A Complete Bibliography of the Journal of Statistical Physics: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 1.09

Title word cross-reference

(d + 1) [BTM^{+}04]. 1 [ABY07, Alb08, BTT07, Hei04, Mas05b, MV03, Mun07, SS09c, SZ03, Too04, WS02a]. 1/2 [Zah02, ABY07, Alb08, DK02, KMSS01, Sla09b]. 2 [Ale01a, AY01, BBL00, BD00, Bir06, BBP02, CN04a, CFN06, Car03, FR00, Fis06a, Fis06b, Fis07, Fis08, FJMR02, FJMR03, Gra09, HL02, HS09b, HK02b, Ito02, Ken04, Kuk04, Luc08a, NH03, Sam05a, Sam05b, Set06]. 2 + 1 [SMT03]. 3 [AC03, Bak06, BSB00, CJS01, Fre06, Rom04, Rom08, SS09e, Sin05]. 3/2 [KCK07]. 4 [Gra09]. _{1} [BFP08]. _{3} [FSG08]. A [Gal06]. A + A → A + S [AFN00]. α [HPWW03]. B_{4} [CM04b]. β = 1 [FM09b, Sin09]. D [CM04c, FH04, TC03, Ber03b, BPS02, BPS03, CM06, FW03, GM07a, HT02, SY03]. d = 1 [BM04c]. D = 2 [RS08]. Δ = \frac{1}{2} [FSZ01]. H [BC04, DMN06, aYL05]. K [ABS01, Kol04, PDV07]. l [Tah09]. L^{q} [AN05]. L_{1} [BIV01]. Z^{2} [DP03]. Z^{d} [Rit06]. N = 4 [GW09]. μ [JH03]. N [Has00, HNS04, HH02, KL06, KL08, O’C01, Wat04, APS09, AVZ00, BJ08, Lan09, OEB01, PIRB03, RdCM04]. O [Wat04]. O(2) [GTZ02]. O(N) [SO02, Ito02, Car06, CSZ00]. p [Fra07b, KM09b, Lou03, Lou08a, MRTZ03, Pan08, Tin03]. p = 1 [BPS02, BPS03]. ϕ^{4} [Hos06, Hol02a]. p ≤ 7 [Häg07]. ± [Fis06a]. ±J [RNVRP04, TPV09, Fis06b, Fis07, Fis08]. Q [AYP07a, EK00, FM09a, JSS05, SvL02, War01]. q > 0 [PS06]. q ≤ 4 [Sea02]. q → 0 [JSS05]. R [BBD08, HKT01, KG00]. R^{2} [BDGSC02]. R^{3} [Sin05]. R^{d} [BP02]. ρ [CPV08]. ρ = 1/2 [Set06]. S [BC02e, Noz04]. sl2 [DFM01]. SLE(κ, ρ) [Kyt06]. SU(1, 1) [BR02]. su(N) [BdGM01]. t′ [LW07, Woj06, LW07, Woj06]. T_{c} [Leg03]. τ_{2}(t_{q}) [Bax04b, Bax06]. U [LW07]. ŝℓ(2) [BP01a]. X [Tor04]. x^{6} + αx^{2} [Suz01]. Y [Tor04]. Z [AV87, AV03, AYP07a, AYP07b]. Z^{2} [CK03b, FF04]. Z^{3} [MS08a]. Z^{d} [BCS07, FK02, PS06]. Z^{d} × Z^{+} [CY09].

-Actions [CK03b]. -Alkane [PIRB03]. -Alkane/Methanol [PIRB03]. -Approach [BIV01]. -Body [HH02]. -Clique [PDV07]. -D [Ale01a, AY01, AC03, BTT07, FJMR02, FJMR03, MV03, SZ03, Too04, WS02a]. -Dependent [AYP07a]. -Dimensional [Ber03b, BBP02, BPS02, BPS03, FW03, GM07a, HT02, Set06, CM04c]. -Dimensions [BTM^{+}04]. -Electron [RdCM04]. -Ferrofluid [GTZ02]. -Invariant [AYP07a, AYP07b]. -Laplacian [Lou08a, Lou03]. -Local [BBD08]. -Matrices [ABS01]. -Model [SvL02]. -Models [CSZ00]. -Nary [Kol04]. -Order [Zah02]. -particle [Lan09, KL06, KL08, O’C01]. -particles [BJ08]. -Percolation [CPV08]. -Point [APS09]. -solutions [AN05]. -Spin [Fra07b, KM09b, MRTZ03, Pan08]. -Spins [Tin03]. -State [CJS01, JSS05]. -Switched [EK00]. -Theorem [BC04]. -Theorems [DMN06]. -Theories [Hol02a]. -Trinomial [War01]. -Vector [AVZ00]. -Wave [Noz04, Tah09].

0-521-41944-1 [Dor02].

1/2 [Cau02b]. 100th [Ano04m, Leb09b]. 101st [Leb09c]. 107 [BPS03]. 108 [Cau02b]. 11 [Dor02]. 117 [Bax06]. 1999 [Mas02a, Pod01b, Pod01a].

2 [Kla04]. 2000 [GCU02, Han02, Opp01, Rit01, Rod02, Wei01, Mas00]. 2001 [Bra02, Hol02a, Opp02a, Pod02b]. 2002 [Dag03b, Kad03b, KK02b, Mas03a, Opp02b, Opp03a, Opp03b, Pod03, Pod04b, Rit03, Shl04, Spo03, Wei02, bA03]. 2003 [Bog04, Dag03a, Leb04b, Nol04, Opp04c, Opp04a, Opp04d, Pod04a, Por04, Rub04, Taq04, Wes04]. 2004 [Leb04a, Leb05b]. 2005 [Leb05c, Leb06b]. 2006 [Leb06c, Leb07b]. 2007 [Leb07c]. 2d [Woj03, BHPH01, KM01, OT03].

3/4 [KK02b]. 3D [Fre03]. 3D-Perturbations [Fre03]. 3rd [Wit03].

47 [AV03].

70th [Ano03m, Ano03n].

80th [Ano03o, Ano04m]. 82nd [Ano00x]. 83rd [Ano00y]. 84th [Ano01u]. 85th [Ano02n, Ano01v]. 86 [Ano01d]. 86th [Ano03m]. 87th [Ano02n]. 88th [Ano03n]. 89th [Ano03o].

90th [Ano04m]. 91st [Leb04a]. 92nd [Leb05b]. 93rd [Leb05c]. 94th [Leb06b]. 95th [Leb06c]. 96th [Leb07b]. 97th [Leb07c]. 98th [Leb08a]. 99th [Leb08b].

= [TC03].

A-D-E [BP01a]. A. [bA09]. A.S.S. [Tal07]. A3 [SB01a]. ABC [ACL^{+}09]. Abelian [ABS01, Pri00]. ABF [JKO^{+}01]. above [HI05]. Abrupt [BBF09]. Absence [BZ08a, MP09, O’C01]. Absolute [DD04]. Absolutely [BSB00]. Absorbing [BFGM02, DM02, SD08]. Abstract [Pro07, Pro09a]. Accelerated [AER05]. Acceleration [BR01]. Accelerator [FGR03]. According [Rei08]. Account [Opp05]. Accuracy [BGGM04, JGB^{+}03]. Acoustic [Ewo01, Naj04, Naj07]. Across [DP04]. Actin [HMO07, JAWC08]. action [Kom07]. Actions [BFKT09, CK03b]. Activation [LB09]. Active [MK07, Ric00, Wes04]. Actomyosin [KC03]. Adaptation [CRVV09]. Adaptive [DSC02, Shl07]. Addendum [Bax06, KP08b, Pro09a]. Additive [CL00, MM05a, Rom08]. Adhesion [KSSM07, Sui00]. Adiabatic [MG02, NS04, NKC00, Pia00]. Adjusted [Rei08]. Adsorbate [KVM^{+}00]. Adsorbates [HM00]. Adsorbing [FT07, MT04, vRR01]. Adsorption [HSA00, OEB01, SS00b, Sud09]. Advanced [Che06, Wei05]. Advection [AG05, Cha09, KMG07, PhD07, YBMS06]. Advective [SKT09]. Advice [Rue02]. Aerodynamic [LLMS02]. Affine [MB03]. against [Bla08]. Agents [Wes04]. Agglomeration [Git07]. Aggregation [FJM01b, MKB00, MM00, Sui00]. Aharonov [dOP08]. Airy [BFP08, BP08b, KT09, MC05, PS02c, Sos00, Wid04]. Akira [Pod03]. AKLT [XKHK08]. Albeverio [Pod07c]. Algebra [DFM01]. Algebraic [BG02, Bax08a, ND08]. Algebras [GI08, MR04, MR05, Nam04]. Algorithm [Cau02a, Cau02b, GW00, Ken02, Ken07, SD05, SD23]. Algorithmical [Too00]. Algorithms [Lia06, MM00, RVMO07]. Aligned [HRS03]. Alkane/Methanol [PIRB03]. All-Atom [SO08]. Alleles [HPK^{+}07]. Alloy [FSG08]. Almost [Blo01, CdO02, FLR03, FFN09, JM00, KL06, KL08]. Along [JS01b, TAL06]. Alphabet [JMU05, Mor07]. Alphabets [MPD09]. Alternate [Cir02]. Alternative [And05, Bao01, Pre07]. Always [IK09a]. among [Fuk08]. Amount [EMC09]. Amplified [MNNS07]. Amplifier [ADLM01]. Amplitude [SS00a, Sea02]. Amplitudes [AKR01, MN02]. Analog [BMA02]. Analogs [BR05b]. Analogues [Kül03]. Analyses [HK02a]. Analysis [AFFS06, AL01, AS03, BCO07, Cai05, Dom03a, EMZ06, FR00, GMW02, HL02, HT02, JW04, JY05, KZ07, LS04a, LR06, OEB01, Paj07, Rom08, ST02, SS09e, Tid01a, Tid01b, Tid04, Yar08]. Analytic [BJ02, CM04b, FZ03, FJM01a, GS03b, JL00b, RK00, Rut01, Suz03, Urr08]. Analytical [BW07a, Lou08b, Rei05]. Analyticity [AS03, BPS02, FPS07, FP04, Mat02, PS06, BPS03]. Analyzing [VB08]. Anchoring [NOB00]. Andersen [Li07]. Anderson [Bel04, BHS07, BK01a, BMP07, Che05, CGK09, CMN06, Eps06, KRT00, KM06, Nak06, Suw09, WZ09]. Angle [Ber03b, Cer06]. Angle-Bending [Ber03b]. Angular [Ale01b, AG09, Fou06, FG08b]. Anharmonic [AVZ00, ALS06b, BLLO09, BCDM07, KP07, KP08b, LS04a, MM05a, RS09, Spo06b, Spo06c]. Animals [Jen01, vRR01]. Anisotropic [HS08, MS02]. Anisotropy [DJ07, FSZ01, NOS05]. Annealed [CS09c, MNP08]. Annihilation [RT08a]. Anniversary [Ano04m]. Annulus [BT04a, Car06]. Anomalies [Man03, SBG03]. Anomalous [Dog00a, DM06c, EK05a, MSVE07, Sam05a, YS05, Wei09]. Anosov [SJW07]. Ansatz [Bax02, MD01, Nep03, PM06]. Ant [Boo01]. Anthology [Opp04d]. Anti [Aus00, BR04, Li06]. Anti-de [Li06]. Anti-Symmetry [BR04]. Antiferromagnet [CJS01, MN00]. Antiferromagnetic [BL08c, JS01a, JSS03, JS06, SS01, SS09b]. Anton [Ton08a]. Anywhere [Fra07a]. Aperiodic [BG01a]. apologies [Lan06]. Application [BEK^{+}07, CSP02, CSN02, Gal06, Jaf03, KP06, Lev03, MNBO09, PM05, SBG03, Tak09a, Yoo07, Bal00, KC03, Mez01]. Applications [AKR01, AM04, AL07, BD04a, BC02c, CKKC00, DR06b, Fre08b, GR06, GC08a, Har08, KN04, KPT05, LLMS02, Nol04, Opp05, Opp06b, Opp06c, Pod07a, Roe03, Ruz00, SSLG^{+}00, Tia04, Yan03, vZC06, Wei02, Wei09]. Applied [BBDR05, HPCH00, ISS07, Mas09a]. Approach [AH03, AMV08, AK02, AS01, BGX01, BR05a, Bar04, BLR08, BCC09, BIV01, Bou09, BZ00a, CT07, CGG06, CL02, CM04a, CP04, CLM05, DMB01, ES09, FJ05, Git08, GAD06, GS07c, GM03a, HKW07, HMO07, Jus01, KS07, Lee05, Lou08b, Mas09b, MV07, Par08, PdKV09, PPR00, Pod08, Pre03, Pre07, PS03b, RR01, Rot06, RF02, Sha01, TC08a, Tao01, WS02a, Zab08, ZQS^{+}05, vK04, dSI04, Pod07b, Sza06]. Approximate [CS02, FM02a]. Approximation [BGGM04, CH06, Fre04, HS09b, JGB^{+}03, KHV00, KRT00, PFK06, Rom04, SW08, Suw09, dARBY04]. Approximations [Cou06, DZ01, GM03b, KMVE03, RT08a, RK00, Tid01a]. Arbitrary [MPD09, PS01, RRS^{+}00]. Arc [CLD05]. Arcsine [Aki08]. Arctic [WMDB02]. Area [Bal00, CZ03, GS00, MC05, MO09a]. Area-Preserving [Bal00, GS00]. Argument [Ric04]. Argumentation [CRVV09]. Arimondo [Shl04]. Arise [CRVV09]. Array [Pro09b]. Arrest [FZS^{+}00, TB04]. Arrhythmias [LKMGG07]. Arrowsmith [KK00b]. Art [Buc05, Yng04]. Artificial [AVBM06]. ascension [ATH06]. ASEP [NAS02, TW08, TW09]. Aspect [Asa00, Fis08]. Aspect-Ratio [Fis08]. Aspects [LKMGG07]. Assignment [Hod01]. Assignments [ALM07]. Assimilation [OT03]. Associated [FM01c, HKT01, KM00, ND08, Shi06a]. Associating [HGST00]. Associative [GL01a, KHV00]. Astigmatism [Bun00]. Asymmetric [AB09, AD02, ALS09, BK08, BB03, BD06c, DLS03, DEL04, DELO05, DP07c, ED04, GM05, Hei04, IS07, JNH^{+}09, KHV00, KK00a, KG00, Lou05, MPS02, MK03, PPH03, RS05, Sam03, SA02, Sch00, Sep01, Set06, Ste04, Uel04, Woj03, ZS02]. Asymptotic [AS01, BFGM02, BCT03, CRT00, CS03a, CS03b, CGV02, Dog00b, EMO08, Els00, FM08, FKLM08, HK02a, HY04, JS01b, JY05, LP08, MVW08, MMR08b, Naj04, NP01, PT04, Tid01a, Var08b, VD05, Ast08]. Asymptotics [ALR07, BC03, BM02b, CMS04, CF03, CCL04, Fuk08, McD01, Piv03, SPVV00, Var08a, Wid04, BM04a]. Asynchronous [RVMO07]. Asynchronously [APL04]. Atmosphere [CD09b]. Atmospheric [CLD05]. Atom [BGX01, FGR03, SO08]. Atomic [OP04, PD05, Yip06]. Atomic-Level [Yip06]. Atoms [ABY07, Alb08, Ano00e, BDM09, BR04, FMS04, Kie09a]. Attached [BD00]. Attenuation [SSE05]. Attracted [DDH08]. Attracting [SLA02]. Attractive [AV03, Bak03, Ess03, FZS^{+}00, HM00, LV02, TTK01, Zha00a, vdHKK02, AV87]. Attractor [CLT07, FKM05, Lem00]. Attractors [MS01]. Atypical [Riv04]. AuCu [FSG08]. Author [Ano00a, Ano01a, Ano02a, Ano03a, Ano04a]. Auto [MNBO09]. Auto-Bispectral [MNBO09]. Autocatalysis [Liu09]. Automata [APL04, CNS08, EK02, EK04, Gác01, HKT01, Kon02a, Kon02b, KM00, MB02, Piv03, TB08, Too08]. Automated [VST09]. Automaton [BKNS01, BF05, SK07]. Autonomous [DMN06, GS01]. Avalanche [PPH03, SSD00]. Average [BZ04, FST06, JR08, Sch00, ZS02]. Averaged [HPWW03, LNM06, Suw09]. Averages [BD04a, FF00, SD06]. Averaging [Car07, Kül01, Lu06, WB04]. Avoiding [CCF^{+}00, CGJ^{+}05, Cau02a, Cau02b, DW05, HT02, Hat07, HvdHS08, Ken02, Ken04, MS08b, SD06, Ghe09]. Avraham [Wei01]. Award [Ano00i, Ano05a]. Axiom [Gal06]. Axis [SM09]. Axisymmetric [GGZD02]. Azimuthal [GGZD02].

B [Mas09a, Rug08]. B.G.K. [Tid04]. B.V [GCU02]. Background [PN03]. Backscattering [GS08a]. Bacteria [JHA07]. Bacteriophage [JHA07]. Bacteriophage-Mediated [JHA07]. Bak [Ban05, MZ02a]. Baker [KS00]. Balance [Cañ07, KP05, Kuk06, LL08b, LL09b]. Ballistic [FJM01b, HP08, Pen01, Pen08]. Band [Bao04, HI04, HI05]. Band-Passing [Bao04]. Bands [HvdHMO09]. Barbara [Lan06]. Bardeen [Hod01]. Barnsley [bA08]. Barrát [bA09]. Barrier [BJ06]. Barrier-Crossing [BJ06]. Barthélemy [bA09]. Baschnagel [Mas02a]. Based [Bar04, CRVV09, DHB07, KR07, SK06, Wei07]. Bases [WNUK01]. Basic [BC02e]. Basins [TNK04]. Basis [DSC02, Hos00, TJ08]. Bassi [Gri00]. Bath [BCL08, BC02b, CN03, Kup04, LPD08]. Baths [BHRW04, HK02a]. Baxter [AYJP01a, Kas01, McC01]. Baxterise [BM01b]. BBGKY [CS08c]. Be [ALM07, CU03, DM09, JHW09, Kad03b]. Beam [Fan05]. BEC [PVZ05]. Becker [DD06, LM02, Sme08]. Becker-Döring [Sme08]. Becker/ [DD06]. Before [BTV09]. Behavior [AFN00, BBF09, BFGM02, BKLO00, BS09b, BFT02, BCM02, CT07, CM02a, CEO07, DG08, DM08, Dog00b, EMO08, Fin03, FW06, GHO^{+}00, GH04, HP08, HvdHS08, JS01b, Kaw03, KAS03, KBS02, KO02, KMS08, LPE04, Lu00, LW04b, LP08, MDB00, Mas05b, MM06b, MMR08b, Naj04, NP01, PZ06a, Por06, Rue02, Sak01, Sak04b, Sak05, SJ02b, SK07, SY03, Sme08, SZ03, Ste04, Tél06b, TBF05, VD05, Zha00b]. Behaviour [AS01, Ben05, BC09, CGI02, Els00, EGGI01, HK02b, MVW08, Ric02, SJ01b, Vel08]. ben [Wei01]. ben-Avraham [Wei01]. Bending [Ber03b]. Benjamin [Dag03b]. Benoît [Wes09]. Berezinskii [GTZ02]. Berlin [Mas02a, Shl04, Wes04]. Bernoulli [AMP04, TW09]. Bertrand [Mas03a]. Bessel [KIK08, Raj09, Sos00]. Bessis [Häg07, LS04b]. Beta [BM04c]. Bethe [Bax02, EE07, FM01a, FM01b, MD01, MS06a, Nep03, PM06, SSD00, SO00a]. Between [Ano01b, Ber03b, CG00b, FZS^{+}00, GPT02, GL04, JS01b, KP01a, LdOP00, LPdO01, LPE04, LR06, MS06b, Nak06, NT04, PIRB03, Rei00, Sak00, Sak06, TV02, ZF09, vEK07, BS07, BL09, RVMO07, Sak07, KG03]. Beyond [ABCM08, Ber07, BM02b, GR00, Hao05, Opp06a, GZG05, CS09c]. BGK [AAP02, BC09, TK02, Zha08]. BGK-Type [AAP02]. Bi [CAD09]. Bi-partite [CAD09]. Bichromatic [OP04]. Bidirectional [GL01a, JNH^{+}09]. Bifurcation [AS06, AN06, EK00, JW04, KMS08]. Bifurcations [AS05, CRV00]. Bilayer [ABY07, Alb08, MS03]. Billiard [BT04a, BM08, Che09a, Che07, LG01, vBD02]. Billiards [BM08, Che06, DP07b, GL04, HS98, LR00, LR02b, MVW08, SL02, TRK03]. Bilocal [CCF^{+}00]. Binary [Bal09, BELM00, BD02, CMSR05, GTA09, GM07a, HD03, Hem00, MR04, MR05, PWC02, PP01, PPR00, SAPD05, ZKD05, BELM06]. Binder [Rap01]. Binding [LL02, WLTH07]. Binomial [BR00b]. Biochemical [KC03, MNNS07]. Biological [Bog04, FFN09, GC08a, Saa07]. Biologists [BA06]. Biology [Wes09, Dag03a, Pod02a]. Biomolecules [Pod07a, SSS07]. Bipartite [Cag02, PM05]. Birkhäuser [Nol04]. Birkhoff [BD04a]. Birkoff [Gal06]. Birth [Ano04m, BBF09, KMS08]. Birthday [Ano02n, Ano03m, Ano03n, Ano03o, Ano04m]. Bispectral [MNBO09]. Bistable [DB00, DB01, GOSG00, MDW09]. Bivariate [DOS01]. Black [Li06, Mag09a]. Bloch [CMS04]. Block [Kül01, XKHK08]. Block-Averaging [Kül01]. Blocking [GS08b]. Blood [OC05]. Blow [Esc07]. Blow-Up [Esc07]. Blue [FMB03]. Blume [BC00, BL08b, GG06, HK02b, MO01, SK07]. BMV [KS08]. Bodies [GT04a]. Body [BDM09, Gal01, HPF^{+}02, HH02, KW00b]. Bogoliubov [BZ00b]. Bohm [dOP08]. Bohmian [DGZ09, GS07b]. Boltzmann [Sha10, Spo06b, Tak10, ADG00, ASB02, Adi04, Ale01b, AG09, Ano00e, Ano00i, Ano05a, AK02, AN00, AN05, AN06, AL07, BGS06, Bar04, BPS06a, BCEP04, BCEP06, BTV09, BCT05, BD06b, BCL08, BHPH01, BC02a, BC02c, BC02d, BCT03, BC03, BGP04, BG06, Cai05, CGR07, CCL09, Cas01, Cas02, Cer05b, Cer06, CD09b, Che05, CS08b, CSP02, CSN02, CM02b, DA03, Erd02, ESY04, EB02, ETB06, FH04, FM01c, Fou06, FG08b, GW00, GM03a, GM03b, HD02, HRK^{+}05, Jia08, JY05, Keh05, KR07, KTH^{+}05, LV01, LT04, Lu00, Lu01, Lu04, Lu05, Lu06, Mas02b, MS02, MMR06, MM06b, NDC02, NSCW04, OC05, PVV02, PhD07, PBFC05, RSMJ^{+}02, RRS^{+}00, Rhe05, SK06, SNC05, Spo06a, Spo06c, SAPD05, SSK02, SFCO02, Tak09a, Tak09b, Tid01b, TKR02, WP02, YHD02, aYL05, Yu06]. Boltzmann [ZQS^{+}05, ZZ06]. Bolyai [Mez01]. Bond [BR06, BPS02, BPS03, FRZB09, PS04b, RS08, Sep01, WP03]. Bonds [HGST00]. Book [Aba00, Ano00c, Ano00d, Ano00b, Ano00e, Ano00f, Ano00g, Ano00h, Ano01b, Ano01c, Bog04, Bra02, Buc05, Cer05a, Dag03b, Dag03a, Dom04, Dor02, GCU02, GC05, Han02, Hol02a, Kad03b, Kan01, Mas00, Mas02a, Mas03a, Mas05a, Mas06, Nol04, Opp01, Opp02a, Opp02b, Opp03a, Opp03b, Opp04c, Opp04a, Opp04d, Opp04b, Opp05, Opp06b, Opp07, Pod01b, Pod01a, Pod02a, Pod02b, Pod03, Pod04b, Pod04a, Pod05, Pod06b, Por04, Rap01, Rit01, Rit03, Rod02, Roe03, Rub04, Rub06a, Sah02, Shl04, Shl05, Spo03, Sza06, Taq04, Wei01, Wei02, Wei05, Wes04, WW05, Wit03, bA03]. Boolean [Cop08, Mec01, ZKBG03a, ZKBG03b]. Boon [Dor02]. Bootstrap [BS09b, Cam05, FS08]. Bordered [VKVT02]. Borders [TNK04]. Borel [PF07]. Born [Cas02, JGB^{+}03, Kie04a]. Bose [BP08a, BZ00b, BNZ02, BD03b, BZ08b, Cso00, DPZ06, GS09, Gra00, JPZ09, LV02, LVZ03, Lee09, Leg03, LS09a, LY01, Lu00, Lu04, Lu05, MV01, OP04, cR04, SKM04, Süt04, Yan03, YY09, Zha00a, Rub04]. Bose-Condensates [Cso00]. Bose-Gas [BZ08b]. Boson [DMP05, PD06]. Bosonic [BFKT09, Got05, Kie09a]. Bosons [BCZ04, Süt03]. Boston [Nol04]. Bottle [CP02]. Bottom [Naj08]. Bouchaud [Mas02a]. Bound [Che07, CE02b, CS09c, LMP05, LL02, Rut08, SO00a, SO00b, YY09]. Bound-Spinons [Rut08]. Boundaries [ALS09, AVBM06, Bah07, BD06c, BCKM00, DDR04, DM02, HD09, JS01c, LMS05, LR00, Lou08b, MT04, MO09b, Tél01, Tid04]. Boundary [Ale01a, AY01, BP01a, BL09, BKM02, CG00b, CAD09, Che09b, CSP02, DELO05, DMB01, DLS09, GW09, HL02, HPWW03, JS08a, JY05, KM01, KM08b, LLSA00, LV02, MMR08a, PS01, SS08, Sab08, SS09a, Sam01a, SJ01a, SS09e, Son09, Sot07, Tid04, WP02, vENS05, vWR05, WDMB05, NOV04]. Boundary-Value [KM08b]. Bounded [AL03, AGL01, BBP02, CMS05, FG08a, JL01]. Bounds [APS09, BR06, BGLT08, CM00, CLT07, FPS07, FL03b, GS07a, Sak07, TV00]. Boussinesq [LW04a]. Bovier [Ton08a]. Bowen [Jia03]. Brain [Mau08]. Branch [CCAD08, Has02]. Branch-Point [CCAD08]. Branched [BI03, KTZ06, SD05, SD23]. Branches [MZ02b]. Branching [BZ09, KZ03b, Mol02, Mol04, Shi09, SD08]. Brankov [Han02]. Bray [GT04b]. Brazil [ATH06, Sai00]. Brazovskii [Shi06b]. Breakage [LW01a]. Breakdown [ADLM01, BEPK^{+}02]. Breaking [AGL01, AR02, BW07a, Fuk00, FW07, GSW07, GR00, JW04, JL01, Luc08a, MPRT^{+}00, Nam04, NS02, SWK04]. Breakup [MVE04]. Breathers [AS01]. Breathing [CLM08]. Brewer [Opp03b]. Brezin [Leb08b]. Bricklayer [Bal04]. Bridges [KIK08]. Brief [Opp05]. Bristol [Roe03]. Broad [KP01a, LdOP00, LPdO01]. Broeck [Kan01]. Broken [BCMP04, De 06, WM07]. Bromberg [Dag03a]. Brownian [CD08, DR05, DEL04, FK00, FVE03, Fel02, FL03a, GJQ06, HR06, HB08, HKW07, KT07, KM09a, Lin08, LM01, MC05, MK04, Ruz00, Shi09, Sin08, Skr03, Sla09a, SW08, Val09b, dG05a, dG05b, Wei02, Wes04]. Brush [Opp04d]. Bryuno [Gen06]. Bucy [MN05]. Building [RF02]. Built [DMP^{+}04]. Bulk [DLR04a, DLR04b, HGST00, PS08a]. Bundles [MK07]. Burgers [ALS06a, AR02, BK03a, CD04, CD09a, CW04, Esc07, Gir03, LRM06, Sim08, Val09b, Val09a, Win02, Yep02]. Burnett [Bob06, Bob08, JS01d].

C [Kan01, Mas09b, Opp02b, Opp04c]. C. [Pod07b]. Cactus [Pre03, Pre07]. Cahn [MPSW00]. Cai [Mas08]. Calculate [FM09b, vZvB02]. Calculation [BD04b, CM04b, FW07, JS04, Sam00]. Calculations [CKKC00, DSC02, FBG02, FRZB09]. Calculus [DW05]. Call [Ano00i]. CAM [SO08]. Cambridge [Aba00, Bra02, Dag03b, Dom04, Dor02, Kan01, Mas00, Mas02a, Mas03a, Opp03a, Pod01b, Pod01a, Pod03, Pod07a, Rap01, Rit01, Rit03, Wei01, Dom04, Pod01a]. ‘Cameo [BK04]. Can [Bak03, BCK00b, DM09, EA07, JHW09, Noz04, DHB07, Fra07a]. Cancer [Kom07]. Cannot [Kad03b]. Canonical [CT04a, CAC02, CETT05, EMZ06, KP01b, Sak00]. Cantor [LdO03]. Capacitors [EK05c]. Capacity [DD09, GR00, SM06]. Capel [BC00, GG06, HK02b, MO01]. Capillary [BLM03, RSMJ^{+}02, dSWP05]. Carbon [BVBF01]. Cardiac [LKMGG07]. Carlo [Ano00c, Rap01, AER05, BLM03, BW07b, BR01, CG00a, FPS00, GC08a, GM00, HRA^{+}04, HGST00, Ken04, Lia06, RT08a, RNVRP04, SD05, SD23, SW08, Vil06a, WS02b]. Cascade [Mol02, Mol04]. Cascades [Arg07, Bak06, CO08a, ZKBG03a]. Cascades.Part [ZKBG03b]. Case [BD06a, BP02, CT04a, Cir02, FT06, Lee09, Pol00, RGRTS06]. Cases [Häg07]. Casimir [BM05b, HB00, SHD03, SK06]. Cassie [DT00]. Castiglione [Opp09]. Catalytic [OBB03, SSK02, ZK00]. Cauchy [AG09, MMR06, NNR09]. Causality [Luc08b]. Cavity [BBLS05, BP09, De 04, GK09a, MP03]. Cayley [GTA09, GS08b, Pat07a, Roz06, Roz08]. CCM [FRZB09]. Celebrating [Ano02n, Ano03m, Ano03n, Ano03o, Ano04m]. Celebration [Sim09]. Cell [DHB07, EE07, Hol07, MV03, YCCN07, LB09, KSSM07]. Cells [HC08]. Cellular [APL04, BKNS01, BF05, CNS08, DHB07, EK02, EK04, Gác01, HKT01, Kon02a, Kon02b, KM00, MNNS07, Piv03, SK07, TB08, Too08]. Censored [DLP09]. Center [Leb07b, Leb07c]. Centers [Nak06]. Central [CC09, Shi09, WW01, HMH05]. Centrally [dBR08]. Cercignani [GC08a]. Certain [CB04, Fuk00, HJ05, HO09, Sos02, Win02]. Chaichian [Roe03]. Chain [AP03, BDM09, BD00, Ber08, FK04, FK05, FSZ01, HWB01, Jac09, JK04, KMSS01, Mez01, Nep03, OBB03, PAY09, PFK06, Rut08, VB08, Yar08, BA01].

Chain-Connections [FK05]. Chains [ALS06b, BS08, BBF09, DVE07, DR06b, EEDJ00, FM05b, GJQ06, LDNG08, LS04a, PD05, Rác00, SC09, WM07, FM05a]. Chandra [FEDZ07]. Changes [CPC03, LNM06]. Channel [PS03a]. Channels [DD09, EE07, SM06]. Chaos [CP04, CG00c, DC00, DC01, Gal09, GG00, KS00, Pod01a, RT08a, Ric00, ZK00, Opp09]. Chaotic [AR00, AG05, BGGZ06, BCMP04, GFH08, GPM^{+}05, JBKN01, JGB^{+}03, LR00, Mac02, Mas07, WS02a, vER07, bA03]. Chapman [Bob06, HD09]. Character [FG00, War01]. Characterization [BZ09, CP04, dFCS06, DB01, LZ04, SS00b]. Characterizations [GK06]. Charge [BVBF01, DJ07, FM08, JS01b, Jan03, JT04, JS08b, LFB04, LWL00, Sam03, Sam06, Sam09, Tél06b]. Charge-Asymmetric [Sam03]. Charge-Stripe [DJ07]. Charged [AR02, HRS03, MNP08, SLA02]. Charges [BCD^{+}04, Sam05a, Tél06b]. Chazottes [Rug08]. Chemical [Ano01c, BD05, CW06, DDB^{+}00, GP04, Kap06, Mal05, Buc07]. Chemically [SSS07]. Chemisorption [EC07]. Chemistry [Dag03a]. Chemists [BA06, Dag03b]. Chemomechanical [LL08b, LL09b]. Chiral [AYJP01a, Bax00, Bax03, Bax05, Bax08b, Bax09, FM08, ND08]. Choice [CS02]. Chordal [Ken07]. Chromatic [JS01a, JSS03, JS08a, SS01]. Circadian [GG00]. Circuits [BMN07]. Circular [BM01a, SSH06a]. Clarendon [Wei02]. Class [BS07, CEMM09, CC00b, DP03, Fou06, GTW01, Kon02a, KMS08, LTWW02, MP09, MS08b, Sak09, TB08, TV02]. Classic [Opp04d]. Classical [AFNvM00, AG09, AC01a, AC01b, Ban05, CR01, Cai04, CDG06, CNV01, DD09, FL03a, Fin03, GL04, Grm08, JS08b, Kie09b, Kie09c, KBS02, LPD08, LMV07, NKC00, PS02a, RC01, RS09, Sam07, SJW07, SO00a, SO00b, SO02, Skr00, Ste04, SST^{+}00, Swe02, Zha08]. Classically [BCMP04]. Classification [BC01, BZ04, BB05, FS07a]. Classify [Cop08]. Clausius [BM01a]. Clerk [Dom04]. Climate [ER00]. Clique [ISS07, PDV07]. Clogging [BH08b]. Close [BCO07, PS00, Rad08a]. Closure [KMVE03]. Clotting [OC05]. Cluster [ABCM03, AV08, Asa00, BC00, BZ00a, Cam01, CK03a, CL06, Dom03a, FBG02, FRZB09, GG02, GG06, HS00, Hei04, PFK06, PS06, vdHJ04]. Clustering [APS09, CM04a, FM07, Gir01, PA00, PA02a, Sal09, SB06]. Clusters [Arg02, BS09a, CZ03, CY09, vDHR06, Pod07a, PM04, Win02, vdHK08]. Co [Opp04c, Opp04d]. Coagulating [HR06]. Coagulation [Cañ07, FG03, GE09, Kol04, LW01a, Nor09]. Coalescence [HNS08]. Coalescing [McD01]. Coarse [BR01, JAWC08, KT06, Opp09]. Coarse-Grained [JAWC08]. Coarse-Graining [KT06]. Coarse-to-Fine [BR01]. Coarsening [NP01]. Codes [PZ06a]. Coding [Sch01]. Coefficient [GJ08c, Lin08, Lou05, Lyb05, Sch00, Tas06]. Coefficients [CdlL05, CS02, CM04c, CM06, Coh09, Coh10, Con04, ECSB00, GA05, GM07a, HJ05, LDNG08, Len00, PS02b, War01]. Coexistence [AB09, BCK00a, DMPV09, EA07, Liu09, PH01]. Coexisting [Rei00]. Cohen [Ano03o, LRB00]. Coherence [Ric00, Aus00]. Coherent [AFHV03, DM02, MPD09, RdCM04, RSV09]. Coins [BM04b]. Colby [Pod04a]. coli} [KMP^{+}07]. Collapse [DH04, Ess03, MNP08, MPT06]. Collapsed [OSv04]. Collapsing [vRR01]. Collatz [RF02]. Collection [Pod07c]. Collective [ES03, Gra00, SBSAD07]. Collide [KIK08]. Colliding [Bou07, ZKBG03a, ZKBG03b]. Colligative [ABC05a, ABC05b]. Collision [Pey09, PhD07]. Collisional [LW01a, Spo06a]. Collisions [Che09a, DR01, Pia01]. Colloidal [SHD03]. Colloids [Git09, Pod05, SAPD05]. Colors [Jan01b, Nie01]. Combinatorial [BCO04, DMB01, HKT01]. Comblike [Zah02]. Combustion [YHD02]. Cometary [FPS04, FS07a, HS09a]. Coming [Sin02]. Commensurate [Yar08]. Comment [BL04, KP01a, LPdO01]. Commentary [Opp04d]. Comments [ER02, Sha10, Tak10]. Common [AMV08]. Communication [Bog09]. Communications [Paj07]. Community [EA07]. Commuting [Dub06]. Compact [Kas01, Par08, Xin09]. Compactness [Lu06]. Comparative [KMVE03]. Comparison [AC03, CG00b, CT04a, GL04, KP01a, LdOP00, LPdO01, ZGA02]. Comparisons [GP04]. Compartments [KL03]. Compass [CO08b]. Compensation [ABY07, Alb08]. Competing [GTA09, MO01, MR04, MR05]. Competition [FZS^{+}00, JHA07]. Complete [Ada01, CY09, EHT00, FM05b, GK06, KM09b]. Completely [FV02]. Completeness [Bax02]. Completing [FM01b]. Complex [dFCS06, GR09, KMVE03, Mac09, Por04, SMS09, Shl05, Shl07, Son09, VB07, bA09]. Complexities [Riv04]. Complexity [Dud07, FW07, Kad03b, Mas00, MM00, SSB04]. Component [AR02, BZ08a, CG00a, FJT03, FT08, FT07, FJM01a, Jan00b, KMST00, LWL00, MT04, ST00, Sam01b, SJ02b, Sam04, SWK04, TM02]. Components [Jan06b, PhD07]. Composite [BM01a, GMW02]. Composites [BM05a]. Compressibility [Rei05, RK05]. Compressible [CG03, GH04, KK02b, KK02c, Mas02b]. Computable [BZ03]. Computation [APL04, BMN09, HL02]. Computational [Jan06a, SC01, Shl07, Suz03, WE07]. Computations [AC03, Arg07, Ken09]. Computer [BCFP05, HPCH00]. Computing [GK09a, Ken08, Rad08b]. Concave [IK09a]. Concavity [DT00, Han06, PG03]. Concentration [Lu05, PdKV09]. Concentrations [ABC05a, ABC05b]. Conceptual [GMW02]. Concise [Dag03b]. Condensates [Cso00, Gra00, MV01, Zha00a]. Condensation [AR02, BLM03, BGH01, BNZ02, DPZ06, EMZ06, FLS07, GSS03, GS08c, GS08d, JPZ09, LVZ03, Leg03, NH03, SM09, SH07, Süt03, Süt04, TAL06, Yan03, Rub04]. Condensations [BZ00b]. Condense [BCZ04]. Condensed [Han02, Opp06b, Rit03, TAL06]. Condition [DN04, DG09b, FW07, IS07, TW09, vENS05]. Conditional [GY05]. Conditioned [KIK08]. Conditions [Ace00, AL03, Ale01a, AY01, BKLO00, BKOS07, BP01a, BLP09, BL09, BKM02, CG00b, Che09b, CSP02, Der08, DELO05, GW09, JY05, KM01, LV02, LL08b, LL09b, NOV04, PS01, PS04b, SDC04, Tid04, WP02, WDMB05].

Conductance [AR00, Kom00a, Kom08]. Conductances [Mat08]. Conduction [BO05, BLLO09, LL00, MNV03]. Conductivity [BM05a, Ber08, FK02, MN04]. Conductor [BCD^{+}04, LB04, LB04]. Conductor/ [LB04]. Conference [Leb07b, Leb07c, Leb08b, Leb08a, Leb09b, Leb09c]. Conference/ [Leb09b]. Configuration [BFPS07, CPC03]. Configurational [KTZ06]. Configurations [BS09a, MNOS04]. Confined [AB09, BMD00, DDB^{+}00, DTY02, KP06, LK09, MT04, MPT06]. Confining [BC09]. Conformal [BLZ01, BP01a, CP02, Dup03, Ken04, Kyt06, LLSA00, Raj09]. Congress [Leb04b]. Conifer [VKVT02]. Conjecture [Bax08b, BB09, Cag02, FLW08, Häg07, JK04, KS08, LS04b]. Conjectures [FKLM08]. Conjugacies [BJR02]. Conjugate [Ban05]. Connected [HL08]. Connectedness [GS04]. Connection [Att00, Ler00]. Connections [BS07, FM05b, FK05, FBC^{+}05, vEK07]. Connectivity [BDGSC02, BCS07, Sak04b, Sak05]. Conservation [KMG07, LR02a, MR01, Mür01, NOV04, Tid04, TV03, WX06]. Conservative [FS06, FNO06, MT08, NOS05, TG00]. Conserved [KK08, NP01, RM00, RS04a]. Conserving [PBFC05]. Considerations [BCEP04, BCEP06, Rei05, TT05]. Consistency [KP01b]. Consistent [AAP02, BG00b, BLL04, BLLO09, GW00, Gri00, HD03, Jac09, KRT00, Rhe05]. Consisting [ABY07, Alb08]. Constant [Bur08, Pan02, PH01]. Constitutive [HS08]. Constrained [AD02, CDA09, Har08, TB04, TBF05]. Constraint [MZ08, Sem08]. Constraints [YS05]. Construction [BV08, BIV01, FPSW01, JKO^{+}01, Kie09c, Rit06, TM08]. Constructive [DZ01, FG00, Roz06]. Contact [Ban05, CY09, DA03, FVE03, Sak01, Sak02, Süt02]. Contact-Interacting [FVE03]. Contained [BDG^{+}09]. Containing [VB07]. Context [CD09b]. Continuity [Ace00, AOT06, BJ02, DD04]. Continuous [BZ08a, BSB00, BDM09, BK02, CT04a, CP01, Cor04, FLM00, GR00, KPT05, LdO03, OT03, Ruz00, Too00, Too07]. Continuously [LD01]. Continuum [BBD04, BDGSC02, CN04a, DMPV08, DMPV09, EY01, FBC^{+}05, HPCH00, KZ07, Lem00, Pen01]. Contour [Roz08]. Contracting [FST06, JR08]. Contractive [BCT05]. Contributions [Ano00l, Ano00o, Ano00r, Ano00u, Ano01g, Ano01j, Ano01m, Ano01p, Ano02d, Ano02g, Ano02j, Ano02m, Ano03c, Ano03f, Ano03i, Ano03l, Ano04d, Ano04g, Ano04i, Ano04l, Ano05d, Ano05g, Riv02, Ano00j, Ano00k, Ano00m, Ano00n, Ano00p, Ano00q, Ano00s, Ano00t, Ano01e, Ano01f, Ano01h, Ano01i, Ano01k, Ano01l, Ano01n, Ano01o, Ano02b, Ano02c, Ano02e, Ano02f, Ano02h, Ano02i, Ano02k, Ano02l, Ano03b, Ano03d, Ano03e, Ano03g, Ano03h, Ano03j, Ano03k, Ano04b, Ano04c, Ano04e, Ano04f, Ano04h, Ano04j, Ano04k, Ano05b, Ano05c, Ano05e, Ano05f]. Control [Mas09a, Pan08, Ric00, SLB00, WS02a]. Controversial [GPL02]. Convection [FN00, Fre06, Lou07]. Convection-reaction-diffusion [Lou07]. Convergence [ADG00, Bah07, BBF09, BF05, BT04b, Blo01, BZ00a, Cañ07, CL03, CCL09, CY09, DN03, DKRS02, Fan04, FPS04, FFN09, GM03b, Li07, Lu05, NNR09, Pey09, Pre05, SW00]. Convergent [Riv02]. Conversion [LLV09]. Convex [JM00, WX06]. Convexity [Blo04]. Cooling [BCT06, MMR06, MM06b]. Cooperative [HSA00, TBF05]. Coordinated [CN01]. Coordinates [ES03]. Copolymer [dHP09]. Copolymers [BG04b, BGLT08, CGG06, CR03]. Copulas [Mas09b]. Core [CR01, Cai04, FHL03, RC01, RS04b, Sam06]. Cores [BP08a]. Corners [TRK03]. Correction [Adi04, AV03, CGJ^{+}05, RT08b, SD23]. Correction-to-Scaling [CGJ^{+}05]. Corrections [CJS01, CSM03, HNO05a, HNO05b, SPVV00, TT05]. Correlated [DD09, LS09a, MDW09, Mes02a, RS08, Tia04, Woj03]. Correlation [BM04a, BHS07, BI03, CS08c, FEDZ07, FM09b, JH03, KW00a, KW02, MM05b, Nag07, PAY09, PG99, PG00, Sam07, Sin09, TTK01, Var08a, ZF09]. Correlations [APS09, BM08, BDLvW08, BCS09, Bud08, BMD00, Che07, CE04, DTP02, DM09, JS01b, KMST00, LFB04, LY07, Lo08, LZ09, Mas00, NOS06, PS01, SJ02a, SJ02b, Sch08]. Corresponding [MR04, MR05]. Corrigendum [KP08b]. Cortex [LK06a]. Coset [FPSW01]. Cosmic [Rub06a]. Couette [AN05, AN06, Gho05, SMRK^{+}02, TTM^{+}01]. Coulomb [AC01a, AC01b, AC04, BMA02, CG00b, CR01, Cai04, Jan00a, JS01b, JS01c, Jan03, JT04, JS04, JS08b, KS02, LS00b, MT04, MM05b, RC01, Sam00, SJ02a, Sam03, Sam05b, Sam07, Tél01, Tél07, TT05]. Coulombic [EN03]. Countable [BI06, JMU05, Mor07]. Countable-Alphabet [Mor07]. Counter [Mon04]. Counter-Example [Mon04]. Counting [CGK09]. Coupled [AE08, BL08a, BBC^{+}01, Erd02, FBG02, FRZB09, FL03a, FW06, JW00, Jus01, LPD08, LLV09, MSVE07, MV01, Rug08, Sam09, SJ01b]. Coupling [AC01b, FV02, GvH07, HD03, Hos06, KK02a, LL08b, LL09b, Mal07, Rhe05, Rot06, Tid01b]. Couplings [BG01a, KP08a, SZ03]. Covariance [Sos02]. Coverings [CC08]. Crash [Por04]. Crawling [MV03]. Created [MM08]. Creation [GS03a]. Critical [ABY07, Alb08, BBK08, BR06, BZ09, Bjö09, BL08c, CN04a, Cam05, CFN06, CJM09, Cam06, Car09, CEO07, DRC04, EMO08, Els00, dSFM02, GD06, Git08, GS01, Han02, Hao05, vDHR06, KS00, KAS03, KBS02, KKS05, LMP05, LPE04, LW04b, Mai03, Man03, MNOS04, MN02, Mis06, MS08a, MM05b, MN00, PDV07, Pod08, Por04, PIRB03, Pri00, Sak01, SS00a, SSLG^{+}00, SHD03, SK07, SY03, SS09d, SDC04, Ste04, Ton07, ZKBG03b, Zha00b, dlLV07, vdHKK02, BL09, HS00, Hol02a]. Critical-Exponent [SS09d]. Criticality [BCO07, BCK00b, CBKM04, KR06, Liu09]. Criticism [Swe04]. Cross [AG09, PIRB03, VTM00, ZF09]. Cross-correlation [ZF09]. Cross-Over [PIRB03]. Crossing [BJ06, KZ03a, Mai03, PH01, SM08]. Crossover [Car03, KAS03, KMSS01, LPE04, PWC02]. Crow [MPD09, PD06]. Crystal [AVZ00, BLL04, EY01, FS03, Luc09b, MS02, NH03]. Crystallization [FY06]. Crystals [BLLO09, DTY02, DKM04, FNP07, Sak04a]. Cube [CLS02]. Cubes [HHB09]. Cubic [AGT07, Dys04, LDNG08, Luc09b, Rid03, SK07, Sum05, Zha00b]. Cubic-Plus-Quartic [LDNG08]. Cumulant [SH05]. Cumulants [BMN09]. Cuprates [RR01]. Curie [CCIL08]. Curious [BL04]. Current [BMN09, BS07, BK08, BDG^{+}06, BD06c, DDR04, DG09b, DG09a, DP07c, MRV02, MH07, Nam04, RS05, Sam09]. Currents [AG07, CCMT09, DDM08, HJ09, VTM00]. Curvature [KK08, Mec01]. Curved [MAOB01, Xin09]. Curves [FF07, Ken08]. Curvilinear [SO08]. Cusps [BM08]. Cut [Cau02a, Cau02b]. Cut-and-Permute [Cau02a, Cau02b]. Cutoff [Ale01b, Cer06, Fou00, FM01c, Fou06, SW08]. Cycle [All08, MDB00]. Cycles [BP08a, DMP05, GRU07, GG00, LJ03, RF02, Tur03]. Cyclic [SF07a]. Cylinder [SWK04]. Cylindrical [SWF07]. Cytoplasm [Hol07]. Cytotoxic [CGS07].

D [Ano03o, BP01a, Han02, Rap01, Rit03, Sin05, Car03, Ale01a, AY01, AC03, Bak06, BTT07, BSB00, BBL00, BD00, Bir06, CN04a, CFN06, FR00, Fis06a, Fis06b, Fis07, Fis08, FJMR02, FJMR03, Fre06, HL02, HS09b, Hei04, HK02b, Ito02, Ken04, Kuk04, Luc08a, Mas05b, MV03, Mun07, NH03, Rom04, Rom08, Sam05a, Sam05b, SS09c, SS09e, SZ03, Too04, WS02a]. D-Fluid [Fre06]. Damping [MG02]. Danchev [Han02]. Daniel [Opp04c, Wei01]. Dark [GG00]. Data [AL01, Dom03a, GOBY06, Jia07, OT03, Sim08]. Dauxois [Shl04]. David [Mau08, Sin02]. Dawn [Pru01]. Days [Ano00z]. Dead [SS00b]. Death [BBF09]. Debye [BMA02, HRS03, TT05]. Debye-Hückel [TT05]. Decay [AV08, BM08, BG02, BCO04, BCT06, BCS07, BCS09, CE04, DMPV08, EX00, GJY04, HMPPMV00, LL09a, Lo08, ND03, PS01, SO00a, Var08a, VMT02]. December [Leb05b, Leb06b, Leb07b]. Decoherence [AE08, Amb06, Pri03]. Decoherent [BG00c]. Decomposition [BGM04, MPSW00]. Decompositions [Jan06b]. Decorrelation [Kra03]. Decrease [BM05a]. Defects [CLK09, MM08, Rit03]. Defined [FKÖ03, HJ05]. Definite [HJ05]. Definition [CCG04, Swe04]. Deformed [HI04]. Degeneracies [GM05]. Degenerate [BT04b, CK00, Rom04]. Degree [GGL09]. Degrees [DRL02, FJMR03, HDMF04, JGB^{+}03, MM02]. Delaunay [BBD08, Der08]. Delay [MDW09, ZKBG03a, ZKBG03b]. Demagnetization [HMG08]. Dembski [Kad03b]. Demichev [Roe03]. Demixing [Hem00, PWC02]. Demonstrated [HDMF04]. Denaturation [Dom05, RG04]. Dendrites [AC03]. Dense [Coh09, Coh10, EN03, Elo08, GW00]. Densities [ALS03, BW07b, BS03, HMH05, RS04a, Sos03]. Density [AC01a, AC01b, AC04, Att00, BG01b, BCEP06, BMD00, CAC02, CC09, CLM07, CR03, Coh09, Coh10, CE02a, DJZ00, DLS02, DG09a, DOS01, DS07, FFN09, GM07a, GPT02, GS09, HDMF04, KS02, KD00, KPvB06, LW01b, Mac09, McD01, MNBO09, Naj04, PZ06a, RS04b, SJ02a, Sam07, SY08, Set06, Tél06b, TTM^{+}01, Wee03, XKHK08]. Density-Dependent [KD00]. Density-Profile [FFN09]. Dependence [ES04, HD03, Kur03, vER07, Nol04]. Dependent [AYJP01b, AYP07a, AYP07b, BGGM04, BR05a, FN00, HDMF04, KD00, LK06b, LR00, LR02b, MS09, PVV02, vWL00]. Depending [Zab06]. Depinning [Pro09b, Ton07, VD01]. Deposition [BS07, Pen01, Pen08, SWK09, TV02]. Depth [BCF^{+}09]. Derivation [ASB02, AGT07, BELM00, BCEP04, BCEP06, BET00, CR03, DR01, Jar00, ND08, Saa07, Tid01b]. Derivatives [FM02b, FKM05]. Derived [BDL06, HMH05]. Describing [FFN09]. Description [AS08, BDG^{+}09, CRT00, DN04, EK05b, FVE03, KZ03b, Roz06]. Descriptions [Lem00]. Design [Kad03b, Mat04]. Destruction [GS01]. Detachments [MK03]. Detailed [Cañ07, Jar00, LDNG08]. Determinant [TW08]. Determinantal [GY05, KT07, Sos00, BS03]. Determinants [BJR02, BLZ01, JK04]. Determination [KK00a, PP01]. Determined [DM09, Fei03]. Determining [KW00a, OT03]. Determinism [LS00a, LS02]. Deterministic [Bla03, Bla05, dBR08, BD05, Cas01, CEMM09, EMMZ06, EGGI01, FGC09, FFN09, Fre03, GPL02, LSSW03, Mür01, RKN00, SD06, Wag00, bA03, Bla08]. Developed [BB09, FJMR02, MDL^{+}03]. Development [FK09a, ICO02, Yu06]. Developments [AM04, Dro00, FM03, FM05a, ONGP01]. Deviation [BBDR05, CLM05, DLS02, DLS03, EHT00, ED04, EMH04, LS00a, LS02, Tin03]. Deviations [AP06, BH08a, BGL05, BD06c, BL08a, BET00, BZ08b, CL00, EK02, EK04, GLM02, LRB05, NR04, Shi06a, Tal07, vWR05]. Diagram [ACL^{+}09, BGOY04, BGLT08, CCIL08, WD01, Woj06]. Diagrams [CM04c, LW07]. Diamagnetism [BD06a]. Diatomic [CK04]. Diblock [CR03]. Dicke [PVZ05]. Dickman [Aba00]. Didier [Por04]. Dielectric [AC01a, AC01b, BEPK^{+}02, HS08, JS01c, JS04, Sam00, Sam01a, Tél01]. Dielectrics [HB00]. Diffeomorphisms [GS00, LSSW03, Wol06]. Differences [ZW04]. Different [Kra03, Lem00, MZ02b, NSCW04]. Differentiability [PSW04]. Differential [Ach08, Cha09, CSN02, DRL02, Gar06, GvH07, HD09, HNS08, JL00a, VEW08]. Diffraction [BH00, BZ08a]. Diffusing [SWF07]. Diffusion [AFN00, ADLM01, BTT07, BR05a, Bar06a, BD06b, BL06, BFP08, BCL^{+}07, CT07, CH06, CDG06, CG00c, Cor05, DMS02, DDB^{+}00, DMR05, DC00, DC01, DM02, Dog00a, DM06c, DG01, EG05, FK00, GJ08a, GJ08c, HMPPMV00, KP06, KS09, KNP07, KD00, Kra03, Kra09, KMP^{+}07, Lin08, LPE04, LK09, Lou03, Lou08a, Lou05, MKB00, MM00, PhD07, PSG06, QQT02, SM08, Sch00, TB04, VB07, dHW04, Lou07, Wei01]. Diffusion-Limited [AFN00, LK09, MM00]. Diffusions [BMW09, CG09, DR05, HP04, Mas09a, PS07a]. Diffusive [Bal00, BDG^{+}09, BCT05, Bla08, BDLvW08, CIS01, DLS03, DG09a, HP08, HS09a, HJ09, MK07, PS03a, RS04a]. Diffusively [JW00]. Diffusivity [BK08]. Dill [Dag03a]. Dilogarithm [Kas01]. Dilute [BBP02, DG08, Lee09, LY01, Pol00, cR04, SB01a, ZJ00, vBD02, vZvB02].

Dilute-Gas [Pol00]. Diluted [CSZ00, De 04, FL03b, GT04b, MRTZ03]. DIMACS [Leb09b]. Dimension [AGT07, BD09, CM02a, Che05, CLT07, DMS02, DTP02, DKK02, DJW07, ECSB00, Gal06, GKR02, GT02, Hir05, HV01, JR08, KKT00, Kon02a, LW04b, MS01, Pri00, RM00, Sim08, SKM04, vdHKK02]. Dimension-Splitting [DMS02]. Dimensional [AFN00, Ada01, AGL01, Akt01, AD02, BH08b, Ber03b, BCD^{+}04, BD03a, BK01a, BDL08, BBC^{+}01, BFT02, BBP02, BCFP05, BET00, BPS02, BPS03, BKL00, BR04, BI03, BK03b, BC02e, Cam05, CGJ^{+}05, Car09, CL07, CM04c, CS04, CS09d, DF02, DJZ00, DDR04, DG09b, DG09a, DM02, DLS09, DFF02, Dun03, FW03, FT08, FT07, FLM00, FJM01a, Fou00, Gal01, GN08, GM07a, Gat00, GS04, GS03b, Gir01, GS00, Got05, GT04a, GSW07, GTZ02, Gry01, GM00, Hao05, HS00, HT02, HvdHMO09, HI04, HPF^{+}02, Jan00b, JS08b, JL00b, KMST00, KS02, KP06, KAS03, KW00b, KMP^{+}07, KPvB06, KP05, Kuk06, KKS05, LRY05, LW07, LFB04, LWL00, LY01, LL00, Lou08b, Luc09b, Mai03, Mal07, MB03, MvWH^{+}01, MO09a, MSVE07, MT04, Nis07, PDvB00]. Dimensional [PS08b, PD05, PS04a, RS04b, Ric02, Riv02, Rom04, Rut01, Sak04a, Sak06, Sak07, SS00a, ST00, Sam01b, SJ02a, SJ02b, Sam03, Sam04, Sam06, SI04, SY03, Süt04, Tao01, Tél01, TM02, Tél07, UC07, WMDB02, WK00, Zhi00, vdHK01, vdHJ04, CJM09, Che09b, CEMM09, DR05, HNOV07, Set06, SKT09, Woj06]. Dimensionality [YFP00]. Dimensions [AH01, BTM^{+}04, BR06, BW07b, CMS05, Cha00, CM04b, CM06, DJ05, DKRS02, Dys05, Gra09, HK01, Hor06, Hos06, Kol08, Lyb05, Man03, Par08, PBFC05, Pro09b, cR04, STV02, SMT03, Urr08, Wat04]. Dimer [Bou07, CC08, DG00, GS08b]. Dimers [CdO02, TW03]. Diode [SJ01c]. Dipole [Dim09, JS01b, KCK07, OP04]. Dipoles [FL03a, Giu09]. Dirac [Lu01, Lu06]. Direct [DA03, KW00a]. Directed [BP08b, HJDRD00, Jan01b, LW04b, Lüb06, OEB01, Ton07, vRR01]. Directional [HGST00]. Dirichlet [BLS04, GF08, KIK08]. Discharge [CLD05]. Discontinuity [CSZ00]. Discontinuous [GS08c, GS08d, TB08, PhD07]. Discrete [Ast08, BZ08a, Ban05, BDM09, BZ03, BCD^{+}04, BTV09, BV08, Cañ07, Car09, CC00a, CC00b, Cor04, FF07, FH04, Fuk00, GJQ06, GTW01, KKT00, LW01a, MCZ06, MZ02a, PS08b, PM06, Pro09b, RRS^{+}00, TKR02, TK02, ALR07]. Discrete-Time [KKT00]. Disk [EFPZ05, JT04, RM08, SSH06a, Tél01, Wag00, vZvB02]. Disks [HPF^{+}02, SST^{+}00]. Disorder [DH08, DT00, FR00, GN08, HMG08, Kom00a, MNP08, PD05, RS08, TPV09]. Disordered [Ber08, BCO07, BL09, CS08b, CG04, DMPV09, De 07, DN03, DM06b, DPZ06, JPZ09, Kül03, KM08b, LM06, Lou08b, Pod06b, Rit01, SCM04, SSE05, Wei01, Zhi00, vBD02, Ton08a, BCK00a]. Disordered/ [BCK00a]. Disorders [Che05]. Dispersal [EA07]. Dispersal-Limitation [EA07]. Dispersing [BM08, Che06]. Dispersion [BP01b, BGL03, Cha09, CGH^{+}03, DKK02, Luc09a, Mat01, YBMS06]. Dispersive [FS05, FS07b]. Displacement [AVZ00, Blo04, SS08, dSWP05]. Displacements [SS09e, vZC06]. Dissipation [Amb06, AEGS04, BDG^{+}04, FW03, GS03a, JH03, LOV04, LK06b, LD01, MSVE07, Nag03, SC09]. Dissipative [BD06a, EL02, EEDJ00, Far05, FS05, FS07b, LW04a, LT04, Mat02, RdCM04, TV00, TE00]. Distance [ALM07, BP01b, PZ06a, SJ02b, Suw09, Tél06b, vdEvdHH08]. Distant [Fei03]. Distinguishable [Nag04, Swe02]. Distributed [CK06, GP08, PS07b]. Distribution [Ace00, BG01a, BG00a, BM00, CZ03, CAC02, DKRS02, FF00, GLTZ06, Has01, Has02, HNOV07, LB01, Led00, Luo01, MC05, MBL08, MO09a, RS05, Ram07, SSD00, Sos02, SST^{+}00, WM07, WK00, vZC06]. Distributional [AG09, Lu04]. Distributions [BR00a, BNZ07, BP07, BR00b, CRVV09, DN04, Fis06b, GMNT05, Man03, MT08, Saa07]. Divergence [HH02, Kie04a, Kie04b]. Divergences [VMT02]. Diversity [Eps06, Gry01]. Divisible [Mag09b]. DLA [BH08b]. DMRG [NAS02]. DNA [Dom05, Hol07, RG04]. Do [BCZ04, GvB02, Sin02]. Dobrushin [Sch06]. Does [Adi04, Boo01, O’C06, Tas06]. Dogs [HNS04]. Domain [Bal01, BK03a, BL09, DTY02, FHAY06, Fis06b, Fis08, GT08, HI04, HI05, NP01, SA02, SH05]. Domains [HL08, LLSA00, MVW08, SSH06b, Tid04, VEW08]. Domany [CN04b, KK00b]. Dominance [CM04c]. Dominate [GvB02]. Dominicis [Pod07b]. Dorfman [Pod01a]. Döring [DD06, LM02, Sme08]. Dorogovtsev [Bog04]. Dots [Jac09, Sha06]. Dotsenko [Rit01]. Double [DB00, KIK08, Sac05]. Double-well [Sac05]. Doubling [KKS05, KMS08]. Doubly [BBS03, HL08]. Doubly-Connected [HL08]. Doukhan [Nol04]. Down [Els00, Hao05, NH03]. Dr [Wit03]. Drift [BCV00, DMR05, KO02, Pia01, PSG06]. Drift-Diffusion [DMR05]. Driven [AH03, ADLM01, BNZ07, BDG^{+}09, BCL08, BCC06, Bla08, CGGP04, CDG06, DLS03, DLS08, ETB06, Gar01, GZ01, GG00, GSW07, IVZ06, JOP06, KMSS01, KM08b, MS05a, MM05a, MB03, MS03, MK07, PDvB00, PS03a, PSG06, RS04a, RK00, Rom08, SJW07, SSS07, SC09, Sid05, Sin08, Sot07, VPB^{+}06, vWR05, EK03]. Driving [Dag03a, Ken08, Lin04]. Drop [GLMSR07]. Droplet [BCK04b, DP03, NH03, PS02c, Rut01]. Droplets [DD06]. Dry [dG05a, dG05b]. Dual [Has01, Sla09b, Yoo07]. Dualities [Kon02a]. Duality [ABS01, Bor02, GKRV09, KW02, Kon02b, Nis07]. Due [BK01a, LFB04, MH07, BM05a]. Dummies [Sha06]. Dust [CD09b]. Dynamic [Ano01c, DFF02, HD03, KCK07, SBG03, WLTH07, YS05]. Dynamical [Adi04, Aki08, AFHV03, AP08, BJR02, BCK00b, Car07, CE04, DJ05, DG00, FF04, FK05, FS07b, dSFM02, FKS08, Gas04, Gas07, GI08, GK02, GSS03, HMS05, Hor06, HAG02, Jia03, Kaw03, Len00, Ler00, MN04, MO01, NS04, TB04, Tur03, Tur06, VTM00, You02, bA03, bA09]. Dynamics [AFN00, Ano01b, AL07, BMW09, BJ08, BCF^{+}09, BBL00, Ber05, BTV09, BGH01, BL08a, Bog07, BM02b, Buc05, CCF^{+}00, CDG06, CMS05, CL02, CLS02, CS08c, CN03, CK06, CE05, CDA09, Cso00, DM08, DGM07a, DGM07b, DLP09, EL02, EEDJ00, Erd02, ES09, EGGI01, ER00, FF07, FPD01, FPS00, FN00, Fra07b, GD06, Gar06, Gar01, GAV00, GFH08, GPM^{+}05, GE09, GM03a, Har08, HNS04, HPCH00, HHB09, HMO07, Iar01, IS07, JAWC08, JW00, KP06, KT06, KK02a, Kaw06, KR06, KVM^{+}00, LL08a, LARvW07, LNT09, Li07, LR02b, Lou08b, Mas07, MS09, MS06a, MDL^{+}03, NOS05, ND03, NKC00, Opp01, PS03a, PA00, PA02a, RSMJ^{+}02, RdCM04, Ric00, RM08, RS06, Rue02, SS04, SA02, Sch07, SCM04, SO08, Spo06c, SSB04, Sui00, TBF05, TB07, TE00]. Dynamics [Vel00, Wei02, WHI02, Wu02, dSWP05, ATH06, Spo06b, Rug08, Buc07, Pod03, Shl04]. Dynamo [AH07, Vin02]. Dyson [Ano04m, Bou07, Luo07, LZ08].

E.Y [Ach08]. Early [Dom03b]. Ecological [JHA07]. Economies [MT08]. Economy [CPT05]. Econophysics [Mas00, Mas03a]. ed [Dom04, Rod02]. Eddy [PS03b]. Edge [Car03, Opp04d, Shc09]. edition [Wit03]. Editor [RT08b]. Edouard [Leb08b]. eds [GC08a, Opp04a, Pod07c, Rug08, Shl04, Wei09]. Edwards [CMN06, KRT00, WP02]. Effect [AH07, BGX01, BK01a, BD01, CT04b, GKR02, GGZD02, Leg03, LM01, MK03, PO07, TK02, Vel00]. Effective [BFKT09, BM05a, FK02, FR00, HMPPMV00, LNM06]. Effects [ADLM01, AMP06, BKM02, EMMZ06, Fei03, HPWW03, JHA07, KTZ06, LK06b, MDW09, NOS05, PDvB00, QAK02, Sam05a, Sid05, TC08a, VTM00, Han02]. Efficiency [Sla09a, SD05, Zha09, SD23]. Efficient [JL00a]. Eigen [MPD09, PD06]. Eigenfunctions [AS03]. Eigenspaces [ND08]. Eigenspectrum [SB01a]. Eigenstates [Süt02]. Eigenvalue [BFP08, CGK09, Has01, LP08, MBL08, NNR09]. Eigenvalue-Counting [CGK09]. Eigenvalues [CAC02, JP01, KM06, Sos02]. Eight [Bax02, Bax04a, BW07b, FM03, FM05a, FM09a, WK04]. Eight-Vertex [Bax02, Bax04a, WK04]. Einstein [Gra00, HE03, JPZ09, KO05, LVZ03, Leg03, Lu00, Lu04, Lu05, MV01, Rub04, Tel06a, Yan03, Zha00a]. Elastic [Ber03b, BC03, Koi07b, LM06, Pro09b, SS08, SS09e]. Elasticity [MP06]. Elastostatics [SS09a]. Electric [EE07, Git09, SSE05, Tél06b]. Electrical [Ber03b, BMN07]. Electrode [MH07]. Electrodynamics [LL02]. Electrolyte [Lev03, MH07, Sam01a, SJ01a, Sam05a, Sam06, Tél06b]. Electrolytes [KHV00]. Electrolytic [TM02]. Electromagnetic [Kie04a, Kie04b]. Electron [AE08, AR00, Bak03, Erd02, LFB04, RdCM04, TJ08, Tia04]. Electronegativity [CW06]. Electronic [DSC02, GPM^{+}05]. Electrons [MP01, Mes02a]. Electrostatic [Sam05b]. Elementary [CRV00, Pan02]. Elements [LP09, Wei05]. Elimination [JGB^{+}03]. Elliott [Ano03n, Dys04, Leb04b, Yng04]. Elliptic [FM09a, GS00]. Elliptical [QLAS02]. Elsevier [GCU02]. Embrechts [Taq04]. Emergent [Spo03]. Emery [BL08b, SK07]. Emil [bA03]. Emulsion [dHP09]. Endomorphisms [AOT06]. Energies [Kie09a, Kie09b, LL02, WK00, Opp02b]. Energy [ADG00, ALS06b, AMT07, AYJP01a, Bar06b, Bax00, Bax03, BNZ07, BBS04, BGP04, BK07, CGR07, CC09, DN03, DMR05, DP04, DLS08, EB02, EGGI01, Far05, Fis06a, FZ03, GK09a, GS07a, GS09, Jab01, JR07, Jan06b, Kur03, LB01, Lee09, LY01, LLV09, NSS04, Nag03, O’C01, Pan02, PVV02, PBFC05, Rác00, Rut08, SY08, Shn03, TG00, TC03, TE00, WE07, Woj09, YY09, ZW04, JR06, Pod07a]. Energy-Conserving [PBFC05]. Energy-Dependent [PVV02]. Energy-Lowering [DN03]. Energy-Transport [DMR05]. Engaged [MKL08]. Engel [Kan01]. Engineering [KLL05]. Engineers [BA06]. Engines [Jan06a]. England [Wei01]. Enhanced [FK00]. Enigmatic [GPL03]. Enriched [JHW09]. Ensemble [AK07, Blo01, Blo04, BB05, CETT05, Has01, NNR09, ND03, VBM06, ZRA04]. Ensembles [Ada01, BS03, CGH^{+}03, CDG07, DMPV08, DGKV07, EHT00, GS08c, GS08d, RS05, Rid04, Shc09, Sin09, Sos03]. Enskog [Bob06, Jia07, PS02b, UC07, Wu09]. Entangled [MBL08]. Entanglement [CCAD08, CAD09, GMW02, GOS08, Hol02b, IK09b, NT04, XKHK08]. Entrainment [GG00]. Entries [PS07b]. Entropic [BGV01, PV02, Sak04a, Sak06, Sak09]. Entropies [TB06]. Entropy [AFHV03, All08, AOT06, AK02, Att00, Bah07, BGH01, BLLO09, BET00, BL08b, BL08c, BMN07, CCAD08, CAD09, DJ05, DP07a, DR03b, DLS07, Dud07, Edw04, ER02, EK06, Fis06a, Fis06b, Fis08, Gar06, Gas04, Gas07, GZ01, Han07, HB08, IVZ06, IK09b, JS01e, Kaw06, Kie09c, KR06, Li06, MR00, MN03, MN05, Mor07, Nag04, Opp02b, Piv03, PG99, PG00, PG03, QQT02, Rad08b, Rob00, RNVRP04, SF07a, Sch01, SSS07, SC09, Swe04, TG00, VTM00, VMT02, ZRA04, ZKD05]. Entropy-Driven [GZ01, IVZ06]. Entropy-Production [VMT02]. Entropy. [Opp04a]. Enumeration [SD05, SD23]. Enumerations [Jen01]. Environment [And08, BCFP05, GJ08b, Pop01, Sac04, Shi09]. Environments [BK09, BK02, CL07, KK02b, KK02c, KO05]. Epitaxial [EY01]. Equality [GJ08a]. Equalization [CW06]. Equation [ADG00, ASB02, Ale01b, AN00, AN05, AN06, BTM^{+}04, Bao00, Bao04, BGS06, Bar04, BLZ01, BK03a, BKLO00, BCEP04, BCEP06, Bir06, BCT05, BCL08, BHPH01, BC02a, BC02c, BC02d, BCT03, BC03, BL06, BGL03, BdMR03, CM00, CR01, Cai04, CS02, CGR07, CCL09, Cas01, Cas02, Cer05b, Cer06, CD04, CD09a, Cha09, CSN02, CKKC00, CW04, CLM05, CD05, DJZ00, DR01, Dog00a, DKRS02, Erd02, ESY04, ETB06, Esc07, FM01a, Fan05, Far02, FS07a, FKS08, Fou00, FM01c, FG08b, GR06, GW00, GS03b, Gra00, GM03a, GK04, HD09, JW04, JH03, Jia07, Jia08, KCT08, KL06, KL08, KR07, Kuk06, KMS08, Lee05, LRM06, LT04, Lou03, Lou08a, Lu00, Lu01, Lu04, Lu05, Lu06, Mat02, Nag03, PO07, PS02b, RC01, Rot06, SS09c, SM08]. Equation [SMT03, SK06, Sha10, Sim08, SV09, Spo06a, Spo06b, Spo06c, SST^{+}00, SFCO02, SW08, TTNK06, Tak09a, Tak09b, Tél07, TKR02, UC07, Val09b, Val09a, WZ09, WW06, Wu09, Yep02, YS00, Yu06, Zah02, ZKBG03a, ZKBG03b, Zha08, Tak10]. Equations [AL03, AV03, AL01, AR02, BDL06, BELM00, BKOS07, BTV09, BD06b, BCG00, BCG01, BC02b, BGP04, BG06, Bob06, Bol09, BKL00, BBS03, Bud08, CT07, Cañ07, CEFM00, CD09b, CE02b, CGV02, DMN06, DRL02, DDM08, Dog00b, EB02, FM01b, FM09a, FH04, FS07a, FJMR02, FG03, Fou06, GJY04, GvH07, HS08, HPWW03, JL00a, JS01d, Kar07, Kas01, KP05, LL08a, LW01a, LM02, LW04a, Los05, Lou07, MS05b, MMR06, MM06b, Ost07, PhD07, PN03, RRS^{+}00, Rom04, Rom08, Sac05, SH05, SEZ05, Str06, Tid01b, VEW08, Wes09, YBMS06, ZZ06, vWL00, AV87, Ach08, GCU02]. Equilibria [BCT05, BC09, CIS01]. Equilibrium [AH03, AP03, BD06a, Bar04, BLR08, BDG^{+}02, BT04b, BDG^{+}04, BDG^{+}06, BCL08, BNZ02, Cañ07, CL03, CCL09, CK00, CP03, CL02, CCMT09, Dom03b, DKRS02, DGZ04, FPS04, FS07a, FNO06, GS03a, GLTZ06, GS07b, GPL03, HKW07, Jaf03, JP02, JMU05, Jus01, KK02a, Lat02, LS04a, Li07, Lu05, Luc08b, LBLB04, Mor07, Opp06a, PVZ05, RY07, Rom08, RS06, SJ01b, TV00, TV02, Uch04, Wag00, Zhi00, BDLvW08, CG09, FGC09, Opp04b, SWK09, WW05, Opp04b]. Equilibriums [Fre01]. Equivalence [Ada01, Bax00, CETT05, EHT00, Rue00b, Sak00]. Equivalent [Bol09, LS04b]. Erased [FLW08, Gra09, HL08]. Erdos [CB04, EMH04, PDV07]. Ergodic [Bla03, GG01, Har08, RB08]. Ergodicity [BCMP04, BDP02, GL04, LW04a, MPT06, MN00, Rom04, Too00, Too04]. Errata [PG00, dG05b]. Erratum [Ano01d, BCG01, BPS03, Cau02b, Coh10, CS09d, GCU03, Gas07, KK02b, Lou08a, NOV04, Pro09a, Sak05, Sod17, Spo06b]. Error [BDM09, RT08b]. Escape [AR06, BBF09, BCC09, HS04, KL09, SSHE06, SSH06a, SSH06b, Ste04, VMT02]. Escape-Rate [VMT02]. Essam [KK00b]. Essay [Kad03b, Taq04]. Essential [MNOS04]. Ester [GC08a]. Estimate [Bou09]. Estimates [BHS07, BDM09, BBP02, BKL00, BM00, CGK09, FJMR02, Ton07, Val00]. Estimating [GOBY06, RS09]. Estimation [AER05, MS01, PS07a]. Estimations [Woj09]. Eternal [BC02a]. Euclidean [AGY02, Con04, KP07, KP08b]. Euler [Bir06, Dub06, FS07a, FMB03, LL08a, Pat07a, SEZ05]. Eulerian [CG09, HPCH00, Kuk04, TV02, TV03]. Evaluation [DRL02, LLMS02, MM05b]. Evaporation [BGH01]. Evaporation/ [BGH01]. Evasion [CGS07]. Even [CM04b]. Event [LB09, Mai03]. Events [KL09, Por04, VEW08, Pod07c]. Every [Too07]. Evidence [BW07a, BFT02, Luc09a, RK05]. Evidences [MPRT^{+}00]. Evolution [BCF^{+}01, Ban05, BdMR03, CLM08, CD09b, Ghe09, HWB01, Iar01, KN04, Ken04, Ken07, Ken09, MZ02a, Nou02, Saa07, Bog04, Sza06]. Evolutions [BBK05, KNK04, LR02a, Yos08]. Evolving [BB03]. Exact [ABCM08, BS07, Ber03b, BP01b, BL09, BC02a, BG06, BD03b, BK03b, CZ03, CSS02, CJSS04, CS08a, CLS02, DLS03, Gat00, Hei04, HPWW03, HJ09, HAG02, KNK04, KMST00, KP06, KL09, KK00a, KW00b, KM09b, MBL08, Mec01, MM05b, PS04a, RS04a, Saa07, Sam03, Sam05a, Sch00, SS09e, Shi06b, Süt02, Urr08, ZS02]. Exactly [BdGM01, BGOY04, EG05, GK02, KMVE03, OBB03, Sam01a, SJ01a, Sam04, Sam05b]. Exactness [HY04]. Examination [LDNG08]. Example [BBK08, FSG08, Man03, Mon04]. Examples [CS04, CS09d, Got05, Süt04]. Exceed [Tas06]. Exchange [AFFS06]. Exchanges [Gal06]. Excitability [GAV00]. Excitable [GOSG00]. Excitation [HI05, PS03a]. Excitations [Iar01, SS08, Sab08, SS09a]. Exclusion [ABS01, ALS09, Bah07, BT04b, BD06c, BDL08, CL07, DF02, DH08, DLS02, DLS03, DEL04, DDR04, DELO05, DG09b, DP07c, ED04, FF04, FF07, GM05, GJ08c, Hei04, IS07, JNH^{+}09, Jun03, KK00a, LOV04, Lou05, Lou08b, MK03, Rit06, SA02, Sep01, Set06, Uch04]. Excursion [MC05]. Excursions [DEL04]. Exhibiting [AG05]. Existence [Bak06, BC09, CP01, Coh09, Coh10, CS04, CS09d, FPS04, FM07, Gho05, GP04, Jia08, Kas02, KK02b, KK02c, MRW02, PS02a, Pol00, Tid01a, Zha08]. Expanding [AP06, CL00, PZ06b]. Expansion [All08, Ast08, Bar06b, Bob06, CR01, Cai04, Cam01, Coh09, Coh10, Fuk03, KPvB06, MS03, PG99, RC01, cR04, SS09e, Str06, dPG04, PG00]. Expansions [ABCM03, ABCM08, AC01b, BZ00a, BC02e, DK02, HPCH00, SM06]. Expectation [FK05, Sal09]. Expected [MCZ06]. Experimental [AG05, PR07, Rác00, SS09d]. Experiments [AFFS06, AC03, SL02]. Explain [EA07]. Explained [Man03]. Explicit [Att00, BCL^{+}07, FS07a, RS08]. Explicitly [BZ03]. Explosion [CGR07]. Exponent [BJ02, DJZ00, dSFM02, JH03, MM05b, Pan02, SJW07, SS09d, vZvB02]. Exponential [AV08, BH08b, BM02b, Che07, DMPV08, JL00a, LL09a, NH03, Ost04, Pan08, Shi06a]. Exponentially [CK06]. Exponentially-Distributed [CK06]. Exponents [BDP02, Cam05, CGJ^{+}05, DR03a, Gam03, HS00, KM09a, KPvB06, LSSW03, LD01, MN00, PDvB00, STV02]. Exports [Kad03a]. Exposed [FW07]. Expression [Dom03a, HPK^{+}07, Mur08]. Expressions [BP01b, DB02, Nag07]. Extended [BCK00a, BC02e, BCM02, CM02a, MDB00, Rug08, UC07, Wit03, JS01a]. Extending [MD01]. Extension [DA03, KK00b]. Extensions [FS06, Grm08, Han06]. External [BR00a, BC09, GHO^{+}00, LRM06, Lin04, LM01, Wee03]. Extinction [KS07, dARBY04]. Extrema [HNOV07]. Extremal [Ast08, Jun03, LKL09]. Extreme [Bur08, Pod07c]. Extreme-Value [Bur08]. Extremes [Mas09b]. Extremum [PG03]. Eye [Hep09]. Eynard [BR05b].

F [Bog04, GCU02, Opp03b, Wes04, bA08]. Faceted [FS03]. Factorised [EMZ06]. Factorization [HKT01]. Factors [CCAD08]. Falcioni [Opp09]. Falicov [DMN00, HK01, LFB04, Mes02b, Woj06]. Falling [GT04a]. Families [CS03b, PTZ03]. Family [BCL^{+}07, DLM05, KP05, LSSW03]. Farey [FKÖ03, FK04, PR07, PFK06]. Fast [Boo01, CT07, CL03, JGB^{+}03, Ken07]. Faster [Ken02]. FCC [LMZ00]. FDLBM [ST02]. Features [GAV00, KMVE03, MNOS04]. February [KK02b]. Feedback [PR07]. Feigenbaum [FKM05]. Fellerian [MRS07]. Fermi [BLP09, BM04c, CS08b, DJ04, Gal01, HY04, LL09a, Lu01, Lu06, OP04]. Fermion [AH01, JS01c, OP04, Shi06a]. Fermionic [AFHV03, SW00, Sal09]. Fernandez [Rug08]. Ferroelectric [BL09]. Ferrofluid [GTZ02]. Ferrofluids [Del05]. Ferromagnet [DG08, Dot06, Rut08]. Ferromagnetic [AYP07a, BIV01, HI04, NSS04, SO00a, SO00b, SO02, TPV09]. Ferromagnetism [FU05]. Feynman [Ost07]. Fiber [DOS01]. Fibers [BM01a]. Fibonacci [NS05]. Fibre [DS07]. Fidelity [LMV07]. Field [AFFS06, AC01a, ADLM01, Bar06a, Bar06b, Bar08, BLZ01, BP01a, BCC06, BL08c, BCZ04, CR07, CS09b, CGT05, CNV01, Con04, CGI02, De 07, DG08, DWMB05, DGR08, DLP09, Dir04, DMP05, Erd02, ES04, ES09, FLW08, FZ03, FT06, FMS04, GHO^{+}00, Gen04, HvdHS08, HKW07, JKO^{+}01, Kad09, KS06, Kie04a, Kie04b, KMSS01, KP04, KM08a, Kül01, Kül03, Kyt06, Lan09, LW07, LPE04, MB03, MS08a, NS02, Nou02, PDvB00, Pat07b, Pey09, RT08c, Rát09, RK00, Roe03, SSD00, Sak06, Sak07, Sak09, Sak01, Sak04b, SW00, Sid05, SSE05, Sla09a, Ste04, SJ01c, WW06, YHD02, ZS02, dARBY04, dPG04, Pod07b, CW04, Sak05]. Field-Driven [PDvB00, RK00]. Fields [AGY02, ALS03, AG05, BY03, BEPK^{+}02, BZ03, BR04, CCAD08, De 04, EE07, Fan04, FG08a, Kor00, Li06, MP01, MN02, Sch06, Sch08, Sos00, Wee03, vdHK08, Pod07b, Git09]. Filament [HMO07, MK07]. Filled [CLS02]. Filling [Riv02]. Film [GMR06]. Films [AB09, GD06, TM02]. Filter [Kra09, MN05]. Filtering [EK06, New08]. Finance [Mas00, Mas05a, Mas08, Mas02a]. Financial [CPP09, Mas02a, Por04, Wit03]. Fine [BR01]. Finetti [Got05]. Finger [FPD01]. Fingering [CT04b]. Finite [Akt01, BJR02, BIV01, BKM02, Bor02, BGM04, CO08a, CT04a, CGGP04, CSP02, CSM03, DP07c, Fis06a, Fis06b, FSZ01, Gam03, GM07b, GGL09, Got05, GT06, Han02, HvdHS08, HNO05b, HKW07, Jan03, JMU05, KPH00, KP01b, KMSS01, Kol08, Kon02a, Kon02b, KM01, Mor07, Nis07, PD05, Pro09b, Rom04, Suw09, TNK04, Ton07, TT05, VST09, ZW04, vdEvdHH08]. Finite-Dimensional [Nis07]. Finite-Range [GM07b, GT06]. Finite-Sampling [ZW04]. Finite-Size [Akt01, BKM02, CGGP04, Fis06a, Fis06b, Han02, KPH00, KP01b, KM01, Ton07, TT05]. Finite-Time [DP07c, Gam03, GGL09]. Finiteness [CT04b]. Finitized [CP02]. First [BBCK04, BCC06, Bra02, Bur08, GM07b, HNO05a, HNO05b, JK01, KC03, Lin04, MM06a, NH03, OvWLH00, PIRB03, Rei00, WLTH07]. First-Order [BBCK04, BCC06, GM07b, NH03, OvWLH00, PIRB03, Rei00, HNO05a, HNO05b]. First-Passage [Bra02, Bur08, KC03, WLTH07]. Fish [Bir07, DM08]. Fisher [Ano03m, CD05, GII08, IK09b, JK04, LW01b, Luo01, Mer03]. Fixation [CF03]. Fixed [AS06, BCF^{+}09, DMB01, FST06, FLS07, Fra07a, Hat07, Kas02, LL09a, PZ06a, TPV09, ABC05a]. Fixed-Boundary [DMB01]. FKG [TTK01]. Flagellar [WLTH07]. Flamm [FT08]. Flat [BW07a, HI04, HI05, Lia06, OSv04]. Flat-Band [HI04, HI05]. Flavors [Jan01b]. Fleas [HNS04]. Flights [CGK^{+}04]. Flip [CGG09, RT08a]. Flip-Annihilation [RT08a]. Flocks [VLCW05]. Flow [ASB02, AS08, Ben05, BGL05, BW07a, Bla03, Bla08, BR01, BCL^{+}07, CMV05, Cer01, CS03a, CG03, DGR08, DR06b, FS07a, Fre03, FMB03, Gar03, GMR06, HRK^{+}05, KTH^{+}05, MN05, OC05, PO07, SJW07, SWK09, SMRK^{+}02, TJ08, TTM^{+}01, TS04, VKVT02, ZQS^{+}05, ZGA02, dSWP05]. Flowers [MP06]. Flowing [HJDRD00]. Flows [AG05, BM08, BD04a, BI06, BD09, BBL00, Bla05, Che07, CSN02, Fan04, FPS04, Gas02, Gen06, HS09a, HD02, Ito02, NDC02, QLAS02, Sab08, SKT09, TAL06, Zha00b]. Floyd [Mas09a]. Fluctuating [BCFP05, Gry01, ICO02, MC05, MB03, dGNPR04]. Fluctuation [AG07, BCG08, BDG^{+}02, BD03a, BGGZ06, BFPS07, GZG05, Jar00, KS07, LOV04, LRB00, LK06b, MMR08a, Nag03, OvWLH00, SC09, Sin08, Sos00, TC07]. Fluctuation-Induced [KS07, OvWLH00]. Fluctuations [AGL01, AVZ00, BS07, Ber05, BDG^{+}06, BGV01, Cam06, CD08, DR03a, DF02, DDM08, DDR04, DELO05, DG09b, DG09a, EEDJ00, Far02, Far05, FS03, FNO06, GGL09, GTW01, JL01, Jan03, JT04, JS08b, Lan09, LV02, LWL00, LP09, MGT07, MDL^{+}03, NV08, PD05, PPR00, Rid03, Rub06b, SA02, SI04, TC08a, TC08b, TV02, Uch04, Vel00, VPB^{+}06, Wei02, ZRA04, dZS04, GC08b, Pod02a]. Fluctuations-Inclusive [PPR00]. Fluid [AC01a, AC01b, AC04, AG05, CRT00, CM02b, GCU02, GH04, ICO02, Jan00a, Jan01a, JT04, LV01, Lyb05, Nad00, NDC02, PWC02, PP01, SEZ05, SFCO02, Wag00, WHI02, dMG04, BELM06, Fre06]. Fluids [Ace00, AH01, BGW01, BELM00, DG01, GC05, HGST00, JS04, JHW09, Keh05, KC02, LS09b, Pod05, RS04b, Rei05, RK05, Sam00, Sam07, SS09d, SAPD05, WBE02, YFP00, dZS04]. Flux [BCKM00, DF02, DP04, EZ04, MCZ06, Rác00]. Flux-Across-Surfaces [DP04]. Foaming [KTH^{+}05]. Fock [ALR07, BGGM04]. Focus [Ler00]. Focussing [Rid03]. Fokker [ASB02, CEFM00, DR01, KL06, KL08, Zah02, Zha08]. Folding [TE00]. Force [Bar06a, Bur08, CDG06, FZS^{+}00, SvL02, WW06]. Forced [BK03a, CD09a, CLM05, MSVE07, Rom04]. Forces [BA06, Ber03b, BM05b, Dag03a, SHD03]. Forcing [Bak06, GFH08, KP05, Kuk06, LW04a, MS05b]. Forests [CS09a, JSS05]. Form [CCAD08, Kar07]. Formalism [AS02, CBKM04, LARvW07, TB06, UC07, VMT02]. Format [Tél07]. Formation [BCK04b, DHB07, Dro00, FPD01]. Formations [VLCW05]. Forms [BLS04, KZ03a, LS04b, Val00]. Formula [BM01a, BCK04b, BEK^{+}07, Gas02, HJ09, IK09b, KK00b, Mag09a]. Formulae [Hir05, KL09]. Formulas [Ast08, BI03, Ito02, Mai03]. Formulated [Rue02]. Formulation [Del05, Fan05, Häg07, KR07, PD06, PV02, WA09]. Fortuin [Arg02]. Forward [HD09]. Foundations [HD02, MPRT^{+}00, Wei09]. Fouque [Fre08a]. Four [Akt01, HWB01, Hos06, Mas03b, MN00, Wat04]. Four-Dimensional [Akt01]. Four-State [HWB01, MN00]. Fourier [BO05, BLL04, KR07, PF07]. Fourth [Lyb05]. FPU [CCG04]. Fractal [AS02, Dys05, HO09, Koz00, Wei07]. Fractal-Based [Wei07]. Fractals [Mis06, SD06, Wei01]. Fraction [CHNO06, FK04, PFK06]. Fractional [AL01, ALS03, Bar06a, BVBF01, Cha09, FK00, FM02b, FKM05, HKW07, Kup04, MK04, Ruz00, YBMS06, Zah02]. Fractions [FKÖ03]. Fracture [BC01, BPS06b]. Fragmentation [Ber03a, BP07, BC01, Cañ07, FG03, GE09, Kol04, MKB00, PD05, RS05]. Frames [Jan06b]. Framework [Cas01, HS08]. Franceschetti [Bog09]. Fredholm [TW08]. Free [AYJP01a, BCG08, Bar06b, BJ08, Bax00, Bax03, BK04, FZ03, GK09a, JS01c, Jan06b, JKO^{+}01, Kad03b, KTH^{+}05, Opp02b, TC03, Woj09, ZW04, CPK07, Iar01]. Free/ [CPK07]. Freedom [DRL02, FJMR03, HDMF04, JGB^{+}03, MM02]. Freeman [Ano04m]. Freezing [MRS07, Sem08]. Frenkel [CM05]. Frequencies [AMP06]. Frequency [Fan05]. Friction [Lin08, Pom05, dG05a, dG05b]. Friendly [KG03]. Frog [LMP05]. Frogs [Pop01]. Frohlinde [Hol02a]. Front [BD01, CK06, Gir03, Lem00, vZvB02]. Frontiers [Pod07c]. Fronts [AS05, Kad06]. Froth [FR00]. Frozen [Rát09]. Fugacity [CR01, Tél07]. Fully [BB09, CN01, FJMR02, MDL^{+}03, PM06]. Fun [Suz01]. Function [Ace00, AK02, Bax03, BM04c, BBCK04, BM00, CS09b, CCL04, FJM01a, GLTZ06, HL02, JS01a, JSS03, JS06, JK01, KIK08, KT09, KW00a, Kom07, Lin08, MC05, Mez01, RR01, SS01, SS09b, Yu06]. Functional [Att00, Bax04b, Bax06, CR03, DLS03, ED04, FM09a, HDMF04, Nep03, Ost04, Ost07, RS04b, Sal09, Suz01, TC03]. Functionals [FH04, Wee03]. Functions [Akt01, Ano01d, AS02, BI03, Car03, CSS02, CJSS04, CS08a, CS08c, Cop08, DSC02, FEDZ07, FM09b, GI08, HMH05, Law09, LP09, MM05b, Nag07, PAY09, PS04a, Ric04, Ruz00, Sam07, Sin09, Too08, YS05, SL02, Weh97]. Fundamental [SF07b]. Fundamentals [Opp06c]. Further [BGX01]. Future [Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g].

G [Ano03o, Fre08a, Han02, Opp04a, Opp04d, Rod02, Wei09]. Gabetta [GC08a]. Gacs [Gra01]. Gain [Kom07]. Gain-of-Function [Kom07]. Gallavotti [LRB00]. Games [MB02, Mie04]. Gamow [Hua09]. Gap [Ale01a, AY01, BG01a, DH08, GF08, LVZ03]. Gapless [HI05]. Gapped [Yar05]. Gaps [DLS08, Kom00b, dlLV07]. Garland [Dag03a]. Garnier [Fre08a]. Gas [AE08, AAP02, ABT03, AL07, Bak03, BTT07, BP02, BTV09, BGH01, BFT02, BDL00, Bro01, BZ00b, BZ08b, BK03b, CGGP04, CC00a, CK00, CS08b, CL02, CLS02, CG00c, Coh09, Coh10, CC00b, DMS02, DM06b, DP07b, Dim09, Dor02, DMP05, DB02, FPS07, FJM01a, GZ01, Gir01, GS09, GT06, KS02, KD00, KPvB06, LRY05, LLMM03, LV02, LB01, Lee09, LY01, LPE04, MPS02, MB02, MM05b, MO09b, Nag03, Nou02, PDvB00, PO07, PY09a, Pol00, PG99, PG00, RKH00, RKN00, cR04, Sak00, Sam01b, SJ02a, Sam03, SS05, SK06, Shn03, Sid05, TAL06, Tél07, TTM^{+}01, TS04, TBF05, YY09, Yep02, ZQS^{+}05]. Gaseous [SK06]. Gases [ABCM03, BDG^{+}06, BCT05, BD02, BG06, BBP02, BMD00, Cer05a, CEMM09, FS07a, GLM02, GC08a, GW00, JM00, LS09a, Mun07, Opp04d, TB04, VPB^{+}06, WW05, vWR05, vZvB02]. Gasket [CCY07, CC08]. Gaskets [HT02, Hat07]. Gauge [KM09b]. Gauss [AS03]. Gaussian [ADLM01, BY03, BKLO00, Bol09, BGM04, CGH^{+}03, CMN06, DR03a, FG08a, FLM00, KO02, LP09, MNBO09, Raw03, Rei08, Sak06, Sak07, Sos00]. Gene [Dom03a, HPK^{+}07, Mur08]. General [BC04, CS04, DD09, FBG02, FM05b, HMS05, McD01, Pro07, Pro09a, QQT02, TT05, UC07, CS09d, SS01, Lan06]. Generalization [Bal09, Bax09, DMSR03, Hat07, RF02]. Generalized [Aki08, BLS04, Bax04b, Bax06, BM01a, BR00b, Blo01, Blo04, BHPH01, Bob08, BL06, CIS01, CGK09, CETT05, DP07b, FK05, GJ08a, GAV00, GS07a, JR06, JR07, KS00, Los05, Luc08b, MVE04, NNR09, Ric04, SMS09, Süt04, Wol06]. Generated [BP07, Cam06, SDBS07]. Generator [vWL00]. Generic [DJW07, Gal06, GT02, Liu09, Luc08a]. Genetic [Eps06, GPM^{+}05]. Genetics [CF03]. Genome [PA00, PA02a]. Gentle [LNT09]. Genus [CSM03]. Geodesic [Mez01]. Geoffrey [Spo03]. Geometric [Der08, JW04, MCG08, SLB00]. Geometrical [Fan05, Mez01]. Geometries [LK09]. Geometry [CFN06, CD08, DMSR00, MP06, Rit03, Sch08, Xin09]. George [Nol04]. Ghirardi [Gri00, Tum06]. Ghirlanda [CG07]. Giacomin [Ton08b]. Giambattista [Ton08b]. Gianfausto [Mas09b]. Giardina [Pod07b]. Gibbs [KP08b, Ada01, BS05, BP02, BBS04, BP07, BCK04b, BMPZ04, CT04a, CRVV09, Cha00, Der08, GT02, HO09, JMU05, KP07, Kur03, LM01, Mor07, MR04, MR05, Roz06, Sak00, Wee03]. Gibbs-Equilibrium [JMU05, Mor07]. Gibbs-Like [BBS04]. Gibbsian [CU03, EL02, ECC03, Kül01, vEK07]. Gibbsianness [DR05, GY05, Kül03, LR02a]. Gilbert [Lan06]. Ginibre [AK07, Rid04]. Ginzburg [EMO08, WBE02]. Ginzburg-Type [WBE02]. Giorgio [Leb08b]. Given [BV08, Lin08]. Glass [Arg07, AK09, Dot04, Eck07, Fis06a, Fis06b, Fis07, Fis08, Git07, GT04b, Has00, ISS07, KRT00, KM09b, Led00, MNS08, MS06a, TB08, WD01, WB04]. Glasses [Bar06b, BCF^{+}09, BK07, CM04a, CGG09, CS09c, De 04, DGM07a, DGM07b, FT06, Fra07b, MPRT^{+}00, NS02, Nis07, Pod02b, VHO09, Pod07b, Pod07a]. Glassy [EGGI01, Gar08, Kaw03, MS06b, TBF05, TB07]. Glauber [BM02b, DLP09, Iar01]. Glazer [Opp02a]. Global [Cer05b, CLT07, GF08, Jia07, Jia08, Li06, MS08a, Wu09, Zha08, Pol00]. GOE [BFP08]. Golden [DJ04, LL09a]. Goodman [Fre08b]. Gordon [CR01, Cai04, MM05b, RC01]. Gordon/ [MM05b]. GOY [BBBF06]. Grad [Kar07]. Gradient [DR05, GJ08b]. Gradients [YCCN07]. Grained [JAWC08]. Graining [KT06, Opp09]. Grand [CT04a, CAC02]. Granular [AR09, AL07, BNZ07, BCT05, BC02b, BCT03, Cer01, Cer05a, DG01, DB02, GM07a, Rad08a, SvL02, TTM^{+}01, TS04, Vil06b, VPB^{+}06, WBE02]. Graph [CB04, KM09b, PDV07, WP03]. Graphical [Sch08]. Graphs [AV08, BMA02, BG00a, BG02, BZ09, Bjö09, BK04, CS03a, CS03b, CE02a, CDA09, EMH04, GR06, KP08a, PTZ03, PS04b, PM05, RT08c, Riv04, Tel06a, Tur03, Tur06, Wu00, vdEvdHH08]. Graphs-with [RT08c]. Gravitating [Gir01]. Gravitational [Fel02, HH02, MPT06]. Gravity [GT06]. Grazing [DR01]. Green [DB02, RR01, Yu06]. Greven [Opp04a]. Grid [Rhe05]. Grids [CSN02, HL02]. Griffiths [BG00b, BL08b, Dot06, SK07]. Gross [JW04]. Ground [CdlL05, FRZB09, GS09, HK01, Has00, HI04, HI05, JM00, Kie09a, Kie09b, KW00b, Lee09, LY01, RNVRP04, Roz06, SY08, Sig09, Woj03, Yar05, YY09, dlLV07]. Ground-State [RNVRP04]. Group [BR01, Cop08, DRC04, FM08, GAD06, HT02, Hat07, Ito02, MTT08, MS08a, RS08, Rob00, Sha01, Shi06b, Rub09a]. Groups [FEDZ07, Piv03]. Growing [GGL09]. Growth [BR00a, Bal01, dBR08, CM00, DHB07, DLS08, EY01, FPD01, GTW01, Gry01, Mez01, MS02, Pen08, PD05, PS04a, SI04, Yos08]. Guerra [CG07, FT06, Tal07]. Guest [JS08b, Sam06, Tél06b, Sam05a]. Guide [Bra02, Gra01, KN04, Rap01, Opp02a]. Guided [MM08]. Gumbel [vdHK08]. Gyroscopically [BHRW04].

H [GCU02, Mas00, Pod04b, Pod04a, Pod07c, Rod02]. Haar [TM08]. Haar-like [TM08]. Hagen [Hol02a, Wit03]. Half [BGS06, CO08b, GS04, Ken08, Riv02, SS08, SS09a, SI04, SS09e]. Half-Filling [Riv02]. Half-Line [GS04]. Half-Plane [CO08b, SS08, SS09a]. Half-Space [BGS06, SI04, SS09e]. Hall [ABS01, FPSW01, Kom00a, Kom00b, Kom08]. Halves [LY07]. Hamiltonian [ALR07, BL08a, Car07, EMMZ06, FS07b, Har08, Jar00, KK02a, Ler00, RY07, Val00, dOP08]. Hamiltonians [Mas05a, BSB00, O’C01]. Hamming [PZ06a]. Handbook [BA06]. Hanson [Mas09a]. Hard [AC04, BC04, BW07b, BP08a, CR01, Cai04, CK04, CM04b, CM04c, CM06, EFPZ05, FPS07, HGST00, HPF^{+}02, HHB09, IVZ06, Jia08, LT04, Lyb05, MMR06, MM06b, Pan02, RC01, RS04b, Rei00, RM08, Sak07, Sam06, Sot07, Urr08, Wag00, vZvB02]. Hard-Core [CR01, RC01, Sam06]. Hard-Disk [EFPZ05, RM08, Wag00]. Hard-Rods [IVZ06]. Hard-Sphere [Pan02]. Hardcore [Der08]. Hardness [CW06]. Hardy [BLS04]. Harish [FEDZ07]. Harish-Chandra [FEDZ07]. Harman [Dom04]. Harmonic [Ber08, BLL04, DR06a, DKM04, EZ04, FNO06, Leg03, RD08, Sak04a, Yan03]. Harmonically [Sin08]. Harnack [BDM09]. Harness [FNP07]. Harris [TTK01]. Hartree [BGGM04]. Hartwig [IK09b, JK04]. Hausdorff [DJ05, DKK02, GT02, HV01, JR08, MS01, Sim08]. Havlin [Wei01]. HCP [LMZ00]. Heads [BM04b]. Healing [CKSZ06, KSSM07]. Heat [BKOS07, BO05, BGL05, BHRW04, BLLO09, CN03, CEMM09, DR06a, EZ04, HK02a, Jan06a, Kup04, LL00, MRV02, MNV03, RD08, SC09, SNC05]. Heat-Type [BKOS07]. Heated [TS04]. Heating [LK06b]. Heavy [BK01a, MP01]. Height [GTW01, MC05, MN00]. Heisenberg [FL03a, PS02a]. Helmholtz [San05]. Hemmer [SSLG^{+}00]. Hénon [DLM05]. Herbut [Git08, Pod08]. Hermitian [CCM00, DJZ00, Has01, KS08, PS08a]. Heterogeneous [DMSR03, DDMSR04, Sah02, ZK00]. Heteropolymer [dHW04]. Hexagon [LG01]. Hexagonal [HN03]. Hexagons [Luc08a]. Hidden [All08, AC05, GKRV09, ZKD05]. Hidetoshi [Pod02b]. Hiemer [Mon04]. Hierarchical [BG02, dFCS06, Hos06, Kom07, MCG08, Rei05, RK05, VHO09, Wat04, WW01]. Hierarchies [LS00b]. Hierarchy [CS08c, JBKN01]. High [ASB02, BGP04, BC02e, CLM07, CKKC00, CGT05, DG08, DGM07a, DGM07b, DWMB05, DLR04a, EB02, FRZB09, Fuk03, GS09, GT04b, HS00, Iar01, MS03, PZ06a, Sak06, Sak07, Sam07, SO00a, SO00b, SO02, SPVV00, WDMB05, YFP00, vdHJ04, RC01]. High-Dimensional [HS00, vdHJ04]. High-Energy [BGP04]. High-Order [FRZB09]. High-Temperature [BC02e, CKKC00, Sam07, RC01]. High-Velocity [ASB02]. Higher [ALS03, BBS03, CSM03, Dup03, Kol08, Len00, Sch08]. Higher-Order [ALS03]. Highly [KHV00]. Hilbert [Got05, Yoo07]. Hill [Leb07b, Leb07c]. Hilliard [MPSW00]. Histogram [KP01a, Lia06, LdOP00, LPdO01]. Historical [Opp04d]. Histories [BG00c, ER00]. History [Dom03b]. Hitting [Gal06]. Hohenberg [CE02b]. Hold [O’C06]. Hölder [Ruz00]. Holds [CY09]. Hole [Li06]. Holes [Dys05, HV01]. Holomorphicity [Car09]. Holstein [PA02b]. Homework [Hod01]. Homoclinic [Ler00]. Homogeneous [BCT06, BCZ04, CGR07, CS08b, FNP07, FS08, Fou06, FG08b, LVZ03, LMP05, Los05, Lu01, ZZ06, Sch08]. Homogenization [Fre06, HP04, Ach08]. Homology [Arg02]. Homopolymers [Per02]. Honeycomb [Bou07, CS08a]. Honoring [Leb08b]. Hookean [TE00]. Hoover [MMR08a, RKH00]. Hopf [EK00]. Hopping [BP08a, BD03b, CK06, DJ07, DPZ06, Woj03]. Horizon [SV07]. Host [dARBY04]. Hove [Var08b]. Howard [Ano02n]. Hubbard [BP08a, BD03b, DPZ06, HI04, HI05, LW07, Mas05b, PA02b, Riv02, SY08, Uel04, Woj03, Woj09]. Hückel [HRS03, TT05]. Hund [FU05]. Husimi [Luo01]. Hybrid [CPK07, CdO07]. Hydrodynamic [ATH06, AV87, AN00, AN05, BELM00, BCG00, BCG01, CLD05, EG00, Kol04, LCS08, MO09b, Mür01, PhD07, PS03b, PY09b, Sep01, Tid01b, Tid04, TV03, GC08b, AV03]. Hydrodynamics [Bob08, DR03b, Grm08, Kuk04, NSCW04, Sai00, SH07, UC07, ZGA02, Dor02]. Hydrogen [ABCM08, BGX01]. Hydrolysis [LLV09]. Hyper [Sum05]. Hyper-Cubic [Sum05]. Hyperbolic [AP08, AMP04, BD04a, BD09, Bob06, CO08a, CZ09, CNV01, Esc07, HV01, Var08a, Wol06, Wu00]. Hyperbolicity [Bun00, BDP02, SH05]. Hypergeometric [Ric04]. Hyperscaling [Sak02]. Hypersphere [Fra07a]. Hyperspheres [BW07b]. Hyperspherical [CG00b]. Hypoelliptic [HP04]. Hypothesis [BGGZ06, ECC03, Tal07]. Hysteresis [Bla05]. Hysteretic [SSD00].

I.M [Wei09]. Ice [Dys04, MTS04]. Ideal [Bak03, BP02, CL02, CLS02, FS07a, JS01c, Sam01a, Tél01, PY09a]. Ideas [ISS07, Pey09]. Identification [GR09]. Identifying [Cas02]. Identities [BM04c, CG07, CGG09, DRC04, War01]. IFS [FST06, JR08]. Igor [Git08, Pod08]. II [Roe03, ABC05b, AVZ00, AC01b, AR02, BM04a, BCEP06, BR04, Cas02, DP07a, DLR04b, DKRS02, EK04, FM05a, Fan05, HNO05b, JS01a, JR07, JS04, JKO^{+}01, Mes02b, MM06b, MR05, RC01, RK05, SSH06a, Sos03, Tak09b, WDMB05, Wu00, ZKBG03b, GCU02]. III [Dom04, Cai04, JSS03, KG03, SSH06b]. Ill [Rue02]. Ill-Formulated [Rue02]. Images [LK06a]. Immersed [Sam05a, Sam06]. Immune [CGS07]. Imperfect [BZ00b, BZ08b]. Implementation [CSP02, Ken02]. Implications [JH03]. Imports [Kad03a]. Imprecise [Gal08]. Improved [BR06, LMP05, Sla09a, FPS07]. Impurities [FT07]. Impurity [Sac04]. In-Depth [BCF^{+}09]. Inaccuracy [ZW04]. Incidence [BG01b]. Incipient [HS00, vdHJ04]. Including [Opp06b]. Inclusive [PPR00]. Incommensurate [Yar08]. Incomplete [FM01a, SD05, SD23]. Incompressible [CSN02, Fre03, Fre06, FPSW01, SKT09]. Increase [Bak03, BM05a]. Increasing [Jan06b, TV00]. Independent [Ano01d, FLS07, Mol02, SS05, Weh97]. Independent-Set [SS05]. Index [Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, DM06c]. Individual [DHB07, Kad06]. Individual-Cell [DHB07]. Induce [Noz04]. Induced [BG04a, BCC09, BMD00, Dog00a, FW03, FHAY06, Fre01, KS07, NV08, OvWLH00, Pia01, YCCN07]. Inductive [BZ00a]. Inelastic [BLR08, BCT06, BCL08, BCG00, BC03, BGP04, EB02, ETB06, Gar03, GA05, MMR06, MM06b, PT04]. Ineleastic [BCG01]. Inequalities [BDM09, BGH01, BGP04, ES04, Hol02b, MS06b, Sak02, TTK01]. Inequality [BLS04, GII08, HT01, Wee03]. Inequivalence [BB05]. Inertia [Bao00]. Inertial [SWK09, TC08a]. Infinite [Aki08, BZ09, BW07a, BP08a, BBP02, BD03b, CGR07, CRV01, DR05, Dim09, DLR04a, DLR04b, DPZ06, Dot04, GL04, HS00, Has00, HNO05a, Jab01, MO01, MSVE07, PS04b, Rei05, RK05, SV07, vdHJ04, Son09]. Infinite-Dimensional [MSVE07, DR05]. Infinite-Range [DPZ06, Has00]. Infinite-Range-Hopping [BP08a, BD03b]. Infinite-Type [BZ09]. Infinitely [BCM02, CM02a, CMS05, GS00, Mag09b]. Inflated [MMR08b]. Influenced [GMR06]. Influences [HPK^{+}07]. Information [AC05, Bog09, EMC09, EA04, GOBY06, GII08, HB08, KT06, Luo01, Luo07, LZ08, MN05, Pod02b]. Informational [LZ04]. Inhibition [BBD08]. Inhomogeneous [CL07, FJM01b, GJ08a, Lou03, Lou07, Lou08a]. Initial [BGGM04, BKLO00, BKOS07, BLP09, BFPS07, DG09b, DG09a, IS07, Pri03, Sim08, SDC04, TW09, Val09b, Val09a, WX06, Win02]. Initialization [Cai05]. Injected [Far02]. Injection [VPB^{+}06]. Insensitivity [Kom00a].

Instabilities [Bob06, DTP02]. Instability [GGZD02, Wag00]. Instanton [DW05]. Institute [Roe03]. Insulator [Noz04, LB04]. Insulator/ [LB04]. Integer [MP09, WMDB02]. Integrable [AK07, AR00, AG09, BP01a, CCAD08, Cso00, LS00b, RS05, WNUK01]. Integral [AL07, Bao00, CKKC00, MK04, TC08a, Tao01, YS00]. Integrals [AM04, Dub06, FEDZ07, Mas05a, Roe03, Ruz00, Wit03]. Integrated [Naj04]. Integration [AK07, Bao04, Sal09]. Intelligence [Kad03b]. Intelligent [Kad03b]. Intensity [GY05]. Interacting [AS05, Ano00h, BM04c, BCM02, CMS05, DFF02, FVE03, GKRV09, GS09, Kol04, Kol07, KP07, KP08b, LS02, LVZ03, LS09b, MP01, Skr03, Sla09a, Too00, Yar05, dSI04, vWR05]. Interaction [BT04a, BELM00, BP02, BFT02, CNS08, CGI02, CGG09, Dot04, FL03a, FZS^{+}00, GT06, Mas03b, SK06, Skr00, Var08b, Yan03]. Interaction-Flip [CGG09]. Interactions [BGMM03, BBDR05, BDM09, BCC06, BCG00, BC03, BGP04, BB05, CS04, CS09d, ES09, FdL06, Fuk00, GTA09, GM07b, HM00, JM00, Jia08, Kom00a, Kom00b, Kon02b, Lan09, Leg03, MR04, MR05, Per02, Pro07, Roz06, Süt02, TC03, WNUK01, dMG04, BCG01, Pro09a, Shl04]. Interactive [Emc04, New08]. Interest [Mas05a]. Interface [BLM03, Dir04, DFF02, GPT02, GG02, Pen08, PIRB03, VD01, dGNPR04, dHW04, dSWP05, vR01, BELM06]. Interfaces [BG04b, BGLT08, CdlL05, CGG06, CK03a, DMN00, Dun03, Gry01, MC05, MH07, MM05c, RK00, dMG04]. Interfacial [KM08a, KM07]. Interferencing [MV01]. Intermediate [Tid01a]. Intermittency [AMP06, BBT03, DR03a, JH03, MDL^{+}03, ND03]. Intermittent [TG02]. Internal [HDMF04, MM00]. International [Leb04b]. Internet [Bog04, Sza06]. Interpolating [Bar08]. Interpretation [CW04]. Intersecting [FR07, OEB01]. Intersection [KM09a]. Interval [AR06, ACL^{+}09, Gal06]. Intracellular [GS07c]. Intrachain [Per02]. Intrinsic [CD09a]. Intrinsically [CK03b]. Introduction [AFG07, Ano00b, Dag03b, DGM04, EvBC02, Mas00, Pod01a, Shl07, Vil06a, Wei05, Rit01]. Invariance [AR02, BM08, Ken04, LLSA00, Mat08, PS02c]. Invariant [AYP07a, AYP07b, BF05, Bir06, CLM08, DD04, EL02, EG00, Has02, Hor06, JBKN01, JR08, KK02b, KK02c, PSW04, Too07]. Invariants [BCO04, BV08, Len00, Spo06a]. Invasion [JHA07, dARBY04]. Invasively [Kie04b]. Inversion [Che09a, SB01a]. Investigation [BCC09, NSCW04]. Inviscid [CD04, Nad00, Sim08, Win02]. Ion [EE07]. Ionization [FMS04]. Ionized [ABCM03]. Ions [Mes02a]. Irreducible [Bar06b]. Irregular [KP04]. Irreversibility [Bro01, Gas04, Gas07]. Irreversible [HPCH00, KZ03b]. ISBN [Dor02]. Ising [Akt01, ABY07, Alb08, AD02, Ale01a, AY01, AYJP01b, AYP07a, AYP07b, BG01a, Bar08, Bax08a, BCF^{+}09, BCO07, Bjö09, BG09, BM02a, BBC^{+}01, BL08c, BC02e, CdlL05, Cam01, CZ03, CAD09, CR05, CP02, Cir02, CGT05, CSM03, Cou06, DM06a, DLP09, Dot06, Fis06a, Fis06b, Fis07, Fis08, FZ03, GHO^{+}00, GOS08, HL02, HRA^{+}04, HvdHS08, HN03, Kas02, KAS03, KZ07, KM01, LLSA00, LPE04, LW01b, Mas03b, MPS02, ND08, ONGP01, PAY09, RS08, RK00, RNVRP04, Roz06, Rut01, Rut08, SSD00, SS00a, SZ03, TPV09, WK00, Wu00, Wu02, Zhi00, vENS05]. Ising-Like [ND08]. Ising-Type [BBC^{+}01]. Islands [GS00, KVM^{+}00]. Isoenergetic [Rue00b]. Isokinetic [Rue00b, ZRA04]. Isolated [AN05, KL06, KL08, Too02]. Isospectral [GS04]. Isotherm [San05]. Isotropic [Hem00, Hos00, Lu00, Lu04, Yar08]. Isserlis [MNBO09]. Issue [Vil06a]. Iterated [Ano01d, BBS03, CN01, Weh97]. Iteration [Pre05]. Itinerant [MP01].

J [Aba00, AV03, Ano01d, Bog04, BPS03, Cau02b, Fre08a, GCU02, Han02, Opp02a, Opp02b, Pod01a, Pod04b]. J. [Fre08a, GC08b, Hod01, Rub09a, Rug08]. J.-P. [Fre08a]. J.-R. [Rug08]. J.V [GC08b]. Jacobi [KIK08]. James [Dom04]. Jamming [Rit06, JS08c, TB07]. Jams [GG01]. Jan [Ano01d]. Janossy [BS03, Sos03]. Jarzynski [GJ08a]. Jean [Dor02, Mas02a]. Jean-Philippe [Mas02a]. Jean-Pierre [Dor02]. Jellium [Dys04]. Jentsch [Pod07c]. John [Hod01]. Joint [Kül03]. Jörg [Mas02a]. Joseph [Fre08b]. Journal [Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Bax06, KK02b, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g]. Jr [GCU02]. Judged [ALM07]. July [Leb04b]. Jump [BMW09, EK05b, Mas09a, Zab06]. Jump-Diffusions [Mas09a]. Jumps [Cha09, Hor06, SB01b]. Jürg [Sim09]. Justification [dOP08]. Justin [Rub09a].

K-SAT [LP01]. Kac [AVE07, BLR08, BCM02, BPS05, CS02, ER07, Fou00, Fra07b, GR06, GF08, Kül01, Kup04, Ost07, PT04, SW08, WW06]. Kadanoff [Opp01]. Kalman [MN05]. Kampen [Rod02]. Kantz [Pod07c]. Kaplan [ECSB00]. Kardar [BTM^{+}04, SS09c, SMT03]. Karhunen [SS09e]. Karhunen-Loéve [SS09e]. Kasteleyn [Arg02]. Kauffman [QAK02]. Kawasaki [Ber05]. Kayser [GCU02]. Kazantsev [Vin02]. Keller [Opp04a, HS09b, HvdHMO09]. Ken [Dag03a]. Kernel [BP08b, Kra09, LCS08, SO00a, Shi06a, Yoo07]. Kernels [Str08]. Kerner [Git07]. Khruslov [Ach08]. Kikuchi [Pre05]. Killing [CGS07]. Kimball [DMN00, HK01, LFB04, Mes02b, Woj06]. Kimura [MPD09, PD06]. Kinds [BNZ02, Edw04]. Kinesin [MvW03]. Kinetic [AL01, AMT07, BDL06, Bal00, BELM00, BCG00, BCG01, BV08, BM02a, CT07, CIS01, Cer05a, CPT05, Del05, DR01, Dog00a, FPS04, FZS^{+}00, GC08a, GW00, GP04, HR06, HS09a, HD02, KCK07, KL06, KL08, LS09b, MT08, Nou02, NOV04, Opp04d, Pan02, PT06, PO07, Pol00, SNC05, SB01b, Str06, TJ08, Tid01b, Tid04, WW05, ZK00, ZGA02]. Kinetically [TB04, TBF05]. Kinetics [Ano01c, Buc07, FSG08, Kup04, MR01]. Kingdom [Bra02]. Kinzel [KK00b]. Kirkpatrick [CC09, CGT05, Led00, Pan08, WB04]. Klages [Wei09]. Klein [CP02]. Kleinert [Hol02a, Wit03]. Klinkenberg [PO07]. Knots [OSv04]. Knotted [Hol02c]. Knudsen [ABT03]. Kohn [PN03]. Kolmogorov [CD05, HD09, Kor00]. Kondo [LK06b, Mal07]. Kontorova [CM05]. Korteweg [BKLO00, Gar01]. Kosterlitz [GTZ02, MM05b]. Kotecký [Sch06]. Kotecký-Shlosman [Sch06]. Kottegoda [Mas09b]. KPZ [ALS06a, PS04a]. Kraichnan [AH07, EX00, Fan04, FJMR02, GH04, KMG07, Vin02]. Kramers [Fre04, Luc08b, Ste04, TTNK06]. Kronig [Luc08b]. Kubo [DB02].

L [Spo03]. L. [Gar08]. lacunary [HO09]. Ladder [BGOY04]. Ladders [BdGM01]. Lagrangian [KK02b, BCG08, BZ03, CGH^{+}03, CG09, HPWW03, HPCH00, KK02c, MDL^{+}03, PS03b]. Lagrangian-Averaged [HPWW03]. Laguerre [DGKV07]. Laguerre-Type [DGKV07]. Lamellipodium [JAWC08]. Laminations [dlLV07]. Landau [Rap01, BR04, EMO08, FJMR03, GM03b, WBE02]. Landauer [KN03]. Landscape [BCF^{+}01, FW07]. Landscapes [HNOV07, Pod07a, Yip06]. Lane [BKNS01, JNH^{+}09]. Langevin [Bao00, Bao04, BL06, EK03, Far02, Gra00, JAWC08, Mag09b]. Langton [Boo01]. Laplacian [Lou08a, FPD01, Lo08, Lou03]. Large [Ano01d, AP06, BH08a, BBDR05, BGL05, BS09a, BD06c, BL08a, BET00, BZ08b, CL00, CE02a, CGV02, CLM05, DM08, DLS02, DLS03, EK02, EK04, EHT00, ED04, EMH04, GLM02, Has00, KNST09, LS00a, LS02, LRB05, NR04, Par08, PS03b, SJ02b, SM06, Shi06a, Tal07, Tin03, Too02, Too08, VEW08, VB07, Wat04, vWR05, Weh97]. Large- [Has00]. Large-Distance [SJ02b]. Large-N [SM06]. Large-Time [CGV02]. Largest [BFP08, NNR09, Sos02, vZvB02]. Lars [Ano04m]. Laser [EK00, ZF09]. Last [BJ06, FR07]. Late [Hod01]. Lateral [HM00, QLAS02]. Lattice [Aba00, AFN00, Ano00g, AK02, Asa00, BP01a, BDG^{+}06, BD02, BL08a, BCKM00, Bro01, BMD00, Cai05, CPK07, CGGP04, CJS01, CSS02, CS03a, CJSS04, CS08a, CS09b, CL06, CSP02, CSN02, CM02b, CRV01, CN01, DMS02, Del05, DM06b, DLS07, DA03, DH04, EE07, FK02, FR07, GLMSR07, GZ01, Giu09, GT06, GM00, HD02, HN03, HRK^{+}05, IVZ06, JS01a, JSS03, JM00, Jen01, Jia03, JY05, KPT05, Keh05, KTH^{+}05, KMP^{+}07, LV01, LRY05, LRB05, LPE04, Liu09, MO09a, Mas02b, MB02, MS02, MZ02b, MMR08b, MS08a, MGT07, MO09b, Mun07, NOS06, Nag03, NDC02, NSCW04, ONGP01, OC05, PVV02, PhD07, PBFC05, PG99, PG00, RSMJ^{+}02, RS09, Rhe05, SSD00, Sak00, Sak06, Sak07, SS01, SS09c, SO00a, SO00b, SO02, SS05, Sea02, SK07, Ser06, Shn03, SNC05]. Lattice [Sid05, Skr00, Spo06b, Spo06c, SAPD05, SSK02, SFCO02, TPV09, TB04, TBF05, TK02, WP02, Wu02, WP03, YHD02, Yar05, Yep02, aYL05, Zha00b, ZQS^{+}05, vWR05, Dor02].

Lattice-Boltzmann [DA03, HRK^{+}05, LV01, Mas02b, NDC02, RSMJ^{+}02, Rhe05]. Lattice-Gas [DMS02, LPE04, PG99, PG00, Sak00, Yep02]. Lattice-Gases [BD02]. Lattices [BBC^{+}01, BK01b, BC02e, CS09a, CNV01, CSM03, DR06a, Elo08, Fin03, Jus01, LL00, LMZ00, Luc09b, MS06a, Par08, PVV02, Pre07, RNVRP04, RD08, Rug08, SJ01b, Sum05]. Laudatio [Leb04b]. Law [Aki08, Ano01d, BZ03, BO05, BLL04, DMSR03, DT00, GPL04, Lin08, Man03, ND03, O’C06, Sod09, Sod17, Too08, WA09, Weh97]. Laws [BLP09, KMG07, LRM06, MR01, NOV04, SF07b, SV07, Tid04, TV03, WX06]. Lax [Mas08]. Layer [ABT03, CRT00]. Layered [Ewo01, Fre08a, Sak04a]. Lead [Fra07a, Fre03]. Leading [MZ02b]. Learn [BCK00b, DHB07]. Learning [Bal09, Kan01]. Least [Kie04b]. Leaves [SS00b]. Lebesgue [CK03b, FW03]. Lebesgue-Measure [FW03]. Leblond [Pod04b]. Lecture [Pod01a]. Lectures [WW05]. Lee [Opp02b, BG01a, Car03, Jaf03]. Lees [WP02]. Legacy [Kie04a]. Lemma [SS05]. Length [FM05a, Fra07a, MS06b, San05, Too04, Zab06]. Lengths [Che05]. Leo [Opp01]. LERW [BBK08]. Lesne [Opp09]. Let [BM01b]. Letter [RT08b]. Letters [Dom04]. Leuzzi [Gar08]. Lev [Rub04]. Level [BG00a, BG04a, Gen04, MS05a, Yip06]. Levels [BR04, EGGI01, NSS04]. Lévy [Bar04, CGK^{+}04, DM06c, EK05b, Kol07, MB06, PS08b, CD09a, EK03]. Lévy-Driven [EK03]. Lévy-Process [CD09a]. Liapunov [CE04]. Library [BC02e]. Lie [Nam04]. Lieb [Ano03n, ES04, Häg07, Han06, Leb04b, RS09, Tia04, Yng04]. Liesegang [Dro00, HvdHMO09]. Life [Mer03, Shl07]. Lifschitz [FJMR03]. Lifshitz [CGV02, LM02, Naj07, Naj08, Vel00, Zhi00]. Light [GG00]. Like [BBS04, Bjö09, BG06, FK00, KL03, LR02b, MNP08, Mis06, ND08, PT06, Rue02, SLA02, Tid04, TM08]. Like-Charged [SLA02]. Limit [AGY02, And08, AN00, BTT07, BGW01, BJ08, BFP08, BM02b, CR07, Cam05, CFN06, Car07, CD09b, CC09, Che05, CS08b, CLD05, CG04, CS09c, CEO07, DVE07, DR01, DGR08, Dim09, Dir04, Dog00a, EA04, FF04, FP04, GTW01, GPL02, HR06, HS00, HMH05, HS09a, JSS05, KS02, Kie09c, Kol04, Kuk04, KM00, Lan09, LP01, LCS08, Lou03, Lou08a, LS09b, MDB00, Mür01, NNR09, Ost04, Ost07, Pen01, Pey09, Rid03, Rit06, Rue00b, Sac05, Sam07, Ser06, Shi09, Süt03, SV07, Tid04, TV03, WW01, Win02, vdHK08]. Limitation [EA07]. Limitations [HDMF04, vWL00]. Limited [AFN00, LK09, MM00]. Limiter [WS02a]. Limiting [BR00a, BKLO00, dBR08, DGM07a, DGM07b, GW00]. Limits [Ano00h, AN05, BK02, CJM09, CEFM00, CS03a, CS03b, Cso00, Fan05, GJ08b, GJ08c, JMU05, Kra03, Lou07, Mor07, Tid01b, Var08b, WX06]. Lindblad [Bud08]. Line [BL09, BCKM00, BC02e, CMN06, DS07, FVE03, GLMSR07, GS04, HC08, MM08, NOB00, PTZ03, PH01, Sui00, WP03]. Line-Graphs [PTZ03]. Linear [AL07, ADLM01, AVBM06, BMW09, Bar04, BKLO00, BKOS07, BM00, BMN07, DR01, Erd02, ECSB00, GZG05, JOP06, KP06, KL06, KL08, LT04, Pod06b, Ryc08, Skr00, Skr03, SD05, SD23, TG02, Yos08]. linearity [Eps06]. Linearized [CT07, Sha10, Tak09a, Tak09b, Tak10]. Link [Spo06b, Spo06c]. Linked [Cam01]. Linked-Cluster [Cam01]. Liouville [BHPH01, CNV01, Dog00b]. Liouvilleans [JP01]. Lipid [GS03b]. Liquid [CK04, DTY02, DRC04, DD06, GPT02, HD03, KW00a, Mas05b, PIRB03, SMRK^{+}02, SBG03, dMG04]. Liquid-Vapor [dMG04]. Liquids [Bol09, FM08, Git09]. Lisbon [Leb04b]. List [Ano01q]. Littlefield [Kad03b]. Lobachevskij [Mez01]. Local [AS06, BBD08, BFT02, BK07, DR03a, GJY04, LR02a, LCS08, RY07, SS05, SDBS07, Too08, Tah09]. Locality [Ber07]. Localization [And05, BLM03, BG04b, BK01b, CdO07, Che05, DG00, Eps06, FLM00, FKS08, GK06, GL04, GM00, KBS02, KP08a, Nak06, WZ09, Yip06]. Localized [dHP09]. Locally [Too00]. Locating [PF07]. Locking [Bar06b, KM08b]. Loéve [SS09e]. Loewner [BBK05, Ghe09, GK04, KNK04, Ken04, Ken07, Ken08, Ken09]. Log [FJM01a, Sam01b]. Logistic [FJ05]. Lognormal [Ost04, Ost07]. Long [BBDR05, BELM00, BCC06, BP01b, BDLvW08, BP08a, BC09, BB05, BCS07, BCM02, CT07, CM02a, CMS05, DMP05, Erd02, FF00, Fin03, FW06, Gar01, Giu09, HvdHS08, KBS02, Lu00, MP01, Mis06, Naj04, Nol04, PDvB00, RS08, SS04, Shl04, Skr03, Sme08, SZ03, Tél06b, Tor04, Tur03, WNUK01, WZ09, dZS04]. Long- [HvdHS08, Tél06b]. Long-Distance [BP01b]. Long-Range [BBDR05, BCC06, CMS05, MP01, Nol04, RS08, Shl04, Skr03, WNUK01, Mis06]. Long-Ranged [dZS04]. Long-Term [FF00]. Long-Time [BC09, Fin03, FW06, Gar01, KBS02, Lu00, SS04, SZ03]. Long-Time-Tail [PDvB00]. Longest [AMV08]. Loop [DFM01, FLW08, Gra09, HL08, KM07, Law09]. Loop-Erased [FLW08, Gra09, HL08]. Loops [CN04a]. Loopwise [Cai04]. Loose [AR09, CM04c, OSv04]. Lorentz [BHPH01, BBP02, BDL00, BK03b, CG00c, CEMM09, KD00, KPvB06, LLMM03, PDvB00, RKH00, RKN00, SV07]. Lorenz [JW00, Luc09a]. Loss [CRV01, FM05b, KT06, Kom07]. Loss- [Kom07]. Lotka [MGT07]. Lovász [SS05]. Low [ABCM08, AH01, AMP06, BM04a, BCEP06, BW07b, BFT02, BM02b, CF03, CLM07, CR05, CHNO06, DLR04b, GM07a, KS02, Lee05, SY08, TTM^{+}01, CR01]. Low-Density [KS02]. Low-lying [Lee05]. Low-Mass [BFT02]. Low-Temperature [ABCM08, BM02b, CR05]. Low-Volume-Fraction [CHNO06]. Lower [BR06, BK01a, CLT07, MS01]. Lowering [DN03]. Lowest [Shn03]. Löwner [KN04]. LRO [Skr00]. LSW [HNO05a, HNO05b, NP01, Vel08]. Ltd [Roe03]. Ludwig [Ano00e]. Lunch [Kad03b]. Luttinger [Ano01q, FM08, Gal01, Hod01, Mas05b, Mat01, Sha01]. Lyapunov [BJ02, BDP02, DR03a, DJZ00, EG00, EFPZ05, Gam03, GvB02, HPF^{+}02, IK09a, KPvB06, PDvB00, Pan02, RM08, SJW07, STV02, TDM02, Wag00, vZvB02]. lying [Lee05].

M [Dom04, Han02, Hod01, Mas03a, Opp02a, Opp09, Pod04a, Roe03, Shl04, Wei02, bA09]. M. [Bog09]. Machlup [Sin08, TC07]. Macroions [MNP08].

Macromolecular [SO08]. Macromolecules [Hem00, Mez01]. Macrophysics [Spo03]. Macroscopic [BDG^{+}02, BDG^{+}09, DMN06, LS00a, LS02]. Made [Opp04c]. Maejima [Taq04]. Magic [Rei00]. Magnetic [BLS04, BEPK^{+}02, BCZ04, BR04, CS09b, CGT05, ES04, FZ03, HY04, KMSS01, LW07, Sac04, dPG04]. Magnetism [AFFS06]. Magnetization [CSZ00, ER07, EMO08, Kas02, Nic01, PH01]. Magnetohydrodynamic [AH07]. Magnetostriction [SZ04]. Magnetotransport [BHPH01]. Magnets [FBG02, FRZB09]. Majda [CLM08]. Major [Lan06]. Majority [dSFM02]. Majority-Vote [dSFM02]. Makoto [Taq04]. Man [Ano00e]. Management [Mas02a]. Mandelbrot [Mol02, SM01]. Manifold [DW05, GS04]. Manifolds [MAOB01]. Mantegna [Mas00]. Many [AH03, CMS05, CEMM09, Gal01, GS00, HC08, HPF^{+}02, KW00b, PA02b, TDM02]. Many-Body [HPF^{+}02, KW00b]. Many-Particle [AH03, CEMM09, TDM02]. Many-Polaron [PA02b]. Many-Sided [HC08]. Map [AS03, Bal00, BBC^{+}01, FJ05, Jus01, Rug08, SJ01b, TG00, TG02]. Maps [AP06, Bal00, Dys05, FW03, GK04, Hao05, Hat07, JBKN01, KS00, KKS05, PZ06b, Var08a, HV01]. Marc [Mas02a]. Marchenko [Ach08]. Market [CPT05]. Markets [CPP09, Por04, Por06, Wit03]. Markov [Ada01, All08, BMW09, BI06, BS08, CU03, DVE07, DZ01, DR06b, FGC09, FM01c, GJQ06, KS09, Kol07, LARvW07, MvW03, Rom08, Sch01, Sch06, Sch08, SC09, VB08, ZKD05]. Markovian [AL03, Bao04, BL06, Bud08, CD04, CDA09, DD09, Fin03, Gas04, Gas07, MR01, PM05, vWL00]. Marro [Aba00]. Martingales [BBK05]. Mass [BTT07, BFT02, CD09b, HD03, Jab01, Kom07, LL02, Mür01, NP01, RM00, Sak07, Sui00]. Mass-action [Kom07]. Massive [CAD09, CL02, CLS02]. Massless [Sak09]. Master [AL03, DDM08, KCT08, Los05, vWL00]. Mastroprieto [Rub09b]. Matching [Cag02, CR05, FKLM08, Mun07]. Material [BC02b, Cer01, Ram07]. Materials [BCT03, GAD06, Vil06b, Sah02]. Mathematical [Aus00, GMW02, HS08, Hep09, HvdHMO09, JP02, Leb04b, WE07, Yng04, Zwe05, dOP08, Ton08a]. Mathematician [Mau08]. Mathematics [Vil06b]. Mating [Nor09]. Matrices [AFNvM00, AK07, ABS01, BG01b, CAC02, CS09b, CCM00, ES07, FM09a, FM09b, HKT01, HJ05, JS01a, JSS03, JS06, LP08, LP09, PS07b, SS01, SS09b, Shc09, Sin09, Sod09, Sos02, Sod17]. Matrix [AS06, Bax04b, Bax06, Bel04, BFP08, BP08b, BEK^{+}07, CGI02, CDG07, EFM09, FW07, LB04, Luo01, LP09, MZ02b, Nag07, NNR09, PS08a, RS05, Ric04, SO00b, SO02, SM06, Str08, WS02b, XKHK08, dSI04]. Matter [AR09, Han02, Mat01, Pom05, Rad08a, Rit03, Opp06b]. Mattis [Opp04c]. Maximal [BL08c, vDHR06, MC05, Zha00b, vdHK08]. Maximization [Att00]. Maximum [BET00, EK06, Kaw06, MCZ06, MK04]. Maxwell [BCT06, BC02d, BG06, FM01c, Gar03, GA05, Jan00a, Dom04]. Maxwellian [BD03a, BC09, CL03]. May [Leb04a, Leb05c, Leb06c, Leb07c, CRVV09, CU03]. Mayer [BMA02]. Mazo [Wei02]. McKean [GR06]. Mean [Bar06b, Bar08, BCC06, CR07, CGI02, De 07, DG08, DGR08, DLP09, DMP05, ES09, FT06, HvdHS08, Kad09, KHV00, KS06, Kül03, KC03, Lan09, LW07, LPE04, Nad00, NS02, Pey09, RT08c, Rát09, Sak01, Sak04b, Sla09a, Suw09, ZS02, dARBY04, dPG04, AC01a, CW04, DWMB05, Sak05]. Mean-Field [BCC06, CR07, DGR08, HvdHS08, KS06, Lan09, LPE04, Sak01, Sak04b, Sla09a, ZS02, AC01a, DWMB05, Sak05]. Mean-Motion [Nad00]. Mean-Spherical [KHV00]. Means [DN04]. Measurable [JR08]. Measure [Aki08, BBS04, CLM08, EFM09, FW03, HS09b, Hor06, JBKN01, KK02b, KK02c, Law09, LTWW02, Mol04, Rei08, Too07, ZZ06]. Measurement [DGZ09]. Measures [BS05, BF05, BCK04a, Bir06, BCL^{+}07, CLM08, CT04a, Cha00, CU03, DD04, EL02, FLR03, GT02, HO09, JR08, Jia03, JL00b, Jun03, KP07, KP08b, Kül03, KM00, LM01, Mec01, Mol02, MR04, MR05, RS04a, Roz06, Sak00, Str08, Wol06, You02, vEK07]. Measuring [BVBF01]. Mechanical [BCC09, HK02a, KNST09, MH07, Par08]. Mechanically [SSS07]. Mechanics [Ano00x, Ano00y, Ano01u, Ano01v, Ano02n, Ano03m, Ano03n, Ano03o, Ano04m, BD06a, BBK05, CP03, CMSR05, Dag03b, Dom03b, Dud07, DGZ09, ES03, Gal08, GS07b, GR09, HPCH00, HB00, JP02, Kan01, KNP07, Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Leb07b, Leb07c, Leb08a, Leb08b, Leb09b, Leb09c, Luc08b, Mas02a, MV07, New08, Opp02a, Opp04c, PV02, Pod01a, Raj09, Roe03, Rue00a, Rue02, SC01, Spo03, Swe02, Tah09, Wit03, dlLV07, vK04, Ton08a, Opp03b, Opp09, Opp07]. Mechanism [Son09, SBSAD07]. Mechanisms [CRVV09, FK09a]. Medal [Leb04b]. Media [BNZ07, Ewo01, Fan05, GOSG00, PO07, RS04b, Sam09, SvL02, VD01, Fre08a, Pod06b]. Mediated [JHA07]. Medium [BMP07, CS08b, Cor05, MO09b, dHW04]. Meester [Bog09]. Meeting [Ano02n, Ano03m, Ano03n, Ano03o, Ano04m, Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Ano00x, Ano00y, Ano01u, Ano01v]. Mehta [BR05b]. Meijer [Pod04b, Rod02]. Melt [MS02]. Melvin [Mas08]. Membrane [EE07, HS04, SWF07]. Membranes [Koi07a]. Memories [GL01a]. Memory [EMMZ06, FM05b]. MEMS [NDC02]. Mendes [Bog04]. Merits [HDMF04]. Meshkov [GGZD02]. Meshwork [Koi07a]. Mesoscopic [CPP09, DDMSR04, GS07c, KPT05, KLL05, WHI02, dMG04]. Messy [Rue02]. Metal [Noz04, SJ01a]. Metallic [FU05]. Metals [MN04]. Metastabilities [KS07]. Metastability [BM02b, BdMR03, Cir02, CN03, CNS08, GS01, MNOS04]. Metastable [BS09b, HMG08, LPD08, MO01, TNK04]. Methanol [PIRB03]. Method [AK02, Asa00, BBLS05, Bol09, CSP02, DA03, EK06, FBG02, FRZB09, FM09b, Fuk03, GK09a, JHW09, LB04, Lo08, MAOB01, MP03, NAS02, Pan08, PF07, Pre05, Rhe05, Roz08, Saa07, Tia04, WS02b, YHD02, ZQS^{+}05]. Methodology [Suz03]. Methods [Ano00c, KLL05, LdOP00, VB08, WE07, KP01a, LPdO01, Opp03a]. Metric [San05]. Metrics [BCT05, BCT06]. Micellar [Mez01]. Michael [Ano03m, bA08]. Michele [Mas09b]. Microcanonical [CETT05, KPH00, KP01b, Kie09c]. Microchannels [ZQS^{+}05]. Microchip [EK00]. Microflow [SNC05]. Microflows [SSK02]. Microgravity [AC03]. Micromagnetics [KPT05]. Microphase [CR03]. Microscopic [Bal01, BR05a, BCF^{+}09, BO05, BD01, BM05b, CG00c, DC00, JS04, Mal05, MDB00, Sam00, Zab08]. Microswimmers [PY09b]. Microwave [FHAY06, SL02]. Microwave-Induced [FHAY06]. Migration [QLAS02, SBSAD07]. Mills [GW09]. Milne [AN00]. Mimicking [GAV00]. Min [KMP^{+}07, Mas08]. Min-Protein [KMP^{+}07]. Minimal [BP01a, CdO07]. Minimum [BDG^{+}04, MBL08, PZ06a]. Minus [KM01]. mis [Por06]. Missing [Jan06b]. Mixed [AY01, BdGM01, BGOY04, BCS07, BCS09, EHT00, LB04, MNBO09, NKC00, RR01]. Mixed-Range [BCS09]. Mixed-Transfer-Matrix [LB04]. Mixing [AP08, BK09, DKRS02, FM05b, GGZD02, KO05, PS06]. Mixture [CC00a, YS00, BELM06]. Mixtures [AAP02, ABT03, BG06, CMSR05, CC00b, FHL03, FV02, GCU02, Gar03, GA05, GM07a, Hem00, OP04, PP01, PPR00, PIRB03, SK06]. Mobility [DG01, KO05]. Möbius [CP02, PM04]. Mode [FV02, GvH07, HD03, KK02a, KM08b, RM08, Yip06]. Mode-Coupling [FV02]. Mode-Locking [KM08b]. Model [ADG00, Akt01, AD02, Ale01a, AY01, AAP02, AVE07, AN00, ADLM01, AYJP01a, AYP07b, ACL^{+}09, BR00a, BJ00, Bal01, Bal04, Ban05, BG01a, BBBF06, Bar08, BLR08, Bax00, Bax03, Bax04b, Bax05, Bax06, Bax08a, Bax08b, Bax09, BKNS01, Bel04, BHS07, BBS04, BG04a, BO05, BBD04, BBD08, BGL05, BD03a, Bir07, BCT06, BCK00a, BK01a, Bjö09, BG09, Bla03, BL09, BHPH01, BM02a, BP08a, BMP07, BCKM00, BKM02, BC09, BC00, Bou07, BL08b, BL08c, BM00, BNZ02, BD03b, Các08, CKSZ06, CLM08, CN04b, Cam01, Car06, CAD09, CRVV09, Cha09, CSS02, CJSS04, CS08a, CS09b, CC09, CCIL08, CR05, Che05, CP02, CK04, Cir02, CGK09, CGT05, CPV08, CLT07, CMN06, CPT05, CSM03, Cou06, DMN00, DW05, DM06a, DM08, DFM01, DM09, DMP^{+}04, DLP09]. Model [DLR04a, DLR04b, DSZ07, DMP05, DPZ06, Dot04, ER07, EG05, EX00, FM01a, FM03, FM05a, FM09a, Fan04, FPS04, dSFM02, Gal01, GLMSR07, GRU07, GTA09, Gat00, GH04, GJQ06, GAV00, GS07a, GS03b, GPM^{+}05, GG00, GG06, GOS08, GK02, GT04b, GR00, HRA^{+}04, HT01, HK01, HD03, HS09a, Has02, Hei04, HDMF04, HWB01, HvdHS08, HvdHMO09, HI04, HI05, HN03, HK02b, HAG02, HMG08, ICO02, Iof02, JSS05, JR07, JL01, JAWC08, Kas02, KS06, KPT05, KK02a, Kaw03, KCK07, KRT00, KAS03, KK08, KW02, KM06, KZ07, KK00b, KM09b, KTH^{+}05, KM01, KMVE03, KMP^{+}07, KR06, Kül01, KMG07, Kur03, LLSA00, LW07, LMP05, Led00, LFB04, LRB00, LPE04, Liu09, LK06b, LR06, LW01b, LD01, MvW03, MKB00, Mal07, MO01, MCG08, Mas02b]. Model [Mas05b, Mat01, MPS02, Mec01, MZ02a, Mes02b, MM06a, MS02, MV03, MZ08, MO09b, Nad00, NAS02, Nak06, NP01, ND08, Nor09, ONGP01, OBB03, Pan08, PS02a, Pat07b, PAY09, PdKV09, PBFC05, PG99, PG00, Pri00, PS06, PA02b, PVZ05, PT04, PY09b, QAK02, RS08, RRS^{+}00, RGRTS06, Riv02, Roz06, Rut01, SSD00, Sak00, SS00a, Sam01a, SJ01a, SI04, Sch07, SMS09, Sea02, SK07, SY03, SY08, SEZ05, Shi06b, SNC05, Sig09, SvL02, SB01b, SZ03, SLA02, Suw09, TTK01, TM02, Tid01b, Tid04, TPV09, Tor04, TK02, Tum06, TE00, Uel04, VD01, Vel00, Vel08, Wat04, WD01, WK00, Woj03, Woj06, Woj09, WB04, WK04, XKHK08, Yan03, Yep02, Zab06, Zah02, ZKBG03a, ZKBG03b, Zhi00, ZF09, ZJ00, dARBY04, dGNPR04]. Model [dMG04, vENS05, vSC00, JR06]. Model-Based [CRVV09]. Modeling [BEPK^{+}02, GS07c, Hol07, KPT05, KTH^{+}05, VTM00, YCCN07, PhD07, Shl05]. Modelling [CPP09, Emc04, GAD06]. Models [Aba00, AH03, Ano00g, AMT07, Arg02, AYJP01b, AYP07a, BZ08a, BS07, Bax02, Bax04a, BP01a, BCO07, BV08, BIV01, dBR08, BCFP05, BZ00a, BCS07, BCS09, BR04, BD05, BC02e, CdlL05, CZ03, CCAD08, CM05, CF03, CK03a, CSZ00, CL06, CG04, CC00b, Cor04, DMS02, DMPV08, DMPV09, De 07, DMR05, DGM07a, DGM07b, DP03, DH04, DHB07, EGGI01, FF04, FKÖ03, FK09a, FR07, FT06, GZ01, GG02, GM07b, GTW01, GP04, HK02a, HD02, vDHR06, Hos06, Ito02, JS01a, JSS03, JS06, JKO^{+}01, KKT00, Koi07a, Kol04, Kom07, Kül03, Kup04, LRY05, LZGM04, LL02, LY07, MP01, Mal05, Mas03b, MT08, MRTZ03, MGT07, MR04, MR05, MPD09, New00, NSCW04, PT06, PD06, PS08a, Pey09, PS02b, Pro07, Pro09a, RM00, RY07, Ric02, RG04, Saa07]. Models [SD09, SS01, SS09b, Sam05b, SCM04, Shl07, Sla09b, Tid01a, TBF05, TV02, Wu00, Wu02, aYL05, dlLV07, Git07, Ton08b]. Moderate [BW07b]. Moderately [Coh09, Coh10]. Modern [Git08, Lan06, Opp05, Pod08, Roe03]. Modes [EG00, EFPZ05, FGR03, Gra00, HPF^{+}02, OT03, Raw03]. Modifiable [Fei03]. Modified [Kar07, LLMM03, Lu00, Lu01, Sme08]. Modular [HMH05, KZ03a]. Modulated [GTA09]. Modulation [DB00]. Molecular [Ano00b, Ano00f, Buc05, CLK09, Har08, HPCH00, HHB09, Jan06a, KL03, KC03, LNT09, Li07, LJ03, LL08b, LL09b, LLV09, MKL08, Pod07a, Ric00, YFP00, Zha09, Dag03a]. Molecule [SKM04]. Molecules [BD03a, BC02d, CL03, FM01c, SWF07, vK02, KIBG^{+}06]. Moment [BDL06, BC02b, BGP04, DR03b, JS01b, KMST00, Kar07, Str06]. Moments [BG01b, Lin04, Mec01]. Momentum [Mür01, PO07]. Monatomic [FS07a]. Monday [Leb07b, Leb07c]. Monolayer [GS03b, Pen01]. Monopole [Li06]. Monotone [FV02, GI08]. monotonic [ATH06]. Monotonicity [CG04, Kie09a]. Monte [Rap01, AER05, BLM03, BW07b, BR01, CG00a, FPS00, GM00, HRA^{+}04, HGST00, Ken04, Lia06, RT08a, RNVRP04, SD05, SD23, SW08, WS02b, Ano00c]. Morphology [BPS06b]. Mortal [LK09]. Mossotti [BM01a]. Most [ER00, Par08]. Motility [MKL08]. Motion [Bao01, Bir07, BdMR03, CL02, CD08, DTY02, FL03a, GPL02, HB08, KT07, KM09a, LNM06, MK04, MG02, Nad00, Ruz00, Wei02, dG05a, dG05b, BELM06]. Motions [DM06c, FK00, GF08, Kol08, Shi09]. Motivated [AFFS06]. Motor [KC03, LJ03, WLTH07]. Motors [CLK09, GJQ06, KL03, LL08b, LL09b, LLV09, MKL08, Sla09a, WE07, Zha09].

Moussa [Häg07, LS04b]. Move [Boo01]. Moves [SDBS07]. Moving [CO08a]. Moyal [Fan05]. Multi [AS02, BZ08a, BBS04, De 04, DHB07, EA04, FF07, FH04, FLM00, Iar01, KK00a, Kon02b, Mas02b, Sak04a, dHW04]. Multi- [FH04]. Multi-Cellular [DHB07]. Multi-Component [BZ08a]. Multi-Dimensional [FLM00]. Multi-Fractal [AS02]. Multi-Information [EA04]. Multi-Interface [dHW04]. Multi-Layered [Sak04a]. Multi-Overlap [De 04]. Multi-particle [Iar01]. Multi-Scale [BBS04]. Multi-Species [KK00a]. Multi-Speed [Mas02b]. Multi-State [Kon02b]. Multi-Type [FF07]. Multibaker [TG00]. Multibody [BP02]. Multicanonical [KP01a, LdOP00, LPdO01]. Multicolor [DH08]. Multicolour [DM06b]. Multicomponent [MPSW00]. Multicritical [Aha03]. Multidimensional [GJ08a, LRM06, Tid04]. Multifractal [BS08, BB09, Ber03a, Hos00, Man03, Ost04, PR07]. Multifractality [CNV01, Dup03, FM02b, FKM05]. Multifractals [Ost07, TB06]. Multigrid [BR01, TKR02]. Multigrid-Solver [TKR02]. Multilayer [DR06b, HSA00]. Multilinear [LP08]. Multiparameter [KKS05]. Multiparticle [PG99, PG00]. Multiphase [HD02]. Multiple [AN05, Bal04, BBK05, CEO07, KM09a, Mur08, ZKBG03a]. Multiplicative [Bao01, BR05a, BBT03, Piv03, SM06]. Multipliers [CE04, Suz01]. Multiscale [DSC02, DVE07, GAD06, PS07a, SO08]. Multiscaling [BC01]. Multispecies [EFM09]. Multitype [GZ01]. Multivariate [SM06]. Murad [Nol04]. Music [Zwe05]. Mutation [CF03]. Mutations [Kom07]. Mutual [GOBY06]. Mutually [CMS05]. My [Mer03, Sin02, Swe04]. Myxococcus [SBSAD07].

N [Bog04, Han02, Mas00, Rod02, SM06]. N.T [Mas09b]. Nagle [Swe04]. Nano [KC02]. Nano- [KC02]. Nanomagnet [Mat04]. Nanoscale [FL06, HM00, KVM^{+}00]. Nanoscopically [AB09]. Nanosystems [Uma09]. Nanotubes [BVBF01]. Narrow [SSHE06, SSH06a, SSH06b, KIBG^{+}06]. Nary [Kol04]. Natural [Aus00, LK06a, Pre05, Rue00a]. Naturally [CRVV09]. Nature [DJ04, JPZ09, Kül01, Luc08a, Mas09b, Pod07c]. Navier [Sin05, Bak06, BKL00, Che09b, CLM05, DGR08, DLS09, FJMR02, GM07a, HPWW03, KP05, Kuk06, LL08a, Mat02, MS05b, Rom04, Rom08]. Near [AC01a, AC01b, AC04, BG04b, CFN06, CJM09, DA03, DOS01, EMO08, HGST00, Jia07, LD01, FN00, Ler00, Wu09]. Near-Critical [CFN06, CJM09]. Near-Planar [DOS01]. Near-Vacuum [Jia07]. Nearest [BBD04, CL07, DF02, DH08, DJ07, MPS02, Skr00]. Nearest-Neighbor [BBD04, DH08, DJ07, Skr00]. Nearly [Ale01a, EK00]. Necklace [DDH06]. necks [KIBG^{+}06]. Need [Adi04]. Needles [HRS03]. Negative [BR00b, CM04c, GK02, Hos06, Lee09, Man03, O’C01, Riv04]. Neighbor [BBD04, CL07, DH08, DJ07, MPS02, Skr00, TC03]. Neighborhood [CEO07, SDBS07]. Neighbors [DF02]. Nelson [Rit03]. Nematode [MV03]. Net [TW03]. Netherlands [GCU02]. Nets [Bog04]. Network [AH03, AG07, BB03, GPM^{+}05, HAG02, LL08b, LL09b, OEB01]. Networks [Ano01b, Ber03b, BK04, Bog04, Bog07, Bog09, CPK07, dFCS06, DOS01, DS07, EZ04, FFN09, GGL09, LB04, LJ03, MNV03, Paj07, RSV09, SMS09, VB07, VB08, bA09]. Neumann [Cas01, Cas02, GI08, MR04, MR05]. Neural [HAG02]. Neuron [GAV00, GR00]. Neutral [EA07, HK01, LB01]. Never [KIK08]. Newtonian [OC05]. Next [TC03]. Nieuwenhuizen [Gar08]. Nilpotent [Piv03]. Ninth [CM06]. Nishimori [Pod02b, CMN06]. NJ [Opp04d]. NLS [AGT07]. No [FG08a, MRV02, Kad03b]. Noise [AMP06, Bao01, Bao04, BR05a, BG04a, BCC09, BD01, CL00, DD09, FW03, Fan05, FJ05, Fre01, FNO06, FG00, GMR06, HAG02, JHA07, KK08, LNM06, MM05a, QAK02, Rom04, Rom08, SM06, SPVV00, Val09a, ZF09, vK04]. Noise-Induced [BG04a, BCC09]. Noises [MDW09]. Noisy [Ber08, Gas02, Mas06]. Nominations [Ano00i]. Non [AL03, AH03, AP06, ABS01, AP03, Bao04, BP02, BDG^{+}02, BDG^{+}04, BDG^{+}06, BC02d, BDLvW08, BL06, Bou07, Bud08, CCMT09, CG09, CSN02, CCM00, Coh09, Coh10, CS04, CS09d, DR03a, DLM05, DJZ00, Eps06, FGC09, FNP07, Fin03, FR07, FM01c, FP04, FdL06, GS03a, GS01, Has01, vDHR06, HV01, HO09, Jaf03, JP02, Jan06b, Kas01, KZ03b, KTZ06, Kül03, LPD08, Lat02, LCS08, Luc08b, MRS07, Mas03b, MZ02a, MR01, Naj08, NKC00, Opp04b, OC05, OEB01, PS07b, PY09a, RB08, SWK09, SJ01b, Sch08, Sig09, SSH06b, SW08, Tah09, TTK01, Too04, Too08, Val00, WNUK01, WW05, Yar05, vEK07, ATH06, Rub09b]. Non-Abelian [ABS01]. Non-Adiabatic [NKC00]. Non-Analyticity [FP04]. Non-Autonomous [GS01]. Non-Colliding [Bou07]. Non-Compact [Kas01]. Non-Critical [vDHR06]. Non-Cutoff [SW08]. Non-Equilibrium [AH03, AP03, BDG^{+}02, BDG^{+}04, CCMT09, GS03a, Jaf03, JP02, Lat02, Luc08b, SJ01b, BDLvW08, CG09, FGC09, Opp04b, SWK09, WW05]. Non-Ergodic [RB08]. Non-Ergodicity [Too04]. Non-Existence [CS04, CS09d]. Non-Gaussian [DR03a]. Non-Hermitian [CCM00]. Non-Homogeneous [FNP07, Sch08]. Non-Hyperbolic [HV01]. Non-Ideal [BP02, PY09a]. Non-Increasing [Jan06b]. Non-Interacting [Yar05]. Non-Intersecting [FR07, OEB01]. Non-lacunary [HO09]. Non-Lifshitz [Naj08]. Non-linearity [Eps06]. Non-Local [LCS08, Too08, Tah09]. Non-Markovian [BL06, MR01]. non-monotonic [ATH06]. Non-Perturbative [Rub09b]. Non-Random [KZ03b, KTZ06]. Non-Relativistic [Sig09]. Non-Resonant [Val00]. Non-Rigidity [DLM05]. Non-Smooth [SSH06b]. Non-Summable [FdL06]. Non-Symmetric [WNUK01]. Non-Symmetrically [PS07b]. Non-Triviality [MZ02a]. Non-Uniform [CSN02]. Non-Uniformly [AP06]. Non-Universality [Mas03b]. Nonchaotic [LRB00]. Noncolliding [KT07]. Noncompactly [Vel08]. Nonconcave [TB06]. Noncondensable [TAL06]. Nonconservative [PT06]. Nonequilibrium [Ano00g, BMN09, BMW09, BD06a, BDG^{+}09, BDL00, DP07b, EMMZ06, ER02, Far05, HE03, HMG08, KNST09, LY07, Liu09, MKB00, OvWLH00, Pod01a, RS06, Rue00a, Skr03, TC07, TC08a, TC08b, Wag00, dZS04, Aba00, Opp07].

Nonequivalence [EHT00, GS08c, GS08d]. Nonergodicity [MB06]. Nonexistence [aYL05]. Nonextensivity [Tor04]. Noninteracting [LS00a]. Nonlinear [AN05, AN06, BCEP04, BCEP06, BD04b, CMS04, DRL02, Dor02, EEDJ00, ER00, FH04, FKS08, Gar03, GK02, KL06, KL08, Kol07, LNM06, Li09, LL00, Luc09a, MRW02, MN02, New08, Sac05, TTM^{+}01, VB07, WZ09, WW01, ZW04]. Nonlocal [BdMR03, Dir04, dMG04]. Nonprimitive [LdO03]. Nonrenormalizability.Part [Kla04]. Nonrigidity [BS08]. Nonsimple [CN04a]. Nontrivial [DVE07, PS04b]. Nonuniformly [Var08a]. Nonuniversality [CHNO06]. Nonzero [HY04]. Normal [Bar06a, DOS01, LPD08, LB04, LS09b, MN04, Süt04, Val00]. Nos [KK02b]. Nosé [MMR08a, RKH00]. Nosé-Hoover [MMR08a, RKH00]. Note [Cir02, DC01, ECSB00, FT06, Iof02, JP01, Kie09b, Pre03, Sos02, vBD02]. Notes [LS09b, Luo07, Pod01a]. Notion [vE00]. Novel [OP04, VBM06]. Nucleation [DD06, HMG08, MVE04, NOS05, Shn03, vB03]. Nucleotide [LLV09]. Number [Che09a, CE02a, CPV08, FBG02, FLS07, FK05, HPK^{+}07, Sam07, Sos00]. Numbers [Ano01d, Too08, Weh97]. Numerical [Bao04, BW07a, BFT02, CG00b, CGG06, Che09b, Ghe09, GGZD02, HK02a, JL00a, Ken09, MPRT^{+}00, QAK02, ST02, WDMB05, RK05]. nut [ATH06]. Nuts [Sai00].

Oasis [FN00]. Obeying [MB06]. Obituary [Ano01r]. Objects [ICO02]. Observable [LZ09]. Observables [BC02e, DGZ04]. Observational [Mas03a]. Observations [Dom03b]. Obstacles [Fuk08, HS98]. Occupation [GL01b, Set06]. Occupation-Time [Set06]. Occurrence [CR05]. Octagonal [DMB01, PVV02]. Octahedron [WMDB02]. Odd [BKOS07, DKRS02, FM05a, FM09b, Lyb05, Sin09, WK04]. Odd-Order [BKOS07]. ODE [Bir07]. Off [AMP06, BBK08, KMP^{+}07]. Off-Critical [BBK08]. Off-Lattice [KMP^{+}07]. Ohmic [LPD08]. Old [Aha03]. On-Line [BC02e]. One [AFN00, AGT07, Ada01, AGL01, AD02, AVBM06, BK08, BH08b, BK01a, BCK00b, BP01b, BBC^{+}01, BFT02, BCFP05, BDL00, BR04, BK03b, CM02a, CG00a, CL07, CEMM09, Cor04, CS04, CS09d, DK02, DTP02, DF02, DJZ00, DDR04, DG09b, DG09a, DM02, DFF02, Dun03, FJT03, FT08, FJM01a, Fou00, Gal01, GN08, GKR02, GS04, GS03b, Gir01, GT04a, GSW07, Gry01, GM00, Hao05, HvdHMO09, HNOV07, HI04, KMST00, KKT00, Kon02a, KW00b, LVZ03, LRB05, LWL00, Lou08b, MvWH^{+}01, PS08b, PD05, PS04a, RM00, RS04b, Sam01b, Sam04, SWK04, SI04, SKM04, Süt04, Tao01, Zhi00, vdHK01, vdHKK02]. One- [HNOV07]. One-Component [CG00a, FJT03, FT08, FJM01a, KMST00, LWL00, Sam04, SWK04]. One-Dimensional [AFN00, Ada01, AGL01, AD02, BK01a, BBC^{+}01, BFT02, BCFP05, BR04, BK03b, CS04, CS09d, DJZ00, DDR04, DFF02, Dun03, FJM01a, Fou00, GN08, GS04, GS03b, Gir01, GT04a, GSW07, GM00, Hao05, HvdHMO09, HI04, KW00b, Lou08b, MvWH^{+}01, PS08b, PD05, PS04a, RS04b, Sam01b, SI04, Tao01, Zhi00, vdHK01, CEMM09]. One-Particle [LVZ03]. One-Phase [LRB05]. One-Step [AVBM06]. One-Time [BP01b]. Onsager [Ano04m, Gal09, Len00, PSW09, SK06, Sin08, TC07, TV03, vK02]. Onuki [Pod03]. Open [Bah07, BD06c, CY09, CP04, CPV08, DJ04, DLS02, DLS03, DDR04, DELO05, DLS07, FS06, JOP06, LMS05, Lou08b, MS09, MO09b, Pri03]. Open-System [CP04]. Opening [HS04]. Operator [AL07, Ast08, Bud08, DG00, GI08, Lat02, Too07]. Operators [BLZ01, BPS06a, BJ02, Cam06, DZ01, DGZ04, ES07, GK06, GS04, JL00b, LdO03, LTWW02, Naj04, Naj08]. Opinion [Sin02]. Oppenheim [Nol04]. Optical [OP04]. Optics [Fan05, FGR03, Fre08b]. Optimal [AER05, BR01, CS02, DMSR00, ES04, MTT08, WS02a]. Optimality [BCF^{+}01]. Optimization [FL03b, Rob00]. Options [Mas05a]. Orbit [BG04a, Mal07]. Orbital [HH02]. Orbits [TDM02]. Order [ALS03, Ano01b, BK08, Bax05, BKOS07, BBCK04, BCC06, BBS03, CM06, Dud07, FRZB09, FSG08, Giu09, GM07b, GTZ02, JK01, Kas02, KM07, LKL09, Len00, LS09b, MM06a, NH03, OvWLH00, PP01, PIRB03, Rei00, Rid04, Sch08, Skr03, SPVV00, VD05, WB04, Xin09, YY09, HNO05a, HNO05b, Zah02]. Ordered [BCK00a, DMPV09, RD08]. Ordering [FSG08, NSS04, KIBG^{+}06]. Orders [MP01]. Organization [Gác01]. Organized [BCK00b, CBKM04, KR06, SDC04]. Orientation [QLAS02, Sid05]. Oriented [BR06, CPV08, Sak02]. Origin [BK04, BM05b, CCM00, RY07, dZS04]. Originating [FZS^{+}00]. Ornstein [Ben05, EK05b, EK05c, Fan04, HK02b, KRT00, KL06, KL08, MM05a, TM08]. Orthogonal [AFNvM00, Blo01, Blo04, Bor02, CGI02, CDG07, DGKV07, FEDZ07, Ryc08, Shc09]. Ortiz [GC08b]. Oscillations [Fre01, GPT02, KMP^{+}07, MNNS07, ZK00]. Oscillator [Amb06, GG00, GvH07, LDNG08, MM05a, Mas06]. Oscillators [BD04b, BHRW04, FW06, FNO06, JW00, KP07, LS04a, MSVE07, Skr00, Skr03, KP08b]. Oscillatory [CDG06, CL02]. Osculating [Ess03]. Ostwald [CHNO06, NV08]. Other [BBC^{+}01, De 07, Man03, Rit06, Roe03, Sos00, ZW04]. Over-Populated [EB02]. Overcoming [Kla04]. Overcomplete [TC03]. Overdamped [BMW09]. Overlap [De 04, Pan08, Tin03]. Overlapping [HGST00]. Overlaps [Bar06b, Led00, MS01]. Oxford [Bog04, Opp02a, Pod02a, Pod02b, Pod04a, Rub04].

P [Dom04, Opp01, Pod04b, Rap01, Roe03]. P. [Fre08a, Opp09, Pod02a]. Pacemakers [KNP07]. Packed [CM04c]. Packet [SLB00]. Packets [KS09]. Packing [AR09, BY03, Elo08, FZS^{+}00, PS00, Rad08a]. pages [Sah02]. Pair [FL03a, KMST00, LM01, Pro07, Pro09a, Skr00, Yoo07]. Pairing [Pan02]. Papanicolaou [Fre08a]. Paper [Gra01]. Papers [Dom04, Opp04d, Rod02]. Papon [Pod04b]. Parabolic [BK01a, BMP07, GS08a]. Paraboloid [FT08]. Paradox [FK09b]. Parallel [CN03, HHB09, Mac02, MM00, PM06].

Paramagnetic [TPV09]. Parameter [Bax05, FSZ01, Mol02, Mol04, PP01, PS07a, VD05, WB04]. Parameters [BP08b, Dud07, LNM06]. Paretian [EK08]. Parisi [BTM^{+}04, SS09c, SMT03, Leb08b, Tal07]. Parking [DFK08, FK09b, Rit06]. Parrondo [MB02]. Part [MMR06, FF04, MM06b, SSHE06, SSH06a, SSH06b, ZKBG03a]. Partial [Ach08, BDP02, DRL02, PSW04, Tah09, VEW08]. Partial-wave [Tah09]. Partially [ABCM03, BCK00a, MNP08]. Particle [AGL01, AGY02, AH03, AV03, Ano00h, BZ08a, BS07, BTT07, BT04b, BCL08, BDL00, Bur08, CM02a, CR07, CEMM09, EK06, GKRV09, GJ08c, GK02, GSW07, GM03b, HS09b, HPCH00, HRK^{+}05, IS07, KT06, KL06, KL08, KG00, LV01, LPD08, LVZ03, Lin08, LR02b, MRS07, MD01, Mür01, O’C01, PS03b, RM00, SJ02b, SEZ05, SMRK^{+}02, Sin08, TDM02, Too04, TC03, WNUK01, dSI04, vZC06, AV87, Iar01, Lan09]. Particles [And08, BCM02, CO08a, CMS05, CD09b, CM02b, DGR08, DA03, Fel02, FG03, GP08, GvB02, GH04, HR06, HNS08, Kol04, Lu00, Lu01, Lu04, Lu05, Lu06, MZ02b, Nag04, QLAS02, Ram07, SWK09, Sos00, Swe02, BJ08, Wes04]. partite [CAD09]. Partition [BBCK04, CSS02, CJSS04, CS08a, CS09b, HL02, HMH05, JS01a, JSS03, JS06, JK01, Law09, Mez01, SS01, SS09b]. Partition-Function [JS01a, JSS03, JS06, SS01, SS09b]. Partitions [BP07, RRS^{+}00, Str08, WMDB02]. Parts [GCU02]. Passage [BJ06, BG04a, Bra02, Bur08, Các08, FR07, KC03, Lin04, WLTH07]. Passing [Bao04]. Passive [Ben05, BCL^{+}07, EX00, Fan04, KO02, KP04, KMVE03, Wie00]. Passively [EK00]. Past [BW07a]. Pasta [BLP09]. Path [AM04, Mas05a, Mez01, Shn03, TC08a, Tao01, Roe03, Wit03]. Path-Integral [Tao01]. Pathogen [dARBY04]. Paths [Bou07, CPV08, EVE06, FM07, FR07, HT02, Hat07, Hol02c, LM01, Tur03]. Pattern [DHB07, Fei03, GR09, SC01, vdHK08]. Patterning [FL06]. Patterns [CR05, DDB^{+}00, Dro00, FHAY06, HM00, LK06a, Shi06b, Mas03a]. Paul [Mas02a, Nol04, Rod02, Taq04]. Pauli [BJ00]. PDE [Con04, RT08c]. PDEs [DD04, EL02]. PDFs [CLM08]. Pedestrian [Jus01]. Peel [dGNPR04]. Pelagic [Bir07]. Penetrable [FHL03]. Pentagrid [AYP07b]. Perceptron [Bal09]. Percolating [Ber03b]. Percolation [AV08, BR06, BS09b, BDGSC02, BP08b, BPS02, BPS03, BCS07, CN04a, CN04b, Cam05, CFN06, CJM09, CZ03, CLM07, CPV08, CRV01, CN01, DM06a, FS08, FR07, FBC^{+}05, GG02, HS00, HvdHS08, HJDRD00, vDHR06, Hol02b, Hol02c, HN03, Jan01b, JS08c, LMZ00, LW04b, Lüb06, MNS08, Mai03, Mis06, PDV07, PS02a, PY09a, PS04b, PM04, RT08c, Rát09, Sak02, Sak04b, Sak05, SJ01c, TB07, Too02, WP03, dHP09, vdHJ04]. Perfect [LB04]. Performance [Paj07]. Period [BD04b, KKS05, KMS08]. Period-Doubling [KMS08]. Periodic [Bao01, BG01a, BG04a, BDL00, BFPS07, BC09, CdlL05, CG00b, CdO02, CG00c, CCL04, DB00, GJY04, Gen04, HL02, HP04, HP08, HK01, JL00b, KD00, Koh01, PN03, RKN00, TDM02, GJQ06]. Periodically [DLS08, Hua09, PSG06, SC09]. Permeable [ALS09]. Permute [Cau02a, Cau02b]. Perpetually [LBLB04]. persistence [Aus00]. Persistent [DM08]. Perspective [AS06, Ton08a]. Perthame [Wes09]. Perturbation [BCDM07, Gen06, GvH07, MS05a, TV02, Val00]. Perturbations [AK09, CB04, Gar01, Naj04, PS08b, WX06, Yar05, Fre03]. Perturbative [BCO07, LS04a, Rub09b]. Perturbed [DMP05, Hua09]. Petrovsky [CD05]. Pfaffian [AK07, BR05b, Nag07, Sos03]. Phase [Aba00, AB09, AH03, Ano00g, AV08, ACL^{+}09, BGOY04, BM05a, BBD04, BBCK04, BCC06, BG09, BM02a, BGLT08, BDLvW08, BCKM00, BB05, BMPZ04, CPC03, CP01, CCIL08, CS04, CS09d, DM06a, Dir04, EN03, FSG08, FG08a, FKÖ03, FHL03, FS08, GTA09, GZ01, GM07b, Has00, HAG02, HMG08, IVZ06, Jaf03, JK01, JNH^{+}09, Jus01, Kad09, KCT08, Kas02, KCK07, KK00a, Koi07a, Koi07b, LW07, LRB05, LZGM04, Liu09, MKB00, MB03, MPS02, MH07, Mes02a, Mes02b, MM06a, MTT08, MGT07, MP09, NH03, Opp05, PS02a, PZ06b, PPR00, Pod01b, Pod04b, PM05, RVMO07, RR01, Rei00, RGRTS06, Rot06, TAL06, Tur06, WD01, Woj06, Yos08, dHP09, vE00, Rub09a, Pod03]. Phase-Field [Dir04]. Phases [BCK00a, BL09, DMPV09, DJ07, DRC04, LS09a, Rei00, SLB00]. Phenomena [CLM07, CKKC00, DRC04, Fre08b, GC08a, Han02, KC02, KMSS01, Ler00, Pod08, Git08]. Phenomenon [Bla05]. Pheromone [YCCN07]. Philadelphia [Roe03]. Philippe [Mas02a]. Phonon [AE08, Erd02, Spo06a, Spo06b, Spo06c]. phospholipid [KIBG^{+}06]. Photons [GC08a]. Phys [AV03, Ano01d, BPS03, Cau02b]. Physical [BGX01, CS04, CS09d, HJDRD00, MV03, Wol06, dZS04, vZC06]. Physicist [Hep09, Lan06]. Physicists [BA06]. Physics [Ano00a, Ano00z, Ano01a, Ano02a, Ano03a, Ano04a, Bax06, Bog09, BPS06b, DHB07, Han02, KK02b, Leb04b, Mas02a, Mas08, Opp01, Opp02b, Opp06b, Pod01a, Pod02b, Pod04b, Pod06a, RB08, Roe03, Rub06a, SBG03, Sza06, Tao01, VBM06, Wit03, Yng04, Zwe05, Opp03a, Pod04a, Rap01, Rit03, Ano00c, Ano00d, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04b, Ano04c]. Physics [Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g]. Pico [KC02]. Pico-Scale [KC02]. Picture [Mag09b, Rut01, SA02]. Pictures [CG09]. Piece [Pom05]. Piecewise [FGC09, TG02]. Pierre [Dor02]. Pinning [Pro09b]. Pirogov [BBCK04, BMPZ04]. Piscunov [CD05]. Piston [CL02, CLS02, GPL02, GM03a, GPL03, GPL04, MG02, NS04]. Pitaevskii [Rub04, JW04]. Pits [VKVT02]. Pivot [Ken02]. Placed [NOB00]. Planar [BM08, CS09a, CC00b, DOS01, DS07, Fre03, HR06, KW02, KM09a, MTS04, SV07]. Planck [ASB02, CEFM00, DR01, KL06, KL08, Zah02, Zha08].

Planck-BGK [Zha08]. Plane [CO08b, Fis06a, GP08, Ghe09, HC08, JS01b, Ken08, Nic01, SS08, SS09a, Sch06, Sch08, TAL06]. Plants [NS05]. Plasma [ABCM08, CG00a, FJT03, FT08, FT07, Jan00b, KMST00, LWL00, MT04, ST00, SJ02b, Sam04, SWK04, TM02]. Plasmas [EN03, LS00b]. Plate [BW07a]. Plateaus [Kom08]. Plus [Bur08, LDNG08]. PNG [PS02c]. Poincaré [CO08b, Leb04b]. Point [APS09, AS06, CCAD08, CEO07, FST06, Gen04, GY05, Hat07, JS01c, KLS07, LL09a, LB01, Luc08a, Luc09b, MCG08, Mur08, Nou02, NOV04, PDV07, PdKV09, Ryc08, Sak04b, Sak05, Shi06a, Sos00, TPV09, Wei07, DM09]. Point-to-Surface [Sak04b, Sak05]. Points [Aha03, Cam06, Car09, EMO08, GT02, Pro09b, SSLG^{+}00, Sim08, dlLV07]. Pointwise [Hor06]. Poiseuille [QLAS02, TS04, ZGA02]. Poisson [SS04, AS01, CM00, Cou06, EK08, Jab01, Luc08a, Luc09b, MRW02]. Poissonian [Fuk08, GHO^{+}00]. Poland [RG04]. Polarization [YCCN07]. Polaron [PA02b]. Polycondensation [KZ03b, KTZ06]. Polydisperse [BGW01, IVZ06]. Polydispersity [BM05a]. Polydomain [FSG08]. Polyelectrolyte [HRS03]. Polyelectrolytes [vSC00]. Polygon [Ric02]. Polygonal [HS98, Sch06, Sch08]. Polygons [MMR08b, SD06]. Polymer [BZ00a, BI03, CMV05, Ess03, Mez01, Pro07, Pro09a, Ton08b, Wie00, Wit03, vdHKK02, Pod04a]. Polymerization [JAWC08]. Polymers [BPS05, CP03, DR06b, EC07, NOB00, OSv04, Pod05, Pod06b, Pod06a, SD05, SD23, Ton07, WHI02, vdHK01, Git09]. Polynomial [HJ05, JS01a, JSS03, JS08a, SS01, SS05]. Polynomials [AFNvM00, BJR02, BR02, Bor02, CS03a, CS03b, EMO08, FY06, Mac09, Rad08b, Ryc08, SM08, SM09]. Polynuclear [BR00a, SI04]. Pools [MTT08]. Populated [EB02]. Population [CF03, FN00]. Pore [Urr08]. Porous [PO07, RS04b]. Positive [CE02a, DJ05, HJ05, Gra01]. Positivity [DN04, Fou00, HJ05, MR00, Ric04, SW00, Tia04, Tid01a]. Possible [HJDRD00]. Postcritically [BJR02]. Potential [Bak03, BMA02, BJ08, BC09, BJ02, Cam06, CMS05, CGK^{+}04, DB01, Fel02, FKS08, GM03b, HMPPMV00, JS08b, JL00b, KS09, Kap06, Lee09, LRM06, LM01, MB03, Sam05b, San05, SSLG^{+}00, SSE05, Suz01, Tah09, Tél06b]. Potentials [AG09, BCM02, CCL09, CCL04, FLM00, GJY04, HP08, Lu06, PN03, Sak09]. Potters [Mas02a]. Potts [Arg02, AYJP01a, Bax00, Bax03, Bax05, Bax08b, Bax09, BBD04, BCK00a, BKM02, CJS01, CZ03, CSS02, CJSS04, CS08a, CS09b, DMPV08, DMPV09, DMP^{+}04, DLR04a, DLR04b, GTA09, GM07b, HDMF04, JS01a, JSS03, JSS05, JS06, KW02, MN00, ND08, SS01, SS09b, Sea02, ZJ00]. Power [BFKT09, BG06, Far02, Jan06a, Lin08, Man03, ND03, PT06, Sin05, VPB^{+}06]. Power-Law [ND03]. Power-Like [PT06]. PowerFLOW [LLMS02]. pp [Kad03b, Pod03, Rap01, bA03]. pp. [Aba00]. Practical [Ano00b]. Preceding [Zab06]. Precise [WA09]. Predictability [SSB04]. Prediction [EMC09, SC01, ZKBG03b]. Predictions [Ken04]. Preface [Ano00v, Ano00w, Ano01s, Ano01t, Ano06, Ano09a, Ano09b, CEEH03, CGL07, KLL05, Leb02a, Leb02b, Leb02c, Leb03, Leb04c, Leb04d, Leb05a, Leb06a, Leb07a, Leb09a]. Preliminary [PVV02]. Premelting [PSW09]. Presence [Bao01, Cau02a, Cau02b, CLK09, Rác00, SD08, TAL06]. Present [EMC09]. Preservation [Sch01]. Preserving [Bal00, FW03, GS00]. Press [Aba00, Bog04, Bra02, Dag03b, Dom04, Dor02, Mas00, Mas02a, Mas03a, Nol04, Opp02a, Opp03a, Opp04a, Pod01b, Pod02a, Pod02b, Pod03, Pod04a, Por04, Rap01, Rit01, Rit03, Rub04, Spo03, Taq04, Wei01, Wei02, bA03, Kan01]. Pressure [Bak03, CLD05, GK09a, GN03, Jan00a, Jan01a, vBD02]. Pressures [FJT03]. Prewetting [PIRB03]. Primitive [BJ00]. Princeton [Opp04a, Por04, Spo03, Taq04, bA03]. Principle [Adi04, BJ00, BDG^{+}04, DR03b, FLR03, GI08, GII08, Kaw06, Tin03, Yoo07, BK04]. Principles [BM08, BD04a, BET00, BMN07, EHT00, Mat08, PG03]. Priori [TV00]. Probabilistic [BLR08, BKL00, CNS08, CEO07, Dom05, EK02, EK04, Kon02a, Kon02b, Too08]. Probabilities [CF03, DP07c, KKT00, KZ03a, vB03]. Probability [ALM07, AS06, Arg07, BR06, BCT06, BPS02, BPS03, BM00, BCL^{+}07, CLM08, DN04, Gal08, KM00, LMP05, Sak04b, Sak05]. Probable [ER00]. Problem [AG09, AMV08, AN00, BLP09, BGH01, BZ00a, CT04b, DLS09, Dir04, DKM04, FKS08, FM02a, Fre06, Gho05, GPL04, HB00, HH02, ISS07, KM08b, LP01, MV07, MMR06, RF02, Sai00, SDBS07, TJ08, Too00, Wie00]. Problems [BGS06, FL03b, KPT05, Kie04a, Kie04b, MRTZ03, MZ08, Rue02, Sem08, WHI02]. Process [ALS09, BK08, Ban05, BBT03, BP07, BDL08, BFP08, BR05b, Bou07, DF02, DLS02, DLS03, DEL04, DDR04, DELO05, DG09b, DP07c, ED04, FM01c, GM05, GJ08c, GSS03, GS08c, GS08d, IS07, JNH^{+}09, Jun03, KT09, Ken08, KK00a, KL06, KL08, KZ03b, LMS05, Lou08b, Luc08a, MMR06, MM06b, PPH03, PS02c, PM05, QQT02, Raj09, RS05, Rit06, Sak01, Sak02, SA02, Sch00, SH07, Sep01, Shi06a, Sui00, SV07, SM01, TM08, Too04, Wid04, ZS02, ZKD05, RT08a, CD09a]. Processes [All08, AVBM06, Bah07, BMW09, BJ06, BF05, Ber03a, BZ09, BT04b, BDGSC02, BD06c, BGM04, CH06, CL07, CY09, CG00c, DH08, DS07, EK08, FGC09, FF07, FFN09, FNP07, Gas04, Gas07, GJ08a, GY05, GL01b, GS08b, GE09, GM03b, JHA07, KT07, Kol07, KLS07, KC03, LOV04, Luc09b, Mas02a, Mas08, Mas09a, MK03, Mur08, OC05, PM06, PSG06, Roe03, Ryc08, SF07a, Set06, Too00, Uch04, Wei07, bA09, Bra02, Taq04]. Processing [Pod02b]. Produces [Ram07]. Product [Gen06]. Production [BLLO09, BMN07, DP07a, DSZ07, MR00, QQT02, SF07a, SSS07, SC09, TG00, VMT02, ZRA04]. Products [CdO07, Gam03]. Prof [Wit03]. Profile [DLS02, DG09a, Far05, FFN09, KM07, MM05c]. Profiles [AC01a, AC01b, AC04, BMD00, Sep01, Tél06b]. Program [Ano00x, Ano00y, Ano00z, Ano01u, Ano01v, Ano02n, Ano03m, Ano03n, Ano03o, Ano04m, Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Leb07b, Leb07c, Leb08a, Leb08b, Leb09b, Leb09c]. Projected [Rei08]. Projection [CU03, Lat02]. Projection-Operator [Lat02]. Projector [MAOB01]. Proof [And05, BCO04, BCK04b, BCT03, CDG07, Häg07, Jia08, Mas05b, Pan02, Süt03]. Propagating [Ewo01]. Propagation [BGL03, BD01, BK01b, BCDM07, CK06, CE02b, DR05, Fre08a, NOS06, vZvB02]. Propagation-Dispersion [BGL03]. Properties [AR00, ABC05a, ABC05b, ALS06a, BL04, BF05, BD00, BDGSC02, BP01b, Bla03, BCG00, BCG01, BDL00, BFPS07, BCS07, CPK07, Che06, CZ09, De 07, EMH04, EE07, FM05b, FZ03, FJM01a, GL01a, GMW02, GSS03, HC08, Hol02a, Jac09, KS02, LPD08, LLMM03, LR00, Nad00, PS08a, PY09a, PS06, PT04, Ric04, Riv04, ST00, Sam01b, SM08, Spo06b, Spo06c, SB06, SLA02, TG02, Ton07, Val09b, Val09a, ZF09]. Property [AMP04, BCT03]. Proposed [Swe04]. Protein [DSZ07, KMP^{+}07, TE00, WE07]. Protrusion [JAWC08]. Pruned [JHW09]. Pruned-Enriched [JHW09]. Pseudo [BM04b]. Pseudosphere [FJT03, JT04]. Publications [Ano01q, Pod02b]. Publishing [Opp04c, Opp04d, Roe03, Wit03]. Pulled [Kad06]. Pulsating [LBLB04]. Pulsation [EK05a]. Pulse [Ewo01]. Pumped [HJ09]. Pumps [AEGS04]. Purchased [Kad03b]. Pure [BM02a, EHT00, Gen04, MBL08, Rei08]. Pyrochlore [MTS04]. Pythagorean [CO08b].

Q [BLZ01]. Q-Operators [BLZ01]. QED [BP09, Sig09]. QFT [CAD09]. Quadratic [LRM06, SV09, Skr00]. Quadrupolar [KIBG^{+}06]. Quadrupole [KCK07]. Quantitative [CM04a, Pey09]. Quantities [ZW04]. Quantization [PTZ03]. Quantized [Kie04b, Kom00a]. Quantum [AL03, AGY02, AM04, Amb06, AVZ00, AC04, ABS01, AEGS04, BMA02, BR05a, BG00a, BG00b, BGOY04, BPS06a, BSB00, BCEP04, BCEP06, BCC09, Bjö09, BG09, Bud08, CCAD08, CP01, CEFM00, Cas01, Cas02, CM05, CCIL08, DD09, DR03b, DMR05, DJ04, DDM08, DRC04, DLS08, DGZ04, ESY04, ES09, EG05, FBG02, FRZB09, FGR03, FPSW01, GLM02, Gal08, GMW02, GII08, GS07b, GvH07, GL04, Gri00, GOS08, Han02, HMH05, HWB01, HB08, HB00, Hua09, Iof02, Jac09, JP02, JOP06, Jan06b, JK04, KN03, Kas01, Kie09a, KBS02, KP08a, Kom00b, KW00b, KP07, KP08b, LRB05, Len00, Li06, LG01, LL02, LS09b, MP01, MD01, MAOB01, MS09, Mes02a, Mes02b, MTS04, MR01, NOS06, Naj07, NR04, NKC00, O’C06, PdKV09, Pri03, Rác00]. Quantum [Raj09, Roe03, Rue00a, Rut08, Sac04, SW00, Sam09, SJW07, Skr00, SB01b, Suz03, Var08b, WNUK01, Wit03, Yar05, Yep02, ZF09, dPG04, vK04, Mas05a, Pod01b, Spo03]. Quartic [LDNG08, Rad08b, TW03]. Quasi [AV08, AS02, BMP07, CR07, CNV01, Gen04, GS03b, JL00b, LM06, MZ02b, SY03, SD08]. Quasi-Classical [CNV01]. Quasi-Particles [MZ02b]. Quasi-Periodic [Gen04, JL00b]. Quasi-Self-Similar [AS02]. Quasi-Static [LM06]. Quasi-Stationary [BMP07, CR07, SD08]. Quasi-Transitive [AV08]. Quasicrystals [CdO07]. Quasifreeness [CS08b]. Quasilocality [FLR03]. Quasiparticle [DK02]. Quasiparticles [RR01]. Quasiperiodic [AYJP01b, AYP07a, BJ02]. Quasispecies [PD06]. Quasistationarity [KL09]. Qubit [LZ09]. Quenched [BCK04a, Lat02, Mat08, SD06, Tin03].

Queuing [RSV09]. Quickly [Mac02].

R [Aba00, Bog09, GCU02, Mas09b, Pod01a, Pod04a, Rit03]. R. [Git07, Rug08, Wei09]. Race [vB03]. Radial [Ace00, Ghe09]. Radiation [Nou02, Sam09]. Radiative [Tid01a]. Radons [Wei09]. Raise [dGNPR04]. RaMOSt [De 06]. Random [AFNvM00, AK07, AS06, And08, AL01, ALS03, Ano01d, AR06, AR09, Ast08, BH00, BL04, Bak06, BH08a, BS05, BY03, BG01b, BB03, BM04b, BKLO00, BKOS07, Bel04, BBT03, BP07, BM01a, BZ04, BZ09, BS09a, BR02, BG04b, BBP02, BCFP05, BP08b, BC00, BK09, BM00, BCL^{+}07, BEK^{+}07, BK01b, Bur08, CPK07, CS09a, CMV05, CK03a, CL06, CL07, CS08b, CCM00, Con04, CDA09, Cor05, CDG07, CB04, CK03b, Cro09, DG00, DDH08, De 04, DFK08, DMB01, DWMB05, Dot06, ES07, Els00, EMH04, FST06, FY06, FF04, FLW08, FLS07, FLM00, FKS08, FM09b, Fra07a, FW07, Gam03, GRU07, Gar01, GS08a, Gas04, Gas07, GK06, GS07a, GG02, GF08, GJ08b, GG06, GMNT05, HL08, HT01, HC08, HNOV07, HMS05, Hor06, JR07]. Random [JS01e, KS09, Ken08, KP08a, Kol08, KO05, Koz00, KPvB06, KZ03b, KTZ06, KP05, Kuk06, Kül01, Kur03, LSSW03, LW04a, LB04, LY07, LMV07, LK09, Lou08b, Luc09b, LP08, LP09, Mac09, MCZ06, MBL08, MO09a, Mas08, Mat08, McD01, MTT08, Mol04, MZ08, MO09b, Nag07, Naj04, Naj07, Naj08, NNR09, OBB03, PDvB00, Pat07b, PS07b, Pop01, PS06, Pro09b, Rad08a, RS08, RS05, RT08c, Raw03, SS08, Sab08, SS09a, SSD00, SM08, SM09, Sch00, Sem08, SM06, SS09e, Shc09, Shi09, SD08, Sod09, Sod17, Sos00, SZ03, SS00b, Sud09, Tao01, Tel06a, Ton07, Tur06, VD01, WZ09, WW01, WX06, Wei05, WDMB05, Win02, Yar08, YBMS06, Zab06, Zab08, ZS02, ZJ00, vENS05, vEK07, vdEvdHH08, vdHK01, vdHKK02, vdHK08, JR06, Weh97]. Random [Bog09, Pod07b, Ton08b, Sah02]. Random-Cluster [BC00, GG02, GG06]. Random-Field [SSD00]. Randomly [CO08b, CD09a, CLM05, Ewo01, Fre08a, MVW08, NOB00, Rei08, VPB^{+}06]. Randomness [Ano01b]. Range [BK08, BBDR05, BJ08, BELM00, BCC06, BIV01, BDLvW08, BP08a, BB05, BCS07, BCS09, BD03b, BGM04, CMS05, CG04, Cro09, CS04, CS09d, DPZ06, Dot04, Fuk00, Giu09, GM07b, GT06, GSS03, GS08c, GS08d, Has00, HvdHS08, Kon02a, Kon02b, LMS05, MP01, MPRT^{+}00, Naj04, Nol04, RS08, RGRTS06, SH07, Shl04, Skr03, Tor04, Uch04, WNUK01, Mis06, PM05]. Ranged [NS02, dZS04]. Rank [Tah09]. Rank-two [Tah09]. Rapid [Kra03]. Rapidly [LS09a]. Rare [KL09, LB09, VEW08]. Rarefied [BGH01, CD09b, GLM02]. Rashba [Mal07]. Ratchet [BCV00]. Ratchets [LJ03]. Rate [AE08, BK08, BCC09, Bud08, CGR07, JH03, Li07, NNR09, Ste04, SWF07, VMT02, Yos08]. Rates [AP08, BT04b, BCT06, DSZ07, KL09, KS07, Gra01, Mas05a]. Ratio [CD09b, Fis08, HD03, Sea02, vdHK08]. Ratios [SS00a]. Rayleigh [BTT07, BFT02]. Rayleigh-Gas [BTT07]. Reacting [Pol00]. Reaction [AFN00, Ano01c, BD06b, Buc07, CH06, CG00c, DDB^{+}00, Fre06, HMPPMV00, KNP07, KMP^{+}07, LPE04, Lou03, Lou08a, SWF07, Wei01, Lou07].

Reaction-Convection [Fre06]. Reaction-Diffusion [DDB^{+}00]. Reactions [BD05, GS07c, GP04, LK09, OBB03, ZK00]. Reactive [BD06b, HM00, SSK02]. Reactivity [CW06]. Reader [Gra01]. Real [AK07, DOS01, KS07, Mac09, SSLG^{+}00, SM08, SM09, VHO09]. Real-Space [KS07]. Realism [BG00b, BG00c, Gri00]. Realizability [KLS07]. Realization [HJDRD00]. Receiving [Leb04b]. Receptor [HS04]. Reciprocal [SK06]. Reciprocity [Gal09, PSW09]. Recurrence [STV02, SV07, Var08a]. Recurrent [Li09, SV09]. Recursion [Ito02]. Recursive [Pre03]. Recycling [GS03a]. Redner [Bra02]. Reduced [Kaw06]. Reducibility [Gen06]. Reduction [Bax08a, BGV01, BI03, FPS00, KP06]. Reductions [AC05]. Reference [Jan06b, Rei05, RK05]. Refined [War01]. Reflecting [DM02, MVW08]. Reflection [AKR01, Tia04]. Refocusing [Ewo01]. Regarding [Nag04]. Regime [Bal00, BCEP06, BCK04b, BMPZ04, DLR04a, DLR04b, JKO^{+}01, Mag09a, O’C06, SH07, SD08, TT05, dHP09]. Regimes [LR06, Tid01a, ZKBG03a]. Region [FPS07, GK06, GT04b, LRB05, PWC02]. Regions [Too02]. Regression [Bud08]. Regular [BS09b, BD02, CPK07, LP09, Par08, Rue02, Sim08, VB07]. Regular-Lattice [CPK07]. Regularization [JS01d, Mat02]. Regulation [FK09a]. Reiss [Ano02n]. Related [AH03, BM08, BF05, CE02b, vDHR06, PS08a, Rug08, WMDB02]. Relation [BCG08, Cor04, GZG05, HE03, JS01b, KO05, Li09, MMR08a, NT04, Rei00, SB01a, Tel06a]. Relations [Bax04b, Bax06, Ber03b, KW02, KP05, Kuk06, Luc08b, Luc09a, LZ04, Nep03, SJW07, SK06, Suz01, TV03]. Relationship [SS00b]. Relativistic [BDL06, CK00, Jia07, Jia08, Sig09, Tum06, Wu09]. Relax [VHO09]. Relaxation [AMT07, BP09, Fin03, GPL03, HS08, HH02, JS01d, KT09, LBLB04, MO09b, NAS02, SvL02, Zhi00]. Relevance [Sai00]. Relevant [DRL02]. Reliability [CS03b]. Reliable [Gác01]. Remark [BL08c, BCS09, Rue00b, vE00]. Remarks [Bax09, CGV02, Fre04, Kuk06]. Renewal [GL01b, PSG06]. Renormalization [BR01, CM05, Cop08, DLM05, DRC04, FM08, GAD06, HT02, Hat07, Ito02, KMG07, KMS08, LL02, MS08a, Oon03, RS08, Rob00, Rub09a, Sam06, Sha01, Shi06b, WS02a, Opp01, Rub09b]. Renormalized [FLR03, SJ01c, SFCO02, Tél06b]. Renyi [CB04, Hor06, EMH04, PDV07]. Renzo [Mas09b]. Repair [CGS07]. Repeated [Var08b]. Repeatedly [WM07]. Repellers [Dys05, HV01, HO09]. Repelling [FST06]. Repetition [CR05]. Replica [De 06, FL03b, FW07, GR00, MPRT^{+}00, NS02, Pan08, RS04b, Rit01]. Reply [BG00b, Gri00, LPdO01, Tak10]. Representation [AL07, Bol09, BC00, EFM09, GG06, KNST09, MN00, TW08]. Representations [BFKT09, LL08b, LL09b, Sch08]. Representing [AGY02]. Reproducing [Yoo07]. Reproductive [Kad06]. Repulsion [Nak06, Sak04a, Sak06, Sak09]. Repulsive [JM00, SS05]. Rescaled [BKLO00]. Research [Kad03b]. Researchers [Wei05]. Reservoirs [BLL04, BLLO09, Jac09]. Residence [Bar06a]. Residual [BL08b, BL08c]. Resistance [FHAY06]. Resistivity [KN03]. Resistor [SJ01c]. Resolution [KVM^{+}00]. Resonance [DB01, DM02, FG00, MDW09, MS09, VST09, ZF09]. Resonances [DP04, RSV09, Sig09, SL02]. Resonant [GFH08, Val00]. Resources [GS03a]. Respect [DT00]. Respond [SS09e]. Response [BMW09, Ber03b, GZG05, GK02, JOP06, LM06, Luc08b, Luc09a, SSD00, Swe04]. Restitution [Tas06]. Restricted [BJ00, DMPV08, KK08, McD01]. Restructuring [ZK00]. Result [And08, FM07, KMST00, KMS08, Pol00]. Results [Aha03, ABCM08, AD02, BH00, BBBF06, Bax00, BM02a, CZ03, CLS02, EHT00, GR06, GP04, GSW07, HD03, JS01a, PVV02, PAY09, Sam05a, SS09e, Urr08]. Resummation [Gen06]. Retrieval [GL01a]. Retrodiction [EMC09]. Returns [CCM00]. Reversal [Fre08a, MN03]. reverse [ATH06]. Reversed [Ewo01, Gas04, Gas07]. Reversibility [QQT02]. Reversible [Bro01, CEFM00, CNS08, Jun03, PdKV09]. Reversible-Work [PdKV09]. Review [Aba00, Ano00c, Ano00d, Ano00b, Ano00e, Ano00f, Ano00g, Ano00h, Ano01b, Ano01c, Bog04, Bra02, Buc05, Cer05a, Dag03b, Dag03a, Dom04, Dor02, GCU02, GC05, Han02, Hol02a, Kad03b, Kan01, Mas00, Mas03a, Mas05a, Mas06, Nol04, Opp01, Opp02a, Opp02b, Opp03a, Opp03b, Opp04c, Opp04a, Opp04d, Opp04b, Opp05, Opp06b, Opp07, Pod01b, Pod01a, Pod02a, Pod02b, Pod03, Pod04b, Pod04a, Pod05, Pod06b, Por04, Rap01, Rit01, Rit03, Rod02, Roe03, Rub04, Rub06a, Sah02, Shl04, Shl05, Spo03, Sza06, Taq04, Wei01, Wei02, Wei05, Wes04, WW05, Wit03, bA03]. Reviews [Mas02a]. Revised [PS02b]. Revision [Hod01]. Revisited [Bax04a, LZ08, SA02, Sha01, SM01]. Revivals [SLB00]. Revolution [McC01]. RG [Gal01, LL09a]. Rheology [OC05]. Rhombus [WMDB02]. Rich [LSSW03]. Richtmyer [GGZD02]. Riemann [Bax03, MvWH^{+}01, SSH06b]. Rigid [BK02, SST^{+}00]. Rigidity [BBLS05, DMN00, DLM05, GG02, MV07]. Rigorous [AGT07, AD02, BH00, BBBF06, BM00, CE02b, DMSR03, DR01, GSW07, Mas05b, MS06b, Tid01b]. Rimini [Tum06]. Ring [AR02, DP07c, JW00, MD01]. Ripening [CHNO06, NV08]. Rise [RSMJ^{+}02]. Risk [Mas02a]. Risks [Mas02a]. River [Opp04d]. Rivet [Dor02]. RMG [BR01]. Road [vK02]. Robert [Wei02]. Robinson [RS09]. Robust [Var08a, vE00]. Rocking [BCV00]. Rodlike [Hem00]. Rods [IVZ06, SLA02]. Roehner [Mas03a]. Role [DHB07, DGZ04, FG00, Gal01, GN03, KSSM07]. Roman [Kad03b]. Room [Leb07b, Leb07c]. Rooted [Sum05]. Roots [BA01, DFM01, FM01a, FM01b, FM09a, SM08, SM09]. Rosario [Mas00]. Rotating [BK03b, LS09a]. Rotational [FL03a]. Rotor [BT04a]. Rotors [HPF^{+}02]. Rough [DMSR03, MS05b, MM05c]. Roughening [KM08a]. Roughness [Pen08]. Rouse [BD00]. Route [OP04]. Routines [Cai05]. row [FK09b]. Rowlinson [YS00]. Rowlison [Iof02]. Rubinow [HvdHMO09]. Rubinshtein [Pod04a]. Ruelle [Jia03, Arg07, Kie09c, Mau08, Sin02, SL02, vBD02]. Ruffo [Shl04]. Rule [BMPZ04, CN03, DJ04, FU05, Jan00b, KMST00, LL09a]. Rules [LM06, Sam09]. Rutgers [Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Leb07b, Leb07c].

S [Bog04, Han02, Opp02a, Opp04d, Pod01b, Sha10, Shl04, NS02]. S. [Mas09b, Pod07c]. Sabra [CLT07]. Sachdev [Pod01b]. Saddle [BL06, Ler00]. Saddle-Focus [Ler00]. Saffman [CT04b]. Saha [ABCM08]. Salesman [SDBS07]. Salient [GAV00]. Salpeter [SO00a]. Salvatore [Sah02]. Same [Kad09]. Sample [Sos02]. Sampled [Rei08]. Sampling [HRA^{+}04, Har08, Mac02, RS06, ZW04]. Sand [DMP^{+}04, HJDRD00]. Sandpile [LD01, Pri00, SD09, VD05]. Sandro [Rub04]. Santa [Lan06]. Santosh [Buc07]. Sarate [GC08b]. Sarina [Dag03a]. SAT [LP01]. Satisfaction [MZ08, Sem08]. Saturation [Sam05b, ZF09]. Sausage [Fuk08]. Scalar [BCL^{+}07, EX00, Fan04, KMVE03, NOV04, Tid04, WX06]. Scale [BBS04, BK04, DDMSR04, DM08, GPL03, Has02, KC02, Mat02, PS02c, SMT03, CPK07]. Scale-Free [BK04]. Scale-Invariant [Has02]. Scaled [NNR09]. Scales [Lem00, MS05b, MS06b]. Scaling [Akt01, AGY02, Ano00h, BCF^{+}01, BLP09, Ber07, Cam05, CFN06, CJM09, CGGP04, CGJ^{+}05, CJS01, Car03, CEFM00, CM05, CLS02, DTP02, DVE07, Dun03, EB02, Fis06a, Fis06b, Fis08, GS08a, GJ08b, GJ08c, Gra09, Han02, HS00, KPH00, KP01b, KM01, KMG07, KKS05, KMS08, LRM06, LR06, LW04b, MN02, MDL^{+}03, Nad00, OEB01, Pod02a, PS04a, Ric02, Str06, TV02, WX06, YS05]. Scattered [GP08]. Scatterers [BK03b]. Scattering [LCS08, RKN00, Tah09, Wag00]. Scenarios [Yip06]. Scheme [AC01a, Bao00, HRA^{+}04, Har08]. Schemes [Rhe05, Rit06, Tal07]. Scheraga [RG04]. Schlögl’s [Liu09]. Schnakenberg [AG07]. Scholes [Mag09a]. Schramm [BBK05, Ghe09, Ken07, Ken09]. Schrödinger [AS01, Ast08, BLZ01, Cam06, CCL04, DG00, DJZ00, FKS08, GJY04, GK06, GS04, Lee05, LdO03, LTWW02, MRW02, Rid03, Rot06, Sac05, SS04, WZ09]. Schrödinger’s [LZ04]. Schulte [Hol02a]. Schur [BR05b]. Schwabl [Opp03b]. Schwarzschild [Li06]. Schweitzer [Wes04]. Schwinger [MPD09, PD06]. Science [Dag03a, Dor02, GCU02, GAD06, KLL05, Pod02b, Pod07a]. Scientific [Dom04, Han02, Hol02a, Opp01, Opp02b, Opp04c, Opp04d, Rod02, Wit03]. Screened [ABCM03, CG00b]. Screening [BK01a, NV08]. Search [SDBS07]. Second [BS07, FSG08, FK09b, GPL04, Hir05, JK01, Liu09, MM06a, SSLG^{+}00, YY09]. Second-Order [FSG08]. Second-row [FK09b]. Section [AG09]. Sedimentation [Fel02, HNS08]. Seeded [dBR08]. Seen [PO07]. Segel [HS09b]. Segregating [BELM00]. Segregation [Uel04, ATH06]. Seiringer [Häg07]. Seismic [ZKBG03a]. Selected [Rod02]. Selection [Fei03, Lem00, RS04a]. Selective [BG04b, BGLT08, CGG06]. Self [AH03, AS02, BGMM03, BDG^{+}09, BCK00b, BC02c, BC02d, BCT03, BC03, BLL04, BLLO09, CCF^{+}00, CGJ^{+}05, Cau02a, Cau02b, CBKM04, CGH^{+}03, CNS08, DW05, DSC02, EX00, Gác01, Ghe09, Gir01, GMNT05, HD03, Has01, HT02, Hat07, HvdHS08, Jac09, Ken02, Ken04, KRT00, Kon02b, KR06, MB03, MM06b, MS08b, PT06, Sak09, Sla09b, SDC04, SD06, Vel08, WB04, Taq04]. Self-Adaptive [DSC02]. Self-Affine [MB03]. Self-Averaging [WB04]. Self-Avoiding [CCF^{+}00, CGJ^{+}05, Cau02a, Cau02b, DW05, HT02, Hat07, HvdHS08, Ken02, Ken04, MS08b, SD06, Ghe09]. Self-Consistent [BLL04, BLLO09, Jac09, KRT00]. Self-Contained [BDG^{+}09]. Self-Driven [AH03]. Self-Dual [Has01, Sla09b]. Self-Duality [Kon02b]. Self-Gravitating [Gir01]. Self-Interaction [CNS08]. Self-Interactions [BGMM03]. Self-Organization [Gác01]. Self-Organized [BCK00b, CBKM04, KR06, SDC04]. Self-potentials [Sak09]. Self-Similar [BC02c, BC02d, BCT03, BC03, CGH^{+}03, EX00, GMNT05, MM06b, Vel08]. Self-Similarity [PT06]. Semi [ALS09, Bao00, BCO04, BW07a, DLR04a, DLR04b, Son09]. Semi-Infinite [BW07a, DLR04a, DLR04b, Son09]. Semi-Integral [Bao00]. Semi-Invariants [BCO04]. Semi-Permeable [ALS09]. Semiclassical [CMS04, RR01, Sac05]. Semiconductor [TJ08]. Semiconductors [ADG00, GC08a]. Semiflexible [CP03]. Semigroups [Kol07]. Sengers [GCU02, GC08b]. Sensitive [CD08]. Sensitivity [SDC04]. Separable [Tah09]. Separating [SV09]. Separation [CR03, FM08, FHL03, JNH^{+}09, MB03]. Sequence [HWB01]. Sequences [WW01]. Sequential [GK09a, HSA00, Lou08b, SS00b, Sud09]. Series [AER05, Aus00, BFKT09, Cas02, Dor02, Fuk03, Gen06, Han02, KIK08, MB06, MM02, PF07, Sin05]. Sessile [GLMSR07]. Set [Bor02, SS05]. Sets [BS05, BP08b, CdO07, DJ05, DM09, GOBY06, GT02, KM00]. Setting [AN05]. Seven [BW07b]. Several [HNOV07, TV03]. Sewell [Spo03]. Shade [VHO09]. Sham [PN03]. Shape [Bal01]. Shapes [dBR08, DDH06, DP03, Dun03, GK04, NH03]. Sharipov [Tak10]. Sharp [BELM06, BG09, BCKM00, CF03, CLT07, Dir04]. Sharpness [AV08]. Shear [ASB02, Ben05, BM00, Cer01, DR06b, Gar03, HRK^{+}05, Keh05, SMRK^{+}02, Yip06]. Shear-Thinning [Keh05]. Sheets [BGMM03]. Shell [CLT07, BBS04, FR00]. Sherrington [CGT05, CC09, Led00, Pan08, WB04]. Shifts [BI06]. Shlomo [Wei01]. Shlosman [Sch06]. Shock [CC00a, Gir03, RS04a]. Shocks [Bal01, Bal04, BTV09, PS03a]. Short [BJ08, CG04, CS04, CS09d, Fuk00, LR02a, MPRT^{+}00, NS02, RGRTS06, Tél06b, vB03, vZC06]. Short-Distance [Tél06b]. Short-Range [Fuk00, MPRT^{+}00]. Short-Ranged [NS02]. Short-Time [vZC06]. Should [ALM07, Sin02]. Show [BM04b]. Shrinking [DLS08]. Sided [HC08]. Sidney [Bra02]. Sierpinski [CCY07, CC08, HT02]. Sign [BBC^{+}01]. Signaling [FFN09]. Signals [BBS04, SB06]. Signature [MNS08, Rác00]. Silo [BFGM02]. Similar [AS02, BC02c, BC02d, BCT03, BC03, CGH^{+}03, EX00, GMNT05, MM06b, Taq04, Vel08]. Similarity [LMZ00, PT06]. Simiu [bA03]. Simple [BL04, BB03, Bla03, BDL08, Bol09, BZ00a, CPT05, DF02, DLS02, DG09b, FT06, Gar03, GJQ06, GAV00, GM03a, LP01, Lou05, Man03, MK03, MV03, MG02, Pol00, SK07, SLA02, TW03, Opp04c]. Simple-Quartic [TW03]. Simplest [BD05]. Simplicity [KM06, SC01]. Simplified [Tid01b]. Simplifying [DMS02]. Simulate [CSN02]. Simulating [Bao01, BBL00, Ken07, LB04]. Simulation [Ano00b, Bao00, BW07b, CK04, DA03, GS07c, HPCH00, HRK^{+}05, ICO02, JHW09, Keh05, LB09, MDB00, RSMJ^{+}02, SSK02, SD05, SD23, SW08, WHI02, YHD02, Buc05]. Simulations [BCFP05, CG00a, CG00b, LV01, NH03, NDC02, OC05, PVV02, PS03b, Rap01, RNVRP04, VKVT02, YS05]. Simultaneous [MM06a]. Sinai [And05, BBCK04, BMPZ04, Jia03, Rue02, SL02, vBD02]. Sinai-Ruelle-Bowen [Jia03]. Sine [CR01, Cai04, RC01, Sos00, MM05b]. Sine-Gordon [CR01, Cai04, RC01]. Singapore [Hol02a, Opp01, Opp02b, Opp04c, Wit03]. Single [Ale01a, BBS04, CH06, GR00, HHB09]. Single-Site [CH06]. Single-Speed [HHB09]. Single-Spin [Ale01a]. Single-Time [BBS04]. Singular [BZ08a, BJ08, BNZ07, GvH07, HD09, LdO03, Lou03, Lou07, Lou08a]. Singularities [BGGZ06, CZ09, FMB03, PF07]. Singularity [BS05, Car03, Dot06, Fis07, FG08b, SMT03]. Site [CH06, DH08, WP03]. Sitter [Li06]. Situations [KKS05]. Six [Bax02, Bax04a, BW07b, BL09, DFM01]. Six-Vertex [BL09, DFM01]. Sixth [KMST00, Ano00z]. Sixth-Moment [KMST00]. Size [Akt01, AV08, BKM02, CGGP04, CSM03, Fis06a, Fis06b, FM09b, FSZ01, Han02, Hei04, KPH00, KP01b, KM01, Mez01, MPD09, Sin09, Ton07, TT05, VST09]. Sizes [SSD00]. Skeletons [Koi07b]. Skew [AFNvM00, Gen06]. Skew-Product [Gen06]. Slab [DM06a]. Slater [SS04]. SLE [HL08, Law09]. SLEs [BBK08, Dub06]. Sliding [MTS04]. Slow [AP08, BdMR03, CL03, Sep01]. Slowing [Els00, Hao05, NH03]. Slowing-Down [Els00]. Slowly [TV00, VHO09]. Slyozov [LM02, Vel00, CGV02]. Small [AK09, BMN09, Bak06, CD09b, Che05, DN04, FG03, Fre01, HS98, HS04, Hos06, Lee09, LR06, LM01, MS05b, MK04, New00, SWF07, Ano01b]. Small-World [LR06]. Smart [Ram07]. Smectic [Xin09]. Smoluchowski [Fre04]. Smooth [AS03, HPCH00, Rue02, SSH06b]. Sneppen [Ban05, MZ02a]. Snurnikov [Mon04]. Soap [TM02]. Social [BK04, Shl07]. society [Pod07c]. Soft [AG09, CCL09, ETB06, GM03b, HPF^{+}02, Lu06, PS02a, TRK03, Yip06]. Soft-Mode [Yip06]. Sokolov [Wei09]. Solid [CM02b, DGR08, EC07, ICO02, Koz00, PdKV09, WMDB02]. Solids [DD06, LM06, PSW09]. Soliton [LS00b, MM05b]. Solitons [DTP02, Gar01]. Sølna [Fre08a]. Soluble [Has02]. Solution [AYJP01a, BKLO00, BW07a, BCL08, BL09, BD03b, Fou00, Hei04, HB00, HAG02, JL00a, Jia07, KM09b, MCZ06, Mez01, MPD09, PD06, PM06, Saa07, Sam03, SV09, Tid01a, Wu09]. Solutions [AL03, ABC05a, ABC05b, Ale01b, AG09, Ano01c, Bak06, BKOS07, BC02a, BC02c, BC02d, BCT03, BG06, BBS03, BK03b, CM00, CS02, CGR07, Cer05b, Cer06, CD04, CD09a, EB02, FS07a, FM02a, FV02, GJY04, Gho05, HS09b, HPWW03, HRS03, KNK04, Lem00, Lev03, LS00b, Lu00, Lu01, Lu04, Lu06, MRTZ03, MM06b, Rom08, Sin05, Vel08, ZZ06, Zha08, ZS02, AN05]. Solvable [BdGM01, EG05, JL01, KMVE03, OBB03, Sam01a, SJ01a, Sam04, Sam05b, Sea02, TM02, VD01]. Solvation [CKKC00]. Solve [Lee05, YBMS06]. Solved [BGOY04, GK02]. Solver [TKR02]. Some [AM04, Ale01b, BH00, BBBF06, Bax09, BCEP04, BCEP06, BCG00, BCG01, BM02a, BBP02, BCFP05, Che09a, CGV02, DD04, Dom03b, DLS08, Fre04, GR06, HMS05, ISS07, KL09, LR00, Naj08, Pey09, PS02b, PS04b, Rue02, Sod09, Sod17, TTK01, TV00, Val09a, Var08a, aYL05, dlLV07, RY07]. Soret [GKR02]. Sornette [Por04]. Sources [AC05, BR00a]. Space [ALR07, BGS06, BBT03, BK09, CPC03, CK03b, FF00, FLM00, GTW01, KCT08, KS07, PS08b, Rad08b, RRS^{+}00, SI04, SS09e, YBMS06]. Space-Discrete [PS08b]. Space-Fractional [YBMS06]. Spaces [CO08a, Got05, Mis06, Yoo07]. Spacing [BG00a]. Spanning [CS09a, CCY07, JSS05]. Sparse [BS09a, Sod09, Sod17]. Spatial [AR02, CRV00, FK09a, Kom07, Mie04, Sch07, VEW08, dARBY04]. Spatially [CGR07, Fou06, FG08b, Lou03, Lou07, Lou08a, Lu01, MDB00, Rug08, ZZ06]. Spatio [CP04, MGT07]. Spatio-Temporal [CP04, MGT07]. Spatiotemporal [GOBY06]. Special [Vil06a]. Species [ALS09, EA07, KK00a, MZ02b]. Specifications [DN04]. Specified [Kad03b]. Speckle [Fre08b]. Spectra [BD09, BSB00, Ber03a, CdO02, CP02, Hir05, TDM02]. Spectral [Ale01a, AY01, AL01, ALS03, ALS06a, AS03, BG01b, BLZ01, Cam06, DH08, DJW07, EK05b, FS05, FLM00, GF08, GM05, HMH05, JL00b, Kom00b, KZ07, LVZ03, Naj08, TG02]. Spectrum [ALR07, Ast08, BA01, DLS08, GvB02, Gen04, IK09a, Kor00, LdO03, LTWW02, MZ02b, ND08, O’C01, PR07, Rut08, SO00b, SO02]. Speculation [Mas03a]. Speed [BLR08, BFS09, CE02b, FM02a, HHB09, Kol08, Mas02b]. Sperm [MV03]. Sphere [FPS07, Pan02, Rei00]. Spheres [BC04, CK04, CM04b, CM04c, CM06, ETB06, LT04, Lyb05, MMR06, MM06b, Pol00, Sot07, Urr08]. Spherical [De 07, DGM07a, DGM07b, FT06, GGZD02, HB00, KHV00, KS06, Koi07a, Koi07b, MCZ06, MCG08, Pat07b, RGRTS06, Urr08]. Spikes [MH07]. Spin [APS09, ABY07, Alb08, Ale01a, Arg07, AK09, Bar06b, BdGM01, BGOY04, BCF^{+}09, BCMP04, BK07, BCS09, BC02e, CM04a, CGI02, CGG09, CS09c, DK02, De 04, De 07, DN03, DGM07a, DGM07b, EGGI01, FM08, FBG02, FK04, FK05, Fis06a, Fis06b, Fis07, Fis08, FL03b, FT06, Fra07b, FSZ01, GHO^{+}00, GT04b, Has00, ISS07, Ito02, JK04, KCT08, KPT05, KCK07, KRT00, KMSS01, KM09b, Led00, Li06, MNS08, Mal07, MPRT^{+}00, Mas03b, MRTZ03, MZ02b, MP09, MPD09, NR04, NS02, Nis07, Pan08, Pod02b, PFK06, PS01, Rác00, RdCM04, Rut08, SO00a, SO00b, SO02, SY03, SCM04, Ser06, Sla09b, Tia04, VHO09, Wat04, WD01, WB04, Yar08, Pod07b]. Spin- [ABY07, Alb08, KCK07, KMSS01, BC02e]. Spin-1 [Yar08]. Spin-Chain [Rut08]. Spin-Charge [FM08]. Spin-Glass [Has00, KRT00]. Spin-Orbit [Mal07]. Spin-Reflection-Positivity [Tia04]. Spinless [LFB04, Mes02a].

Spinning [Pom05]. Spinodal [MPSW00, MVE04]. Spinons [Rut08]. Spins [FBG02, FG08a, Fin03, Tin03, XKHK08]. Spiral [HI04, Sum05]. Splitting [DMS02, GS01]. Splittings [PSW04]. Spontaneous [AR02, DH04, FL06, GSW07, Nam04, Nic01, RSV09]. Spread [FBC^{+}05]. Spread-Out [FBC^{+}05]. Springer [Mas02a, Opp03b, Pod04b, Sah02, Shl04, Wes04]. Square [Ace00, ADLM01, CJS01, CSS02, CN01, DB00, DB01, Dys04, JS01a, MO09a, ONGP01, SS01, Suw09, TPV09, WP03]. Square-Lattice [JS01a, SS01, TPV09]. Square-Well [Ace00]. Squares [HHB09, KS08]. SRB [AOT06, Wol06, You02]. Stability [AS01, HMS05, MRW02, Mie04, NSCW04, ST02, Uma09, Woj03, Wu09, Zha00a]. Stabilized [BHRW04]. Stable [Aki08, Bar04, BCK04a, BDP02, Eps06, Fre01, Lem00, Sim08]. Stadium [LR02b]. Stadium-Like [LR02b]. Standard [Bal00, Sig09]. Stanley [Mas00]. Star [Bjö09]. Star-Like [Bjö09]. Stars [Kie09a]. Stat [AV03, Ano01d, BPS03]. State [AVE07, BR05a, BP02, Bol09, CR01, Cai04, CJS01, DLS02, FHAY06, GCU02, GS03b, GS09, Has00, HNS04, HWB01, HI05, JSS05, Kie09a, Kie09b, Kon02b, Kur03, LPD08, Lee09, LY01, MBL08, MN00, MO09b, NS02, OvWLH00, RC01, RS04a, RNVRP04, Saa07, SD09, iST06, SY08, Sig09, SST^{+}00, TC08b, Tél07, Yar05, dlLV07, Gar08]. State-Dependent [BR05a]. States [ABS01, AR02, AP03, BGGM04, BDG^{+}02, BDG^{+}04, BCT06, BDL00, BNZ02, CdlL05, CR07, DK02, DJZ00, Dun03, ER02, EMZ06, Far05, FRZB09, FPSW01, Gat00, Got05, HK01, HE03, HI04, JM00, JMU05, KS00, KNST09, KW00b, Lee05, LY07, LZ09, MO01, MRW02, Mor07, MKL08, MPD09, Naj04, PA02b, Rei08, Rei05, RK05, Rom08, Roz06, Rue00a, RSV09, SO00a, SO00b, TNK04, TC07, TE00, Uma09, Var08b, VMT02, Woj03, dZS04]. States-Exclusion [ABS01]. Static [CLK09, HDMF04, LM06, SBG03]. Statics [Opp01]. Stationary [AN05, AN06, AR02, Bah07, Bak06, BFGM02, BDG^{+}02, BDG^{+}04, BW07a, BMP07, BDL00, CR07, DLS03, Dun03, EFM09, Far05, Gat00, GSS03, LPD08, MRW02, PS04a, Rot06, Sak00, SD08, ZF09]. Statist [Cau02b]. Statistical [AR00, Ano00a, Ano00c, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00x, Ano00y, Ano00z, Ano01a, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01u, Ano01v, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Asa00, BD06a, BBK05, Bax06]. Statistical [BD00, BP01b, Bog09, BPS06b, CD09a, CP03, Che06, CZ09, CMSR05, Dag03a, Dom03b, Dud07, ES03, FVE03, FKÖ03, FJMR02, GR09, HC08, HB00, JP02, KK02b, KTZ06, Kuk04, KMG07, Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Leb07b, Leb07c, Leb08a, Leb08b, Leb09b, Leb09c, Luc08b, Mas02a, MH07, MV07, New08, Opp03a, Opp04b, Opp06c, Opp07, Opp09, Par08, PV02, Pod01a, Pod06a, PG99, PA00, PA02a, Rap01, RB08, Rit01, Roe03, Rue00a, Rue02, SBG03, Swe02, Sza06, Tah09, Tao01, Val09b, Val09a, VBM06, WW05, dlLV07, PG00, Ton08a, Dag03b, Kan01, Opp01, Opp02a, Opp03b, Opp04c, Opp06b, Pod02b, Rub06a]. Statistically [BR01]. Statistics [ABS01, AG05, BJ06, Bar04, Bar06a, BK07, Bur08, Các08, DS07, EGGI01, GL01b, GGL09, Hos00, LP08, MB06, MvWH^{+}01, Rid04, Ryc08, SKT09, TE00, WLTH07, Wit03, Pod06b]. Status [SF07b]. Steady [AN00, AP03, DLS02, ER02, EMZ06, HE03, KNST09, LY07, MT08, MO09b, OvWLH00, RS04a, Saa07, SD09, iST06, TC07, TC08b, VMT02]. Steady-State [RS04a]. Steep [CGK^{+}04]. Stefan [Dir04]. Steiner [Git09]. Stell [SSLG^{+}00]. Step [AVBM06, DG09b, DG09a, FS03, GMNT05, IS07, Roz06, TW09]. Sticky [GH04]. Stieltjes [Ruz00]. Stochastic [Ale01a, AY01, AS08, Ano01c, Bak06, BBBF06, BBS04, BGL05, BDG^{+}06, BL08a, BLL04, BLLO09, BKL00, Các08, CKSZ06, CF03, Cha09, CN03, CD05, DD04, DN03, DB00, DB01, DM02, DOS01, DS07, DKK02, EL02, ES07, EK05c, FF07, FK09a, FM07, FG00, GOSG00, GvH07, HS09b, Has02, HN03, JL00a, KN04, KT06, Ken04, KZ07, KMP^{+}07, KZ03b, KC03, LR02a, LNT09, LKMGG07, MS05a, Mas09a, MS05b, MDW09, Mie04, MGT07, MKL08, Mur08, Pia00, PS02b, PM06, RY07, Roe03, RF02, SD09, Sak00, Sch07, SCM04, SS09e, SZ03, TC08b, VEW08, Vel00, VD05, Yos08, ZF09, bA03, Mas02a]. Stochastically [BCK04a, MSVE07]. Stochasticity [EK03, EA07, Fre03]. Stochastics [Mat02]. Stock [Por04]. Stokes [Bak06, BCV00, BKL00, Che09b, CLM05, DGR08, DLS09, FJMR02, GM07a, HPWW03, KP05, Kuk06, LL08a, Mat02, MS05b, Rom04, Rom08, Sab08, Sin05, Suz01]. Stokes- [HPWW03]. Stokesian [CS08c]. Storage [GR00]. Stored [EMC09]. Strands [BGMM03]. Strange [EZ04]. Strategy [MTT08]. Strength [ES04, JK01, WM07]. Stress [Jan01a, SMRK^{+}02]. Stretched [Che07]. Stretching [AG05, SKT09]. Strict [Fou00, HT01]. String [Mun07, Pro09b]. Stringari [Rub04]. Strip [CS03a, CP02, PM04]. Stripe [DJ07]. Striped [Shi06b]. Stripes [LFB04]. Strips [CSS02, CJSS04, CS08a, CS09b, MP06]. Strong [Ano01d, BEPK^{+}02, BR04, CCL09, FG08b, GKR02, Gen04, MG02, NT04, TPV09, Weh97, Skr00]. Strong-Disorder [TPV09]. Strongly [HJ09, LS09a, Mes02a, Tia04, Woj03]. Structural [De 07, Dot04]. Structure [AK07, BG01a, BS09a, Bog07, CS09b, DSC02, DOS01, FS07b, FJM01a, KW00a, KK00a, Piv03, PA00, PA02a, SC01, Sot07, SDBS07, Sza06, YS05]. Structured [Pod05, Pod05]. Structures [AP08, De 04, GOSG00, Rei00, Rub06a]. Students [Wei05]. Studies [BLM03, Che09b, QAK02, Suz03, WDMB05]. Study [BD06a, BLR08, CD09b, Dom05, Ghe09, GGZD02, GM00, HGST00, HvdHMO09, Jaf03, JS08c, KMVE03, Mas03a, MH07, OP04, RS04b]. Studying [WE07]. Subadditive [Han07]. Subadditivity [NT04].

Subcritical [AV08, dHP09]. Subcubes [MZ08]. Subdiffusion [Mag09b]. Subdiffusive [BSB00, Mag09a]. Subextensive [Fis07]. Subgraphs [VB07]. Subharmonicity [Bou09]. Subject [Bur08, LNM06, PS08b]. Subordination [EK05b]. Subsequence [AMV08]. Subshifts [JMU05, Mor07]. Substrate [DMSR00]. Substrates [DMSR03, Xin09]. Such [CS04, CS09d]. Sufficient [PS04b]. Sullivan [Lan06]. Sum [CS02, Jan00b, KMST00, LM06, Sam09]. Summable [FdL06]. Sums [CL03, Gal06, KS08]. Sunday [Leb07b, Leb07c]. Super [GW09]. Superadditivity [Luo07, LZ08]. Superbosonization [BEK^{+}07]. Superconducting [RR01]. Superconductivity [Hod01, Noz04]. Superconductors [Rub06b]. Supercritical [Too02]. Superdiffusive [Ben05, CEMM09, KO02]. Superdiffusivity [LRY05, SS09c, Set06]. Superfluidity [OP04]. Superfractals [bA08]. Superintegrable [Bax08b, Bax09, ND08]. Superradiance [PVZ05]. Supersaturated [Shn03]. Superstable [BZ08b, CMS05]. Supersymmetric [GW09, MS08a]. Supersymmetry [Has00, TTNK06]. Support [CM00]. Supported [Vel08]. Supports [ZZ06]. Sure [Blo01, FFN09]. Surface [BCG08, Bax03, Cau02b, CMSR05, CMN06, CG03, Dir04, DLR04a, DLR04b, FPD01, GK09a, HNOV07, HK02b, KTH^{+}05, Lev03, LD01, Sak04b, Sak05, Sam01a, SJ01a, SK06, TJ08, ZK00, dMG04]. Surfaces [BPS06b, Cau02a, DDMSR04, DP04, EC07, Koi07b, SSH06b, SSK02, ZJ00]. Surfactants [Pod05, Pod05]. Surmise [GT08]. Survey [BGS06, vdHK01]. Survival [KKT00, Opp02a, Sak04b, Sak05]. Susceptibilities [AYP07a]. Susceptibility [AYJP01b, AYP07b, JS04, ONGP01, Sam00]. Suspended [DA03, ICO02, SAPD05]. Suspension [BI06]. Suspensions [HRK^{+}05, LV01, SHD03, SMRK^{+}02]. Swelling [Mez01]. Swift [CE02b]. Switched [EK00]. Switching [WLTH07]. Symbolic [FF00]. Symmetric [DF02, DLS02, DDR04, DG09b, DB00, KM09b, WNUK01]. Symmetrically [PS07b]. Symmetries [AH01, GKRV09, LK06a]. Symmetrized [FR07]. Symmetry [AGL01, BW07a, BBC^{+}01, BR04, De 06, DFM01, DWMB05, Fei03, Fuk00, FW07, GSW07, GR00, JW04, JL01, Lou05, Luc08a, MPRT^{+}00, NS02, SWK04, Tak09a, Tak09b, WDMB05, Sha10, Tak10]. Symmetry-Breaking [BW07a]. Symplectic [CDG07, DGKV07, FEDZ07]. Synapses [Bal09]. Synaptic [CRVV09, HS04]. Synchronisation [PY09b]. Synchronization [DB00]. Synchronous [RVMO07]. System [Aki08, ABY07, Alb08, AS05, And08, AK09, AR02, AS01, Bak06, BDL00, BCDM07, CM02a, CP01, Che09b, CP04, DLS03, DB00, DLS09, Dir04, FNO06, GSW07, HR06, HS09b, HMPPMV00, HMS05, Hua09, IVZ06, Jab01, Jaf03, JS01b, KCT08, KNP07, KL06, KL08, Lat02, Ler00, Luc09a, Mac09, MRW02, MSVE07, MS03, McD01, MDW09, Mür01, NS04, PS03a, RS04a, Rei00, SS04, Sin05, Sme08, TDM02, Tin03, UC07]. Systematic [AC01b, KPvB06, ZW04]. Systems [APS09, Ada01, AGL01, AGY02, AH03, AFHV03, AP08, AV03, Ano00h, AMP04, Asa00, BDL06, BMN09, BMA02, BBDR05, BM04c, BDG^{+}09, BD06b, BCC06, BCK00b, Bla08, BDLvW08, BCMP04, BB05, BCM02, CR07, CG00b, CR01, Cai04, CCG04, Car07, CEFM00, CDG06, CLM07, Che09a, CZ09, CE04, CGI02, CS04, CS09d, DJ05, DK02, DDB^{+}00, DN03, DJ04, DLS07, DG09a, DHB07, DLS08, EG00, EFPZ05, EMMZ06, EK03, ETB06, FK05, FS05, FS06, FS07b, FM07, FL03b, FdL06, FPSW01, Gal06, Gal01, GHO^{+}00, GS03a, Gen04, GMW02, GKRV09, GLTZ06, GJ08b, GT08, GS01, GR09, Han02, HMH05, Har08, HNS04, HNO05a, HNO05b, HPF^{+}02, Hor06, HJ09, JPZ09, JOP06, JS01c, Jan03, JS08b, Jia03, KPT05, Kom00b, KNST09, KW00b, KG00, KM08b, LS00a]. Systems [LS02, Lan09, LNM06, LVZ03, LARvW07, LRB05, MRS07, MPSW00, MS05a, MN04, MDB00, MNNS07, MMR08a, MT04, MS09, MPT06, MZ02b, MK07, MS06b, MP09, NOS06, NR04, NKC00, Paj07, Pan02, Pri03, PS01, RC01, RM00, RS09, RdCM04, RM08, SJW07, SSLG^{+}00, SO00a, SO00b, SO02, Shl04, Shl07, SSE05, Skr00, Skr03, SDC04, Suz03, Swe02, TC08b, Tao01, Tél01, Tia04, TT05, TV00, TV03, TC03, Var08b, VTM00, WNUK01, WS02a, Yar05, You02, dPG04, dSI04, vEK07, AV87, Ton08a, Bog09, GC08a, Por04, Rit01, Rug08, Wei01, bA03, Ano00b, Shl05].

T [Shl04, LB09]. T-cell [LB09]. T.M [Gar08]. Tableaux [KG03]. Tagged [BTT07, BT04b, GJ08c, IS07]. Tail [ASB02, KW00a, MM06b, PDvB00, Zhi00]. Tails [BM04b, BK01a, BGP04, BG06, BM00, EB02, KBS02, Naj07, Naj08, PT06]. Takahashi [BMD00]. Takata [Sha10]. Talagrand [FT06]. Tamas [Pod02a]. Tanaka [Opp03a]. Taqqu [Nol04]. Targeted [EK03]. TASEP [BFPS07, BFS09, DPP09, EFM09, Son09]. Taxonomy [Oon03]. Taylor [AN06, CT04b, PF07]. Technique [Bao01]. Techniques [Bar08, BBDR05, SM06, vZvB02]. Temperature [ABCM08, BM04a, BCF^{+}09, BM02b, BC02e, CN04b, CCG04, CKKC00, CGT05, DG08, DGM07a, DGM07b, DP03, DLR04a, DLR04b, DKM04, EGGI01, Fuk03, Gir01, GT04b, HY04, HN03, Iar01, JMU05, KMSS01, Kur03, LM06, MS03, MP03, Mor07, RC01, SSD00, Sam07, SO00a, SO00b, SO02, Sot07, Wu02, vER07, CR05]. Temperatures [ABY07, Alb08, TNK04]. Temporal [CP04, MGT07]. Tension [BTM^{+}04, CRV01, Dir04, DT00, FPD01, GLMSR07, HK02b, Lev03, Sam01a, SJ01a, dMG04]. Tensions [CMSR05]. Tensor [Jan00a, Jan01a, JS04, Sam00]. Tenth [CM06]. Term [FF00, Lou03, Lou08a]. Terms [CMN06, HB00, ZF09]. Ternary [MR04, MR05]. Tessellations [Der08, HC08, Luc08a, Luc09b]. Test [BCG08]. Testing [AVE07, MTT08]. Tests [Ken04]. Their [BC02c, CS03a, CS03b, DM09, FM09a, FS06, PTZ03, Pey09, Tid01b, BR05b]. Them [You02]. Theorem [AK07, AG07, BC04, BGGZ06, BR05b, Bud08, Car07, CY09, CS04, CS09d, DP04, Han06, Jar00, LRB00, LK06b, MNBO09, Mon04, San05, Sch06, Shi09, Sin08, aYL05, vZC06, HMH05]. Theorems [CC09, CEO07, DMN06, FT06, GTW01, Pen01, SC09, TC07, WW01, vdHK08]. Theoretic [FK05]. Theoretical [AFFS06, Dro00, HMO07, MPRT^{+}00, QAK02]. Theories [AKR01, BP01a, Kad09, NS02, Ste04, Hol02a, WW05]. Theory [ABCM08, AM04, AG07, AVE07, Ast08, Bal00, BLZ01, Bel04, BDG^{+}02, BD03a, BBCK04, BCK00b, BP08b, BEK^{+}07, CR01, Cai04, CIS01, Cer05a, CR03, CW06, CNV01, Con04, CDG07, DWMB05, DD06, DGZ04, EY01, EVE06, EA07, FLW08, FM05b, FS05, FGR03, FJMR02, FZ03, FV02, Fre08b, GW09, GC08a, GZG05, GG01, HY04, HD02, HDMF04, Hod01, HNO05a, HNO05b, HRS03, Jaf03, JP02, JOP06, Kie04a, Kie04b, KM08a, KTZ06, KC03, Kyt06, LS00a, LS02, Lia06, Luc08b, MS09, MS08a, MVE04, Opp04d, PD06, PO07, Pol00, Pru01, RC01, Rei05, RK05, Rit01, Roe03, Rub06b, SS01, SW00, SO08, Sin08, SJ01c, Str06, TJ08, TC07, Tid01a, Val00, WBE02, WW05, YS00, ZGA02, ATH06, Pod07b, Gal01, Han02, Mas02a, Nol04]. There [DMSR00]. Thermal [Bao04, Ber08, BD02, BC02b, BP09, FMS04, GLTZ06, GPL03, GMR06, PVV02, PD05, PBFC05, TK02, dZS04, Ano00d, Opp02b]. Thermally [JOP06]. Thermodynamic [BGW01, Car07, CBKM04, CG04, CS09c, DTP02, EA04, FF04, GPL02, HD02, KS02, Kie09c, LARvW07, RY07, Rid03, Rit06, Rue00b, SF07a, ST00, Sam01b, Ser06, SLA02, Süt03, TB06, Uma09]. Thermodynamical [AMV08]. Thermodynamics [ABCM08, Ano00f, BDG^{+}09, CCMT09, Dag03a, FGC09, FK04, GPL04, Jan06b, Koh01, Lev03, LL08b, LL09b, Mal05, Opp06a, Opp06c, PG03, QQT02, SF07b, iST06, Shl04, TC08b, WW05, Gar08, O’C06, WA09, Opp04b]. ThermoElectric [Jac09]. Thermostat [LNT09, Li07]. Thermostated [MMR08a, RKH00, Wag00]. Thermostating [RKN00, VTM00]. Thermostats [BL08a, DP07a, Gal09, Rue00b]. Thermostatted [BDL00, Lat02, WW06]. Theta [KIK08]. Thickness [DOS01]. Thickness-Density [DOS01]. Thin [AB09, DS07, GD06, GMR06]. Thin-Film [GMR06]. Thinking [GK09b]. Thinning [Keh05]. Third [O’C06, WA09]. Thirring [ES04]. Thomas [HY04]. Thomson [BCK04b]. Thouless [GTZ02, MM05b]. Three [BD03a, BC02e, CMS05, DM09, Ess03, KAS03, KMP^{+}07, Luc09b, MD01, SKT09, Urr08, WMDB02]. Three-Dimensional [BD03a, BC02e, KAS03, KMP^{+}07, Luc09b, WMDB02, SKT09]. Three-Particle [MD01]. Three-point [DM09]. Threshold [PS04b]. Thresholds [LMZ00]. Tiles [Nie01]. Tilings [BH00, DMB01, DWMB05, WMDB02, WDMB05]. Time [AER05, Aus00, BJ06, BGGM04, Bar06a, BH08b, BBT03, BBS04, BP01b, BC09, BK09, BCM02, Các08, CM02a, Che07, CGV02, CCL04, DP07c, Erd02, Ewo01, FF00, FS05, Fin03, FN00, FW06, Fre08a, Gal06, GJY04, Gam03, Gar01, Gas04, Gas07, GJQ06, GL01b, GGL09, GE09, GTW01, GPL03, HDMF04, HKW07, KKT00, KBS02, KO05, Kra03, KC03, LR02a, Lin04, LK06b, LR00, LR02b, Lu00, MN03, MCZ06, MNOS04, MB06, MDW09, MM02, MS09, MS06b, PDvB00, PM06, QQT02, SS04, Set06, SMT03, Sme08, SZ03, Too00, WZ09, WLTH07, vB03, vWL00, vZC06, CT07]. Time-Decay [GJY04]. Time-Dependent [BGGM04, FN00, LK06b, LR00, LR02b, MS09, vWL00]. Time-Mixing [KO05]. Time-Periodic [BC09, CCL04, GJY04, GJQ06]. Time-Reversal [MN03]. Time-Reversed [Ewo01, Gas04, Gas07]. Time-Reversibility [QQT02]. Times [CF03, Cha00, Mag09b, NAS02, Zab06]. Timescales [Fin03]. Timestepping [JL00a]. Toda [AKR01, EEDJ00]. Toda-Type [AKR01]. Toeplitz [JK04]. Tom [Sim09]. Tomoyasu [Opp03a]. Tonchev [Han02]. Tool [GOBY06]. Topological [BS08, Eck07, Rad08b]. Topology [CPC03, CCMT09, MN04, RGRTS06, Xin09]. Torquato [Sah02]. Torus [Arg02, CM00, FW03, PM04, SS00a]. Total [NP01, Ric04]. Totally [BK08, GM05, Sep01]. Toy [KK02a, MZ08]. Trace [Gas02]. Tracer [Ben05, GvB02, KO02, KP04, Sch00, TB04]. Tracers [KO05]. Tracking [ER00]. Tracy [Sod17, Sod09]. Traffic [AS08, BKNS01, Bla03, Bla05, GG01, KL03, LZGM04, VST09]. Trafficking [HS04, Hol07]. Trajectories [FFN09, LKL09, Mac02]. Trajectory [MS08a]. Trans [Opp03b]. Transduction [WE07]. Transfer [Bax04b, BBS04, CS09b, JS01a, JSS03, JS06, LB04, MZ02b, SS01, SS09b, SO00b, SO02, SNC05, Tid01a, Bax06]. Transfer-Matrix [MZ02b]. Transform [KR07, PF07]. Transformations [CL00, Gam03]. Transient [AFN00, KS00]. Transition [AV08, AVE07, BBLS05, BBD04, BG09, BG04b, CN04b, CP01, EVE06, Ess03, FG08a, FKÖ03, FPS00, GN08, GM07b, GS08c, GS08d, HD03, HN03, HAG02, IVZ06, Jaf03, Kas02, Koi07a, Koi07b, LR06, MNS08, MKB00, Mes02a, Mes02b, MS06a, Noz04, OvWLH00, OEB01, Pod03, RVMO07, Rei00, RG04, SZ04, Ton07, TB08, Uma09, WS02b, Yar08, Git07]. Transitions [Aba00, AH03, Ano00g, BBCK04, BCC06, BDLvW08, BBC^{+}01, BB05, CPC03, CS04, CS09d, DM06a, DLR04a, DLR04b, EN03, ER00, FSG08, FS08, GZ01, JK01, Jus01, Kad09, KCK07, LZGM04, MRS07, MH07, MM06a, MTT08, MGT07, MP09, NH03, Opp05, PZ06b, PPR00, PM05, RGRTS06, Rub09a, TNK04, Tur06, Yos08, ZKBG03b, bA03, vE00, CS09d, Pod01b, Pod04b]. Transitive [AV08, Sla09b]. Translation [AGL01, EG00, JL01, SWK04]. Translation-Invariant [EG00]. Translational [AR02]. Transmission [BP01b]. Transport [ADG00, AFHV03, ALS06b, AMT07, AEGS04, BJ08, BPS06a, BSB00, CPK07, CdO02, CdO07, CMV05, CLK09, Coh09, Coh10, CEMM09, Cor05, DMR05, DR06a, EMMZ06, EE07, ECSB00, Gar03, GA05, GM07a, Gat00, GK09b, GK02, Jac09, KC02, KP04, Kor00, LPD08, LLMM03, LDNG08, PS08b, PS02b, RD08, Tid01a, Wei09, Wes09, GC08a]. Transports [TG00]. Transverse [PAY09]. Trap [Leg03, MCZ06]. Trapped [Cso00, Süt03]. Traps [Gra00, NOB00, OP04, Süt04]. Traveling [SDBS07]. Travelling [Bla08, CO08b, CD05]. Treated [Hos00]. Treatment [EE07, Sla09a]. Tree [BCO04, DFK08, GTA09, GS08b, Pat07a, PR07, Roz06]. Trees [BH08a, BS09b, CCY07, DJW07, FS08, Jen01, LMP05, Pre03, Roz08, Sud09, Sum05, VHO09, vR01]. Trend [CK00, TV00]. Triangle [HT01, RT08c, SST^{+}00]. Triangles [CE02a]. Triangular [CJSS04, CL06, CRV01, CSM03, MN00, Wu02, JSS03]. Triangular-Lattice [JSS03]. Triangulated [Koi07b]. Triangulations [CE05]. Trickiness [Sam07]. Tricritical [CEO07, EMO08, Lüb06, PIRB03]. Tridiagonal [CCM00]. Trigonometric [FY06]. Trinomial [War01]. Trivial [CJM09, KK02a]. Triviality [Hos06, MZ02a, Wat04]. Trotter [Suw09]. trough [Bar08]. Truncated [CS02, Sal09]. Truncation [FdL06]. Trusted [Ano00e]. Tsori [Git09]. Tube [KL03]. Tube-Like [KL03]. Tubular [MVW08]. Tuesday [Leb07b, Leb07c]. Tug [MKL08]. Tug-of-War [MKL08]. tune [Lan06]. Tunnelling [MNOS04]. Turbulence [ALS06a, AH07, BCG08, BBT03, BBS04, BB09, Ber07, BET00, CLT07, CG03, ECC03, FK00, FJMR02, FJMR03, Gir03, GN03, Hos00, JH03, KMVE03, MDL^{+}03, OT03, PS03b, SFCO02, Win02, YS05]. Turbulent [Ben05, Fan05, Kra03, KMG07, SWK09, SB06]. Turning [DM08]. Turtle [vB03]. Twice [GP08]. Twist [CCAD08, GK09b]. Two [AR02, ALS09, Bax00, BKNS01, BR06, BDM09, BG04a, BCD^{+}04, BM05a, BDL08, BP08b, BFS09, BET00, BKL00, BNZ02, BC02e, Cam05, CJM09, CGJ^{+}05, Car09, Che09b, DM09, DLS09, DKM04, Fan05, FT08, FT07, Gat00, Gen04, GS00, GP04, GTZ02, GPL03, HK01, HNS04, HJ05, HNOV07, HPF^{+}02, Jan00b, JS08b, JNH^{+}09, KMST00, KS02, KIK08, Kra03, KG03, KPvB06, KP05, Kuk06, KM07, KKS05, LRY05, LW07, LFB04, LWL00, LY01, LL00, Liu09, LZ09, Mai03, MS05a, Mal07, MB03, MO09a, MT04, MRTZ03, PDvB00, PBFC05, PS03a, RS04a, cR04, Ric02, Riv02, Roz06, Rut01, Sak04a, Sak06, Sak07, SS00a, ST00, Sam01b, SJ02a, SJ02b, Sam03, Sam04, Sam06, TJ08, Tél01, TM02, Tél07, UC07, Urr08, WK00, Woj06, vEK07, Tah09]. Two- [BC02e]. Two-Body [BDM09]. Two-Channel [PS03a]. Two-Component [AR02, FT07, Jan00b, MT04, ST00, Sam01b, SJ02b, TM02]. Two-Dimensional [BDL08, BET00, BKL00, Cam05, CGJ^{+}05, Car09, FT08, FT07, Gat00, GS00, GTZ02, HPF^{+}02, Jan00b, JS08b, KMST00, KS02, KPvB06, KP05, Kuk06, KKS05, LFB04, LWL00, LY01, LL00, Mai03, MB03, MO09a, MT04, PDvB00, Ric02, Rut01, SS00a, ST00, SJ02a, SJ02b, Sam03, Sam04, Sam06, Tél01, TM02, Tél07, UC07, WK00, CJM09, Che09b, HNOV07, Woj06]. Two-Frequency [Fan05]. Two-Lane [BKNS01, JNH^{+}09]. Two-Level [BG04a, Gen04]. Two-Loop [KM07]. Two-Phase [Liu09]. Two-point [DM09]. Two-Qubit [LZ09]. Two-State [HNS04]. Two-Step [Roz06]. Two-Surface [TJ08]. Two-Temperature [DKM04]. Two-Time-Scale [GPL03]. Type [AKR01, AS06, AAP02, AN06, BKOS07, BZ09, BBC^{+}01, DGKV07, DR01, Dir04, FF07, JMU05, Kol04, Kol07, Mor07, PVZ05, TTK01, WBE02]. Typical [BS09a]. Typicality [Rei08].

U [Git09]. Uhlenbeck [Ben05, EK05b, EK05c, Fan04, KL06, KL08, MM05a, TM08]. Ulam [BLP09]. Ultracold [OP04]. Ultrametric [De 06]. Ultrametricity [CM04a].

Unbounded [APS09, BK03a, MVW08, PS01, SZ03]. Uncertainty [Gar06, GI08, GII08, LZ04]. Unconventional [DRC04]. Uncorrelated [BGGM04]. Uncoupled [HNS04]. Understand [SBG03]. Understanding [BGX01, SSLG^{+}00]. Undriven [BNZ07]. Unfolding [CRV00]. Uniaxial [KCT08]. Unidirectional [FK00]. Unidirectionally [SJ01b]. Unified [MCZ06]. Uniform [AR06, CSN02, CNV01, Elo08, LdO03]. Uniformly [AP06, Fra07a, GP08, PZ06b]. Unifying [Pru01]. Uniqueness [Bak06, BP02, FM05b, Fou06, FG08b, Gho05, GS07b, Mol04, Yar05, DMPV08]. Unitary [BP01a]. United [Bra02]. Unity [BA01, DFM01, FM01a, FM01b, FM09a, Tas06]. Universal [CZ03, CETT05, Fin03, FG00, LW04b, SS00a, Sea02]. Universality [BM05b, CDG07, DLM05, DGKV07, FPSW01, JS01c, KKS05, LLSA00, Mas03b, MVE04, PS08a, Shc09, Sos02, vdEvdHH08]. University [Aba00, Bog04, Bra02, Dag03b, Dom04, Dor02, Kan01, Leb04a, Leb05b, Leb05c, Leb06c, Leb06b, Leb07b, Leb07c, Mas00, Mas02a, Mas03a, Opp02a, Opp03a, Opp04a, Pod01b, Pod01a, Pod02a, Pod02b, Pod03, Pod04a, Por04, Rap01, Rit01, Rit03, Rub04, Spo03, Taq04, Wei01, bA03]. Unmix [GC05]. Unoriented [vdHJ04]. Unravelings [DDM08]. Unreliable [FBC^{+}05]. Unsolvability [Too00]. Unstable [BG04a]. Unusual [CJS01]. Upadhyay [Buc07]. Update [PM06]. Updating [APL04, Cir02, CN03, RVMO07]. Uphill [vB03]. Upper [CE02b, LMP05, LW04b, Pri00, SJW07, YY09]. Urban [VB08]. Use [PVV02]. Used [JHW09]. Using [Bou09, CK04, Cop08, DSC02, JL00a, LS00a, LS02, Mas09b, NAS02, SLB00, YBMS06].

V [GCU02, Pod07c]. V.A. [Ach08]. Vacancy [PdKV09, TW03]. Vacuum [Jia07, Wu09]. Valence [FRZB09]. Valence-Bond [FRZB09]. Validations [FKLM08]. Validity [BMN07]. Value [Bjö09, Bur08, KM08b]. Valued [HS09b]. Values [BZ09, FK05, Mis06, MK04, Sal09, SS09d]. Vanishing [BM04c, ABC05b]. Vapor [GPT02, TAL06, dMG04]. Vapour [PIRB03]. Variability [SSB04, SS00b]. Variable [GJ08c, GMNT05, Hei04, Too04]. Variables [Ano01d, FLS07, MNBO09, Sem08, WW01, Weh97]. Variance [BK08, FPS00, Set06, vdEvdHH08]. Variation [Asa00]. Variational [AS01, BD04a, BD04b, FLR03, GS07a, Hir05, Pre03, Pre07, Yoo07]. Various [LR06]. Varying [LD01]. Vector [AVZ00, BCS09, Kor00]. Vectors [Hua09, JS01e]. Velocities [Win02]. Velocity [ASB02, BZ03, BTV09, CM00, CO08a, CC00a, CGH^{+}03, CC00b, Cor04, DGZ09, FH04, HKW07, KK02b, KK02c, KP04, Lu05, MDL^{+}03, Pia01, RS09, RRS^{+}00, RK00, Sab08, Val09b, Val09a]. Verena [Hol02a]. Verhulst [Các08]. Verification [Rut01]. Verlag [Pod04b, Sah02, Shl04]. Version [Liu09, Tum06]. Versions [Law09]. Versus [DJ07, GG00, LSSW03, Ric00]. Vertex [Bax02, Bax04a, BL09, DFM01, FM03, FM05a, FM09a, KP08a, WK04]. Vertical [EK00]. Very [DN04, Lu06]. Vesicles [Car03, KIBG^{+}06]. Vespignani [bA09]. Via [Sal09, All08, Bak06, BJR02, BET00, BCM02, CL03, CMS05, EMO08, FL03a, Has00, HPCH00, HMG08, JS01d, Ryc08, Tah09]. Viana [GT04b]. Vicious [DPP09, KG03]. Vicsek [Pod02a]. Victor [Rit01]. Video [PR07]. Vieri [Rub09b]. View [BCF^{+}09, MCG08, Nou02, NOV04]. Viewed [FS06]. Views [Ano01c]. Villani [Häg07, LS04b]. Violation [LK06b]. Virial [CM04c, CM06, Lyb05, cR04]. Virtual [JAWC08]. Virtues [vWL00]. Virus [Hol07, Sch07]. Visco [LM06]. Visco-Elastic [LM06]. Viscosity [LOV04]. Viscous [BBL00, CRT00, CT04b]. Visual [LK06a]. Vlasov [CM00, Jab01, Lan09]. Void [DS07]. Vol [Dom04, Han02, KK02b, Roe03]. Volterra [MGT07]. Volume [CT04a, CSP02, CHNO06, Dim09, GII08, MO01]. Volumes [DN04]. Voronoi [DSC02, Luc08a, Luc09b, SEZ05]. Vortex [BBL00, FM07]. Vortices [BDL08, ER07, LB01]. Vote [dSFM02]. Voter [SMS09]. Vries [BKLO00, Gar01]. vs [BLP09, EK05b, NS02, Pre03]. Vulpiani [Opp09].

W [Fre08b, Opp03b]. Waals [BA06, FP04, Kaw03]. Waerden [Cag02]. Wagner [LM02, Vel00]. Waiting [Cha00, Mag09b, Zab06]. Walk [And05, AR06, CCM00, Cro09, DDH08, Fra07a, HvdHS08, Ken04, MCZ06, MvWH^{+}01, McD01, Pat07a, Raw03, SD08, Wei05, Yar08, Zab06]. Walker [DM08]. Walkers [DPP09, Ess03, FVE03, KG03, LK09]. Walks [BL04, BZ04, BZ09, BCFP05, BK09, BK02, CPK07, CCF^{+}00, CGJ^{+}05, Cau02a, Cau02b, Els00, FF04, FLW08, Ghe09, Gra09, GMNT05, HL08, Ken02, Koz00, MO09a, Mat08, MS08b, OEB01, SD06, Tel06a, YBMS06, Zab08]. Wall [AC01a, AC01b, AC04, BL09, BMD00, DDH08, DT00, DFF02, Fis06b, Fis08, HGST00, HI04, HI05, JS01b, Pia00, SA02, SWK09, SD08]. Wall-Induced [BMD00]. Walls [AB09, BFGM02, KG03, Sak06, Sak07]. War [MKL08]. Ward [BM04c, DRC04]. Wark [Opp02a]. Warnecke [Opp04a]. Water [SBG03, VKVT02]. Watermelon [OEB01]. Wave [CCL04, DKRS02, FK00, GLTZ06, HM00, KS09, Lem00, Noz04, SLB00, Tah09, Fre08a]. Wave-Like [FK00]. Wavefront [FM02a]. Waveguide [Naj07]. Waveguides [AR00]. Waves [CMS04, CC00a, CD05, Fan05, MM08, Naj07, WX06]. Wavevector [AYJP01b, AYP07b]. Wavevector-Dependent [AYJP01b, AYP07b]. Weak [Bao00, Blo01, BKM02, Cer05b, Cer06, DN04, DJ07, DGZ09, FJ05, KBS02, KR07, Lu06, MG02, Pro09b, Rot06, TAL06, Yar05, AC01b]. Weak-Coupling [AC01b]. Weakly [AE08, ALS06b, AMP04, BZ00b, BZ08b, CMS04, CS08b, DDH08, DELO05, ED04, Erd02, FW06, GS09, JR08, LS09b, MS08b, RB08, Spo06b, Spo06c, vWR05]. Weber [Tum06]. Wegner [Bou09]. Wehr [Ano01d]. Wei [Mas08]. Weierstrass [SM01]. Weighted [BZ09]. Weiss [AFFS06, CCIL08]. Well [Ace00, CGK^{+}04, DB01, Sac05]. Wells [DB00]. Wetting [BLM03, DMSR00, DDMSR04, GN08, PIRB03]. Which [Ram07, Rue02, You02]. White [Bao01, CL00, Fan05, FG00, GCU02, Val09a]. White-Noise [Fan05, Val09a]. Whitening [GOBY06]. Who [Ano00e]. Whole [Ghe09]. Whole-Plane [Ghe09]. Widom [Dag03b, Sod17, Iof02, Sod09, YS00]. Width [MM05c]. Widths [Kom08].

Wiener [Fuk08]. Wigner [Fan05, GT08, Han07, Luo07, LZ08, PS07b]. Wild [CS02, CL03]. Wilkens [Shl04]. William [Kad03b]. Winding [HL08, PM04]. Windings [FF04]. Winterbottom [BIV01]. Wires [KN03]. with/against [Bla08]. Without [Ale01b, Cer06, FPD01, Fou00, FM01c, Fou06, Kie04a, Kie04b, MRV02, Kad03b]. Witten [Lo08]. WKB [KS07]. Wolfang [Mas02a]. Work [PdKV09, San05, Sin08, TC08a]. Workshop [Leb09b]. World [Dys04, Han02, Hol02a, LR06, New00, Opp01, Opp02b, Opp04c, Opp04d, Rod02, Wit03]. Worlds [Ano01b]. Worm [Hos00]. Would [Rue02]. Wound [CKSZ06, KSSM07]. Wulff [DDH06]. WWW [Bog04].

xanthus [SBSAD07]. xiii [Rap01]. XORSAT [MRTZ03]. Xu [Mas08]. xv [Aba00]. XXZ [BA01, FM01a, FSZ01, KMSS01, Nep03]. XY [AP03]. Xylem [VKVT02].

Y. [Git09]. Yanase [Luo07, LZ08, Han07]. Yang [BG01a, Car03, GW09, Jaf03]. Yangian [BGX01]. Yasha [Rue02]. Years [Sin02]. Yeast [YCCN07]. York [Dag03a, Opp03b, Pod04b, Sah02]. Yorke [ECSB00]. Young [DMSR03, KG03]. Yukawa [CG00a, Jan01a].

Zernike [HK02b, KRT00]. Zero [BTM^{+}04, BK08, BTT07, BP01b, CN04b, DJ05, DG08, DP04, DP03, EGGI01, FHAY06, Gir01, GSS03, GS08c, GS08d, HN03, JMU05, JL00b, LM06, LMS05, LTWW02, MP03, Mor07, PdKV09, PM05, SSD00, SM08, SH07, Sot07, Uch04, Wu02, vER07]. Zero-Dimensional [JL00b]. Zero-Energy [DP04]. Zero-Entropy [DJ05]. Zero-Point [PdKV09]. Zero-Range [GS08c, GS08d, LMS05, SH07, PM05]. Zero-Range-Exclusion [Uch04]. Zero-Resistance [FHAY06]. Zero-Temperature [CN04b, HN03, Mor07, Sot07, Wu02]. Zeroes [JK01, LW01b]. Zeros [BG01a, BBCK04, JS01a, JSS03, JS06, KT09, Mac09, SS01, SS09b]. Zeta [SL02]. Zeta-functions [SL02]. Zhang [BTM^{+}04, KR06, SS09c, SMT03]. Zinn [Rub09a]. Zone [LR06]. Zwanzig [AVE07, Kup04].

References

Andries:2002:CBT

[AAP02] Pierre Andries, Kazuo Aoki, and Benoit Perthame. A consistent BGK-type model for gas mixtures. Journal of Statistical Physics, 106(5–6):993–1018, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

             /link.springer.com/article/10.1023/A%3A1014033703134.

Albano:2009:PCN

[AB09] Ezequiel V. Albano and Kurt Binder. Phase coexistence in nanoscopically thin films confined by asymmetric walls. Journal of Statistical Physics, 135(5–6):991–1008, June 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9710-8.

Abad:2000:BRB

[Aba00] Enrique Abad. Book review: Nonequilibrium Phase Transitions in Lattice Models J. Marro and R. Dickman, Cambridge University Press, Cambridge, 1999, xv + 327 pp. Journal of Statistical Physics, 100(3–4):797–800, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018687828373.

Alexander:2005:CPSa

[ABC05a] Kenneth S. Alexander, Marek Biskup, and Lincoln Chayes. Colligative properties of solutions: I. Fixed concentrations. Journal of Statistical Physics, 119(3–4):479–507, May 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3016-2.

Alexander:2005:CPSb

[ABC05b] Kenneth S. Alexander, Marek Biskup, and Lincoln Chayes. Colligative properties of solutions: II. Vanishing concentrations. Journal of Statistical Physics, 119(3–4):509–537, May 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3017-1.

Alastuey:2003:SCE

[ABCM03] A. Alastuey, V. Ballenegger, F. Cornu, and Ph. A. Martin. Screened cluster expansions for partially ionized gases. Journal of Statistical Physics, 113(3–4):455–503, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026064617421.

Alastuey:2008:ERT

[ABCM08] A. Alastuey, V. Ballenegger, F. Cornu, and Ph. A. Martin. Exact results for thermodynamics of the hydrogen plasma: Lowtemperature expansions beyond Saha theory. Journal of Statistical Physics, 130(6):1119–1176, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9464-0.

Ardonne:2001:NAQ

[ABS01] Eddy Ardonne, Peter Bouwknegt, and Kareljan Schoutens. Non-Abelian quantum Hall states-exclusion statistics, K-matrices, and duality. Journal of Statistical Physics, 102(3–4):421–469, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004878231034.

Aoki:2003:KLG

[ABT03] Kazuo Aoki, Claude Bardos, and Shigeru Takata. Knudsen layer for gas mixtures. Journal of Statistical Physics, 112(3– 4):629–655, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023876025363.

Albayrak:2007:CCT

[ABY07] E. Albayrak, T. Bulut, and S. Yilmaz. Critical and compensation temperatures of the Ising bilayer system consisting of spin-1/2 and spin-1 atoms. Journal of Statistical Physics, 127 (5):967–983, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9311-3.

Aqua:2001:DPCa

[AC01a] Jean-Noël Aqua and Françoise Cornu. Density profiles in a classical Coulomb fluid near a dielectric wall. I. Mean-field scheme. Journal of Statistical Physics, 105(1–2):211–243, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1012238211895.

Aqua:2001:DPCb

[AC01b] Jean-Noël Aqua and Françoise Cornu. Density profiles in a classical Coulomb fluid near a dielectric wall. II. Weak-coupling systematic expansions. Journal of Statistical Physics, 105(1–2): 245–283, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012290228733.

Altundas:2003:CDD

[AC03] Y. B. Altundas and G. Caginalp. Computations of dendrites in 3-D and comparison with microgravity experiments. Journal of Statistical Physics, 110(3–6):1055–1067, March 2003.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022140725763.

Aqua:2004:DPQ

[AC04] J.-N. Aqua and F. Cornu. Density profiles in a quantum Coulomb fluid near a hard wall. Journal of Statistical Physics, 115(3–4):997–1036, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000022372.82047.
6a.

Ay:2005:RHI

[AC05] Nihat Ay and James P. Crutchfield. Reductions of hidden information sources. Journal of Statistical Physics, 120(3–4):659–684, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-6797-4.

Acedo:2000:CCR

[Ace00] L. Acedo. Continuity conditions for the radial distribution function of square-well fluids. Journal of Statistical Physics, 99 (3–4):707–723, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018631109833.

Achdou:2008:VMK

[Ach08] Yves Achdou. V.A. Marchenko and E.Y. Khruslov: Homogenization of Partial Differential Equations. Journal of Statistical Physics, 133(4):797–799, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9633-9.

Ayyer:2009:PDA

[ACL^{+}09] A. Ayyer, E. A. Carlen, J. L. Lebowitz, P. K. Mohanty, D. Mukamel, and E. R. Speer. Phase diagram of the ABC model on an interval. Journal of Statistical Physics, 137(5–6): 1166–1204, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9834-x.

Aldous:2002:AOD

[AD02] David Aldous and Persi Diaconis. The asymmetric onedimensional constrained Ising model: Rigorous results. Journal of Statistical Physics, 107(5–6):945–975, June 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015170205728.

Adams:2001:CEG

[Ada01] Stefan Adams. Complete equivalence of the Gibbs ensembles for one-dimensional Markov systems. Journal of Statistical Physics, 105(5–6):879–908, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013505328774.

Abdallah:2000:CBE

[ADG00] N. Ben Abdallah, L. Desvillettes, and S. Génieys. On the convergence of the Boltzmann equation for semiconductors toward the energy transport model. Journal of Statistical Physics, 98(3– 4):835–870, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018635827617.

Adib:2004:DBP

[Adi04] Artur B. Adib. Does the Boltzmann principle need a dynamical correction? Journal of Statistical Physics, 117(3–4): 581–597, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-3454-2.

Asselah:2001:DEB

[ADLM01] A. Asselah, P. Dai Pra, J. L. Lebowitz, and Ph. Mounaix. Diffusion effects on the breakdown of a linear amplifier model driven by the square of a Gaussian field. Journal of Statistical Physics, 104(5–6):1299–1315, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010470231689.

Adami:2008:RDE

[AE08] Riccardo Adami and László Erdős. Rate of decoherence for an electron weakly coupled to a phonon gas. Journal of Statistical Physics, 132(2):301–328, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9561-8.

Avron:2004:TDQ

[AEGS04] J. E. Avron, A. Elgart, G. M. Graf, and L. Sadun. Transport and dissipation in quantum pumps. Journal of Statistical Physics, 116(1–4):425–473, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037245.45780.e1.

Alexander:2005:AMC

[AER05] Francis J. Alexander, Gregory L. Eyink, and Juan M. Restrepo. Accelerated Monte Carlo for optimal estimation of time series. Journal of Statistical Physics, 119(5–6):1331–1345, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-3770-1.

Albert:2006:MWE

[AFFS06] C. Albert, L. Ferrari, J. Fröhlich, and B. Schlein. Magnetism and the Weiss exchange field — a theoretical analysis motivated by recent experiments. Journal of Statistical Physics, 125 (1):77–124, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9120-0.

Alber:2007:I

[AFG07] Mark Alber, Erwin Frey, and Ray Goldstein. Introduction. Journal of Statistical Physics, 128(1–2):1–3, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9331-z.

Alicki:2003:CTD

[AFHV03] R. Alicki, M. Fannes, B. Haegeman, and D. Vanpeteghem. Coherent transport and dynamical entropy for fermionic systems. Journal of Statistical Physics, 113(3–4):549–574, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026020802400.

Abad:2000:ODL

[AFN00] Enrique Abad, Harry L. Frisch, and Gregoire Nicolis. Onedimensional lattice dynamics of the diffusion-limited reaction A + A → A + S: Transient behavior. Journal of Statistical Physics, 99(5–6):1397–1407, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018649108442.

Adler:2000:CSO

[AFNvM00] M. Adler, P. J. Forrester, T. Nagao, and P. van Moerbeke. Classical skew orthogonal polynomials and random matrices. Journal of Statistical Physics, 99(1–2):141–170, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018644606835.

Arratia:2005:SSF

[AG05] P. E. Arratia and J. P. Gollub. Statistics of stretching fields in experimental fluid flows exhibiting chaotic advection. Journal of Statistical Physics, 121(5–6):805–822, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8664-8.

Andrieux:2007:FTC

[AG07] David Andrieux and Pierre Gaspard. Fluctuation theorem for currents and Schnakenberg network theory. Journal of Statistical Physics, 127(1):107–131, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9233-5.

Alonso:2009:DCS

[AG09] Ricardo J. Alonso and Irene M. Gamba. Distributional and classical solutions to the Cauchy Boltzmann problem for soft potentials with integrable angular cross section. Journal of Statistical Physics, 137(5–6):1147–1165, December 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9873-3.

Aizenman:2001:BFT

[AGL01] M. Aizenman, S. Goldstein, and J. L. Lebowitz. Bounded fluctuations and translation symmetry breaking in one-dimensional particle systems. Journal of Statistical Physics, 103(3–4):601– 618, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010397401128.

Adami:2007:RDC

[AGT07] Riccardo Adami, François Golse, and Alessandro Teta. Rigorous derivation of the cubic NLS in dimension one. Journal of Statistical Physics, 127(6):1193–1220, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9271-z.

Albeverio:2002:REQ

[AGY02] Sergio Albeverio, Hanno Gottschalk, and Minoru W. Yoshida. Representing Euclidean quantum fields as scaling limit of particle systems. Journal of Statistical Physics, 108(1–2):361–369, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015456005923.

Anderson:2001:SFF

[AH01] Philip W. Anderson and F. Duncan M. Haldane. The symmetries of fermion fluids at low dimensions. Journal of Statistical Physics, 103(3–4):425–428, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010324912515.

Aldana:2003:PTS

[AH03] Maximino Aldana and Cristián Huepe. Phase transitions in selfdriven many-particle systems and related non-equilibrium models: A network approach. Journal of Statistical Physics, 112 (1–2):135–153, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023675519930.

Arponen:2007:DEK

[AH07] Heikki Arponen and Peter Horvai. Dynamo effect in the Kraichnan magnetohydrodynamic turbulence. Journal of Statistical Physics, 129(2):205–239, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9399-5.

Aharony:2003:ONR

[Aha03] Amnon Aharony. Old and new results on multicritical points. Journal of Statistical Physics, 110(3–6):659–669, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022103717585.

Ansumali:2002:EFA

[AK02] Santosh Ansumali and Iliya V. Karlin. Entropy function approach to the lattice Boltzmann method. Journal of Statistical Physics, 107(1–2):291–308, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014575024265.

Akemann:2007:ISG

[AK07] Gernot Akemann and Eugene Kanzieper. Integrable structure of Ginibre’s ensemble of real random matrices and a Pfaffian integration theorem. Journal of Statistical Physics, 129(5–6): 1159–1231, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9381-2.

Arguin:2009:SPS

[AK09] Louis-Pierre Arguin and Nicola Kistler. Small perturbations of a spin glass system. Journal of Statistical Physics, 135(5– 6):1167–1180, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9694-4.

Akimoto:2008:GAL

[Aki08] Takuma Akimoto. Generalized arcsine law and stable law in an infinite measure dynamical system. Journal of Statistical Physics, 132(1):171–186, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9544-9.

Ahn:2001:ARA

[AKR01] Changrim Ahn, Chanju Kim, and Chaiho Rim. Applications of reflection amplitudes in Toda-type theories. Journal of Statistical Physics, 102(3–4):385–419, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004826214195.

Aktekin:2001:FSS

[Akt01] N. Aktekin. The finite-size scaling functions of the fourdimensional Ising model. Journal of Statistical Physics, 104 (5–6):1397–1406, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010457905088.

Anh:2001:SAF

[AL01] V. V. Anh and N. N. Leonenko. Spectral analysis of fractional kinetic equations with random data. Journal of Statistical Physics, 104(5–6):1349–1387, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010474332598.

Aissani:2003:CBS

[AL03] A. Aissani and K. Lendi. Conditions for bounded solutions of non–Markovian quantum master equations. Journal of Statistical Physics, 111(5–6):1353–1362, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023064518885.

Arlotti:2007:IRL

[AL07] Luisa Arlotti and Bertrand Lods. Integral representation of the linear Boltzmann operator for granular gas dynamics with applications. Journal of Statistical Physics, 129(3):517–536, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9402-1.

Albayrak:2008:CCT

[Alb08] E. Albayrak. Critical and compensation temperatures of the Ising bilayer system consisting of spin-1/2 and spin-1 atoms. Journal of Statistical Physics, 130(4):829–830, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9330-0.

Alexander:2001:SGDa

[Ale01a] Kenneth S. Alexander. The spectral gap of the 2-D stochastic Ising model with nearly single-spin boundary conditions. Journal of Statistical Physics, 104(1–2):59–87, July 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010301525937.

Alexandre:2001:SSB

[Ale01b] Radjesvarane Alexandre. Some solutions of the Boltzmann equation without angular cutoff. Journal of Statistical Physics, 104 (1–2):327–358, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010317913642.

Allahverdyan:2008:EHM

[All08] Armen E. Allahverdyan. Entropy of hidden Markov processes via cycle expansion. Journal of Statistical Physics, 133(3): 535–564, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9613-0.

Abe:2007:HSD

[ALM07] Sumiyoshi Abe, Bernhard Lesche, and Jens Mund. How should the distance of probability assignments be judged? Journal of Statistical Physics, 128(5):1189–1196, September 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9344-7.

Albeverio:2007:SHF

[ALR07] Sergio Albeverio, Saidakhmat N. Lakaev, and Tulkin H. Rasulov. On the spectrum of an Hamiltonian in Fock space. Discrete spectrum asymptotics. Journal of Statistical Physics, 127 (2):191–220, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9240-6.

Anh:2003:HOS

[ALS03] V. V. Anh, N. N. Leonenko, and L. M. Sakhno. Higher-order spectral densities of fractional random fields. Journal of Statistical Physics, 111(3–4):789–814, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022898131682.

Anh:2006:SPB

[ALS06a] V. V. Anh, N. N. Leonenko, and L. M. Sakhno. Spectral properties of Burgers and KPZ turbulence. Journal of Statistical Physics, 122(5):949–974, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9009-3.

Aoki:2006:ETW

[ALS06b] Kenichiro Aoki, Jani Lukkarinen, and Herbert Spohn. Energy transport in weakly anharmonic chains. Journal of Statistical Physics, 124(5):1105–1129, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9171-2.

Ayyer:2009:TSA

[ALS09] Arvind Ayyer, Joel L. Lebowitz, and Eugene R. Speer. On the two species asymmetric exclusion process with semi-permeable boundaries. Journal of Statistical Physics, 135(5–6):1009–1037, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9724-2.

Albeverio:2004:SND

[AM04] S. Albeverio and S. Mazzucchi. Some new developments in the theory of path integrals, with applications to quantum theory. Journal of Statistical Physics, 115(1–2):191–215, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019836.37663.d9.

Ambegaokar:2006:DDQ

[Amb06] Vinay Ambegaokar. Dissipation and decoherence in a quantum oscillator. Journal of Statistical Physics, 125(5–6):1183– 1192, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8018-6.

Arbieto:2004:BPW

[AMP04] Alexander Arbieto, Carlos Matheus, and Maria José Pacifico. The Bernoulli property for weakly hyperbolic systems. Journal of Statistical Physics, 117(1–2):243–260, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044058.99450.c9.

Aumaitre:2006:ELF

[AMP06] Sébastien Aumaître, Kirone Mallick, and François Pétrélis. Effects of the low frequencies of noise on on–off intermittency. Journal of Statistical Physics, 123(4):909–927, May 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9116-9.

Aoki:2007:KRM

[AMT07] Kazuo Aoki, Peter Markowich, and Shigeru Takata. Kinetic relaxation models for energy transport. Journal of Statistical Physics, 127(2):287–312, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9273-x.

Amsalu:2008:TAL

[AMV08] Saba Amsalu, Heinrich Matzinger, and Marina Vachkovskaia. Thermodynamical approach to the longest common subsequence problem. Journal of Statistical Physics, 131(6):1103–1120, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9533-z.

Arkeryd:2000:MPH

[AN00] L. Arkeryd and A. Nouri. On the Milne problem and the hydrodynamic limit for a steady Boltzmann equation model. Journal of Statistical Physics, 99(3–4):993–1019, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018655815285.

Arkeryd:2005:SNB

[AN05] L. Arkeryd and A. Nouri. The stationary nonlinear Boltzmann equation in a Couette setting with multiple, isolated L^{q}-solutions and hydrodynamic limits. Journal of Statistical Physics, 118(5– 6):849–881, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2708-3.

Arkeryd:2006:TCT

[AN06] L. Arkeryd and A. Nouri. On a Taylor–Couette type bifurcation for the stationary nonlinear Boltzmann equation. Journal of Statistical Physics, 124(2–4):401–443, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8008-8.

Andreoletti:2005:APL

[And05] Pierre Andreoletti. Alternative proof for the localization of Sinai’s walk. Journal of Statistical Physics, 118(5–6):883–933, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2122-x.

Andreoletti:2008:LRS

[And08] Pierre Andreoletti. A limit result for a system of particles in random environment. Journal of Statistical Physics, 131 (2):235–246, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9497-z.

Anonymous:2000:AIJ

[Ano00a] Anonymous. Author index for journal of statistical physics (2000). Journal of Statistical Physics, 101(5–6):1137–1150, December 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018859608667.

Anonymous:2000:BRBc

[Ano00b] Anonymous. Book review: A Practical Introduction to the Simulation of Molecular Systems. Journal of Statistical Physics, 99 (3–4):1053–1055, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018668218011.

Anonymous:2000:BRBa

[Ano00c] Anonymous. Book review: Monte Carlo Methods in Statistical Physics. Journal of Statistical Physics, 98(1–2):503–505, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018651611707.

Anonymous:2000:BRBb

[Ano00d] Anonymous. Book review: Thermal Physics. Journal of Statistical Physics, 99(3–4):1051–1052, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018616201173.

Anonymous:2000:BRL

[Ano00e] Anonymous. Book review: Ludwig Boltzmann: The man who trusted atoms. Journal of Statistical Physics, 98(5–6): 1429–1432, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018640401191.

Anonymous:2000:BRM

[Ano00f] Anonymous. Book review: Molecular thermodynamics. Journal of Statistical Physics, 98(5–6):1425–1427, March 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018688317121.

Anonymous:2000:BRN

[Ano00g] Anonymous. Book review: Nonequilibrium phase transitions in lattice models. Journal of Statistical Physics, 98(5–6): 1417–1418, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018684116212.

Anonymous:2000:BRS

[Ano00h] Anonymous. Book review: Scaling limits of interacting particle systems. Journal of Statistical Physics, 98(5–6):1419–1424, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018636300282.

Anonymous:2000:CNB

[Ano00i] Anonymous. Call for nominations for the 2001 Boltzmann award. Journal of Statistical Physics, 99(3–4):1071, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018624402990.

Anonymous:2000:FCBa

[Ano00j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 98(1–2):511–512, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018655712616.

Anonymous:2000:FCBb

[Ano00k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 98(3–4):971–972, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018644528759.

Anonymous:2000:FCBc

[Ano00l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 98(5–6):1433–1434, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018692418029.

Anonymous:2000:FCBd

[Ano00m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 99(1–2):627–628, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018692023231.

Anonymous:2000:FCBe

[Ano00n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 99(3–4):1073–1074, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018676419828.

Anonymous:2000:FCBf

[Ano00o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 99(5–6):1429–1430, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018605326188.

Anonymous:2000:FCBg

[Ano00p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 100(1–2):489–490, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018772716913.

Anonymous:2000:FCBh

[Ano00q] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 100(3–4):803–804, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018715729885.

Anonymous:2000:FCBi

[Ano00r] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 100(5–6):1179–1182, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1018732317577.

Anonymous:2000:FCBj

[Ano00s] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 101(1–2):711, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1026677417714.

Anonymous:2000:FCBk

[Ano00t] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 101(3–4):933, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026679307759.

Anonymous:2000:FCBl

[Ano00u] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 101(5–6):1151–1152, December 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026673424597.

Anonymous:2000:Pa

[Ano00v] Anonymous. Preface. Journal of Statistical Physics, 100(1–2): 1–2, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018619123888.

Anonymous:2000:Pb

[Ano00w] Anonymous. Preface. Journal of Statistical Physics, 101(1– 2):1–2, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026467624528.

Anonymous:2000:PSMa

[Ano00x] Anonymous. Program of the 82nd Statistical Mechanics Meeting. Journal of Statistical Physics, 99(3–4):1057–1065, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018620302081.

Anonymous:2000:PSMb

[Ano00y] Anonymous. Program of the 83rd Statistical Mechanics Meeting. Journal of Statistical Physics, 100(5–6):1173–1178, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018731629283.

Anonymous:2000:PSS

[Ano00z] Anonymous. Program of the Sixth Statistical Physics Days. Journal of Statistical Physics, 99(3–4):1067–1070, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018672318920.

Anonymous:2001:AIJ

[Ano01a] Anonymous. Author index for journal of statistical physics (2001). Journal of Statistical Physics, 105(5–6):945–956, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017455530700.

Anonymous:2001:BRB

[Ano01b] Anonymous. Book review: Small Worlds. The Dynamics of Networks Between Order and Randomness. Journal of Statistical Physics, 102(1–2):369–370, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026585314580.

Anonymous:2001:BRS

[Ano01c] Anonymous. Book review: Stochastic and dynamic views of chemical reaction kinetics in solutions. Journal of Statistical Physics, 102(5–6):1447–1448, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1017385317289.

Anonymous:2001:ESL

[Ano01d] Anonymous. Erratum on ‘A Strong Law of Large Numbers for Iterated Functions of Independent Random Variables,’ Jan Wehr, J. Stat. Phys. 86:1373 (1997). Journal of Statistical Physics, 104(3–4):901, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017339609610. See [Weh97].

Anonymous:2001:FCBa

[Ano01e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 102(1–2):371–372, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026635609790.

Anonymous:2001:FCBb

[Ano01f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 102(3–4):1083–1084, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017365722166.

Anonymous:2001:FCBc

[Ano01g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 102(5–6):1449, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017352500611.

Anonymous:2001:FCBd

[Ano01h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 103(1–2):409–410, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017354107340.

Anonymous:2001:FCBe

[Ano01i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 103(3–4):645–646, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017356601230.

Anonymous:2001:FCBf

[Ano01j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 103(5–6):1155–1156, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017330102034.

Anonymous:2001:FCBg

[Ano01k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 104(1–2):487–488, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017347020370.

Anonymous:2001:FCBh

[Ano01l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 104(3–4):903–904, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017391601031.

Anonymous:2001:FCBi

[Ano01m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 104(5–6):1409–1410, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1017461013531.

Anonymous:2001:FCBj

[Ano01n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 105(1–2):411–412, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017354715529.

Anonymous:2001:FCBk

[Ano01o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 105(3–4):721, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017418322579.

Anonymous:2001:FCBl

[Ano01p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 105(5–6):957–958, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017452814770.

Anonymous:2001:LPL

[Ano01q] Anonymous. Luttinger’s publications list. Journal of Statistical Physics, 103(3–4):641–643, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017372818875.

Anonymous:2001:O

[Ano01r] Anonymous. Obituary. Journal of Statistical Physics, 103(3– 4):413–415, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017320711606.

Anonymous:2001:Pa

[Ano01s] Anonymous. Preface. Journal of Statistical Physics, 102(3–4): 373–374, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017317713287.

Anonymous:2001:Pb

[Ano01t] Anonymous. Preface. Journal of Statistical Physics, 103(3– 4):411–412, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017325827536.

Anonymous:2001:PSMa

[Ano01u] Anonymous. Program of the 84th Statistical Mechanics Meeting. Journal of Statistical Physics, 103(5–6):1147–1153, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017310126922.

Anonymous:2001:PSMb

[Ano01v] Anonymous. Program of the 85th Statistical Mechanics Meeting. Journal of Statistical Physics, 105(1–2):405–410, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017321831459.

Anonymous:2002:AIJ

[Ano02a] Anonymous. Author index for journal of statistical physics (2002). Journal of Statistical Physics, 109(5–6):1125–1138, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020442926264.

Anonymous:2002:FCBa

[Ano02b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 106(1–2):405–406, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017461514114.

Anonymous:2002:FCBb

[Ano02c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 106(3–4):855–856, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017413028207.

Anonymous:2002:FCBc

[Ano02d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 106(5–6):1255–1257, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017421209689.

Anonymous:2002:FCBd

[Ano02e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 107(1–2):597–598, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017439802190.

Anonymous:2002:FCBe

[Ano02f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 107(3–4):943–944, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017481004763.

Anonymous:2002:FCBf

[Ano02g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 107(5–6):1305, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017408028927.

Anonymous:2002:FCBg

[Ano02h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 108(1–2):371–372, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017470302349.

Anonymous:2002:FCBh

[Ano02i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 108(3–4):719–720, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017465022900.

Anonymous:2002:FCBi

[Ano02j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 108(5–6):1303–1304, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019712104553.

Anonymous:2002:FCBj

[Ano02k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 109(1–2):351–352, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1019998719962.

Anonymous:2002:FCBk

[Ano02l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 109(3–4):921–922, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1020471409426.

Anonymous:2002:FCBl

[Ano02m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 109(5–6):1139–1140, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020494910335.

Anonymous:2002:PSM

[Ano02n] Anonymous. Program of the 87th Statistical Mechanics Meeting Celebrating the 85th Birthday of Howard Reiss. Journal of Statistical Physics, 109(1–2):345–349, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019965204550.

Anonymous:2003:AIJ

[Ano03a] Anonymous. Author index for journal of statistical physics (2003). Journal of Statistical Physics, 113(5–6):889–900, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027324913870.

Anonymous:2003:FCBa

[Ano03b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 110(1–2):461–464, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021058212231.

Anonymous:2003:FCBb

[Ano03c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 110(3–6):1523–1524, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1022162217780.

Anonymous:2003:FCBc

[Ano03d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 111(1–2):503–504, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022209626946.

Anonymous:2003:FCBd

[Ano03e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 111(3–4):1025–1026, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022861508794.

Anonymous:2003:FCBe

[Ano03f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 111(5–6):1405–1406, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023042831547.

Anonymous:2003:FCBf

[Ano03g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 112(1–2):435–436, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023656410361.

Anonymous:2003:FCBg

[Ano03h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 112(3–4):887–888, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023844513068.

Anonymous:2003:FCBh

[Ano03i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 112(5–6):1221–1222, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1024612924881.

Anonymous:2003:FCBi

[Ano03j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 113(1–2):387, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1025791313728.

Anonymous:2003:FCBj

[Ano03k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 113(3–4):635–636, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1026081121056.

Anonymous:2003:FCBk

[Ano03l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 113(5–6):901–902, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027373024634.

Anonymous:2003:PSMa

[Ano03m] Anonymous. Program of the 86th Statistical Mechanics Meeting Celebrating the 70th Birthday of Michael E. Fisher. Journal of Statistical Physics, 110(3–6):1515–1522, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022125717102.

Anonymous:2003:PSMb

[Ano03n] Anonymous. Program of the 88th Statistical Mechanics Meeting Celebrating the 70th Birthday of Elliott Lieb. Journal of Statistical Physics, 111(5–6):1397–1403, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023072820702.

Anonymous:2003:PSMc

[Ano03o] Anonymous. Program of the 89th Statistical Mechanics Meeting Celebrating the 80th Birthday of E. G. D. Cohen. Journal of Statistical Physics, 113(1–2):381–386, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025791100815.

Anonymous:2004:AIJ

[Ano04a] Anonymous. Author index for journal of statistical physics (2004). Journal of Statistical Physics, 117(5–6):1073–1087, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5732-4.

Anonymous:2004:FCBa

[Ano04b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 114(1–2):535–536, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003355.16175.cd.

Anonymous:2004:FCBb

[Ano04c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 114(3–4):1181–1182, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012707.94583.6c.

Anonymous:2004:FCBc

[Ano04d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 114(5–6):1629–1631, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000014282.38647.39.

Anonymous:2004:FCBd

[Ano04e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 115(1–2):715–716, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000020056.35273.31.

Anonymous:2004:FCBe

[Ano04f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 115(3–4):1147–1148, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022594.27245.b6.

Anonymous:2004:FCBf

[Ano04g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 115(5–6):1769–1772, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028249.92254.5e.

Anonymous:2004:FCBg

[Ano04h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 116(1–4):1199–1200, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037395.29242.20.

Anonymous:2004:FCBh

[Ano04i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 116(5–6):1707–1708, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/B%3AJOSS.0000041913.24232.7f.

Anonymous:2004:FCBi

[Ano04j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 117(1–2):385–386, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000044206.36476.34.

Anonymous:2004:FCBj

[Ano04k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 117(3–4):783–784, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-004-9165-x.

Anonymous:2004:FCBk

[Ano04l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 117(5–6):1089–1090, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5733-3.

Anonymous:2004:PSM

[Ano04m] Anonymous. Program of the 90th Statistical Mechanics Meeting Celebrating the 80th Birthday of Freeman Dyson and the 100th Anniversary of the Birth of Lars Onsager. Journal of Statistical Physics, 115(3–4):1139–1145, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022576.38044.55.

Anonymous:2005:BA

[Ano05a] Anonymous. Boltzmann award (2004). Journal of Statistical Physics, 118(1–2):371–372, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2511-1.

Anonymous:2005:FCBa

[Ano05b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 118(1–2):369–370, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2510-2.

Anonymous:2005:FCBb

[Ano05c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 118(3–4):793–794, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-2513-7.

Anonymous:2005:FCBc

[Ano05d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 118(5–6):1273–1274, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-3484-9.

Anonymous:2005:FCBd

[Ano05e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 119(1–2):457–458, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-4962-4.

Anonymous:2005:FCBe

[Ano05f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 119(3–4):951–952, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-6322-9.

Anonymous:2005:FCBf

[Ano05g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 119(5–6):1423–1424, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-7343-0.

Anonymous:2006:P

[Ano06] Anonymous. Preface. Journal of Statistical Physics, 124(2– 4):269, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9119-6.

Anonymous:2009:Pa

[Ano09a] Anonymous. Preface. Journal of Statistical Physics, 134 (5–6):807, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9735-z.pdf.

Anonymous:2009:Pb

[Ano09b] Anonymous. Preface. Journal of Statistical Physics, 135 (5–6):787, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9763-8.pdf.

Alves:2006:CSE

[AOT06] José F. Alves, Krerley Oliveira, and Ali Tahzibi. On the continuity of the SRB entropy for endomorphisms. Journal of Statistical Physics, 123(4):763–785, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9059-1.

Aschbacher:2003:NES

[AP03] Walter H. Aschbacher and Claude-Alain Pillet. Non-equilibrium steady states of the XY chain. Journal of Statistical Physics, 112(5–6):1153–1175, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1024619726273.

Araujo:2006:LDN

[AP06] V. Araújo and M. J. Pacifico. Large deviations for non-uniformly expanding maps. Journal of Statistical Physics, 125(2):411–453, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9183-y.

Alves:2008:SRM

[AP08] José F. Alves and Vilton Pinheiro. Slow rates of mixing for dynamical systems with hyperbolic structures. Journal of Statistical Physics, 131(3):505–534, May 2008. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9482-6.

Adachi:2004:CAU

[APL04] Susumu Adachi, Ferdinand Peper, and Jia Lee. Computation by asynchronously updating cellular automata. Journal of Statistical Physics, 114(1–2):261–289, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003112.54283.ac.

Abdesselam:2009:CBP

[APS09] Abdelmalek Abdesselam, Aldo Procacci, and Benedetto Scoppola. Clustering bounds on n-point correlations for unbounded spin systems. Journal of Statistical Physics, 136(3):405–452, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9789-y.

Akguc:2000:CSP

[AR00] Gursoy B. Akguc and L. E. Reichl. Conductance and statistical properties of chaotic and integrable electron waveguides. Journal of Statistical Physics, 98(3–4):813–834, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018683710779.

Arndt:2002:SBT

[AR02] Peter F. Arndt and Vladimir Rittenberg. Spontaneous breaking of translational invariance and spatial condensation in stationary states on a ring. II. The charged system and the twocomponent Burgers equations. Journal of Statistical Physics, 107 (5–6):989–1013, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015101506637.

Antal:2006:EUR

[AR06] T. Antal and S. Redner. Escape of a uniform random walk from an interval. Journal of Statistical Physics, 123(6):1129–1144, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9139-2.

Aristoff:2009:RLP

[AR09] David Aristoff and Charles Radin. Random loose packing in granular matter. Journal of Statistical Physics, 135(1):1–23, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9722-4.

Arguin:2002:HFK

[Arg02] Louis-Pierre Arguin. Homology of Fortuin–Kasteleyn clusters of Potts models on the torus. Journal of Statistical Physics, 109(1– 2):301–310, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019979326380.

Arguin:2007:SGC

[Arg07] Louis-Pierre Arguin. Spin glass computations and Ruelle’s probability cascades. Journal of Statistical Physics, 126(4–5): 951–976, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9207-7.

Arriola:2001:VAS

[AS01] Enrique Ruíz Arriola and Juan Soler. A variational approach to the Schrödinger–Poisson system: Asymptotic behaviour, breathers, and stability. Journal of Statistical Physics, 103(5– 6):1069–1105, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010369224196.

Aouidi:2002:MFF

[AS02] Jamil Aouidi and Mourad Ben Slimane. Multi-fractal formalism for quasi-self-similar functions. Journal of Statistical Physics, 108(3–4):541–590, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1015729908167.

Antoniou:2003:ASE

[AS03] I. Antoniou and S. A. Shkarin. Analyticity of smooth eigenfunctions and spectral analysis of the Gauss map. Journal of Statistical Physics, 111(1–2):355–369, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022217410549.

Amann:2005:BSI

[AS05] A. Amann and E. Schöll. Bifurcations in a system of interacting fronts. Journal of Statistical Physics, 119(5–6):1069–1138, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-4405-2.

Albers:2006:PLB

[AS06] D. J. Albers and J. C. Sprott. Probability of local bifurcation type from a fixed point: A random matrix perspective. Journal of Statistical Physics, 125(4):885–921, November 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9232-6.

Alperovich:2008:SDT

[AS08] Timur Alperovich and Alexandros Sopasakis. Stochastic description of traffic flow. Journal of Statistical Physics, 133(6): 1083–1105, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9652-6.

Asada:2000:NSA

[Asa00] Hiromu Asada. A new statistical aspect of the cluster variation method for lattice systems. Journal of Statistical Physics, 98(3– 4):621–637, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018615223074.

Acedo:2002:DHV

[ASB02] L. Acedo, A. Santos, and A. V. Bobylev. On the derivation of a high-velocity tail from the Boltzmann–Fokker–Planck equation for shear flow. Journal of Statistical Physics, 109(5–6): 1027–1050, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020424610273.

Astrauskas:2008:ETS

[Ast08] A. Astrauskas. Extremal theory for spectrum of random discrete Schrödinger operator. I. Asymptotic expansion formulas. Journal of Statistical Physics, 131(5):867–916, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9519-x.

Alam:2006:HTR

[ATH06] Meheboob Alam, L. Trujillo, and H. J. Herrmann. Hydrodynamic theory for reverse Brazil nut segregation and the nonmonotonic ascension dynamics. Journal of Statistical Physics, 124(2–4):587–623, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9078-y.

Attard:2000:EDF

[Att00] Phil Attard. The explicit density functional and its connection with entropy maximization. Journal of Statistical Physics, 100 (1–2):445–473, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018668502023.

Ausloos:2000:ACA

[Aus00] Marcel Ausloos. (anti)coherence and (anti)persistence in natural and mathematical time series. Journal of Statistical Physics, 101 (1–2):707, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026486710705.

Andjel:1987:HEA

[AV87] Enrique Daniel Andjel and Maria Eulália Vares. Hydrodynamic equations for attractive particle systems on Z. Journal of Statistical Physics, 47(1–2):265–288, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009046. See cor-

rection [AV03].

Andjel:2003:CHE

[AV03] Enrique Daniel Andjel and Maria Eulália Vares. Correction to ‘Hydrodynamic Equations for Attractive Particle Systems on Z,’ J. Stat. Phys. 47:265 (1987). Journal of Statistical Physics, 113(1–2):379–380, October 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025739208508. See [AV87].

Antunovic:2008:SPT

[AV08] Ton’ci Antunović and Ivan Veselić. Sharpness of the phase transition and exponential decay of the subcritical cluster size for percolation on quasi-transitive graphs. Journal of Statistical Physics, 130(5):983–1009, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9459-x.

Azaele:2006:LOS

[AVBM06] Sandro Azaele, Igor Volkov, Jayanth R. Banavar, and Amos Maritan. Linear one-step processes with artificial boundaries. Journal of Statistical Physics, 125(2):491–511, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9158-z.

Ariel:2007:TTS

[AVE07] G. Ariel and E. Vanden-Eijnden. Testing transition state theory on Kac–Zwanzig model. Journal of Statistical Physics, 126 (1):43–73, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9165-0.

Angelescu:2000:QVA

[AVZ00] N. Angelescu, A. Verbeure, and V. A. Zagrebnov. Quantum nvector anharmonic crystal II: Displacement fluctuations. Journal of Statistical Physics, 100(5–6):829–851, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018715125648.

Alexander:2001:SGDb

[AY01] Kenneth S. Alexander and Nobuo Yoshida. The spectral gap of the 2-D stochastic Ising model with mixed boundary conditions. Journal of Statistical Physics, 104(1–2):89–109, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010382316368.

Au-Yang:2001:BSF

[AYJP01a] Helen Au-Yang, Bai-Qi Jin, and Jacques H. H. Perk. Baxter’s solution for the free energy of the chiral Potts model. Journal of Statistical Physics, 102(3–4):471–499, February 2001.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004830415104.

Au-Yang:2001:WDS

[AYJP01b] Helen Au-Yang, Bai-Qi Jin, and Jacques H. H. Perk. Wavevector-dependent susceptibility in quasiperiodic Ising models. Journal of Statistical Physics, 102(3–4):501–543, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004882431942.

Yong:2005:NTS

[aYL05] Wen an Yong and Li-Shi Luo. Nonexistence of H theorem for some lattice Boltzmann models. Journal of Statistical Physics, 121(1–2):91–103, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-5958-9.

Au-Yang:2007:DSF

[AYP07a] Helen Au-Yang and Jacques H. H. Perk. Q-dependent susceptibilities in ferromagnetic quasiperiodic Z-invariant Ising models. Journal of Statistical Physics, 127(2):265–286, April 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9213-9.

Au-Yang:2007:WDS

[AYP07b] Helen Au-Yang and Jacques H. H. Perk. Wavevector-dependent susceptibility in Z-invariant pentagrid Ising model. Journal of Statistical Physics, 127(2):221–264, April 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9212-x.

Braak:2001:SXC

[BA01] Daniel Braak and Natan Andrei. On the spectrum of the XXZ– Chain at roots of unity. Journal of Statistical Physics, 105(3–4): 677–709, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012236111393.

ben-Avraham:2003:BRB

[bA03] Daniel ben Avraham. Book review: Chaotic Transitions in Deterministic and Stochastic Dynamical Systems. Emil Simiu, 224 pp., Princeton University Press, 2002. Journal of Statistical Physics, 110(1–2):459–460, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021047501904.

Ben-Avraham:2006:BWF

[BA06] Daniel Ben-Avraham. Van der Waals Forces: A Handbook for Biologists, Chemists, Engineers, and Physicists. Journal of Statistical Physics, 123(3):709–710, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9137-4.

ben-Avraham:2008:MFB

[bA08] Daniel ben Avraham. Michael F. Barnsley: Superfractals. Journal of Statistical Physics, 130(1):203–204, January 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9441-7.

ben-Avraham:2009:BMB

[bA09] Daniel ben Avraham. A. Barrát, M. Barthélemy, A. Vespignani: Dynamical Processes on Complex Networks. Journal of Statistical Physics, 135(4):773–774, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9761-x.

Bahadoran:2007:CES

[Bah07] C. Bahadoran. On the convergence of entropy for stationary exclusion processes with open boundaries. Journal of Statistical Physics, 126(4–5):1069–1082, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9229-1.

Baker:2003:CAP

[Bak03] George A. Baker, Jr. Can an attractive potential increase the pressure of an ideal electron gas? Journal of Statistical Physics, 110(3–6):971–979, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022132523945.

Bakhtin:2006:EUS

[Bak06] Yuri Bakhtin. Existence and uniqueness of stationary solutions for 3D Navier–Stokes system with small random forcing via stochastic cascades. Journal of Statistical Physics, 122(2): 351–360, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8014-x.

Balescu:2000:KTA

[Bal00] R. Balescu. Kinetic theory of area-preserving maps. Application to the standard map in the diffusive regime. Journal of Statistical Physics, 98(5–6):1169–1234, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018619829416.

Balazs:2001:MSS

[Bal01] Márton Balázs. Microscopic shape of shocks in a domain growth model. Journal of Statistical Physics, 105(3–4):511–524, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012271624597.

Balazs:2004:MSB

[Bal04] Márton Balázs. Multiple shocks in bricklayers’ model. Journal of Statistical Physics, 117(1–2):77–98, October 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044060.25344.58.

Baldassi:2009:GLP

[Bal09] Carlo Baldassi. Generalization learning in a perceptron with binary synapses. Journal of Statistical Physics, 136(5):902– 916, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9822-1.

Bandt:2005:DEM

[Ban05] Christoph Bandt. The discrete evolution model of Bak and Sneppen is conjugate to the classical contact process. Journal of Statistical Physics, 120(3–4):685–693, August 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5965-x.

Bao:2000:SIS

[Bao00] Jing-Dong Bao. Semi-integral scheme for simulation of Langevin equation with weak inertia. Journal of Statistical Physics, 99 (1–2):595–602, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018665211378.

Bao:2001:AST

[Bao01] Jing-Dong Bao. Alternative simulating technique for periodic motion in the presence of multiplicative white noise. Journal of Statistical Physics, 102(1–2):361–368, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026533330510.

Bao:2004:NIN

[Bao04] Jing-Dong Bao. Numerical integration of a non–Markovian Langevin equation with a thermal band-passing noise. Journal of Statistical Physics, 114(1–2):503–513, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003118.62044.b7.

Barkai:2004:SEB

[Bar04] Eli Barkai. Stable equilibrium based on Lévy statistics:a linear Boltzmann equation approach. Journal of Statistical Physics, 115(5–6):1537–1565, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000028068.93241.
3a.

Barkai:2006:RTS

[Bar06a] E. Barkai. Residence time statistics for normal and fractional diffusion in a force field. Journal of Statistical Physics, 123 (4):883–907, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9109-8.

Barra:2006:IFE

[Bar06b] Adriano Barra. Irreducible free energy expansion and overlaps locking in mean field spin glasses. Journal of Statistical Physics, 123(3):601–614, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-9006-6.

Barra:2008:MFI

[Bar08] Adriano Barra. The mean field Ising model trough interpolating techniques. Journal of Statistical Physics, 132(5):787– 809, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9567-2.

Baxter:2000:ETR

[Bax00] R. J. Baxter. Equivalence of the two results for the free energy of the chiral Potts model. Journal of Statistical Physics, 98(3– 4):513–535, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018621421257.

Baxter:2002:CBA

[Bax02] R. J. Baxter. Completeness of the Bethe ansatz for the six and eight-vertex models. Journal of Statistical Physics, 108(1–2):1– 48, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015437118218.

Baxter:2003:RSC

[Bax03] R. J. Baxter. The Riemann surface of the chiral Potts model free energy function. Journal of Statistical Physics, 112(1–2):1–26, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023611702183.

Baxter:2004:SEV

[Bax04a] R. J. Baxter. The six and eight-vertex models revisited. Journal of Statistical Physics, 116(1–4):43–66, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037215.07702.93.

Baxter:2004:TMF

[Bax04b] R. J. Baxter. Transfer matrix functional relations for the generalized τ_{2}(t_{q}) model. Journal of Statistical Physics, 117(1– 2):1–25, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000044062.64287.b9. See

addendum [Bax06].

Baxter:2005:OPC

[Bax05] R. J. Baxter. The order parameter of the chiral Potts model. Journal of Statistical Physics, 120(1–2):1–36, July 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5534-3.

Baxter:2006:ATM

[Bax06] R. J. Baxter. Addendum to: ‘Transfer Matrix Functional Relations for the Generalized τ_{2}(t_{q}) Model, Journal of Statistical Physics 117, 1–25 (2004)’. Journal of Statistical Physics, 123(3): 705, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9106-y. See [Bax04b].

Baxter:2008:ARI

[Bax08a] R. J. Baxter. Algebraic reduction of the Ising model. Journal of Statistical Physics, 132(6):959–982, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9587-y.

Baxter:2008:CSC

[Bax08b] R. J. Baxter. A conjecture for the superintegrable chiral Potts model. Journal of Statistical Physics, 132(6):983–1000, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9588-x.

Baxter:2009:SRG

[Bax09] R. J. Baxter. Some remarks on a generalization of the superintegrable chiral Potts model. Journal of Statistical Physics, 137(5– 6):798–813, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9778-1.

Bauer:2003:SAE

[BB03] Michel Bauer and Denis Bernard. A simple asymmetric evolving random network. Journal of Statistical Physics, 111(3–4):703– 737, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022842013935.

Bouchet:2005:CPT

[BB05] F. Bouchet and J. Barré. Classification of phase transitions and ensemble inequivalence, in systems with long range interactions. Journal of Statistical Physics, 118(5–6):1073–1105, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-2059-0.

Benzi:2009:FDT

[BB09] Roberto Benzi and Luca Biferale. Fully developed turbulence and the multifractal conjecture. Journal of Statistical Physics, 135(5–6):977–990, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9738-9.

Barbato:2006:SRR

[BBBF06] D. Barbato, M. Barsanti, H. Bessaih, and F. Flandoli. Some rigorous results on a stochastic GOY model. Journal of Statistical Physics, 125(3):677–716, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9203-y.

Boldrighini:2001:ITO

[BBC^{+}01] C. Boldrighini, L. A. Bunimovich, G. Cosimi, S. Frigio, and A. Pellegrinotti. Ising-type and other transitions in onedimensional coupled map lattices with sign symmetry. Journal of Statistical Physics, 102(5–6):1271–1283, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004892312745.

Biskup:2004:PFZ

[BBCK04] M. Biskup, C. Borgs, J. T. Chayes, and R. Kotecký. Partition function zeros at first-order phase transitions: Pirogov– Sinai theory. Journal of Statistical Physics, 116(1–4):97–155, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037243.48527.e3.

Bertin:2004:PTN

[BBD04] Etienne Bertin, Jean-Michel Billiot, and Rémy Drouilhet. Phase transition in the nearest-neighbor continuum Potts model. Journal of Statistical Physics, 114(1–2):79–100, January 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003056.84984.8b.

Bertin:2008:LDI

[BBD08] Etienne Bertin, Jean-Michel Billiot, and Rémy Drouilhet. Rlocal Delaunay inhibition model. Journal of Statistical Physics, 132(4):649–667, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9565-4.

Barre:2005:LDT

[BBDR05] Julien Barré, Freddy Bouchet, Thierry Dauxois, and Stefano Ruffo. Large deviation techniques applied to systems with longrange interactions. Journal of Statistical Physics, 119(3–4):677– 713, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-3768-8.

Barrera:2009:ACE

[BBF09] J. Barrera, O. Bertoncini, and R. Fernández. Abrupt convergence and escape behavior for birth and death chains. Journal of Statistical Physics, 137(4):595–623, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9861-7.pdf.

Bauer:2005:MSL

[BBK05] Michel Bauer, Denis Bernard, and Kalle Kytölä. Multiple Schramm–Loewner evolutions and statistical mechanics martingales. Journal of Statistical Physics, 120(5–6):1125–1163, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-005-7002-5.

Bauer:2008:LEC

[BBK08] Michel Bauer, Denis Bernard, and Kalle Kytölä. LERW as an example of off-critical SLEs. Journal of Statistical Physics, 132 (4):721–754, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9569-0.

Bene:2000:SFV

[BBL00] J. Bene, S. Bröcheler, and H. Lustfeld. Simulating 2D flows with viscous vortex dynamics. Journal of Statistical Physics, 101(1– 2):567–577, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026483521802.

Barre:2005:CMR

[BBLS05] J. Barré, A. R. Bishop, T. Lookman, and A. Saxena. The cavity method for the rigidity transition. Journal of Statistical Physics, 118(5–6):1057–1071, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2709-2.

Boldrighini:2002:SED

[BBP02] C. Boldrighini, L. A. Bunimovich, and A. Pellegrinotti. Some estimates for 2-dimensional infinite and bounded dilute random Lorentz gases. Journal of Statistical Physics, 109(3–4): 729–745, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020470800432.

Brown:2003:SDH

[BBS03] Becky A. Brown, A. Ray Brown, and Michael F. Shlesinger. Solutions of doubly and higher order iterated equations. Journal of Statistical Physics, 110(3–6):1087–1097, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022144826671.

Benzi:2004:GLM

[BBS04] Roberto Benzi, Luca Biferale, and Mauro Sbragaglia. A Gibbs-like measure for single-time, multi-scale energy transfer in stochastic signals and Shell model of turbulence. Journal of Statistical Physics, 114(1–2):137–154, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003107.88526.bf.

Benzi:2003:ITM

[BBT03] Roberto Benzi, Luca Biferale, and Federico Toschi. Intermittency in turbulence: Multiplicative random process in space and time. Journal of Statistical Physics, 113(5–6):783–798, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027360619182.

Bouabci:2000:RCR

[BC00] M. B. Bouabci and C. E. I. Carneiro. Random-cluster representation for the Blume–Capel model. Journal of Statistical Physics, 100(5–6):805–827, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018723327466.

Bershadskii:2001:CMF

[BC01] A. Bershadskii and Emily S. C. Ching. Classification of multiscaling in fracture and fragmentation. Journal of Statistical Physics, 104(1–2):49–57, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010349409099.

Bobylev:2002:EES

[BC02a] A. V. Bobylev and C. Cercignani. Exact eternal solutions of the Boltzmann equation. Journal of Statistical Physics, 106(5– 6):1019–1038, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014085719973.

Bobylev:2002:MEG

[BC02b] A. V. Bobylev and C. Cercignani. Moment equations for a granular material in a thermal bath. Journal of Statistical Physics, 106(3–4):547–567, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013754205008.

Bobylev:2002:SSSa

[BC02c] A. V. Bobylev and C. Cercignani. Self-similar solutions of the Boltzmann equation and their applications. Journal of Statistical Physics, 106(5–6):1039–1071, March 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014037804043.

Bobylev:2002:SSSb

[BC02d] A. V. Bobylev and C. Cercignani. Self-similar solutions of the Boltzmann equation for non–Maxwell molecules. Journal of Statistical Physics, 108(3–4):713–717, August 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015790226822.

Butera:2002:LLE

[BC02e] P. Butera and M. Comi. An on-line library of extended hightemperature expansions of basic observables for the spin-S Ising models on two- and three-dimensional lattices. Journal of Statistical Physics, 109(1–2):311–315, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019995830014.

Bobylev:2003:SSA

[BC03] A. V. Bobylev and C. Cercignani. Self-similar asymptotics for the Boltzmann equation with inelastic and elastic interactions. Journal of Statistical Physics, 110(1–2):333–375, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021031031038.

Bednorz:2004:GTH

[BC04] Adam Bednorz and Bogdan Cichocki. General H-theorem for hard spheres. Journal of Statistical Physics, 114(1–2):327–360, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003114.38280.e4.

Bosi:2009:BME

[BC09] Roberta Bosi and Maria J. Cáceres. The BGK model with external confining potential: Existence, long-time behaviour and time-periodic Maxwellian equilibria. Journal of Statistical Physics, 136(2):297–330, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9782-5.

Biskup:2006:MFD

[BCC06] Marek Biskup, Lincoln Chayes, and Nicholas Crawford. Meanfield driven first-order phase transitions in systems with longrange interactions. Journal of Statistical Physics, 122(6): 1139–1193, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8072-0.

Bhattacharya:2009:INI

[BCC09] Satyabrata Bhattacharya, Sudip Chattopadhyay, and Jyotipratim Ray Chaudhuri. Investigation of noise-induced escape rate: A quantum mechanical approach. Journal of Statistical Physics, 136(4):733–750, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9802-5.

Berkenbusch:2004:DCT

[BCD^{+}04] Marko Kleine Berkenbusch, Isabelle Claus, Catherine Dunn, Leo P. Kadanoff, Maciej Nicewicz, and Shankar C. Venkataramani. Discrete charges on a two dimensional conductor. Journal of Statistical Physics, 116(5–6):1301–1358, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041741.27244.ac.

Butta:2007:PPA

[BCDM07] Paolo Buttà, Emanuele Caglioti, Sara Di Ruzza, and Carlo Marchioro. On the propagation of a perturbation in an anharmonic system. Journal of Statistical Physics, 127(2):313–325, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9278-0.

Benedetto:2004:SCD

[BCEP04] D. Benedetto, F. Castella, R. Esposito, and M. Pulvirenti. Some considerations on the derivation of the nonlinear quantum Boltzmann equation. Journal of Statistical Physics, 116(1–4):381–410, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037205.09518.3f.

Benedetto:2006:SCD

[BCEP06] D. Benedetto, F. Castella, R. Esposito, and M. Pulvirenti. Some considerations on the derivation of the nonlinear quantum Boltzmann equation II: The low density regime. Journal of Statistical Physics, 124(2–4):951–996, August 2006. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-9010-x.

Banavar:2001:SOL

[BCF^{+}01] Jayanth R. Banavar, Francesca Colaiori, Alessandro Flammini, Amos Maritan, and Andrea Rinaldo. Scaling, optimality, and landscape evolution. Journal of Statistical Physics, 104(1–2):1– 48, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010397325029.

Belletti:2009:DVM

[BCF^{+}09] F. Belletti, A. Cruz, L. A. Fernandez, A. Gordillo-Guerrero, M. Guidetti, A. Maiorano, F. Mantovani, E. Marinari, V. Martin-Mayor, J. Monforte, A. Muñoz Sudupe, D. Navarro, G. Parisi, S. Perez-Gaviro, J. J. Ruiz-Lorenzo, and et al. An in-depth view of the microscopic dynamics of Ising spin glasses at fixed temperature. Journal of Statistical Physics, 135(5– 6):1121–1158, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9727-z.

Boldrighini:2005:CSS

[BCFP05] C. Boldrighini, G. Cosimi, S. Frigio, and A. Pellegrinotti. Computer simulations for some one-dimensional models of random walks in fluctuating random environment. Journal of Statistical Physics, 121(3–4):361–372, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-7012-3.

Bobylev:2000:SPK

[BCG00] A. V. Bobylev, J. A. Carrillo, and I. M. Gamba. On some properties of kinetic and hydrodynamic equations for inelastic interactions. Journal of Statistical Physics, 98(3–4):743–773, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018627625800.

Bobylev:2001:ESP

[BCG01] A. V. Bobylev, J. A. Carrillo, and I. M. Gamba. Erratum on ‘On Some Properties of Kinetic and Hydrodynamic Equations for Ineleastic Interactions’. Journal of Statistical Physics, 103 (5–6):1137–1138, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010325409175.

Bandi:2008:TFR

[BCG08] M. M. Bandi, J. R. Cressman, Jr., and W. I. Goldburg. Test of the fluctuation relation in Lagrangian turbulence on a free surface. Journal of Statistical Physics, 130(1):27–38, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9355-4.

Biskup:2000:CPD

[BCK00a] M. Biskup, L. Chayes, and R. Kotecký. Coexistence of partially Disordered/ ordered phases in an extended Potts model. Journal of Statistical Physics, 99(5–6):1169–1206, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018680520737.

Blanchard:2000:WCO

[BCK00b] Ph. Blanchard, B. Cessac, and T. Krüger. What can one learn about self-organized criticality from dynamical systems theory? Journal of Statistical Physics, 98(1–2):375–404, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018639308981.

Bianchi:2004:SSQ

[BCK04a] Alessandra Bianchi, Pierluigi Contucci, and Andreas Knauf. Stochastically stable quenched measures. Journal of Statistical Physics, 117(5–6):831–844, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-5707-5.

Biskup:2004:PGT

[BCK04b] Marek Biskup, Lincoln Chayes, and Roman Kotecký. A proof of the Gibbs–Thomson formula in the droplet formation regime. Journal of Statistical Physics, 116(1–4):175–203, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037209.36990.eb.

Borgs:2000:SPB

[BCKM00] C. Borgs, J. T. Chayes, C. King, and N. Madras. Sharp phase boundaries for a lattice flux line model. Journal of Statistical Physics, 98(5–6):1075–1113, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018611627599.

Bronski:2007:EFP

[BCL^{+}07] Jared C. Bronski, Roberto Camassa, Zhi Lin, Richard M. McLaughlin, and Alberto Scotti. An explicit family of probability measures for passive scalar diffusion in a random flow. Journal of Statistical Physics, 128(4):927–968, August 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9316-y.

Bisi:2008:ESI

[BCL08] Marzia Bisi, José A. Carrillo, and Bertrand Lods. Equilibrium solution to the inelastic Boltzmann equation driven by a particle bath. Journal of Statistical Physics, 133(5):841–870, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9636-6.

Butta:2002:LTB

[BCM02] Paolo Buttà, Emanuele Caglioti, and Carlo Marchioro. On the long time behavior of infinitely extended systems of particles interacting via Kac potentials. Journal of Statistical Physics, 108 (1–2):317–339, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015451905014.

Borgonovi:2004:BEC

[BCMP04] F. Borgonovi, G. L. Celardo, M. Maianti, and E. Pedersoli. Broken ergodicity in classically chaotic spin systems. Journal of Statistical Physics, 116(5–6):1435–1447, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041745.62340.00.

Bertini:2004:CPT

[BCO04] Lorenzo Bertini, Emilio N. M. Cirillo, and Enzo Olivieri. A combinatorial proof of tree decay of semi-invariants. Journal of Statistical Physics, 115(1–2):395–413, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019813.58778.bf.

Bertini:2007:PAD

[BCO07] Lorenzo Bertini, Emilio N. M. Cirillo, and Enzo Olivieri. Perturbative analysis of disordered Ising models close to criticality. Journal of Statistical Physics, 126(4–5):987–1006, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9214-8.

Braga:2007:DPC

[BCS07] Gastão A. Braga, Leandro M. Ciolleti, and Rémy Sanchis. Decay properties of the connectivity for mixed long range percolation models on Z^{d}. Journal of Statistical Physics, 129(3): 587–591, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9347-4.

Braga:2009:RDC

[BCS09] Gastão A. Braga, Leandro M. Cioletti, and Rémy Sanchis. A remark on the decay of correlations for mixed-range spin vector models. Journal of Statistical Physics, 136(1):195–198, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9777-2.

Bobylev:2003:PAP

[BCT03] A. V. Bobylev, C. Cercignani, and G. Toscani. Proof of an asymptotic property of self-similar solutions of the Boltzmann equation for granular materials. Journal of Statistical Physics, 111(1–2):403–417, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022273528296.

Bisi:2005:CMB

[BCT05] M. Bisi, J. A. Carrillo, and G. Toscani. Contractive metrics for a Boltzmann equation for granular gases: Diffusive equilibria. Journal of Statistical Physics, 118(1–2):301–331, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-8785-5.

Bisi:2006:DRP

[BCT06] M. Bisi, J. A. Carrillo, and G. Toscani. Decay rates in probability metrics towards homogeneous cooling states for the inelastic Maxwell model. Journal of Statistical Physics, 124(2– 4):625–653, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9035-9.

Bena:2000:SDR

[BCV00] I. Bena, M. Copelli, and C. Van den Broeck. Stokes’ drift: A rocking ratchet. Journal of Statistical Physics, 101(1–2): 415–424, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026466206162.

Briet:2004:DBC

[BCZ04] Philippe Briet, Horia D. Cornean, and Valentin A. Zagrebnov. Do bosons condense in a homogeneous magnetic field? Journal of Statistical Physics, 116(5–6):1545–1578, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041748.02351.07.

Benichou:2000:SPA

[BD00] Olivier Bénichou and Jean Desbois. Statistical properties of the 2D attached Rouse chain. Journal of Statistical Physics, 101(3– 4):921–931, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026433823078.

Brunet:2001:EMN

[BD01] Éric Brunet and Bernard Derrida. Effect of microscopic noise on front propagation. Journal of Statistical Physics, 103(1– 2):269–282, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004875804376.

Blaak:2002:RBT

[BD02] Ronald Blaak and David Dubbeldam. Regular binary thermal lattice-gases. Journal of Statistical Physics, 108(1–2):283–315, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015499820944.

Bezandry:2003:FTT

[BD03a] P. H. Bezandry and T. Diagana. Fluctuation theory for a three-dimensional model of Maxwellian molecules. Journal of Statistical Physics, 110(3–6):1375–1395, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022169532123.

Bru:2003:ESI

[BD03b] J.-B. Bru and T. C. Dorlas. Exact solution of the infinite-range-hopping Bose–Hubbard model. Journal of Statistical Physics, 113(1–2):177–196, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025774821712.

Barreira:2004:BAH

[BD04a] Luis Barreira and Paulo Doutor. Birkhoff averages for hyperbolic flows: Variational principles and applications. Journal of Statistical Physics, 115(5–6):1567–1603, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028069.64945.65.

Benguria:2004:VCP

[BD04b] Rafael Benguria and M. Cristina Depassier. Variational calculation of the period of nonlinear oscillators. Journal of Statistical Physics, 116(1–4):923–931, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037219.42798.f7.

Bunimovich:2005:DMS

[BD05] Leonid A. Bunimovich and Mark F. Demers. Deterministic models of the simplest chemical reactions. Journal of Statistical Physics, 120(1–2):239–252, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-5254-8.

Bandyopadhyay:2006:DDC

[BD06a] Malay Bandyopadhyay and Sushanta Dattagupta. Dissipative diamagnetism — a case study for equilibrium and nonequilibrium statistical mechanics. Journal of Statistical Physics, 123 (6):1273–1284, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9114-y.

Bisi:2006:RBE

[BD06b] M. Bisi and L. Desvillettes. From reactive Boltzmann equations to reaction–diffusion systems. Journal of Statistical Physics, 124 (2–4):881–912, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8075-x.

Bodineau:2006:CLD

[BD06c] T. Bodineau and B. Derrida. Current large deviations for asymmetric exclusion processes with open boundaries. Journal of Statistical Physics, 123(2):277–300, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9048-4.

Barreira:2009:DSH

[BD09] Luis Barreira and Paulo Doutor. Dimension spectra of hyperbolic flows. Journal of Statistical Physics, 136(3):505–525, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9790-5.

Bertini:2002:MFT

[BDG^{+}02] L. Bertini, A. De Sole, D. Gabrielli, G. Jona-Lasinio, and C. Landim. Macroscopic fluctuation theory for stationary nonequilibrium states. Journal of Statistical Physics, 107(3–4):635– 675, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014525911391.

Bertini:2004:MDP

[BDG^{+}04] L. Bertini, A. De Sole, D. Gabrielli, G. Jona-Lasinio, and C. Landim. Minimum dissipation principle in stationary nonequilibrium states. Journal of Statistical Physics, 116(1–4): 831–841, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000037220.57358.94.

Bertini:2006:NEC

[BDG^{+}06] L. Bertini, A. De Sole, D. Gabrielli, G. Jona-Lasinio, and C. Landim. Non equilibrium current fluctuations in stochastic lattice gases. Journal of Statistical Physics, 123(2):237–276, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9056-4.

Bertini:2009:TNT

[BDG^{+}09] L. Bertini, A. De Sole, D. Gabrielli, G. Jona-Lasinio, and C. Landim. Towards a nonequilibrium thermodynamics: A selfcontained macroscopic description of driven diffusive systems. Journal of Statistical Physics, 135(5–6):857–872, June 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9670-4.

Batchelor:2001:ESM

[BdGM01] M. T. Batchelor, J. de Gier, and M. Maslen. Exactly solvable su(N) mixed spin ladders. Journal of Statistical Physics, 102(3– 4):559–566, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004886500083.

Blanchard:2002:CPC

[BDGSC02] Ph. Blanchard, G. Dell’Antonio, D. Gandolfo, and M. Sirugue-Collin. Connectivity properties of continuum percolation processes on R^{2}. Journal of Statistical Physics, 106(1–2):1–22, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013157427257.

Bonetto:2000:PSN

[BDL00] F. Bonetto, D. Daems, and J. L. Lebowitz. Properties of stationary nonequilibrium states in the thermostatted periodic Lorentz gas i: The one particle system. Journal of Statistical Physics, 101 (1–2):35–60, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026414222092.

Bagland:2006:MSD

[BDL06] Véronique Bagland, Pierre Degond, and Mohammed Lemou. Moment systems derived from relativistic kinetic equations. Journal of Statistical Physics, 125(3):621–659, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9173-0.

Bodineau:2008:VTD

[BDL08] T. Bodineau, B. Derrida, and Joel L. Lebowitz. Vortices in the two-dimensional simple exclusion process. Journal of Statistical Physics, 131(5):821–841, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9518-y.

Bodineau:2008:LRC

[BDLvW08] T. Bodineau, B. Derrida, V. Lecomte, and F. van Wijland. Long range correlations and phase transitions in non-equilibrium diffusive systems. Journal of Statistical Physics, 133(6):1013– 1031, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9647-3.

Benguria:2009:HID

[BDM09] R. Benguria, J. Dolbeault, and R. Monneau. Harnack inequalities and discrete–continuous error estimates for a chain of atoms with two-body interactions. Journal of Statistical Physics, 134 (1):27–51, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9662-4.

Butta:2003:SMM

[BdMR03] Paolo Buttà, Anna de Masi, and Emanuele Rosatelli. Slow motion and metastability for a nonlocal evolution equation. Journal of Statistical Physics, 112(3–4):709–764, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023832210342.

Burns:2002:PHL

[BDP02] K. Burns, D. Dolgopyat, and Ya. Pesin. Partial hyperbolicity, Lyapunov exponents and stable ergodicity. Journal of Statistical Physics, 108(5–6):927–942, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019779128351.

Bunder:2007:SFA

[BEK^{+}07] J. E. Bunder, K. B. Efetov, V. E. Kravtsov, O. M. Yevtushenko, and M. R. Zirnbauer. Superbosonization formula and its application to random matrix theory. Journal of Statistical Physics, 129(5–6):809–832, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9405-y.

Bellissard:2004:RMT

[Bel04] Jean Bellissard. Random matrix theory and the Anderson model. Journal of Statistical Physics, 116(1–4):739–754, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037246.61440.6c.

Bastea:2000:BFL

[BELM00] S. Bastea, R. Esposito, J. L. Lebowitz, and R. Marra. Binary fluids with long range segregating interaction. I: Derivation of kinetic and hydrodynamic equations. Journal of Statistical Physics, 101(5–6):1087–1136, December 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026481706240.

Bastea:2006:SIM

[BELM06] Sorin Bastea, Raffaele Esposito, Joel L. Lebowitz, and Rossana Marra. Sharp interface motion of a binary fluid mixture. Journal of Statistical Physics, 124(2–4):445–483, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9040-z.

Benabou:2005:SBP

[Ben05] Gaël Benabou. Superdiffusive behaviour of a passive Ornstein– Uhlenbeck tracer in a turbulent shear flow. Journal of Statistical Physics, 121(3–4):319–341, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-7003-4.

Ben-Ezra:2002:MDB

[BEPK^{+}02] Y. Ben-Ezra, Yu. V. Pershin, Yu. A. Kaplunovsky, I. D. Vagner, and P. Wyder. Modeling of the dielectric breakdown under strong magnetic fields. Journal of Statistical Physics, 106(3– 4):653–662, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013714523664.

Berestycki:2003:MSF

[Ber03a] Julien Berestycki. Multifractal spectra of fragmentation processes. Journal of Statistical Physics, 113(3–4):411–430, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026060516513.

Bergman:2003:ERB

[Ber03b] David J. Bergman. Exact relations between elastic and electrical response of d-dimensional percolating networks with anglebending forces. Journal of Statistical Physics, 111(1–2):171–199, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022205007823.

Bernardin:2005:FKD

[Ber05] Cèdric Bernardin. Fluctuations for Kawasaki dynamics. Journal of Statistical Physics, 119(3–4):827–852, May 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3769-7.

Bershadskii:2007:BSL

[Ber07] Alexander Bershadskii. Beyond scaling and locality in turbulence. Journal of Statistical Physics, 128(3):721–739, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9322-0.

Bernardin:2008:TCN

[Ber08] Cédric Bernardin. Thermal conductivity for a noisy disordered harmonic chain. Journal of Statistical Physics, 133(3): 417–433, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9620-1.

Boucher:2000:DME

[BET00] Christopher Boucher, Richard S. Ellis, and Bruce Turkington. Derivation of maximum entropy principles in two-dimensional turbulence via large deviations. Journal of Statistical Physics, 98(5–6):1235–1278, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018671813486.

Belitsky:2005:IMC

[BF05] Vladimir Belitsky and Pablo A. Ferrari. Invariant measures and convergence properties for cellular automaton 184 and related processes. Journal of Statistical Physics, 118(3–4):589–623, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8822-4.

Barros:2002:ABS

[BFGM02] Saulo R. M. Barros, Pablo A. Ferrari, Nancy L. Garcia, and Servet Martínez. Asymptotic behavior of a stationary silo with absorbing walls. Journal of Statistical Physics, 106(3–4): 521–546, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013702220938.

Balaban:2009:PSR

[BFKT09] Tadeusz Balaban, Joel Feldman, Horst Knörrer, and Eugene Trubowitz. Power series representations for bosonic effective actions. Journal of Statistical Physics, 134(5–6):839–857, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9634-8.

Bornemann:2008:APL

[BFP08] Folkmar Bornemann, Patrik L. Ferrari, and Michael Prähofer. The Airy_{1} process is not the limit of the largest eigenvalue in GOE matrix diffusion. Journal of Statistical Physics, 133(3): 405–415, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9621-0.

Borodin:2007:FPT

[BFPS07] Alexei Borodin, Patrik L. Ferrari, Michael Prähofer, and Tomohiro Sasamoto. Fluctuation properties of the TASEP with periodic initial configuration. Journal of Statistical Physics, 129(5–6):1055–1080, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9383-0.

Borodin:2009:TST

[BFS09] Alexei Borodin, Patrik L. Ferrari, and Tomohiro Sasamoto. Two speed TASEP. Journal of Statistical Physics, 137(5–6):

936–977, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9837-7.

Boldrighini:2002:NEL

[BFT02] C. Boldrighini, S. Frigio, and D. Tognetti. Numerical evidence for the low-mass behavior of the one-dimensional Rayleigh gas with local interaction. Journal of Statistical Physics, 108(3– 4):703–712, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015738209984.

Barra:2000:LSD

[BG00a] F. Barra and P. Gaspard. On the level spacing distribution in quantum graphs. Journal of Statistical Physics, 101(1–2): 283–319, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026495012522.

Bassi:2000:CQR

[BG00b] Angelo Bassi and GianCarlo Ghirardi. Consistent quantum realism: A reply to Griffiths. Journal of Statistical Physics, 99 (5–6):1427, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018653209350.

Bassi:2000:DHR

[BG00c] Angelo Bassi and GianCarlo Ghirardi. Decoherent histories and realism. Journal of Statistical Physics, 98(1–2):457–494, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018647510799.

Barata:2001:DGS

[BG01a] João C. A. Barata and Pedro S. Goldbaum. On the distribution and gap structure of Lee–Yang zeros for the Ising model: Periodic and aperiodic couplings. Journal of Statistical Physics, 103 (5–6):857–891, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010332500031.

Bauer:2001:RIM

[BG01b] M. Bauer and O. Golinelli. Random incidence matrices: Moments of the spectral density. Journal of Statistical Physics, 103 (1–2):301–337, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004879905284.

Barra:2002:ADH

[BG02] Felipe Barra and Thomas Gilbert. Algebraic decay in hierarchical graphs. Journal of Statistical Physics, 109(3–4):777– 801, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020427018179.

Berglund:2004:NIP

[BG04a] Nils Berglund and Barbara Gentz. On the noise-induced passage through an unstable periodic orbit I: Two-level model. Journal of Statistical Physics, 114(5–6):1577–1618, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013966.54504.da.

Bodineau:2004:LTR

[BG04b] Thierry Bodineau and Giambattista Giacomin. On the localization transition of random copolymers near selective interfaces. Journal of Statistical Physics, 117(5–6):801–818, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5705-7.

Bobylev:2006:BEM

[BG06] A. V. Bobylev and I. M. Gamba. Boltzmann equations for mixtures of Maxwell gases: Exact solutions and power like tails. Journal of Statistical Physics, 124(2–4):497–516, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9044-8.

Bjornberg:2009:PTQ

[BG09] J. E. Björnberg and G. R. Grimmett. The phase transition of the quantum Ising model is sharp. Journal of Statistical Physics, 136(2):231–273, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9788-z.

Bardos:2004:ATD

[BGGM04] Claude Bardos, François Golse, Alex D. Gottlieb, and Norbert J. Mauser. Accuracy of the time-dependent Hartree– Fock approximation for uncorrelated initial states. Journal of Statistical Physics, 115(3–4):1037–1055, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022381.86923.0a.

Bonetto:2006:CHF

[BGGZ06] F. Bonetto, G. Gallavotti, A. Giuliani, and F. Zamponi. Chaotic hypothesis, fluctuation theorem and singularities. Journal of Statistical Physics, 123(1):39–54, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9047-5.

Bobylev:2001:EIE

[BGH01] A. V. Bobylev, R. Grzhibovskis, and A. Heintz. Entropy inequalities for Evaporation/ condensation problem in rarefied gas dynamics. Journal of Statistical Physics, 102(5–6):1151–1176, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004832126858.

Boon:2003:PDE

[BGL03] Jean Pierre Boon, Patrick Grosfils, and James F. Lutsko. Propagation-dispersion equation. Journal of Statistical Physics, 113(3–4):527–548, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026068718330.

Bertini:2005:LDS

[BGL05] Lorenzo Bertini, Davide Gabrielli, and Joel L. Lebowitz. Large deviations for a stochastic model of heat flow. Journal of Statistical Physics, 121(5–6):843–885, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5527-2.

Bodineau:2008:CSI

[BGLT08] Thierry Bodineau, Giambattista Giacomin, Hubert Lacoin, and Fabio Lucio Toninelli. Copolymers at selective interfaces: New bounds on the phase diagram. Journal of Statistical Physics, 132 (4):603–626, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9579-y.

Brydges:2004:FRD

[BGM04] David C. Brydges, G. Guadagni, and P. K. Mitter. Finite range decomposition of Gaussian processes. Journal of Statistical Physics, 115(1–2):415–449, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019818.81237.66.

Banavar:2003:SIS

[BGMM03] Jayanth R. Banavar, Oscar Gonzalez, John H. Maddocks, and Amos Maritan. Self-interactions of strands and sheets. Journal of Statistical Physics, 110(1–2):35–50, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021010526495.

Batchelor:2004:QPD

[BGOY04] M. T. Batchelor, X.-W. Guan, N. Oelkers, and Z.-J. Ying. Quantum phase diagram of an exactly solved mixed spin ladder. Journal of Statistical Physics, 116(1–4):571–589, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037225.79748.98.

Bobylev:2004:MIH

[BGP04] A. V. Bobylev, I. M. Gamba, and V. A. Panferov. Moment inequalities and high-energy tails for Boltzmann equations with inelastic interactions. Journal of Statistical Physics, 116(5–6): 1651–1682, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000041751.11664.ea.

Bardos:2006:HSP

[BGS06] Claude Bardos, François Golse, and Yoshio Sone. Half-space problems for the Boltzmann equation: A survey. Journal of Statistical Physics, 124(2–4):275–300, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9077-z.

Bodineau:2001:ERF

[BGV01] Thierry Bodineau, Giambattista Giacomin, and Yvan Velenik. On entropic reduction of fluctuations. Journal of Statistical Physics, 102(5–6):1439–1445, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004860900450.

Banerjee:2001:TLP

[BGW01] Shubho Banerjee, R. B. Griffiths, and M. Widom. Thermodynamic limit for polydisperse fluids. Journal of Statistical Physics, 104(3–4):725–752, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1010380705975.

Bai:2001:FUH

[BGX01] Cheng-Ming Bai, Mo-Lin Ge, and Kang Xue. Further understanding of hydrogen atom: Yangian approach and physical effect. Journal of Statistical Physics, 102(3–4):545–557, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004834516013.

Baake:2000:DRT

[BH00] Michael Baake and Moritz Höffe. Diffraction of random tilings: Some rigorous results. Journal of Statistical Physics, 99(1– 2):219–261, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018648707744.

Bakhtin:2008:LDR

[BH08a] Yuri Bakhtin and Christine Heitsch. Large deviations for random trees. Journal of Statistical Physics, 132(3):551–560, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9540-0.

Benjamini:2008:ECT

[BH08b] Itai Benjamini and Christopher Hoffman. Exponential clogging time for a one dimensional DLA. Journal of Statistical Physics, 131(6):1185–1188, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9557-4.

Bobylev:2001:LEG

[BHPH01] A. V. Bobylev, Alex Hansen, J. Piasecki, and E. H. Hauge. From the Liouville equation to the generalized Boltzmann equation for magnetotransport in the 2D Lorentz model. Journal of Statistical Physics, 102(5–6):1133–1150, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004880010020.

Bloch:2004:GSO

[BHRW04] Anthony M. Bloch, Patrick Hagerty, Alberto G. Rojo, and Michael I. Weinstein. Gyroscopically stabilized oscillators and heat baths. Journal of Statistical Physics, 115(3–4):1073–1100, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022367.36305.d3.

Bellissard:2007:CEA

[BHS07] Jean V. Bellissard, Peter D. Hislop, and Günter Stolz. Correlation estimates in the Anderson model. Journal of Statistical Physics, 129(4):649–662, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9409-7.

Brydges:2003:DRF

[BI03] David C. Brydges and John Z. Imbrie. Dimensional reduction formulas for branched polymer correlation functions. Journal of Statistical Physics, 110(3–6):503–518, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022143331697.

Barreira:2006:SFC

[BI06] Luis Barreira and Godofredo Iommi. Suspension flows over countable Markov shifts. Journal of Statistical Physics, 124 (1):207–230, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9140-9.

Biryuk:2006:IME

[Bir06] Andrei Biryuk. On invariant measures of the 2D Euler equation. Journal of Statistical Physics, 122(4):597–616, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8011-0.

Birnir:2007:OMM

[Bir07] Björn Birnir. An ODE model of the motion of pelagic fish. Journal of Statistical Physics, 128(1–2):535–568, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9292-2.

Bodineau:2001:WCF

[BIV01] T. Bodineau, D. Ioffe, and Y. Velenik. Winterbottom construction for finite range ferromagnetic models: An L_{1}-approach. Journal of Statistical Physics, 105(1–2):93–131, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1012277926007.

Baker:2000:PPR

[BJ00] George A. Baker, Jr. and J. D. Johnson. The Pauli principle and the restricted primitive model. Journal of Statistical Physics, 100(1–2):233–242, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018647930248.

Bourgain:2002:CLE

[BJ02] J. Bourgain and S. Jitomirskaya. Continuity of the Lyapunov exponent for quasiperiodic operators with analytic potential. Journal of Statistical Physics, 108(5–6):1203–1218, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019751801035.

Bao:2006:LPT

[BJ06] Jing-Dong Bao and Ying Jia. Last passage time statistics for barrier-crossing processes. Journal of Statistical Physics, 123 (4):861–869, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9082-2.

Barre:2008:FTL

[BJ08] J. Barré and P. E. Jabin. Free transport limit for n-particles dynamics with singular and short range potential. Journal of Statistical Physics, 131(6):1085–1101, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9526-y.

Bjornberg:2009:CVQ

[Bjö09] Jakob E. Björnberg. Critical value of the quantum Ising model on star-like graphs. Journal of Statistical Physics, 135(3):571– 583, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9749-6.

Baladi:2002:DDD

[BJR02] Viviane Baladi, Yunping Jiang, and Hans Henrik Rugh. Dynamical determinants via dynamical conjugacies for postcritically finite polynomials. Journal of Statistical Physics, 108(5–6): 973–993, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019783229260.

Biskup:2001:SED

[BK01a] Marek Biskup and Wolfgang König. Screening effect due to heavy lower tails in one-dimensional parabolic Anderson model. Journal of Statistical Physics, 102(5–6):1253–1270, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004840328675.

Bunimovich:2001:LPR

[BK01b] Leonid A. Bunimovich and Milena A. Khlabystova. Localization and propagation in random lattices. Journal of Statistical Physics, 104(5–6):1155–1171, September 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010409913034.

Bunimovich:2002:WRE

[BK02] Leonid A. Bunimovich and Milena A. Khlabystova. Walks in rigid environments: Continuous limits. Journal of Statistical Physics, 108(5–6):905–925, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019727111513.

Bec:2003:FBE

[BK03a] Jérémie Bec and Konstantin Khanin. Forced Burgers equation in an unbounded domain. Journal of Statistical Physics, 113(5– 6):741–759, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1027356518273.

Bunimovich:2003:ODL

[BK03b] Leonid A. Bunimovich and Milena A. Khlabystova. Onedimensional Lorentz gas with rotating scatterers: Exact solutions. Journal of Statistical Physics, 112(5–6):1207–1218, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1024623827182.

Blanchard:2004:CPO

[BK04] Ph. Blanchard and T. Krüger. The ‘Cameo Principle’ and the origin of scale-free graphs in social networks. Journal of Statistical Physics, 114(5–6):1399–1416, March 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013969.87579.39.

Bovier:2007:LES

[BK07] Anton Bovier and Irina Kurkova. Local energy statistics in spin glasses. Journal of Statistical Physics, 126(4–5):933–949, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9141-8.

Balazs:2008:OCV

[BK08] Márton Balázs and Júlia Komjáthy. Order of current variance and diffusivity in the rate one totally asymmetric zero range process. Journal of Statistical Physics, 133(1):59–78, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9604-1.

Bricmont:2009:RWS

[BK09] Jean Bricmont and Antti Kupiainen. Random walks in space time mixing environments. Journal of Statistical Physics, 134 (5–6):979–1004, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9689-1.

Bricmont:2000:PET

[BKL00] J. Bricmont, A. Kupiainen, and R. Lefevere. Probabilistic estimates for the two-dimensional stochastic Navier–Stokes equations. Journal of Statistical Physics, 100(3–4):743–756, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018627609718.

Beghin:2000:GLB

[BKLO00] Luisa Beghin, Viktoria P. Knopova, Nikolai N. Leonenko, and Enzo Orsingher. Gaussian limiting behavior of the rescaled solution to the linear Korteweg–de Vries equation with random initial conditions. Journal of Statistical Physics, 99(3–4):769–781, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018687327580.

Borgs:2002:FSE

[BKM02] C. Borgs, R. Kotecký, and I. Medved’. Finite-size effects for the Potts model with weak boundary conditions. Journal of Statistical Physics, 109(1–2):67–131, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019931410450.

Belitsky:2001:CAM

[BKNS01] V. Belitsky, J. Krug, E. Jordão Neves, and G. M. Schütz. A cellular automaton model for two-lane traffic. Journal of Statistical Physics, 103(5–6):945–971, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010361022379.

Beghin:2007:SLO

[BKOS07] L. Beghin, Yu. Kozachenko, E. Orsingher, and L. Sakhno. On the solutions of linear odd-order heat-type equations with random initial conditions. Journal of Statistical Physics, 127 (4):721–739, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9309-x.

Baake:2004:CCP

[BL04] Michael Baake and Matthias Löwe. Comment on: Curious properties of simple random walks. Journal of Statistical Physics, 116(5–6):1449–1451, September 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041746.43933.73.

Boilley:2006:NMD

[BL06] David Boilley and Yoann Lallouet. Non-Markovian diffusion over a saddle with a generalized Langevin equation. Journal of Statistical Physics, 125(2):473–489, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9197-5.

Bodineau:2008:LDL

[BL08a] Thierry Bodineau and Raphaël Lefevere. Large deviations of lattice Hamiltonian dynamics coupled to stochastic thermostats. Journal of Statistical Physics, 133(1):1–27, October 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9601-4.

Braga:2008:REB

[BL08b] Gastão A. Braga and Paulo C. Lima. On the residual entropy of the Blume–Emery–Griffiths model. Journal of Statistical Physics, 130(3):571–578, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9457-z.

Braga:2008:RRE

[BL08c] Gastão A. Braga and Paulo C. Lima. A remark on the residual entropy of the antiferromagnetic Ising model in the maximal critical field. Journal of Statistical Physics, 131(6):1189–1193, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9549-4.

Bleher:2009:ESS

[BL09] Pavel Bleher and Karl Liechty. Exact solution of the six-vertex model with domain wall boundary conditions. Critical line between ferroelectric and disordered phases. Journal of Statistical Physics, 134(3):463–485, February 2009. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9688-2.

Blank:2003:EPS

[Bla03] Michael Blank. Ergodic properties of a simple deterministic traffic flow model. Journal of Statistical Physics, 111(3–4):903–930, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022806500731.

Blank:2005:HPD

[Bla05] Michael Blank. Hysteresis phenomenon in deterministic traffic flows. Journal of Statistical Physics, 120(3–4):627–658, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-5959-8.

Blank:2008:TAF

[Bla08] Michael Blank. Travelling with/against the flow. Deterministic diffusive driven systems. Journal of Statistical Physics, 133(4): 773–796, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9637-5.

Bonetto:2004:FLH

[BLL04] Federico Bonetto, Joel L. Lebowitz, and Jani Lukkarinen. Fourier’s law for a harmonic crystal with self-consistent stochastic reservoirs. Journal of Statistical Physics, 116(1–4):783–813, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037232.14365.10.

Bonetto:2009:HCE

[BLLO09] F. Bonetto, J. L. Lebowitz, J. Lukkarinen, and S. Olla. Heat conduction and entropy production in anharmonic crystals with selfconsistent stochastic reservoirs. Journal of Statistical Physics, 134(5–6):1097–1119, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9657-1.

Binder:2003:MCS

[BLM03] Kurt Binder, David Landau, and Marcus Müller. Monte Carlo studies of wetting, interface localization and capillary condensation. Journal of Statistical Physics, 110(3–6):1411–1514, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022173600263.

Blower:2001:ASW

[Blo01] Gordon Blower. Almost sure weak convergence for the generalized orthogonal ensemble. Journal of Statistical Physics, 105(1– 2):309–335, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012294429641.

Blower:2004:DCG

[Blo04] Gordon Blower. Displacement convexity for the generalized orthogonal ensemble. Journal of Statistical Physics, 116(5–6): 1359–1387, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000041742.86859.cd.

Benettin:2009:FPU

[BLP09] G. Benettin, R. Livi, and A. Ponno. The Fermi–Pasta–Ulam problem: Scaling laws vs. initial conditions. Journal of Statistical Physics, 135(5–6):873–893, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9660-6.

Bassetti:2008:PSS

[BLR08] Federico Bassetti, Lucia Ladelli, and Eugenio Regazzini. Probabilistic study of the speed of approach to equilibrium for an inelastic Kac model. Journal of Statistical Physics, 133(4): 683–710, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9630-z.

Balinsky:2004:GHI

[BLS04] Alexander Balinsky, Ari Laptev, and Alexander V. Sobolev. Generalized Hardy inequality for the magnetic Dirichlet forms. Journal of Statistical Physics, 116(1–4):507–521, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037228.35518.ca.

Bazhanov:2001:SDS

[BLZ01] Vladimir V. Bazhanov, Sergei L. Lukyanov, and Alexander B. Zamolodchikov. Spectral determinants for Schrödinger equation and q-operators of conformal field theory. Journal of Statistical Physics, 102(3–4):567–576, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004838616921.

Bronski:2000:RET

[BM00] Jared C. Bronski and Richard M. McLaughlin. Rigorous estimates of the tails of the probability distribution function for the random linear shear model. Journal of Statistical Physics, 98(3– 4):897–915, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018639928526.

Berlyand:2001:GCM

[BM01a] L. Berlyand and V. Mityushev. Generalized Clausius–Mossotti formula for random composite with circular fibers. Journal of Statistical Physics, 102(1–2):115–145, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026512725967.

Boukraa:2001:LB

[BM01b] S. Boukraa and J.-M. Maillard. Let’s baxterise. Journal of Statistical Physics, 102(3–4):641–700, February 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004842717830.

Bodineau:2002:SNR

[BM02a] T. Bodineau and Fabio Martinelli. Some new results on the kinetic Ising model in a pure phase. Journal of Statistical Physics, 109(1–2):207–235, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019939712267.

Bovier:2002:MGD

[BM02b] Anton Bovier and Francesco Manzo. Metastability in Glauber dynamics in the low-temperature limit: Beyond exponential asymptotics. Journal of Statistical Physics, 107(3–4):757–779, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014586130046.

Bach:2004:CLT

[BM04a] Volker Bach and Jacob Schach Møller. Correlation at low temperature: II. Asymptotics. Journal of Statistical Physics, 116 (1–4):591–628, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000037236.34145.20.

Bauke:2004:PRC

[BM04b] Heiko Bauke and Stephan Mertens. Pseudo random coins show more heads than tails. Journal of Statistical Physics, 114(3–4): 1149–1169, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://arxiv.org/abs/

cond-mat/0307138;   http://link.springer.com/article/
10.1023/B%3AJOSS.0000012521.67853.9a.

Benfatto:2004:WIV

[BM04c] G. Benfatto and V. Mastropietro. Ward identities and vanishing of the beta function for d = 1 interacting Fermi systems. Journal of Statistical Physics, 115(1–2):143–184, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000019825.37968.23.

Berlyand:2005:IDE

[BM05a] L. Berlyand and V. Mityushev. Increase and decrease of the effective conductivity of two phase composites due to polydispersity. Journal of Statistical Physics, 118(3–4):481–509, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8818-0.

Buenzli:2005:MOU

[BM05b] Pascal R. Buenzli and Philippe A. Martin. Microscopic origin of universality in Casimir forces. Journal of Statistical Physics, 119 (1–2):273–307, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-1990-4.

Balint:2008:DCI

[BM08] Péter Bálint and Ian Melbourne. Decay of correlations and invariance principles for dispersing billiards with cusps, and related planar billiard flows. Journal of Statistical Physics, 133(3): 435–447, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9623-y.

Ballenegger:2002:QMG

[BMA02] V. Ballenegger, Ph. A. Martin, and A. Alastuey. Quantum Mayer graphs for Coulomb systems and the analog of the Debye potential. Journal of Statistical Physics, 108(1–2):169–211, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015443603197.

Buschle:2000:WID

[BMD00] J. Buschle, P. Maass, and W. Dieterich. Wall-induced density profiles and density correlations in confined Takahashi lattice gases. Journal of Statistical Physics, 99(1–2):273–312, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018652808652.

Bruers:2007:VEP

[BMN07] Stijn Bruers, Christian Maes, and Karel Netocný. On the validity of entropy production principles for linear electrical circuits. Journal of Statistical Physics, 129(4):725–740, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9412-z.

Baiesi:2009:CCC

[BMN09] Marco Baiesi, Christian Maes, and Karel Netocný. Computation of current cumulants for small nonequilibrium systems. Journal of Statistical Physics, 135(1):57–75, April 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9723-3.

Boldrighini:2007:APM

[BMP07] C. Boldrighini, S. Molchanov, and A. Pellegrinotti. Anderson parabolic model for a quasi-stationary medium. Journal of Statistical Physics, 129(1):151–169, October 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9364-3.

Bovier:2004:GPR

[BMPZ04] A. Bovier, I. Merola, E. Presutti, and M. Zahradník. On the Gibbs phase rule in the Pirogov–Sinai regime. Journal of Statistical Physics, 114(5–6):1235–1267, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013970.66907.b9.

Baiesi:2009:NLR

[BMW09] Marco Baiesi, Christian Maes, and Bram Wynants. Nonequilibrium linear response for Markov dynamics, i: Jump processes and overdamped diffusions. Journal of Statistical Physics, 137(5–6):1094–1116, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9852-8.

Bru:2002:ESM

[BNZ02] J.-B. Bru, B. Nachtergaele, and V. A. Zagrebnov. The equilibrium states for a model with two kinds of Bose condensation. Journal of Statistical Physics, 109(1–2):143–176, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019935511359.

Ben-Naim:2007:SED

[BNZ07] E. Ben-Naim and A. Zippelius. Singular energy distributions in driven and undriven granular media. Journal of Statistical Physics, 129(4):677–697, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9411-0.

Bernardin:2005:FLM

[BO05] Cédric Bernardin and Stefano Olla. Fourier’s law for a microscopic model of heat conduction. Journal of Statistical Physics, 121(3–4):271–289, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-7578-9.

Bobylev:2006:ICE

[Bob06] A. V. Bobylev. Instabilities in the Chapman–Enskog expansion and hyperbolic Burnett equations. Journal of Statistical Physics, 124(2–4):371–399, August 2006. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-005-8087-6.

Bobylev:2008:GBH

[Bob08] A. V. Bobylev. Generalized Burnett hydrodynamics. Journal of Statistical Physics, 132(3):569–580, August 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9556-5.

Boguna:2004:BRB

[Bog04] Marian Boguñá. Book review: Evolution of Networks. From Biological Nets to the Internet and WWW. S. N. Dorogovtsev and J. F. F. Mendes, Oxford University Press, Oxford, 2003. Journal of Statistical Physics, 114(5–6):1627–1628, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000014281.61380.75.

Boguna:2007:SDN

[Bog07] Marián Boguñá. The structure and dynamics of networks. Journal of Statistical Physics, 126(2):419–421, January 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9267-8.

Boguna:2009:MFR

[Bog09] Marián Boguñá. M. Franceschetti, R. Meester: Random Networks for Communication. From Statistical Physics to Information Systems. Journal of Statistical Physics, 135(3):585–586, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9740-2.

Bolmatov:2009:ESS

[Bol09] Dima Bolmatov. Equations of state for simple liquids from the Gaussian equivalent representation method. Journal of Statistical Physics, 137(4):765–773, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9874-2.

Boon:2001:HFD

[Boo01] Jean Pierre Boon. How fast does Langton’s ant move? Journal of Statistical Physics, 102(1–2):355–360, January 2001.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026581213671.

Borodin:2002:DOP

[Bor02] Alexei Borodin. Duality of orthogonal polynomials on a finite set. Journal of Statistical Physics, 109(5–6):1109–1120, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020432812090.

Boutillier:2007:NCP

[Bou07] Cédric Boutillier. Non-colliding paths in the honeycomb dimer model and the Dyson process. Journal of Statistical Physics, 129(5–6):1117–1135, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9431-9.

Bourgain:2009:AWE

[Bou09] Jean Bourgain. An approach to Wegner’s estimate using subharmonicity. Journal of Statistical Physics, 134(5–6):969–978, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9729-x.

Behrend:2001:ICB

[BP01a] Roger E. Behrend and Paul A. Pearce. Integrable and conformal boundary conditions for ŝℓ(2) A-D-E lattice models and unitary minimal conformal field theories. Journal of Statistical Physics, 102(3–4):577–640, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004890600991.

Blanco:2001:LDT

[BP01b] R. Blanco and L. Pesquera. Long-distance transmission at zero dispersion: Exact expressions for one-time statistical properties. Journal of Statistical Physics, 104(1–2):167–192, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010357610916.

Belitsky:2002:UGS

[BP02] V. Belitsky and E. A. Pechersky. Uniqueness of Gibbs state for non-ideal gas in R^{d}: The case of multibody interaction.

Journal of Statistical Physics, 106(5–6):931–955, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014029602226.

Berestycki:2007:GDR

[BP07] Nathanaël Berestycki and Jim Pitman. Gibbs distributions for random partitions generated by a fragmentation process. Journal of Statistical Physics, 127(2):381–418, April 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9261-1.

Boland:2008:LCI

[BP08a] G. Boland and J. V. Pulé. Long cycles in the infiniterange-hopping Bose–Hubbard model with hard cores. Journal of Statistical Physics, 132(5):881–905, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9586-z.

Borodin:2008:AKT

[BP08b] Alexei Borodin and Sandrine Péché. Airy kernel with two sets of parameters in directed percolation and random matrix theory. Journal of Statistical Physics, 132(2):275–290, July 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9553-8.

Bruneau:2009:TRQ

[BP09] L. Bruneau and C.-A. Pillet. Thermal relaxation of a QED cavity. Journal of Statistical Physics, 134(5–6):1071–1095, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9656-2.

Braga:2002:ADB

[BPS02] Gastão A. Braga, Aldo Proccaci, and Remy Sanchis. Analyticity of the d-dimensional bond percolation probability around p = 1. Journal of Statistical Physics, 107(5–6):1267–1282, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1015170027110. See erratum [BPS03].

Braga:2003:EAD

[BPS03] Gastão A. Braga, Aldo Procacci, and Remy Sanchis. Erratum on ‘Analyticity of the d-Dimensional Bond Percolation Probability Around p = 1,’ J. Stat. Phys. 107:1267 (2002). Journal of Statistical Physics, 111(1–2):501, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022289931930. See [BPS02].

Butta:2005:KP

[BPS05] Paolo Buttà, Aldo Procacci, and Benedetto Scoppola. Kac polymers. Journal of Statistical Physics, 119(3–4):643–658, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-3771-0.

Bechouche:2006:QTB

[BPS06a] Philippe Bechouche, Frédéric Poupaud, and Juan Soler. Quantum transport and Boltzmann operators. Journal of Statistical Physics, 122(3):417–436, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8082-y.

Bouchbinder:2006:SPF

[BPS06b] Eran Bouchbinder, Itamar Procaccia, and Shani Sela. Statistical physics of fracture surfaces morphology. Journal of Statistical Physics, 125(5–6):1025–1064, December 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9045-7.

Baik:2000:LDP

[BR00a] Jinho Baik and Eric M. Rains. Limiting distributions for a polynuclear growth model with external sources. Journal of Statistical Physics, 100(3–4):523–541, August 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018615306992.

Betancort-Rijo:2000:GNB

[BR00b] J. Betancort-Rijo. Generalized negative binomial distributions. Journal of Statistical Physics, 98(3–4):917–933, February 2000.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018691912596.

Brandt:2001:RMR

[BR01] Achi Brandt and Dorit Ron. Renormalization multigrid (RMG): Statistically optimal renormalization group flow and coarse-to-fine Monte Carlo acceleration. Journal of Statistical Physics, 102(1–2):231–257, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026520927784.

Bleher:2002:RP

[BR02] Pavel Bleher and Denis Ridzal. SU(1, 1) random polynomials. Journal of Statistical Physics, 106(1–2):147–171, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013124213145.

Brummelhuis:2004:ODM

[BR04]  Raymond Brummelhuis and Mary Beth Ruskai.  One-

dimensional models for atoms in strong magnetic fields, II: Antisymmetry in the Landau levels. Journal of Statistical Physics, 116(1–4):547–570, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037229.51177.
6d.

Barik:2005:QSD

[BR05a] Debashis Barik and Deb Shankar Ray. Quantum statedependent diffusion and multiplicative noise: A microscopic approach. Journal of Statistical Physics, 120(1–2):339–365, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-5251-y.

Borodin:2005:EMT

[BR05b] Alexei Borodin and Eric M. Rains. Eynard–Mehta theorem, Schur process, and their Pfaffian analogs. Journal of Statistical Physics, 121(3–4):291–317, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-7583-z.

Belitsky:2006:ILB

[BR06] Vladimir Belitsky and Thomas Logan Ritchie. Improved lower bounds for the critical probability of oriented bond percolation in two dimensions. Journal of Statistical Physics, 122(2):279–302, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8022-x.

Bray:2002:BRB

[Bra02] A. J. Bray. Book review: A Guide to First-Passage Processes. Sidney Redner. Cambridge University Press, United Kingdom, 2001. Journal of Statistical Physics, 106(3–4):853–854, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013783011368.

Broadwell:2001:IRL

[Bro01] James E. Broadwell. Irreversibility in a reversible lattice gas. Journal of Statistical Physics, 103(5–6):1125–1136, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010373325105.

Borodin:2003:JDD

[BS03] Alexei Borodin and Alexander Soshnikov. Janossy densities. I. Determinantal ensembles. Journal of Statistical Physics, 113(3– 4):595–610, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026025003309.

Barral:2005:ISS

[BS05] Julien Barral and Stéphane Seuret. Inside singularity sets of random Gibbs measures. Journal of Statistical Physics, 120(5–6): 1101–1124, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5458-y.

Balazs:2007:ECB

[BS07] Márton Balázs and Timo Seppäläinen. Exact connections between current fluctuations and the second class particle in a class of deposition models. Journal of Statistical Physics, 127 (2):431–455, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9291-3.

Barreira:2008:MNT

[BS08] Luis Barreira and Vitor Saraiva. Multifractal nonrigidity of topological Markov chains. Journal of Statistical Physics, 130 (2):387–412, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9429-3.

Bertoin:2009:STC

[BS09a] Jean Bertoin and Vladas Sidoravicius. The structure of typical clusters in large sparse random configurations. Journal of Statistical Physics, 135(1):87–105, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9728-y.

Biskup:2009:MBB

[BS09b] Marek Biskup and Roberto H. Schonmann. Metastable behavior for bootstrap percolation on regular trees. Journal of Statistical Physics, 136(4):667–676, August 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9798-x.pdf.

Bellissard:2000:SQT

[BSB00] J. Bellissard and H. Schulz-Baldes. Subdiffusive quantum transport for 3D Hamiltonians with absolutely continuous spectra. Journal of Statistical Physics, 99(1–2):587–594, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018613227308.

Balint:2004:RIA

[BT04a] Péter Bálint and Serge Troubetzkoy. Rotor interaction in the annulus billiard. Journal of Statistical Physics, 117(3–4):681– 702, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2274-8.

Bertini:2004:EPD

[BT04b] Lorenzo Bertini and Cristina Toninelli. Exclusion processes with degenerate rates: Convergence to equilibrium and tagged particle. Journal of Statistical Physics, 117(3–4):549–580, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-3453-3.

Bahraminasab:2004:ZTK

[BTM^{+}04] A. Bahraminasab, S. M. A. Tabei, A. A. Masoudi, F. Shahbazi, and M. Reza Rahimi Tabar. Zero tension Kardar– Parisi–Zhang equation in (d + 1)-dimensions. Journal of Statistical Physics, 116(5–6):1521–1544, September 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041747.55551.8b.

Balint:2007:ZML

[BTT07] Péter Bálint, Bálint Tóth, and Péter Tóth. On the zero mass limit of tagged particle diffusion in the 1-D Rayleigh-gas. Journal of Statistical Physics, 127(4):657–675, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9304-2.

Berthelin:2009:DVB

[BTV09] Florent Berthelin, Athanasios E. Tzavaras, and Alexis Vasseur. From discrete velocity Boltzmann equations to gas dynamics before shocks. Journal of Statistical Physics, 135(1):153–173, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9709-1.

Buchete:2005:BRB

[Buc05] Nicolae-Viorel Buchete. Book review: The Art of Molecular Dynamics Simulation. Journal of Statistical Physics, 118(1–2): 365–367, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-8788-2.

Buchete:2007:SKU

[Buc07] Nicolae-Viorel Buchete. Santosh K. Upadhyay. Chemical Kinetics and Reaction Dynamics. Journal of Statistical Physics, 129 (2):407–408, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9418-6.

Budini:2008:OCQ

[Bud08] Adrián A. Budini. Operator correlations and quantum regression theorem in non–Markovian Lindblad rate equations. Journal of Statistical Physics, 131(1):51–78, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9476-9.

Bunimovich:2000:HA

[Bun00] Leonid A. Bunimovich. Hyperbolicity and astigmatism. Journal of Statistical Physics, 101(1–2):373–384, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026405920274.

Burkhardt:2008:FPE

[Bur08] Theodore W. Burkhardt. First-passage and extreme-value statistics of a particle subject to a constant force plus a random force. Journal of Statistical Physics, 133(2):217–230, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9615-y.

Bobylev:2008:CDK

[BV08] A. V. Bobylev and M. C. Vinerean. Construction of discrete kinetic models with given invariants. Journal of Statistical Physics, 132(1):153–170, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9536-9.

Bena:2001:MFC

[BVBF01] Cristina Bena, Smitha Vishveshwara, Leon Balents, and Matthew P. A. Fisher. Measuring fractional charge in carbon nanotubes. Journal of Statistical Physics, 103(3–4):429–440, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010376929353.

Bichsel:2007:SFP

[BW07a] Denis Bichsel and Peter Wittwer. Stationary flow past a semiinfinite flat plate: Analytical and numerical evidence for a symmetry-breaking solution. Journal of Statistical Physics, 127 (1):133–170, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9248-y.

Bishop:2007:MCS

[BW07b] Marvin Bishop and Paula A. Whitlock. Monte Carlo simulation of hard hyperspheres in six, seven and eight dimensions for low to moderate densities. Journal of Statistical Physics, 126(2): 299–314, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9266-9.

Baryshnikov:2003:GFR

[BY03] Yu. Baryshnikov and J. E. Yukich. Gaussian fields and random packing. Journal of Statistical Physics, 111(1–2):443–463, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022229713275.

Bovier:2000:SIA

[BZ00a] Anton Bovier and Milos Zahradník. A simple inductive approach to the problem of convergence of cluster expansions of polymer models. Journal of Statistical Physics, 100(3–4):765–778, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018631710626.

Bru:2000:CBW

[BZ00b] J.-B. Bru and V. A. Zagrebnov. On condensations in the Bogoliubov weakly imperfect Bose gas. Journal of Statistical Physics, 99(5–6):1297–1338, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018692823463.

Bennett:2003:DVF

[BZ03] Curtis D. Bennett and Craig L. Zirbel. Discrete velocity fields with explicitly computable Lagrangian law. Journal of Statistical Physics, 111(3–4):681–701, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022889929864.

Bertacchi:2004:CAR

[BZ04] Daniela Bertacchi and Fabio Zucca. Classification on the average of random walks. Journal of Statistical Physics, 114(3–4):

947–975, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000012513.55697.65.

Baake:2008:ASC

[BZ08a] Michael Baake and Natali Zint. Absence of singular continuous diffraction for discrete multi-component particle models. Journal of Statistical Physics, 130(4):727–740, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-007-9445-3.pdf.

Bru:2008:LDS

[BZ08b] J.-B. Bru and V. A. Zagrebnov. Large deviations in the superstable weakly imperfect Bose-gas. Journal of Statistical Physics, 133(2):379–400, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9593-0.

Bertacchi:2009:CCV

[BZ09] Daniela Bertacchi and Fabio Zucca. Characterization of critical values of branching random walks on weighted graphs through infinite-type branching processes. Journal of Statistical Physics, 134(1):53–65, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9653-5.

Chan:2002:DEG

[CAC02] Garnet Kin-Lic Chan, Paul W. Ayers, and Ernest S. Croot III. On the distribution of eigenvalues of grand canonical density matrices. Journal of Statistical Physics, 109(1–2):289–299, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1019999930923.

Caceres:2008:PTS

[Các08] Manuel O. Cáceres. Passage time statistics in a stochastic Verhulst model. Journal of Statistical Physics, 132(3):487–500, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9554-7.

Castro-Alvaredo:2009:BPE

[CAD09] Olalla A. Castro-Alvaredo and Benjamin Doyon. Bi-partite entanglement entropy in massive QFT with a boundary: the Ising model. Journal of Statistical Physics, 134(1):105–145, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9664-2.

Caglioti:2002:BMV

[Cag02] Emanuele Caglioti. Bipartite matching and Van der Waerden conjecture. Journal of Statistical Physics, 107(3–4):857–867, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014594331863.

Caillol:2004:SGT

[Cai04] Jean-Michel Caillol. Sine-Gordon theory for the equation of state of classical hard–core Coulomb systems. III. Loopwise expansion. Journal of Statistical Physics, 115(5–6):1461–1504, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000028066.25728.cf.

Caiazzo:2005:ALB

[Cai05] Alfonso Caiazzo. Analysis of lattice Boltzmann initialization routines. Journal of Statistical Physics, 121(1–2):37–48, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-7010-5.

Campostrini:2001:LCE

[Cam01] Massimo Campostrini. Linked-cluster expansion of the Ising model. Journal of Statistical Physics, 103(1–2):369–394, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004884006193.

Camia:2005:SLC

[Cam05] Federico Camia. Scaling limit and critical exponents for two-dimensional bootstrap percolation. Journal of Statistical Physics, 118(1–2):85–101, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-004-8778-4.

Camus:2006:SFS

[Cam06] Brice Camus. Spectral fluctuations of Schrödinger operators generated by critical points of the potential. Journal of Statistical Physics, 123(4):811–829, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9113-z.

Canizo:2007:CED

[Cañ07] José A. Cañizo. Convergence to equilibrium for the discrete coagulation–fragmentation equations with detailed balance. Journal of Statistical Physics, 129(1):1–26, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9373-2.

Cardy:2003:CSF

[Car03] John Cardy. Crossover scaling functions for 2 d vesicles, and the Yang–Lee edge singularity. Journal of Statistical Physics, 110 (3–6):519–526, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022195415768.

Cardy:2006:MA

[Car06] John Cardy. The O(n) model on the annulus. Journal of Statistical Physics, 125(1):1–21, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9186-8.

Carati:2007:ATH

[Car07] A. Carati. An averaging theorem for Hamiltonian dynamical systems in the thermodynamic limit. Journal of Statistical Physics, 128(4):1057–1077, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9332-y.

Cardy:2009:DHT

[Car09] John Cardy. Discrete holomorphicity at two-dimensional critical points. Journal of Statistical Physics, 137(5–6):814–824, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9870-6.

Castella:2001:NEQ

[Cas01] F. Castella. From the von Neumann equation to the quantum Boltzmann equation in a deterministic framework. Journal of Statistical Physics, 104(1–2):387–447, July 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010374114551.

Castella:2002:NEQ

[Cas02] F. Castella. From the von Neumann equation to the quantum Boltzmann equation II: Identifying the Born series. Journal of Statistical Physics, 106(5–6):1197–1220, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014098122698.

Causo:2002:CPA

[Cau02a] Maria Serena Causo. Cut-and-permute algorithm for selfavoiding walks in the presence of surfaces. Journal of Statistical Physics, 108(1–2):247–281, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015447704106. See er-

ratum [Cau02b].

Causo:2002:ECP

[Cau02b] Maria Serena Causo. Erratum: ‘Cut-and-Permute Algorithm for Self-Avoiding Walks in the Presence of Surfaces’ J. Statist. Phys. 108(1/2):252 (2002). Journal of Statistical Physics, 109 (1–2):343, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019943220480. See [Cau02a].

Coulomb:2004:CPE

[CB04] Stéphane Coulomb and Michel Bauer. On certain perturbations of the Erdős–Renyi random graph. Journal of Statistical Physics, 115(5–6):1251–1282, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028060.07802.9d.

Cessac:2004:SOC

[CBKM04] B. Cessac, Ph. Blanchard, T. Krüger, and J. L. Meunier. Selforganized criticality and thermodynamic formalism. Journal of Statistical Physics, 115(5–6):1283–1326, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028057.16662.89.

Cercignani:2000:SWD

[CC00a] Carlo Cercignani and Henri Cornille. Shock waves for a discrete velocity gas mixture. Journal of Statistical Physics, 99 (1–2):115–140, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018692522765.

Cornille:2000:CPD

[CC00b] Henri Cornille and Carlo Cercignani. A class of planar discrete velocity models for gas mixtures. Journal of Statistical Physics, 99(3–4):967–991, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018603831215.

Chang:2008:DCS

[CC08] Shu-Chiuan Chang and Lung-Chi Chen. Dimer coverings on the Sierpinski gasket. Journal of Statistical Physics, 131(4):631–650, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9516-0.

Chatterjee:2009:CLT

[CC09] Sourav Chatterjee and Nicholas Crawford. Central limit theorems for the energy density in the Sherrington–Kirkpatrick model. Journal of Statistical Physics, 137(4):639–666, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9865-3.

Cardy:2008:FFB

[CCAD08] J. L. Cardy, O. A. Castro-Alvaredo, and B. Doyon. Form factors of branch-point twist fields in quantum integrable models and entanglement entropy. Journal of Statistical Physics, 130(1): 129–168, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9422-x.

Caracciolo:2000:BDS

[CCF^{+}00] Sergio Caracciolo, Maria Serena Causo, Giovanni Ferraro, Mauro Papinutto, and Andrea Pelissetto. Bilocal dynamics for self-avoiding walks. Journal of Statistical Physics, 100(5–6): 1111–1145, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018719226557.

Carati:2004:DTF

[CCG04] A. Carati, P. Cipriani, and L. Galgani. On the definition of temperature in FPU systems. Journal of Statistical Physics, 115 (3–4):1101–1112, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000022378.52789.b6.

Chayes:2008:PDQ

[CCIL08] Lincoln Chayes, Nicholas Crawford, Dmitry Ioffe, and Anna Levit. The phase diagram of the quantum Curie–Weiss model. Journal of Statistical Physics, 133(1):131–149, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9608-x.

Costin:2004:TAS

[CCL04] O. Costin, R. D. Costin, and J. L. Lebowitz. Time asymptotics of the Schrödinger wave function in time-periodic potentials. Journal of Statistical Physics, 116(1–4):283–310, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000037244.42209.f7.

Carlen:2009:SCE

[CCL09] Eric A. Carlen, Maria C. Carvalho, and Xuguang Lu. On strong convergence to equilibrium for the Boltzmann equation with soft potentials. Journal of Statistical Physics, 135(4):681–736, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9741-1.

Cicuta:2000:NHT

[CCM00] G. M. Cicuta, M. Contedini, and L. Molinari. Non-Hermitian tridiagonal random matrices and returns to the origin of a random walk. Journal of Statistical Physics, 98(3–4):685–699, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018671308053.

Chernyak:2009:NET

[CCMT09] Vladimir Y. Chernyak, Michael Chertkov, Sergey V. Malinin, and Razvan Teodorescu. Non-equilibrium thermodynamics and topology of currents. Journal of Statistical Physics, 137(1): 109–147, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9832-z.

Chang:2007:STS

[CCY07] Shu-Chiuan Chang, Lung-Chi Chen, and Wei-Shih Yang. Spanning trees on the Sierpinski gasket. Journal of Statistical Physics, 126(3):649–667, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-006-9262-0.

Chabanol:2004:MSI

[CD04] Marie-Line Chabanol and Jean Duchon. Markovian solutions of inviscid Burgers equation. Journal of Statistical Physics, 114(1– 2):525–534, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000003120.32992.a9.

Conlon:2005:TWS

[CD05] Joseph G. Conlon and Charles R. Doering. On travelling waves for the stochastic Fisher–Kolmogorov–Petrovsky– Piscunov equation. Journal of Statistical Physics, 120(3–4): 421–477, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5960-2.

Chevalier:2008:BMS

[CD08] C. Chevalier and F. Debbasch. Is Brownian motion sensitive to geometry fluctuations? Journal of Statistical Physics, 131 (4):717–731, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9481-7.

Chabanol:2009:LPI

[CD09a] Marie-Line Chabanol and Jean Duchon. Lévy-Process intrinsic statistical solutions of a randomly forced Burgers equation. Journal of Statistical Physics, 136(6):1095–1104, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9824-z.

Charles:2009:SMR

[CD09b] Frédérique Charles and Laurent Desvillettes. Small mass ratio limit of Boltzmann equations in the context of the study of evolution of dust particles in a rarefied atmosphere. Journal of Statistical Physics, 137(3):539–567, November 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9858-2.

Coolen:2009:CMD

[CDA09] A. C. C. Coolen, A. De Martino, and A. Annibale. Constrained Markovian dynamics of random graphs. Journal of Statistical Physics, 136(6):1035–1067, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9821-2.

Castella:2006:DDC

[CDG06] F. Castella, P. Degond, and Th. Goudon. Diffusion dynamics of classical systems driven by an oscillatory force. Journal of Statistical Physics, 124(2–4):913–950, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9071-5.

Costin:2007:PUO

[CDG07] Ovidiu Costin, Percy Deift, and Dimitri Gioev. On the proof of universality for orthogonal and symplectic ensembles in random matrix theory. Journal of Statistical Physics, 129(5–6): 937–948, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9277-1.

Caffarelli:2005:IGS

[CdlL05] Luis A. Caffarelli and Rafael de la Llave. Interfaces of ground states in Ising models with periodic coefficients. Journal of Statistical Physics, 118(3–4):687–719, February 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8825-1.

Carvalho:2002:STA

[CdO02] Túlio O. Carvalho and César R. de Oliveira. Spectra and transport in almost periodic dimers. Journal of Statistical Physics, 107(5–6):1015–1030, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1015153523475.

Carvalho:2007:HQT

[CdO07] Túlio O. Carvalho and César R. de Oliveira. Hybrid quasicrystals, transport and localization in products of minimal sets. Journal of Statistical Physics, 127(6):1221–1235, June 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9299-8.

Collet:2002:NLG

[CE02a] P. Collet and J.-P. Eckmann. The number of large graphs with a positive density of triangles. Journal of Statistical Physics, 109(5–6):923–943, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020489507547.

Collet:2002:RUB

[CE02b] P. Collet and J.-P. Eckmann. A rigorous upper bound on the propagation speed for the Swift–Hohenberg and related equations. Journal of Statistical Physics, 108(5–6):1107–1124, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1019795631986.

Collet:2004:LMD

[CE04] P. Collet and J.-P. Eckmann. Liapunov multipliers and decay of correlations in dynamical systems. Journal of Statistical Physics, 115(1–2):217–254, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019817.71073.61.

Collet:2005:DT

[CE05] P. Collet and J.-P. Eckmann. Dynamics of triangulations. Journal of Statistical Physics, 121(5–6):1073–1081, December 2005.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-6985-2.

Chertkov:2003:P

[CEEH03] Misha Chertkov, Bob Ecke, Gregory Eyink, and Darryl Holm. Preface. Journal of Statistical Physics, 113(5–6):637–642, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027383632386.

Castella:2000:FPE

[CEFM00] François Castella, László Erdős, Florian Frommlet, and Peter A. Markowich. Fokker–Planck equations as scaling limits of reversible quantum systems. Journal of Statistical Physics, 100 (3–4):543–601, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018667323830.

Collet:2009:SHT

[CEMM09] Pierre Collet, Jean-Pierre Eckmann, and Carlos Mejía-Monasterio. Superdiffusive heat transport in a class of deterministic one-dimensional many-particle Lorentz gases. Journal of Statistical Physics, 136(2):331–347, July 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9783-4.

Costeniuc:2007:MCB

[CEO07] Marius Costeniuc, Richard S. Ellis, and Peter Tak-Hun Otto. Multiple critical behavior of probabilistic limit theorems in the neighborhood of a tricritical point. Journal of Statistical Physics, 127(3):495–552, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9290-4.

Cercignani:2001:SFG

[Cer01] C. Cercignani. Shear flow of a granular material. Journal of Statistical Physics, 102(5–6):1407–1415, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004804815471.

Cercignani:2005:BRK

[Cer05a] Carlo Cercignani. Book review: Kinetic theory of granular gases. Journal of Statistical Physics, 118(5–6):1263–1264, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-2116-8.

Cercignani:2005:GWS

[Cer05b] Carlo Cercignani. Global weak solutions of the Boltzmann equation. Journal of Statistical Physics, 118(1–2):333–342, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-8786-4.

Cercignani:2006:WSB

[Cer06] Carlo Cercignani. Weak solutions of the Boltzmann equation without angle cutoff. Journal of Statistical Physics, 123(4):753– 762, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9115-x.

Costeniuc:2005:GCE

[CETT05] Marius Costeniuc, Richard S. Ellis, Hugo Touchette, and Bruce Turkington. The generalized canonical ensemble and its universal equivalence with the microcanonical ensemble. Journal of Statistical Physics, 119(5–6):1283–1329, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-4407-0.

Cercueil:2003:SAF

[CF03] Alain Cercueil and Olivier François. Sharp asymptotics for fixation times in stochastic population genetics models at low mutation probabilities. Journal of Statistical Physics, 110(1–2): 311–332, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021078914200.

Camia:2006:SLG

[CFN06] Federico Camia, Luiz Renato G. Fontes, and Charles M. Newman. The scaling limit geometry of near-critical 2D percolation. Journal of Statistical Physics, 125(5–6):1155–1171, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-9014-6.

Caillol:2000:MCS

[CG00a] J. M. Caillol and D. Gilles. Monte Carlo simulations of the Yukawa one-component plasma. Journal of Statistical Physics, 100(5–6):933–947, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018727428374.

Caillol:2000:NSS

[CG00b] J. M. Caillol and D. Gilles. Numerical simulations of screened Coulomb systems. A comparison between hyperspherical and periodic boundary conditions. Journal of Statistical Physics, 100(5–6):905–932, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018775311536.

Claus:2000:MCR

[CG00c] I. Claus and P. Gaspard. Microscopic chaos and reaction– diffusion processes in the periodic Lorentz gas. Journal of Statistical Physics, 101(1–2):161–186, October 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026447129361.

Cressman:2003:CFT

[CG03] J. R. Cressman and W. I. Goldburg. Compressible flow: Turbulence at the surface. Journal of Statistical Physics, 113(5–6): 875–883, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1027368920999.

Contucci:2004:MTL

[CG04] Pierluigi Contucci and Sandro Graffi. Monotonicity and thermodynamic limit for short range disordered models. Journal of Statistical Physics, 115(1–2):581–589, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019812.03696.b7.

Contucci:2007:GGI

[CG07] Pierluigi Contucci and Cristian Giardinà. The Ghirlanda– Guerra identities. Journal of Statistical Physics, 126(4–5):

917–931, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9091-1.

Chetrite:2009:ELP

[CG09] Raphaël Chetrite and Krzysztof Gawedzki. Eulerian and Lagrangian pictures of non-equilibrium diffusions. Journal of Statistical Physics, 137(5–6):890–916, December 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9803-4.

Caravenna:2006:NAC

[CGG06] Francesco Caravenna, Giambattista Giacomin, and Massimiliano Gubinelli. A numerical approach to copolymers at selective interfaces. Journal of Statistical Physics, 122(4):799–832, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8081-z.

Contucci:2009:IFI

[CGG09] Pierluigi Contucci, Cristian Giardinà, and Claudio Giberti. Interaction-flip identities in spin glasses. Journal of Statistical Physics, 135(5–6):1181–1203, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9706-4.

Caracciolo:2004:FSS

[CGGP04] Sergio Caracciolo, Andrea Gambassi, Massimiliano Gubinelli, and Andrea Pelissetto. Finite-size scaling in the driven lattice gas. Journal of Statistical Physics, 115(1–2):281–322, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019824.34397.1d.

Chaves:2003:LDG

[CGH^{+}03] Marta Chaves, Krzysztof Gawedzki, Peter Horvai, Antti Kupiainen, and Massimo Vergassola. Lagrangian dispersion in Gaussian self-similar velocity ensembles. Journal of Statistical Physics, 113(5–6):643–692, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1027348316456.

Contucci:2002:MFB

[CGI02] Pierluigi Contucci, Sandro Graffi, and Stefano Isola. Mean field behaviour of spin systems with orthogonal interaction matrix. Journal of Statistical Physics, 106(5–6):895–914, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014073417247.

Caracciolo:2005:CSE

[CGJ^{+}05] Sergio Caracciolo, Anthony J. Guttmann, Iwan Jensen, Andrea Pelissetto, Andrew N. Rogers, and Alan D. Sokal. Correction-to-scaling exponents for two-dimensional self-avoiding walks. Journal of Statistical Physics, 120(5–6):1037–1100, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-7004-3.

Chechkin:2004:LFS

[CGK^{+}04] Aleksei V. Chechkin, Vsevolod Yu. Gonchar, Joseph Klafter, Ralf Metzler, and Leonid V. Tanatarov. Lévy flights in a steep potential well. Journal of Statistical Physics, 115(5–6): 1505–1535, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000028067.63365.04.

Combes:2009:GEC

[CGK09] Jean-Michel Combes, François Germinet, and Abel Klein. Generalized eigenvalue-counting estimates for the Anderson model. Journal of Statistical Physics, 135(2):201–216, April 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9731-3.pdf.

Contucci:2007:P

[CGL07] Pierluigi Contucci, Sandro Graffi, and Joel L. Lebowitz. Preface. Journal of Statistical Physics, 126(4–5):729, March 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9256-y.

Carlen:2007:REI

[CGR07] Eric Carlen, Ester Gabetta, and Eugenio Regazzini. On the rate of explosion for infinite energy solutions of the spatially homogeneous Boltzmann equation. Journal of Statistical Physics, 129 (4):699–723, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9403-0.

Chan:2007:CKI

[CGS07] Cliburn Chan, Andrew J. T. George, and Jaroslav Stark. Cytotoxic Killing and immune evasion by repair. Journal of Statistical Physics, 128(1–2):393–411, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9242-4.

Comets:2005:ISK

[CGT05] Francis Comets, Francesco Guerra, and Fabio Lucio Toninelli. The Ising–Sherrington–Kirkpatrick model in a magnetic field at high temperature. Journal of Statistical Physics, 120(1–2):147– 165, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-5471-1.

Collet:2002:SRL

[CGV02] Jean-François Collet, Thierry Goudon, and Alexis Vasseur. Some remarks on large-time asymptotic of the Lifshitz–Slyozov equations. Journal of Statistical Physics, 108(1–2):341–359, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1015404021853.

Canet:2006:SSA

[CH06] L. Canet and H. J. Hilhorst. Single-site approximation for reaction–diffusion processes. Journal of Statistical Physics, 125 (3):517–531, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9206-8.

Chazottes:2000:DWT

[Cha00] Jean-René Chazottes. Dimensions and waiting times for Gibbs measures. Journal of Statistical Physics, 98(1–2):305–320, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018683024003.

Chakraborty:2009:SDE

[Cha09] Paramita Chakraborty. A stochastic differential equation model with jumps for fractional advection and dispersion. Journal of Statistical Physics, 136(3):527–551, August 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9794-1.

Chen:2005:LLB

[Che05] Thomas Chen. Localization lengths and Boltzmann limit for the Anderson model at small disorders in dimension 3. Journal of Statistical Physics, 120(1–2):279–337, July 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5255-7.

Chernov:2006:ASP

[Che06] N. Chernov. Advanced statistical properties of dispersing billiards. Journal of Statistical Physics, 122(6):1061–1094, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9036-8.

Chernov:2007:SEB

[Che07] N. Chernov. A stretched exponential bound on time correlations for billiard flows. Journal of Statistical Physics, 127(1):21–50, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9293-1.

Chen:2009:INC

[Che09a] Lizhou Chen. Inversion number and collisions in some billiard systems. Journal of Statistical Physics, 137(2):331–342, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9846-6.

Chernov:2009:NST

[Che09b] N. Chernov. Numerical studies of a two-dimensional Navier– Stokes system with new boundary conditions. Journal of Statistical Physics, 135(4):751–761, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9759-4.

Conti:2006:NLV

[CHNO06] Sergio Conti, Andreas Hönig, Barbara Niethammer, and Felix Otto. Nonuniversality in low-volume-fraction Ostwald ripening. Journal of Statistical Physics, 124(1):231–259, July 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9154-3.

Cirillo:2002:NMI

[Cir02] Emilio N. M. Cirillo. A note on the metastability of the Ising model: The alternate updating case. Journal of Statistical Physics, 106(1–2):385–390, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013144717688.

Cercignani:2001:DEG

[CIS01] Carlo Cercignani, Reinhard Illner, and Cristina Stoica. On diffusive equilibria in generalized kinetic theory. Journal of Statistical Physics, 105(1–2):337–352, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1012246513712.

Camia:2009:TCN

[CJM09] Federico Camia, Matthijs Joosten, and Ronald Meester. Trivial, critical and near-critical scaling limits of two-dimensional percolation. Journal of Statistical Physics, 137(1):57–69, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9841-y.pdf.

Cardy:2001:UCS

[CJS01] John Cardy, Jesper Lykke Jacobsen, and Alan D. Sokal. Unusual corrections to scaling in the 3-state Potts antiferromagnet on a square lattice. Journal of Statistical Physics, 105(1–2):25–47, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012273825098.

Chang:2004:EPM

[CJSS04] Shu-Chiuan Chang, Jesper Lykke Jacobsen, Jesús Salas, and Robert Shrock. Exact Potts model partition functions for strips of the triangular lattice. Journal of Statistical Physics, 114(3– 4):763–823, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000012508.58718.83.

Cercignani:2000:TED

[CK00] C. Cercignani and G. M. Kremer. Trend to equilibrium of a degenerate relativistic gas. Journal of Statistical Physics, 98(1– 2):441–456, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018695426728.

Cerny:2003:IRC

[CK03a] J. Cerný and R. Kotecký. Interfaces for random cluster models. Journal of Statistical Physics, 111(1–2):73–106, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022248822844.

Courbage:2003:IRA

[CK03b] M. Courbage and B. Kamiński. On intrinsically random Z^{2}- actions on a Lebesgue space. Journal of Statistical Physics, 112 (1–2):421–427, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023600325382.

Ciccotti:2004:SDL

[CK04] Giovanni Ciccotti and Galina Kalibaeva. Simulation of a diatomic liquid using hard spheres model. Journal of Statistical Physics, 115(1–2):701–714, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019823.05483.ae.

Cohen:2006:FPD

[CK06] Elisheva Cohen and David A. Kessler. Front propagation dynamics with exponentially-distributed hopping. Journal of Statistical Physics, 122(5):925–948, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-9004-8.

Chialvo:2000:AIE

[CKKC00] A. A. Chialvo, P. G. Kusalik, Yu. V. Kalyuzhnyi, and P. T. Cummings. Applications of integral equation calculations to high-temperature solvation phenomena. Journal of Statistical Physics, 100(1–2):167–199, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018639728431.

Callaghan:2006:SMW

[CKSZ06] Thomas Callaghan, Evgeniy Khain, Leonard M. Sander, and Robert M. Ziff. A stochastic model for wound healing. Journal of Statistical Physics, 122(5):909–924, March 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9022-1.

Carmona:2000:LDE

[CL00] S. C. Carmona and A. Lopes. Large deviations for expanding transformations with additive white noise. Journal of Statistical Physics, 98(5–6):1311–1333, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018675914395.

Chernov:2002:DMP

[CL02] N. Chernov and J. L. Lebowitz. Dynamics of a massive piston in an ideal gas: Oscillatory motion and approach to equilibrium. Journal of Statistical Physics, 109(3–4):507–527, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020450228657.

Carlen:2003:FSC

[CL03] Eric A. Carlen and Xuguang Lu. Fast and slow convergence to equilibrium for Maxwellian molecules via wild sums. Journal of Statistical Physics, 112(1–2):59–134, July 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023623503092.

Chayes:2006:RCM

[CL06] L. Chayes and H. K. Lei. Random cluster models on the triangular lattice. Journal of Statistical Physics, 122(4):647–670, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8078-7.

Chayes:2007:ODN

[CL07] Lincoln Chayes and Thomas M. Liggett. One dimensional nearest neighbor exclusion processes in inhomogeneous and random environments. Journal of Statistical Physics, 129(2):193–203, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9397-7.

Choquet:2005:HLA

[CLD05] Isabelle Choquet and Brigitte Lucquin-Desreux. Hydrodynamic limit for an arc discharge at atmospheric pressure. Journal of Statistical Physics, 119(1–2):197–239, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2711-8.

Chai:2009:TMM

[CLK09] Yan Chai, Reinhard Lipowsky, and Stefan Klumpp. Transport by molecular motors in the presence of static defects. Journal of Statistical Physics, 135(2):241–260, April 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9715-3.pdf.

Collina:2005:LDA

[CLM05] R. Collina, R. Livi, and A. Mazzino. Large deviation approach to the randomly forced Navier–Stokes equation. Journal of Statistical Physics, 118(3–4):451–479, February 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8817-1.

Chayes:2007:PPL

[CLM07] L. Chayes, J. L. Lebowitz, and V. Marinov. Percolation phenomena in low and high density systems. Journal of Statistical Physics, 129(3):567–585, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9408-8.

Camassa:2008:EPM

[CLM08] Roberto Camassa, Zhi Lin, and Richard M. McLaughlin. Evolution of the probability measure for the Majda model: New invariant measures and breathing PDFs. Journal of Statistical Physics, 130(2):343–371, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9427-5.

Chernov:2002:SDM

[CLS02] N. Chernov, J. L. Lebowitz, and Ya. Sinai. Scaling dynamics of a massive piston in a cube filled with ideal gas: Exact results. Journal of Statistical Physics, 109(3–4):529–548, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020402312727.

Constantin:2007:SLB

[CLT07] Peter Constantin, Boris Levant, and Edriss S. Titi. Sharp lower bounds for the dimension of the global attractor of the Sabra shell model of turbulence. Journal of Statistical Physics, 127 (6):1173–1192, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9317-x.

Caglioti:2000:BGV

[CM00] Emanuele Caglioti and Carlo Marchioro. Bounds on the growth of the velocity support for the solutions of the Vlasov–Poisson equation in a torus. Journal of Statistical Physics, 100(3–4): 659–677, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018623508809.

Caglioti:2002:LTB

[CM02a] E. Caglioti and C. Marchioro. On the long time behavior of a particle in an infinitely extended system in one dimension. Journal of Statistical Physics, 106(3–4):663–680, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013766507734.

Chopard:2002:LBS

[CM02b] Bastien Chopard and Stefan Marconi. Lattice Boltzmann solid particles in a lattice Boltzmann fluid. Journal of Statistical Physics, 107(1–2):23–37, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014542116996.

Ciliberti:2004:QCA

[CM04a] Stefano Ciliberti and Enzo Marinari. A quantitative clustering approach to ultrametricity in spin glasses. Journal of Statistical Physics, 115(1–2):557–580, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019831.87525.db.

Clisby:2004:ACH

[CM04b] N. Clisby and B. M. McCoy. Analytic calculation of B_{4} for hard spheres in even dimensions. Journal of Statistical Physics, 114(5–6):1343–1361, March 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013959.30878.d2.

Clisby:2004:NVC

[CM04c] N. Clisby and B. M. McCoy. Negative virial coefficients and the dominance of loose packed diagrams for D-dimensional hard spheres. Journal of Statistical Physics, 114(5–6):1361–1392, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000013960.83555.7d.

Catarino:2005:RQS

[CM05] Nuno R. Catarino and Robert S. MacKay. Renormalization and quantum scaling of Frenkel–Kontorova models. Journal of Statistical Physics, 121(5–6):995–1014, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8671-9.

Clisby:2006:NTO

[CM06] Nathan Clisby and Barry M. McCoy. Ninth and tenth order virial coefficients for hard spheres in d dimensions. Journal of Statistical Physics, 122(1):15–57, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8080-0.

Contucci:2006:STN

[CMN06] Pierluigi Contucci, Satoshi Morita, and Hidetoshi Nishimori. Surface terms on the Nishimori line of the Gaussian Edwards– Anderson model. Journal of Statistical Physics, 122(2):303–312, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8020-z.

Carles:2004:SAW

[CMS04] Rémi Carles, Peter A. Markowich, and Christof Sparber. Semiclassical asymptotics for weakly nonlinear Bloch waves. Journal of Statistical Physics, 117(1–2):343–375, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044070.34410.17.

Cavallaro:2005:DIM

[CMS05] G. Cavallaro, C. Marchioro, and C. Spitoni. Dynamics of infinitely many particles mutually interacting in three dimensions via a bounded superstable long-range potential. Journal of Statistical Physics, 120(1–2):367–416, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-5466-y.

Coninck:2005:SMS

[CMSR05] J. De. Coninck, S. Miracle-Solé, and J. Ruiz. On the statistical mechanics and surface tensions of binary mixtures. Journal of Statistical Physics, 119(3–4):597–642, May 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3251-6.

Celani:2005:PTR

[CMV05] A. Celani, S. Musacchio, and D. Vincenzi. Polymer transport in random flow. Journal of Statistical Physics, 118(3–4):531–554, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8820-6.

Cuansing:2001:IFC

[CN01] E. Cuansing and H. Nakanishi. Iterated fully coordinated percolation on a square lattice. Journal of Statistical Physics, 105(3– 4):659–675, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012284027322.

Cirillo:2003:MSD

[CN03] Emilio N. M. Cirillo and Francesca R. Nardi. Metastability for a stochastic dynamics with a parallel heat bath updating rule. Journal of Statistical Physics, 110(1–2):183–217, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021070712382.

Camia:2004:CNL

[CN04a] Federico Camia and Charles M. Newman. Continuum nonsimple loops and 2D critical percolation. Journal of Statistical Physics, 116(1–4):157–173, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037221.31328.75.

Camia:2004:PTZ

[CN04b] Federico Camia and Charles M. Newman. The percolation transition in the zero-temperature domany model. Journal of Statistical Physics, 114(5–6):1199–1210, March 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013965.36344.75.

Cirillo:2008:MRP

[CNS08] Emilio N. M. Cirillo, Francesca R. Nardi, and Cristian Spitoni. Metastability for reversible probabilistic cellular automata with self-interaction. Journal of Statistical Physics, 132(3):431–471, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9563-6.

Comtet:2001:MUH

[CNV01] Alain Comtet, Sergei Nechaev, and Raphaël Voituriez. Multifractality in uniform hyperbolic lattices and in quasi-classical Liouville field theory. Journal of Statistical Physics, 102(1–2): 203–230, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026568810945.

Cammarota:2008:CPM

[CO08a] V. Cammarota and E. Orsingher. Cascades of particles moving at finite velocity in hyperbolic spaces. Journal of Statistical Physics, 133(6):1137–1159, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9648-2.

Cammarota:2008:TRP

[CO08b] V. Cammarota and E. Orsingher. Travelling randomly on the Poincaré half-plane with a Pythagorean compass. Journal of Statistical Physics, 130(3):455–482, February 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9430-x.

Cohen:2009:NED

[Coh09] E. G. D. Cohen. Non existence of a density expansion of the transport coefficients in a moderately dense gas. Journal of Statistical Physics, 137(5–6):878–889, December 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

009-9838-6. See erratum [Coh10].

Cohen:2010:ENE

[Coh10] E. G. D. Cohen. Erratum to: Non existence of a density expansion of the transport coefficients in a moderately dense gas. Journal of Statistical Physics, 138(1–3):551, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9851-9.pdf. See [Coh09].

Conlon:2004:PRC

[Con04] Joseph G. Conlon. PDE with random coefficients and Euclidean field theory. Journal of Statistical Physics, 116(1–4):933–958, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037204.93858.f2.

Coppersmith:2008:URG

[Cop08] S. N. Coppersmith. Using the renormalization group to classify Boolean functions. Journal of Statistical Physics, 130(6): 1063–1085, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9486-2.

Cornille:2004:DVM

[Cor04] H. Cornille. Discrete velocity models and one continuous relation. Journal of Statistical Physics, 114(3–4):977–1014, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012514.21906.4e.

Corngold:2005:TDR

[Cor05] Noel Corngold. Transport and diffusion in a random medium. Journal of Statistical Physics, 120(3–4):521–541, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-6801-z.

Coupier:2006:PAI

[Cou06] D. Coupier. Poisson approximations for the Ising model. Journal of Statistical Physics, 123(2):473–495, April 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9033-y.

Cassandro:2001:EPT

[CP01] Marzio Cassandro and Pierre Picco. Existence of a phase transition in a continuous quantum system. Journal of Statistical Physics, 103(5–6):841–856, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010352720562.

Chui:2002:FCS

[CP02] C. H. Otto Chui and Paul A. Pearce. Finitized conformal spectra of the Ising model on the Klein bottle and Möbius strip. Journal of Statistical Physics, 107(5–6):1167–1205, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015113909363.

Chatterji:2003:SMS

[CP03] Apratim Chatterji and Rahul Pandit. The statistical mechanics of semiflexible equilibrium polymers. Journal of Statistical Physics, 110(3–6):1219–1248, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022109313467.

Cipriani:2004:OSA

[CP04] P. Cipriani and A. Politi. An open-system approach for the characterization of spatio-temporal chaos. Journal of Statistical Physics, 114(1–2):205–228, January 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003110.15959.d3.

Casetti:2003:PTT

[CPC03] Lapo Casetti, Marco Pettini, and E. G. D. Cohen. Phase transitions and topology changes in configuration space. Journal of Statistical Physics, 111(5–6):1091–1123, June 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023044014341.

Candia:2007:TPR

[CPK07] Julián Candia, Paul E. Parris, and V. M. Kenkre. Transport properties of random walks on Scale–Free/ regular-lattice hybrid networks. Journal of Statistical Physics, 129(2):323–333, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9362-5.

Cordier:2009:MMF

[CPP09] Stephane Cordier, Lorenzo Pareschi, and Cyrille Piatecki. Mesoscopic modelling of financial markets. Journal of Statistical Physics, 134(1):161–184, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9667-z.

Cordier:2005:KMS

[CPT05] Stephane Cordier, Lorenzo Pareschi, and Giuseppe Toscani. On a kinetic model for a simple market economy. Journal of Statistical Physics, 120(1–2):253–277, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-5456-0.

Comets:2008:NOP

[CPV08] Francis Comets, Serguei Popov, and Marina Vachkovskaia. The number of open paths in an oriented ρ-percolation model. Journal of Statistical Physics, 131(2):357–379, April 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9506-2.

Caillol:2001:SGT

[CR01] Jean-Michel Caillol and Jean-Luc Raimbault. Sine-Gordon theory for the equation of state of classical hard-core Coulomb systems. I. Low fugacity expansion. Journal of Statistical Physics, 103(5–6):753–776, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010396502815.

Choksi:2003:DDF

[CR03] Rustum Choksi and Xiaofeng Ren. On the derivation of a density functional theory for microphase separation of diblock copolymers. Journal of Statistical Physics, 113(1–2):151–176, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1025722804873.

Ren:2004:VED

[cR04] Hai cang Ren. The virial expansion of a dilute Bose gas in two dimensions. Journal of Statistical Physics, 114(1–2):481–501, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003117.00478.77.

Chazottes:2005:ORM

[CR05] J.-R. Chazottes and F. Redig. Occurrence, repetition and matching of patterns in the low-temperature Ising model. Journal of Statistical Physics, 121(3–4):579–605, November 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7575-z.

Caglioti:2007:QSS

[CR07] E. Caglioti and F. Rousset. Quasi-stationary states for particle systems in the mean-field limit. Journal of Statistical Physics, 129(2):241–263, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9390-1.

Croydon:2009:RWR

[Cro09] David A. Croydon. Random walk on the range of random walk. Journal of Statistical Physics, 136(2):349–372, July 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9785-2.

Cerda:2000:ADV

[CRT00] Enrique Cerda, René Rojas, and Enrique Tirapegui. Asymptotic description of a viscous fluid layer. Journal of Statistical Physics, 101(1–2):553–565, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026411510531.

Coullet:2000:SUE

[CRV00] P. Coullet, E. Risler, and N. Vandenberghe. Spatial unfolding of elementary bifurcations. Journal of Statistical Physics, 101(1– 2):521–541, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026415607690.

Connelly:2001:PLT

[CRV01] Robert Connelly, Konstantin Rybnikov, and Stanislav Volkov. Percolation of the loss of tension in an infinite triangular lattice. Journal of Statistical Physics, 105(1–2):143–171, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1012282026916.

Cessac:2009:HGD

[CRVV09] B. Cessac, H. Rostro, J. C. Vasquez, and T. Viéville. How Gibbs distributions May naturally arise from synaptic adaptation mechanisms. A model-based argumentation. Journal of Statistical Physics, 136(3):565–602, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9786-1.

Carlen:2002:OCC

[CS02] Eric A. Carlen and Francesco Salvarani. On the optimal choice of coefficients in a truncated wild sum and approximate solutions for the Kac equation. Journal of Statistical Physics, 109(1–2): 261–277, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019943813176.

Chang:2003:FPT

[CS03a] Shu-Chiuan Chang and Robert Shrock. Flow polynomials and their asymptotic limits for lattice strip graphs. Journal of Statistical Physics, 112(3–4):815–879, August 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023836311251.

Chang:2003:RPT

[CS03b] Shu-Chiuan Chang and Robert Shrock. Reliability polynomials and their asymptotic limits for families of graphs. Journal of Statistical Physics, 112(5–6):1019–1077, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024663508526.

Cuesta:2004:GNE

[CS04] José A. Cuesta and Angel Sánchez. General non-existence theorem for phase transitions in one-dimensional systems with short range interactions, and physical examples of such transitions. Journal of Statistical Physics, 115(3–4):869–893, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000022373.63640.4e. See erratum [CS09d].

Chang:2008:EPM

[CS08a] Shu-Chiuan Chang and Robert Shrock. Exact Potts model partition functions for strips of the honeycomb lattice. Journal of Statistical Physics, 130(5):1011–1024, March 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9461-3.

Chen:2008:BLQ

[CS08b] Thomas Chen and Itaru Sasaki. Boltzmann limit and quasifreeness for a homogeneous Fermi gas in a weakly disordered random medium. Journal of Statistical Physics, 132(2):329–353, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9560-9.

Cichocki:2008:SDB

[CS08c] Bogdan Cichocki and Krzysztof Sadlej. Stokesian dynamics — the BBGKY hierarchy for correlation functions. Journal of Statistical Physics, 132(1):129–151, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9542-y.

Caracciolo:2009:SFR

[CS09a] Sergio Caracciolo and Andrea Sportiello. Spanning forests on random planar lattices. Journal of Statistical Physics, 135(5– 6):1063–1104, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9733-1.

Chang:2009:SPF

[CS09b] Shu-Chiuan Chang and Robert Shrock. Structure of the partition function and transfer matrices for the Potts model in a magnetic field on lattice strips. Journal of Statistical Physics, 137(4): 667–699, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9868-0.

Contucci:2009:TLS

[CS09c] Pierluigi Contucci and Shannon Starr. Thermodynamic limit for spin glasses. Beyond the annealed bound. Journal of Statistical Physics, 135(5–6):1159–1166, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9676-y.

Cuesta:2009:EGN

[CS09d] José A. Cuesta and Angel Sánchez. Erratum to: General Non-Existence Theorem for Phase Transitions in One-Dimensional Systems with Short Range Interactions, and Physical Examples of Such Transitions. Journal of Statistical Physics, 137(3):593– 594, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-009-9862-6.pdf. See [CS04].

Costa-Santos:2003:FSC

[CSM03] Ruben Costa-Santos and Barry M. McCoy. Finite size corrections for the Ising model on higher genus triangular lattices. Journal of Statistical Physics, 112(5–6):889–920, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1024697307618.

Chew:2002:NDL

[CSN02] Y. T. Chew, C. Shu, and X. D. Niu. A new differential lattice Boltzmann equation and its application to simulate incompressible flows on non-uniform grids. Journal of Statistical Physics, 107(1–2):329–342, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014579125174.

Csordas:2000:ILD

[Cso00] András Csordás. Integrable limits of dynamics in trapped Bosecondensates. Journal of Statistical Physics, 101(1–2):259–272, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026455331178.

Chew:2002:IBC

[CSP02] Y. T. Chew, C. Shu, and Y. Peng. On implementation of boundary conditions in the application of finite volume lattice Boltzmann method. Journal of Statistical Physics, 107(1–2):539–556, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014547712879.

Chang:2002:EPM

[CSS02] Shu-Chiuan Chang, Jesús Salas, and Robert Shrock. Exact Potts model partition functions for strips of the square lattice. Journal of Statistical Physics, 107(5–6):1207–1253, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015165926201.

Chayes:2000:DMD

[CSZ00] Lincoln Chayes, Senya B. Shlosman, and Valentin A. Zagrebnov. Discontinuity of the magnetization in diluted O(n)-models. Journal of Statistical Physics, 98(3–4):537–549, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018659005327.

Cancrini:2004:CFV

[CT04a] Nicoletta Cancrini and Christel Tremoulet. Comparison of finite volume canonical and grand canonical Gibbs measures: The continuous case. Journal of Statistical Physics, 117(5–6): 1023–1046, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-5714-6.

Crowdy:2004:EFS

[CT04b] Darren Crowdy and Saleh Tanveer. The effect of finiteness in the Saffman–Taylor viscous fingering problem. Journal of Statistical Physics, 114(5–6):1501–1536, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013962.78542.33.

Caceres:2007:KAL

[CT07] María J. Cáceres and Giuseppe Toscani. Kinetic approach to long time behavior of linearized fast diffusion equations. Journal of Statistical Physics, 128(4):883–925, August 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9329-6.

Chazottes:2003:PMM

[CU03] J.-R. Chazottes and E. Ugalde. Projection of Markov measures May be Gibbsian. Journal of Statistical Physics, 111(5– 6):1245–1272, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023056317067.

Choquard:2004:MFI

[CW04] Philippe Choquard and Joël Wagner. On the ‘Mean Field’ interpretation of Burgers’ equation. Journal of Statistical Physics, 116(1–4):843–853, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037211.80229.
04.

Cohen:2006:HEE

[CW06] Morrel H. Cohen and Adam Wasserman. On hardness and electronegativity equalization in chemical reactivity theory. Journal of Statistical Physics, 125(5–6):1121–1139, December 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9031-0.

Chen:2009:CCT

[CY09] Xinxing Chen and Qiang Yao. The complete convergence theorem holds for contact processes on open clusters of Z^{d} × Z^{+}. Journal of Statistical Physics, 135(4):651–680, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9756-7.

Cardy:2003:ERU

[CZ03] John Cardy and Robert M. Ziff. Exact results for the universal area distribution of clusters in percolation, Ising, and Potts models. Journal of Statistical Physics, 110(1–2):1–33, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1021069209656.

Chernov:2009:SPH

[CZ09] Nikolai Chernov and Hong-Kun Zhang. On statistical properties of hyperbolic systems with singularities. Journal of Statistical Physics, 136(4):615–642, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9804-3.

Ding:2003:ELB

[DA03] E-Jiang Ding and Cyrus K. Aidun. Extension of the lattice-Boltzmann method for direct simulation of suspended particles near contact. Journal of Statistical Physics, 112(3–4):685–708, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023880126272.

Dagdug:2003:BRBb

[Dag03a] Leonardo Dagdug. Book review: Molecular Driving Forces: Statistical Thermodynamics in Chemistry and Biology. Ken A. Dill and Sarina Bromberg, Garland Science, New York, 2003. Journal of Statistical Physics, 113(5–6):885–887, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1027321005069.

Dagdug:2003:BRBa

[Dag03b] Leonardo Dagdug. Book review: Statistical Mechanics: A Concise Introduction for Chemists. Benjamin Widom, Cambridge University Press, Cambridge, 2002. Journal of Statistical Physics, 112(1–2):433–434, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023604426290.

deAguiar:2004:IEM

[dARBY04] M. A. M. de Aguiar, E. M. Rauch, and Y. Bar-Yam. Invasion and extinction in the mean field approximation for a spatial host–pathogen model. Journal of Statistical Physics, 114(5–6): 1417–1451, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013958.15218.47.

Dhara:2000:SUP

[DB00] Asish K. Dhara and S. R. Banerjee. Synchronization under periodic modulation of symmetric double square Wells in a bistable stochastic system. Journal of Statistical Physics, 99(3–4):799– 812, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018691428489.

Dhara:2001:NCS

[DB01] Asish K. Dhara and S. R. Banerjee. New characterization of stochastic resonance in bistable square potential well. Journal of Statistical Physics, 105(5–6):771–787, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013592910118.

Dufty:2002:GKE

[DB02] James W. Dufty and J. Javier Brey. Green–Kubo expressions for a granular gas. Journal of Statistical Physics, 109(3–4): 433–448, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020489910002.

Boer:2008:LSD

[dBR08] Anne Fey den Boer and Frank Redig. Limiting shapes for deterministic centrally seeded growth models. Journal of Statistical Physics, 130(3):579–597, February 2008. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9450-6.

Dettmann:2000:MCD

[DC00] C. P. Dettmann and E. G. D. Cohen. Microscopic chaos and diffusion. Journal of Statistical Physics, 101(3–4):775–817, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026477605331.

Dettmann:2001:NCD

[DC01] C. P. Dettmann and E. G. D. Cohen. Note on chaos and diffusion. Journal of Statistical Physics, 103(3–4):589–599, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010345417058.

DaPrato:2004:ACI

[DD04] Giuseppe Da Prato and Arnaud Debussche. Absolute continuity of the invariant measures for some stochastic PDEs. Journal of Statistical Physics, 115(1–2):451–468, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019816.42914.19.

Dreyer:2006:BDT

[DD06] Wolfgang Dreyer and Frank Duderstadt. On the Becker/ Döring theory of nucleation of liquid droplets in solids. Journal of Statistical Physics, 123(1):55–87, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9024-z.

Datta:2009:CCQ

[DD09] Nilanjana Datta and Tony Dorlas. Classical capacity of quantum channels with general Markovian correlated noise. Journal of Statistical Physics, 134(5–6):1173–1195, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9726-0.

DeKepper:2000:RDP

[DDB^{+}00] P. De Kepper, E. Dulos, J. Boissonade, A. De Wit, G. Dewel, and P. Borckmans. Reaction-diffusion patterns in confined chemical systems. Journal of Statistical Physics, 101(1–2):495–508, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026462105253.

DeConinck:2006:NWS

[DDH06] Joël De Coninck, François Dunlop, and Thierry Huillet. A necklace of Wulff shapes. Journal of Statistical Physics, 123 (1):223–236, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8019-5.

DeConinck:2008:RWW

[DDH08] Joël De Coninck, François Dunlop, and Thierry Huillet. Random walk weakly attracted to a wall. Journal of Statistical Physics, 133(2):271–280, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9609-9.

Derezinski:2008:FQC

[DDM08] Jan Dereziński, Wojciech De Roeck, and Christian Maes. Fluctuations of quantum currents and unravelings of master equations. Journal of Statistical Physics, 131(2):341–356, April 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9500-8.

DeConinck:2004:WHS

[DDMSR04] Joël De Coninck, Christophe Dobrovolny, Salvador Miracle-Solé, and Jean Ruiz. Wetting of heterogeneous surfaces at the mesoscopic scale. Journal of Statistical Physics, 114(3–4):575–604, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012503.98210.67.

Derrida:2004:CFO

[DDR04] B. Derrida, B. Douçot, and P.-E. Roche. Current fluctuations in the one-dimensional symmetric exclusion process with open boundaries. Journal of Statistical Physics, 115(3–4):717–748, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022379.95508.b2.

DeSanctis:2004:RMO

[De 04] Luca De Sanctis. Random multi-overlap structures and cavity fields in diluted spin glasses. Journal of Statistical Physics, 117(5–6):785–799, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-5704-8.

DeSanctis:2006:UBR

[De 06] Luca De Sanctis. Ultrametric broken replica symmetry RaMOSt. Journal of Statistical Physics, 122(5):857–874, March 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-9008-4.

DeSanctis:2007:SPD

[De 07] Luca De Sanctis. Structural properties of the disordered spherical and other mean field spin models. Journal of Statistical Physics, 126(4–5):817–835, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9167-y.

Derrida:2004:AEP

[DEL04] B. Derrida, C. Enaud, and J. L. Lebowitz. The asymmetric exclusion process and Brownian excursions. Journal of Statistical Physics, 115(1–2):365–382, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019833.35328.b4.

Dellar:2005:LKF

[Del05] Paul J. Dellar. Lattice kinetic formulation for ferrofluids. Journal of Statistical Physics, 121(1–2):105–118, October 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8056-0.

Derrida:2005:FWA

[DELO05] B. Derrida, C. Enaud, C. Landim, and S. Olla. Fluctuations in the weakly asymmetric exclusion process with open boundary conditions. Journal of Statistical Physics, 118(5–6):795–811, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-1989-x.

Dereudre:2008:GDT

[Der08] David Dereudre. Gibbs Delaunay tessellations with geometric hardcore conditions. Journal of Statistical Physics, 131 (1):127–151, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9479-6.

DeMasi:2002:FFO

[DF02] A. De Masi and P. A. Ferrari. Flux fluctuations in the one dimensional nearest neighbors symmetric simple exclusion process. Journal of Statistical Physics, 107(3–4):677–683, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014577928229.

Costa:2006:HCC

[dFCS06] Luciano da Fontoura Costa and Filipi Nascimento Silva. Hierarchical characterization of complex networks. Journal of Statistical Physics, 125(4):841–872, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9130-y.

Dunlop:2002:DOD

[DFF02] F. M. Dunlop, P. A. Ferrari, and L. R. G. Fontes. A dynamic one-dimensional interface interacting with a wall. Journal of Statistical Physics, 107(3–4):705–727, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1014755529138.

Dehling:2008:PRT

[DFK08] H. G. Dehling, S. R. Fleurke, and C. Külske. Parking on a random tree. Journal of Statistical Physics, 133(1):151–157, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-008-9589-9.pdf.

Deguchi:2001:LAS

[DFM01] Tetsuo Deguchi, Klaus Fabricius, and Barry M. McCoy. The sl2 loop algebra symmetry of the six-vertex model at roots of unity. Journal of Statistical Physics, 102(3–4):701–736, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004894701900.

DeBievre:2000:DLR

[DG00] Stephan De Bièvre and François Germinet. Dynamical localization for the random dimer Schrödinger operator. Journal of Statistical Physics, 98(5–6):1135–1148, March 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018615728507.

Dufty:2001:MDG

[DG01] James W. Dufty and Vicente Garzó. Mobility and diffusion in granular fluids. Journal of Statistical Physics, 105(5–6): 723–744, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013545908301.

deGennes:2005:BMD

[dG05a] P.-G de Gennes. Brownian motion with dry friction. Journal of Statistical Physics, 119(5–6):953–962, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

005-4650-4. See erratum [dG05b].

deGennes:2005:EBM

[dG05b] P.-G. de Gennes. Errata: Brownian motion with dry friction. Journal of Statistical Physics, 119(5–6):1421, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

005-7319-0. See [dG05a].

DeSanctis:2008:MFD

[DG08] Luca De Sanctis and Francesco Guerra. Mean field dilute ferromagnet: High temperature and zero temperature behavior. Journal of Statistical Physics, 132(5):759–785, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9575-2.

Derrida:2009:CFOb

[DG09a] Bernard Derrida and Antoine Gerschenfeld. Current fluctuations in one dimensional diffusive systems with a step initial density profile. Journal of Statistical Physics, 137(5–6):978–1000, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9830-1.

Derrida:2009:CFOa

[DG09b] Bernard Derrida and Antoine Gerschenfeld. Current fluctuations of the one dimensional symmetric simple exclusion process with step initial condition. Journal of Statistical Physics, 136(1):1– 15, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9772-7.

Deift:2007:UOS

[DGKV07] P. Deift, D. Gioev, T. Kriecherbauer, and M. Vanlessen. Universality for orthogonal and symplectic Laguerre-type ensembles. Journal of Statistical Physics, 129(5–6):949–1053, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9325-x.

DiCastro:2004:I

[DGM04] Carlo Di Castro, Francesco Guerra, and Fabio Martinelli. Introduction. Journal of Statistical Physics, 115(1–2):3–6, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000020050.81014.86.

Dembo:2007:LDSa

[DGM07a] Amir Dembo, Alice Guionnet, and Christian Mazza. Limiting dynamics for spherical models of spin glasses at high temperature. Journal of Statistical Physics, 126(4–5):781–815, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9228-2.

Dembo:2007:LDSb

[DGM07b] Amir Dembo, Alice Guionnet, and Christian Mazza. Limiting dynamics for spherical models of spin glasses at high temperature. Journal of Statistical Physics, 128(4):847–881, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9239-z.

deGier:2004:RPM

[dGNPR04] Jan de Gier, Bernard Nienhuis, Paul A. Pearce, and Vladimir Rittenberg. The raise and peel model of a fluctuating interface. Journal of Statistical Physics, 114(1–2):1–35, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000003102.81727.fd.

Desvillettes:2008:MFL

[DGR08] Laurent Desvillettes, François Golse, and Valeria Ricci. The mean-field limit for solid particles in a Navier–Stokes flow. Journal of Statistical Physics, 131(5):941–967, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9521-3.

Durr:2004:QER

[DGZ04] Detlef Dürr, Sheldon Goldstein, and Nino Zanghì. Quantum equilibrium and the role of operators as observables in quantum theory. Journal of Statistical Physics, 116(1–4):959–1055, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037234.80916.d0.

Durr:2009:WMV

[DGZ09] Detlef Dürr, Sheldon Goldstein, and Nino Zanghì. On the weak measurement of velocity in Bohmian mechanics. Journal of Statistical Physics, 134(5–6):1023–1032, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9674-0.

Dowker:2004:SCM

[DH04] Fay Dowker and Joe Henson. Spontaneous collapse models on a lattice. Journal of Statistical Physics, 115(5–6):1327–1339, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028061.97843.84.

Dermoune:2008:SGM

[DH08] Azzouz Dermoune and Philippe Heinrich. Spectral gap for multicolor nearest-neighbor exclusion processes with site disorder.

Journal of Statistical Physics, 131(1):117–125, April 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9496-0.

Drasdo:2007:RPG

[DHB07] Dirk Drasdo, Stefan Hoehme, and Michael Block. On the role of physics in the growth and pattern formation of multi-cellular systems: What can we learn from individual-cell based models? Journal of Statistical Physics, 128(1–2):287–345, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9289-x.

denHollander:2009:LPC

[dHP09] F. den Hollander and N. Pétrélis. On the localized phase of a copolymer in an emulsion: Subcritical percolation regime. Journal of Statistical Physics, 134(2):209–241, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-008-9663-3.pdf.

denHollander:2004:DHM

[dHW04] Frank den Hollander and Mario V. Wüthrich. Diffusion of a heteropolymer in a multi-interface medium. Journal of Statistical Physics, 114(3–4):849–889, February 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012510.81452.4a.

Dimock:2009:IVL

[Dim09] J. Dimock. Infinite volume limit for the dipole gas. Journal of Statistical Physics, 135(3):393–427, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9739-8.

Dirr:2004:SPS

[Dir04] Nicolas Dirr. A Stefan problem with surface tension as the sharp interface limit of a nonlocal system of phase-field type. Journal of Statistical Physics, 114(3–4):1085–1113, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012517.33719.9f.

Derezinski:2004:NFG

[DJ04] Jan Dereziński and Vojkan Jaksić. On the nature of Fermi golden rule for open quantum systems. Journal of Statistical Physics, 116(1–4):411–423, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037208.99352.0a.

Dai:2005:HDZ

[DJ05] Xiongping Dai and Yunping Jiang. Hausdorff dimensions of zeroentropy sets of dynamical systems with positive entropy. Journal of Statistical Physics, 120(3–4):511–519, August 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7008-z.

Derzhko:2007:CSP

[DJ07] V. Derzhko and J. Jȩdrzejewski. Charge-stripe phases versus a weak anisotropy of nearest-neighbor hopping. Journal of Statistical Physics, 126(3):467–505, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9245-1.

Durhuus:2007:SDG

[DJW07] Bergfinnur Durhuus, Thordur Jonsson, and John F. Wheater. The spectral dimension of generic trees. Journal of Statistical Physics, 128(5):1237–1260, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9348-3.

Derrida:2000:LED

[DJZ00] Bernard Derrida, Jesper Lykke Jacobsen, and Reuven Zeitak. Lyapunov exponent and density of states of a one-dimensional non–Hermitian Schrödinger equation. Journal of Statistical Physics, 98(1–2):31–55, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018666620368.

Datta:2002:EOQ

[DK02] Nilanjana Datta and Tom Kennedy. Expansions for one quasiparticle states in spin 1/2 systems. Journal of Statistical Physics, 108(3–4):373–399, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015713005441.

Dolgopyat:2002:HDS

[DKK02] D. Dolgopyat, V. Kaloshin, and L. Koralov. Hausdorff dimension in stochastic dispersion. Journal of Statistical Physics, 108(5–6): 943–971, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019731212422.

Dudnikova:2004:TTP

[DKM04] T. V. Dudnikova, A. I. Komech, and N. J. Mauser. On two-temperature problem for harmonic crystals. Journal of Statistical Physics, 114(3–4):1035–1083, February 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012516.89488.20.

Dudnikova:2002:CED

[DKRS02] T. V. Dudnikova, A. I. Komech, N. E. Ratanov, and Y. M. Suhov. On convergence to equilibrium distribution, II. The wave equation in odd dimensions, with mixing. Journal of Statistical Physics, 108(5–6):1219–1253, September 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019755917873.

delaLlave:2007:CPI

[dlLV07] Rafael de la Llave and Enrico Valdinoci. Critical points inside the gaps of ground state laminations for some models in statistical mechanics. Journal of Statistical Physics, 129(1):81–119, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9345-6.

DeCarvalho:2005:RHF

[DLM05] A. De Carvalho, M. Lyubich, and M. Martens. Renormalization in the Hénon family, I: Universality but non-rigidity. Journal of Statistical Physics, 121(5–6):611–669, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8668-4.

Ding:2009:CGD

[DLP09] Jian Ding, Eyal Lubetzky, and Yuval Peres. Censored Glauber dynamics for the mean field Ising model. Journal of Statistical Physics, 137(3):407–458, November 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9859-1.pdf.

Dobrovolny:2004:STSa

[DLR04a] C. Dobrovolny, L. Laanait, and J. Ruiz. Surface transitions of the semi-infinite Potts model I: The high bulk temperature regime. Journal of Statistical Physics, 114(5–6):1269–1302, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000013957.89983.81.

Dobrovolny:2004:STSb

[DLR04b] C. Dobrovolny, L. Laanait, and J. Ruiz. Surface transitions of the semi-infinite Potts model II: The low bulk temperature regime. Journal of Statistical Physics, 116(5–6):1405– 1434, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000041744.83013.5d.

Derrida:2002:LDD

[DLS02] B. Derrida, J. L. Lebowitz, and E. R. Speer. Large deviation of the density profile in the steady state of the open symmetric simple exclusion process. Journal of Statistical Physics, 107 (3–4):599–634, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014555927320.

Derrida:2003:ELD

[DLS03] B. Derrida, J. L. Lebowitz, and E. R. Speer. Exact large deviation functional of a stationary open driven diffusive system: The asymmetric exclusion process. Journal of Statistical Physics, 110 (3–6):775–810, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022111919402.

Derrida:2007:EOL

[DLS07] B. Derrida, J. L. Lebowitz, and E. R. Speer. Entropy of open lattice systems. Journal of Statistical Physics, 126(4–5):1083–1108, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9160-5.

Duclos:2008:EGS

[DLS08] Pierre Duclos, Ondra Lev, and Pavel Stovícek. On the energy growth of some periodically driven quantum systems with shrinking gaps in the spectrum. Journal of Statistical Physics, 130(1):169–193, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9419-5.

Dinaburg:2009:NBP

[DLS09] Efim Dinaburg, Dong Li, and Yakov G. Sinai. A new boundary problem for the two dimensional Navier–Stokes system. Journal of Statistical Physics, 135(4):737–750, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9760-y.

Dhara:2002:CSR

[DM02] Asish K. Dhara and Tapan Mukhopadhyay. Coherent stochastic resonance in one dimensional diffusion with one reflecting and one absorbing boundaries. Journal of Statistical Physics, 107 (3–4):685–703, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014530012299.

DeSantis:2006:SPP

[DM06a] Emilio De Santis and Rossella Micieli. Slab percolation and phase transitions for the Ising model. Journal of Statistical Physics, 122(1):59–72, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-8086-7.

Dermoune:2006:AMD

[DM06b] Azzouz Dermoune and Servet Martinez. Around multicolour disordered lattice gas. Journal of Statistical Physics, 123(1): 181–192, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-9012-8.

Dorea:2006:ADI

[DM06c] Chang C. Y. Dorea and Ary V. Medino. Anomalous diffusion index for Lévy motions. Journal of Statistical Physics, 123(3):685– 698, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9074-2.

Degond:2008:LSD

[DM08] Pierre Degond and Sébastien Motsch. Large scale dynamics of the persistent turning Walker model of fish behavior. Journal of Statistical Physics, 131(6):989–1021, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9529-8.

Deng:2009:HMS

[DM09] Xinghua Deng and Robert V. Moody. How model sets can be determined by their two-point and three-point correlations. Journal of Statistical Physics, 135(4):621–637, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9742-0.

Destainville:2001:FBO

[DMB01] N. Destainville, R. Mosseri, and F. Bailly. Fixed-boundary octagonal random tilings: A combinatorial approach. Journal of Statistical Physics, 102(1–2):147–190, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026564710037.

deMasi:2004:LVI

[dMG04] A. de Masi and T. Gobron. Liquid-vapor interfaces and surface tension in a mesoscopic model of fluid with nonlocal interactions. Journal of Statistical Physics, 115(1–2):643–679, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000019834.85465.a2.

Datta:2000:RIF

[DMN00] Nilanjana Datta, Alain Messager, and Bruno Nachtergaele. Rigidity of interfaces in the Falicov–Kimball model. Journal of Statistical Physics, 99(1–2):461–555, April 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018609126399.

DeRoeck:2006:TMA

[DMN06] Wojciech De Roeck, Christian Maes, and Karel Netocný. Htheorems from macroscopic autonomous equations. Journal of Statistical Physics, 123(3):571–584, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9079-x.

Dinaburg:2004:PMB

[DMP^{+}04] E. Dinaburg, C. Maes, S. Pirogov, F. Redig, and A. Rybko. The Potts model built on sand. Journal of Statistical Physics, 117(1– 2):179–198, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000044067.96085.e2.

Dorlas:2005:LCP

[DMP05] Teunis C. Dorlas, Philippe A. Martin, and Joseph V. Pule. Long cycles in a perturbed mean field model of a Boson gas. Journal of Statistical Physics, 121(3–4):433–461, November 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7582-0.

DeMasi:2008:PMC

[DMPV08] Anna De Masi, Immacolata Merola, Errico Presutti, and Yvon Vignaud. Potts models in the continuum. Uniqueness and exponential decay in the restricted ensembles. Journal of Statistical Physics, 133(2):281–345, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9603-2.

DeMasi:2009:COD

[DMPV09] Anna De Masi, Immacolata Merola, Errico Presutti, and Yvon Vignaud. Coexistence of ordered and disordered phases in Potts models in the continuum. Journal of Statistical Physics, 134 (2):243–306, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9677-x.

Degond:2005:QET

[DMR05] Pierre Degond, Florian Méhats, and Christian Ringhofer. Quantum energy-transport and drift-diffusion models. Journal of Statistical Physics, 118(3–4):625–667, February 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8823-3.

DSouza:2002:DSS

[DMS02] Raissa M. D’Souza, Norman H. Margolus, and Mark A. Smith. Dimension-splitting for simplifying diffusion in lattice-gas models. Journal of Statistical Physics, 107(1–2):401–422, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014587326991.

DeConinck:2000:TOS

[DMSR00] J. De Coninck, S. Miracle-Solé, and J. Ruiz. Is there an optimal substrate geometry for wetting? Journal of Statistical Physics, 100(5–6):981–997, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018745122923.

DeConinck:2003:RGY

[DMSR03] J. De Coninck, S. Miracle-Solé, and J. Ruiz. Rigorous generalization of Young’s law for heterogeneous and rough substrates. Journal of Statistical Physics, 111(1–2):107–127, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022200906915.

DeSantis:2003:CEL

[DN03] Emilio De Santis and Charles M. Newman. Convergence in energy-lowering (disordered) stochastic spin systems. Journal of Statistical Physics, 110(1–2):431–442, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021039200087.

Dachian:2004:DSM

[DN04] S. Dachian and B. S. Nahapetian. Description of specifications by means of probability distributions in small volumes under condition of very weak positivity. Journal of Statistical Physics, 117(1–2):281–300, October 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044069.91072.0b.

Dogbe:2000:ADL

[Dog00a] Christian Dogbé. Anomalous diffusion limit induced on a kinetic equation. Journal of Statistical Physics, 100(3–4):603–632, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018619407900.

Dogbe:2000:ABL

[Dog00b] Christian Dogbé. Asymptotic behavior for the Liouville equations. Journal of Statistical Physics, 99(3–4):873–902, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018647613468.

Domany:2003:CAG

[Dom03a] Eytan Domany. Cluster analysis of gene expression data. Journal of Statistical Physics, 110(3–6):1117–1139, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022148927580.

Domb:2003:SOE

[Dom03b] Cyril Domb. Some observations on the early history of equilibrium statistical mechanics. Journal of Statistical Physics, 110 (3–6):475–496, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022139230789.

Domb:2004:BRS

[Dom04] Cyril Domb. Book review: The Scientific Letters and Papers of James Clerk Maxwell. P. M. Harman, ed. Vol. III, Cambridge University Press, Cambridge. Journal of Statistical Physics, 116(5–6):1703–1706, September 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041912.31649.5a.

Dombry:2005:PSD

[Dom05] Clément Dombry. A probabilistic study of DNA denaturation. Journal of Statistical Physics, 120(3–4):695–719, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-5962-0.

deOliveira:2008:MJA

[dOP08] César R. de Oliveira and Marciano Pereira. Mathematical justification of the Aharonov–Bohm Hamiltonian. Journal of Statistical Physics, 133(6):1175–1184, December 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9631-y.

Dorfman:2002:BRB

[Dor02] J. R. Dorfman. Book review: Lattice Gas Hydrodynamics. Jean-Pierre Rivet and Jean Pierre Boon, Cambridge University Press, 2001, ISBN 0-521-41944-1, Cambridge Nonlinear Science Series 11. Journal of Statistical Physics, 107(1–2):593–595, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014503930626.

Dodson:2001:BNT

[DOS01] C. T. J. Dodson, Y. Oba, and W. W. Sampson. Bivariate normal thickness-density structure in real near-planar stochastic fiber networks. Journal of Statistical Physics, 102(1–2):345–353, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026529229601.

Dotsenko:2004:IRI

[Dot04] Victor Dotsenko. Infinite range interaction model of a structural glass. Journal of Statistical Physics, 115(3–4):823–837, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022368.39876.1c.

Dotsenko:2006:GSR

[Dot06] Victor Dotsenko. Griffiths singularity in the random Ising ferromagnet. Journal of Statistical Physics, 122(2):197–216, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8079-6.

Descombes:2003:DSC

[DP03] Xavier Descombes and Eugène Pechersky. Droplet shapes for a class of models in Z^{2} at zero temperature. Journal of Statistical Physics, 111(1–2):129–169, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022252923753.

DellAntonio:2004:FAS

[DP04] G. F. Dell’Antonio and G. Panati. The flux-across-surfaces theorem and zero-energy resonances. Journal of Statistical Physics, 116(1–4):1161–1180, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037235.96575.
dc.

Dairbekov:2007:EPT

[DP07a] Nurlan S. Dairbekov and Gabriel P. Paternain. Entropy production in thermostats II. Journal of Statistical Physics, 127 (5):887–914, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9301-5.

Deryabin:2007:NGG

[DP07b] Mikhail V. Deryabin and Lev D. Pustyl’nikov. Nonequilibrium gas and generalized billiards. Journal of Statistical Physics, 126 (1):117–132, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9250-4.

Dorlas:2007:FTC

[DP07c] T. C. Dorlas and V. B. Priezzhev. Finite-time current probabilities in the asymmetric exclusion process on a ring. Journal of Statistical Physics, 129(4):787–805, November 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9406-x.

dePasquale:2004:EAM

[dPG04] F. de Pasquale and S. M. Giampaolo. Expansion around the mean field in quantum magnetic systems. Journal of Statistical Physics, 115(1–2):125–142, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019837.56894.8a.

Dorlas:2009:VWT

[DPP09] T. C. Dorlas, A. M. Povolotsky, and V. B. Priezzhev. From vicious walkers to TASEP. Journal of Statistical Physics, 135 (3):483–517, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9748-7.

Dorlas:2006:CDI

[DPZ06] T. C. Dorlas, L. A. Pastur, and V. A. Zagrebnov. Condensation in a disordered infinite-range hopping Bose–Hubbard model. Journal of Statistical Physics, 124(5):1137–1178, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9176-x.

Desvillettes:2001:RDL

[DR01] L. Desvillettes and V. Ricci. A rigorous derivation of a linear kinetic equation of Fokker–Planck type in the limit of grazing collisions. Journal of Statistical Physics, 104(5–6):1173– 1189, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010461929872.

Datta:2003:NGF

[DR03a] Sandip Datta and Ramakrishna Ramaswamy. Non-Gaussian fluctuations of local Lyapunov exponents at intermittency. Journal of Statistical Physics, 113(1–2):283–295, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025783023529.

Degond:2003:QMH

[DR03b] P. Degond and C. Ringhofer. Quantum moment hydrodynamics and the entropy principle. Journal of Statistical Physics, 112(3– 4):587–628, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023824008525.

Dereudre:2005:PGI

[DR05] David Dereudre and Sylvie Roelly. Propagation of Gibbsianness for infinite-dimensional gradient Brownian diffusions. Journal of Statistical Physics, 121(3–4):511–551, November 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7580-2.

Dhar:2006:HTH

[DR06a] Abhishek Dhar and Dibyendu Roy. Heat transport in harmonic lattices. Journal of Statistical Physics, 125(4):801–820, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9235-3.

Dubbeldam:2006:MMC

[DR06b] J. Dubbeldam and F. Redig. Multilayer Markov chains with applications to polymers in shear flow. Journal of Statistical Physics, 125(1):225–243, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9166-z.

DiCastro:2004:RGW

[DRC04] C. Di Castro, R. Raimondi, and S. Caprara. Renormalization group and Ward identities in quantum liquid phases and in unconventional critical phenomena. Journal of Statistical Physics, 115(1–2):91–123, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000019832.91097.0b.

Degenhard:2002:TER

[DRL02] Andreas Degenhard and Javier Rodríguez-Laguna. Towards the evaluation of the relevant degrees of freedom in nonlinear partial differential equations. Journal of Statistical Physics, 106(5–6): 1093–1120, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014041904951.

Droz:2000:RTD

[Dro00] Michel Droz. Recent theoretical developments on the formation of Liesegang patterns. Journal of Statistical Physics, 101(1–2): 509–519, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026489416640.

Dodson:2007:PLP

[DS07] C. T. J. Dodson and W. W. Sampson. Planar line processes for void and density statistics in thin stochastic fibre networks. Journal of Statistical Physics, 129(2):311–322, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9379-9.

DeFabritiis:2002:ESC

[DSC02] G. De Fabritiis, S. Succi, and P. V. Coveney. Electronic structure calculations using self-adaptive multiscale Voronoi basis functions. Journal of Statistical Physics, 107(1–2):159–171, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014558520631.

Filho:2002:DCE

[dSFM02] Abel G. da Silva Filho and F. G. Brady Moreira. Dynamical critical exponent for the majority-vote model. Journal of Statistical Physics, 106(1–2):391–401, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013196701758.

deSanctis:2004:MAI

[dSI04] L. de Sanctis and M. Isopi. On the ‘Matrix Approach’ to interacting particle systems. Journal of Statistical Physics, 115 (1–2):383–393, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000019828.19669.37.

dosSantos:2005:DID

[dSWP05] Luís O. E. dos Santos, Fabiano G. Wolf, and Paulo C. Philippi. Dynamics of interface displacement in capillary flow. Journal of Statistical Physics, 121(1–2):197–207, October 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7001-6.

Dong:2007:TMP

[DSZ07] J. J. Dong, B. Schmittmann, and R. K. P. Zia. Towards a model for protein production rates. Journal of Statistical Physics, 128 (1–2):21–34, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9134-7.

Dunlop:2000:CLC

[DT00] François Dunlop and Krzysztof Topolski. Cassie’s law and concavity of wall tension with respect to disorder. Journal of Statistical Physics, 98(5–6):1115–1134, March 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018663611669.

Dauxois:2002:TIO

[DTP02] Thierry Dauxois, Nikos Theodorakopoulos, and Michel Peyrard. Thermodynamic instabilities in one dimension: Correlations, scaling and solitons. Journal of Statistical Physics, 107(3– 4):869–891, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014546415934.

Denniston:2002:DMC

[DTY02] Colin Denniston, Géza Tóth, and J. M. Yeomans. Domain motion in confined liquid crystals. Journal of Statistical Physics, 107(1–2):187–202, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014562721540.

Dubedat:2006:EIC

[Dub06] Julien Dubédat. Euler integrals for commuting SLEs. Journal of Statistical Physics, 123(6):1183–1218, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9132-9.

Dudko:2007:SME

[Dud07] Olga K. Dudko. Statistical mechanics: Entropy, order parameters, and complexity. Journal of Statistical Physics, 126(2): 429–430, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9263-z.

Dunlop:2003:SSS

[Dun03] François Dunlop. Stationary states and scaling shapes of onedimensional interfaces. Journal of Statistical Physics, 111(1– 2):433–442, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022277629205.

Duplantier:2003:HCM

[Dup03] Bertrand Duplantier. Higher conformal multifractality. Journal of Statistical Physics, 110(3–6):691–738, March 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022107818494.

DeVille:2007:NSL

[DVE07] R. E. Lee DeVille and Eric Vanden-Eijnden. A nontrivial scaling limit for multiscale Markov chains. Journal of Statistical Physics, 126(1):75–94, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9237-1.

David:2005:ICS

[DW05] François David and Kay J. Wiese. Instanton calculus for the self-avoiding manifold model. Journal of Statistical Physics, 120(5–6):875–1035, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-5253-9.

Destainville:2005:RTH

[DWMB05] N. Destainville, M. Widom, R. Mosseri, and F. Bailly. Random tilings of high symmetry: I. Mean-field theory. Journal of Statistical Physics, 120(5–6):799–835, September 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-6989-y.

Dyson:2004:EWS

[Dys04] Freeman Dyson. Elliott’s world: From square ice to cubic jellium. Journal of Statistical Physics, 116(1–4):3–8, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://labs.adsabs.harvard.edu/ui/
abs/2004JSP...116....3D;   http://link.springer.com/
article/10.1023/B%3AJOSS.0000037240.68210.60.

Dysman:2005:FDR

[Dys05] Michelle Dysman. Fractal dimensions for repellers of maps with holes. Journal of Statistical Physics, 120(3–4):479–509, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-5964-y.

Ding:2001:CAM

[DZ01] Jiu Ding and Aihui Zhou. Constructive approximations of Markov operators. Journal of Statistical Physics, 105(5–6): 863–878, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013553227865.

deZarate:2004:POL

[dZS04] José M. Ortiz de Zárate and Jan V. Sengers. On the physical origin of long-ranged fluctuations in fluids in thermal nonequilibrium states. Journal of Statistical Physics, 115(5–6):1341–1359, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028062.57459.52.

Erb:2004:MIT

[EA04] Ionas Erb and Nihat Ay. Multi-information in the thermodynamic limit. Journal of Statistical Physics, 115(3–4):949–976, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022375.49904.ea.

Etienne:2007:NCT

[EA07] Rampal S. Etienne and David Alonso. Neutral community theory: How stochasticity and dispersal-limitation can explain species coexistence. Journal of Statistical Physics, 128(1–2):485– 510, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9163-2.

Ernst:2002:SSI

[EB02] M. H. Ernst and R. Brito. Scaling solutions of inelastic Boltzmann equations with over-populated high energy tails. Journal of Statistical Physics, 109(3–4):407–432, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020437925931.

Erban:2007:CPS

[EC07] Radek Erban and S. Jonathan Chapman. On chemisorption of polymers to solid surfaces. Journal of Statistical Physics, 127 (6):1255–1277, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9224-6.

Eyink:2003:GHT

[ECC03] Gregory L. Eyink, Shiyi Chen, and Qiaoning Chen. Gibbsian hypothesis in turbulence. Journal of Statistical Physics, 113(5– 6):719–740, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1027304501435.

Eckmann:2007:TG

[Eck07] Jean-Pierre Eckmann. A topological glass. Journal of Statistical Physics, 129(2):289–309, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9387-9.

Evans:2000:NKY

[ECSB00] Denis J. Evans, E. G. D. Cohen, Debra J. Searles, and F. Bonetto. Note on the Kaplan–Yorke dimension and linear transport coefficients. Journal of Statistical Physics, 101(1– 2):17–34, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026449702528.

Enaud:2004:LDF

[ED04] C. Enaud and B. Derrida. Large deviation functional of the weakly asymmetric exclusion process. Journal of Statistical Physics, 114(3–4):537–562, February 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012501.43746.cf.

Edwards:2004:NKE

[Edw04] S. F. Edwards. New kinds of entropy. Journal of Statistical Physics, 116(1–4):29–42, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037233.36686.
2f.

Erdem:2007:TPC

[EE07] Riza Erdem and Cesur Ekiz. The transport properties of the cell membrane ion channels in electric fields: Bethe lattice treatment. Journal of Statistical Physics, 129(3):469–481, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9370-5.

Ebeling:2000:NDF

[EEDJ00] W. Ebeling, U. Erdmann, J. Dunkel, and M. Jenssen. Nonlinear dynamics and fluctuations of dissipative Toda chains. Journal of Statistical Physics, 101(1–2):443–457, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026407415248.

Evans:2009:MRS

[EFM09] Martin R. Evans, Pablo A. Ferrari, and Kirone Mallick. Matrix representation of the stationary measure for the multispecies TASEP. Journal of Statistical Physics, 135(2):217–239, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9696-2.

Eckmann:2005:LMH

[EFPZ05] Jean-Pierre Eckmann, Christina Forster, Harald A. Posch, and Emmanuel Zabey. Lyapunov modes in hard-disk systems. Journal of Statistical Physics, 118(5–6):813–847, March 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2687-4.

Eckmann:2000:HLM

[EG00] Jean-Pierre Eckmann and Omri Gat. Hydrodynamic Lyapunov modes in translation-invariant systems. Journal of Statistical Physics, 98(3–4):775–798, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018679609870.

Esposito:2005:ESM

[EG05] M. Esposito and P. Gaspard. Exactly solvable model of quantum diffusion. Journal of Statistical Physics, 121(3–4):463– 496, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-7577-x.

Esposti:2001:SEL

[EGGI01] M. Degli Esposti, C. Giardinà, S. Graffi, and S. Isola. Statistics of energy levels and zero temperature dynamics for deterministic spin models with glassy behaviour. Journal of Statistical Physics, 102(5–6):1285–1313, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004844429584.

Ellis:2000:LDP

[EHT00] Richard S. Ellis, Kyle Haven, and Bruce Turkington. Large deviation principles and complete equivalence and nonequivalence results for pure and mixed ensembles. Journal of Statistical Physics, 101(5–6):999–1064, December 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026446225804.

Erneux:2000:NVH

[EK00] Thomas Erneux and Gregory Kozyreff. Nearly vertical Hopf bifurcation for a passively Q-switched microchip laser. Journal of Statistical Physics, 101(1–2):543–552, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026482609796.

Eizenberg:2002:LDP

[EK02] A. Eizenberg and Y. Kifer. Large deviations for probabilistic cellular automata. Journal of Statistical Physics, 108(5–6): 1255–1280, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019708001944.

Eliazar:2003:LDL

[EK03] Iddo Eliazar and Joseph Klafter. Lévy-Driven Langevin systems: Targeted stochasticity. Journal of Statistical Physics, 111 (3–4):739–768, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022894030773.

Eizenberg:2004:LDP

[EK04] A. Eizenberg and Y. Kifer. Large deviations for probabilistic cellular automata II. Journal of Statistical Physics, 117(5–6):

845–889, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-5708-4.

Eliazar:2005:AP

[EK05a] Iddo Eliazar and Joseph Klafter. Anomalous pulsation. Journal of Statistical Physics, 120(3–4):587–626, August 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5478-7.

Eliazar:2005:LOU

[EK05b] Iddo Eliazar and Joseph Klafter. Lévy, Ornstein–Uhlenbeck, and subordination: Spectral vs. jump description. Journal of Statistical Physics, 119(1–2):165–196, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2710-9.

Eliazar:2005:SOU

[EK05c] Iddo Eliazar and Joseph Klafter. Stochastic Ornstein–Uhlenbeck capacitors. Journal of Statistical Physics, 118(1–2):177–198, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8782-8.

Eyink:2006:MEM

[EK06] Gregory L. Eyink and Sangil Kim. A maximum entropy method for particle filtering. Journal of Statistical Physics, 123(5): 1071–1128, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9124-9.

Eliazar:2008:PPP

[EK08] Iddo Eliazar and Joseph Klafter. Paretian Poisson processes. Journal of Statistical Physics, 131(3):487–504, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9505-3.

E:2002:GDI

[EL02] Weinan E and Di Liu. Gibbsian dynamics and invariant measures for stochastic dissipative PDEs. Journal of Statistical Physics, 108(5–6):1125–1156, September 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019747716056.

Eloranta:2008:DPU

[Elo08] Kari Eloranta. Dense packing on uniform lattices. Journal of Statistical Physics, 130(4):741–755, February 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9448-0.

Elskens:2000:ABC

[Els00] Yves Elskens. Asymptotic behaviour for critical slowing-down random walks. Journal of Statistical Physics, 101(1–2):397–404, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026426524818.

Emch:2004:IM

[Emc04]  Gérard G. Emch.  Interactive modelling.  Journal of

Statistical Physics, 116(1–4):17–28, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037239.45958.99.

Ellison:2009:PRA

[EMC09] Christopher J. Ellison, John R. Mahoney, and James P. Crutchfield. Prediction, retrodiction, and the amount of information stored in the present. Journal of Statistical Physics, 136(6): 1005–1034, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9808-z.pdf.

Engel:2004:LDP

[EMH04] Andreas Engel, Rémi Monasson, and Alexander K. Hartmann. On large deviation properties of Erdős–Rényi random graphs. Journal of Statistical Physics, 117(3–4):387–426, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-2268-6.

Eckmann:2006:MEN

[EMMZ06] Jean-Pierre Eckmann, Carlos Mejía-Monasterio, and Emmanuel Zabey. Memory effects in nonequilibrium transport for deterministic Hamiltonian systems. Journal of Statistical Physics, 123 (6):1339–1360, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9153-4.

Ellis:2008:ABM

[EMO08] Richard S. Ellis, Jonathan Machta, and Peter Tak-Hun Otto. Asymptotic behavior of the magnetization near critical and tricritical points via Ginzburg–Landau polynomials. Journal of Statistical Physics, 133(1):101–129, October 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9606-z.

Evans:2006:CAC

[EMZ06] M. R. Evans, Satya N. Majumdar, and R. K. P. Zia. Canonical analysis of condensation in factorised steady states. Journal of Statistical Physics, 123(2):357–390, April 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9046-6.

Ebeling:2003:CPT

[EN03] W. Ebeling and G. Norman. Coulombic phase transitions in dense plasmas. Journal of Statistical Physics, 110(3–6):861–877, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022120121219.

Epstein:2006:ALN

[Eps06] Charles L. Epstein. Anderson localization, non-linearity and stable genetic diversity. Journal of Statistical Physics, 124(1):25– 46, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9149-0.

Eyink:2000:MPH

[ER00] G. L. Eyink and J. M. Restrepo. Most probable histories for nonlinear dynamics: Tracking climate transitions. Journal of Statistical Physics, 101(1–2):459–472, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026437432570.

Evans:2002:CEN

[ER02] Denis J. Evans and Lamberto Rondoni. Comments on the entropy of nonequilibrium steady states. Journal of Statistical Physics, 109(3–4):895–920, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1020435219996.

ElBouanani:2007:VMK

[ER07] H. El Bouanani and M. Rouleux. Vortices and magnetization in Kac’s model. Journal of Statistical Physics, 128(3):741–770, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9319-8.

Erdos:2002:LBE

[Erd02] László Erdős. Linear Boltzmann equation as the long time dynamics of an electron weakly coupled to a phonon field. Journal of Statistical Physics, 107(5–6):1043–1127, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015157624384.

Edwards:2003:SMC

[ES03] S. F. Edwards and Moshe Schwartz. Statistical mechanics in collective coordinates. Journal of Statistical Physics, 110(3– 6):497–502, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022191214859.

Erdos:2004:MLT

[ES04] László Erdős and Jan Philip Solovej. Magnetic Lieb–Thirring inequalities with optimal dependence on the field strength. Journal of Statistical Physics, 116(1–4):475–506, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037216.45270.1d.

Edelman:2007:RMS

[ES07] Alan Edelman and Brian D. Sutton. From random matrices to stochastic operators. Journal of Statistical Physics, 127(6): 1121–1165, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9226-4.

Erdos:2009:QDM

[ES09] László Erdős and Benjamin Schlein. Quantum dynamics with mean field interactions: a new approach. Journal of Statistical Physics, 134(5–6):859–870, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9570-7.

Escudero:2007:BHB

[Esc07] Carlos Escudero. Blow-up of the hyperbolic Burgers equation. Journal of Statistical Physics, 127(2):327–338, April 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9276-7.

Essam:2003:TAO

[Ess03] John W. Essam. Three attractive osculating walkers and a polymer collapse transition. Journal of Statistical Physics, 110(3–6): 1191–1207, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022105112559.

Erdos:2004:QBE

[ESY04] László Erdős, Manfred Salmhofer, and Horng-Tzer Yau. On the quantum Boltzmann equation. Journal of Statistical Physics, 116(1–4):367–380, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037224.56191.
ed.

Ernst:2006:BED

[ETB06] M. H. Ernst, E. Trizac, and A. Barrat. The Boltzmann equation for driven systems of inelastic soft spheres. Journal of Statistical Physics, 124(2–4):549–586, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9062-6.

Ernst:2002:I

[EvBC02] M. H. Ernst, H. van Beijeren, and E. G. D. Cohen. Introduction. Journal of Statistical Physics, 109(3–4):355–359, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020481608184.

E:2006:TTT

[EVE06] Weinan E and Eric Vanden-Eijnden. Towards a theory of transition paths. Journal of Statistical Physics, 123(3):503–523, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-9003-9.

Ewodo:2001:RTR

[Ewo01] Jean Ndzié Ewodo. Refocusing of a time-reversed acoustic pulse propagating in randomly layered media. Journal of Statistical Physics, 104(5–6):1253–1272, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010466130781.

Eyink:2000:SSD

[EX00] Gregory L. Eyink and Jack Xin. Self-similar decay in the Kraichnan model of a passive scalar. Journal of Statistical Physics, 100 (3–4):679–741, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018675525647.

E:2001:CTE

[EY01] Weinan E and Nung Kwan Yip. Continuum theory of epitaxial crystal growth. I. Journal of Statistical Physics, 104(1–2):221– 253, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010361711825.

Eckmann:2004:SHF

[EZ04] J.-P. Eckmann and E. Zabey. Strange heat flux in (an)harmonic networks. Journal of Statistical Physics, 114(1–2):515–523, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003119.91989.48.

Fannjiang:2004:CPS

[Fan04] Albert C. Fannjiang. Convergence of passive scalar fields in Ornstein–Uhlenbeck flows to Kraichnan’s model. Journal of Statistical Physics, 114(1–2):115–135, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003106.02538.ae.

Fannjiang:2005:WNG

[Fan05] Albert C. Fannjiang. White-noise and geometrical optics limits of Wigner–Moyal equation for beam waves in turbulent media II: Two-frequency formulation. Journal of Statistical Physics, 120 (3–4):543–586, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5961-1.

Farago:2002:IPF

[Far02] Jean Farago. Injected power fluctuations in Langevin equation. Journal of Statistical Physics, 107(3–4):781–803, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014538214117.

Farago:2005:EPF

[Far05] Jean Farago. Energy profile fluctuations in dissipative nonequilibrium stationary states. Journal of Statistical Physics, 118(3– 4):373–405, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-8814-4.

Franceschetti:2005:CPU

[FBC^{+}05] Massimo Franceschetti, Lorna Booth, Matthew Cook, Ronald Meester, and Jehoshua Bruck. Continuum percolation with unreliable and spread-out connections. Journal of Statistical Physics, 118(3–4):721–734, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-8826-0.

Farnell:2002:CCM

[FBG02] D. J. J. Farnell, R. F. Bishop, and K. A. Gernoth. Coupled cluster method calculations of quantum magnets with spins of general spin quantum number. Journal of Statistical Physics, 108 (3–4):401–428, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015769622279.

Friedli:2006:TSN

[FdL06] S. Friedli and B. N. B. de Lima. On the truncation of systems with non-summable interactions. Journal of Statistical Physics, 122(6):1215–1236, March 2006. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-005-8023-9.

Ferrer:2007:CFH

[FEDZ07] A. Prats Ferrer, B. Eynard, P. Di Francesco, and J.-B. Zuber. Correlation functions of Harish-Chandra integrals over the orthogonal and the symplectic groups. Journal of Statistical Physics, 129(5–6):885–935, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9350-9.

Feigenbaum:2003:PSD

[Fei03] Mitchell J. Feigenbaum. Pattern selection: Determined by symmetry and modifiable by distant effects. Journal of Statistical Physics, 112(1–2):219–275, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023683721747.

Felderhof:2002:SBP

[Fel02] B. U. Felderhof. Sedimentation of Brownian particles in a gravitational potential. Journal of Statistical Physics, 109(3–4): 483–493, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020446127749.

Fan:2000:DLT

[FF00] Ai-Hua Fan and De-Jun Feng. On the distribution of long-term time averages on symbolic space. Journal of Statistical Physics, 99(3–4):813–856, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018643512559.

Fayolle:2004:DWR

[FF04] Guy Fayolle and Cyril Furtlehner. Dynamical windings of random walks and exclusion models. Part I: Thermodynamic limit in Z^{2}. Journal of Statistical Physics, 114(1–2):229–260, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003111.88829.9d.

Fayolle:2007:SDD

[FF07] Guy Fayolle and Cyril Furtlehner. Stochastic dynamics of discrete curves and multi-type exclusion processes. Journal of Statistical Physics, 127(5):1049–1094, June 2007. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9286-0.

Fernandez:2009:DPP

[FFN09] Roberto Fernández, Luiz R. Fontes, and E. Jordão Neves. Density-profile processes describing biological signaling networks: Almost sure convergence to deterministic trajectories. Journal of Statistical Physics, 136(5):875–901, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-009-9819-9.

Fulinski:2000:UCS

[FG00] A. Fuliński and P. F. Góra. Universal character of stochastic resonance and a constructive role of white noise. Journal of Statistical Physics, 101(1–2):483–493, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026463523619.

Fournier:2003:SPC

[FG03] Nicolas Fournier and Jean-Sébastien Giet. On small particles in coagulation–fragmentation equations. Journal of Statistical Physics, 111(5–6):1299–1329, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1023060417976.

Ferrari:2008:NPT

[FG08a] Pablo A. Ferrari and Sebastian P. Grynberg. No phase transition for Gaussian fields with bounded spins. Journal of Statistical Physics, 130(1):195–202, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9423-9.

Fournier:2008:USH

[FG08b] Nicolas Fournier and Hélène Guérin. On the uniqueness for the spatially homogeneous Boltzmann equation with a strong angular singularity. Journal of Statistical Physics, 131(4):749–781, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9511-5.

Faggionato:2009:NET

[FGC09] A. Faggionato, D. Gabrielli, and M. Ribezzi Crivellari. Nonequilibrium thermodynamics of piecewise deterministic Markov processes. Journal of Statistical Physics, 137(2):259–304, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9850-x.

Fishman:2003:TQA

[FGR03] Shmuel Fishman, Italo Guarneri, and Laura Rebuzzini. A theory for quantum accelerator modes in atom optics. Journal of Statistical Physics, 110(3–6):911–943, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022176306198.

Feldman:2004:NFM

[FH04] Mikhail Feldman and Seung-Yeal Ha. Nonlinear functionals of multi-D discrete velocity Boltzmann equations. Journal of Statistical Physics, 114(3–4):1015–1033, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012515.85916.2a.

Finkler:2006:DPM

[FHAY06] Ilya Finkler, Bertrand I. Halperin, Assa Auerbach, and Amir Yacoby. Domain patterns in the microwave-induced zero-resistance state. Journal of Statistical Physics, 125(5–6):1093–1107, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9037-7.

Finken:2003:PSP

[FHL03] R. Finken, J.-P. Hansen, and A. A. Louis. Phase separation of penetrable core mixtures. Journal of Statistical Physics, 110(3– 6):1015–1037, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022136624854.

Fine:2003:ULT

[Fin03] Boris V. Fine. Universal long-time relaxation on lattices of classical spins: Markovian behavior on non–Markovian timescales. Journal of Statistical Physics, 112(1–2):319–327, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023639906726.

Fisch:2006:FSSa

[Fis06a] Ronald Fisch. Finite-size scaling in the energy–entropy plane for the 2D ± Ising spin glass. Journal of Statistical Physics, 125(3): 777–792, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9164-1.

Fisch:2006:FSSb

[Fis06b] Ronald Fisch. Finite-size scaling of the domain wall entropy distributions for the 2D ±J Ising spin glass. Journal of Statistical Physics, 125(3):793–803, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9223-7.

Fisch:2007:SSI

[Fis07] Ronald Fisch. Subextensive singularity in the 2D ±J Ising spin glass. Journal of Statistical Physics, 128(4):1113–1124, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9336-7.

Fisch:2008:ARS

[Fis08] Ronald Fisch. Aspect-ratio scaling of domain wall entropy for the 2D ±J Ising spin glass. Journal of Statistical Physics, 130 (3):561–569, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9436-4.

Fogedby:2005:WNA

[FJ05] Hans C. Fogedby and Mogens H. Jensen. Weak noise approach to the logistic map. Journal of Statistical Physics, 121(5–6): 759–778, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5457-z.

Forrester:2001:APS

[FJM01a] P. J. Forrester, B. Jancovici, and D. S. McAnally. Analytic properties of the structure function for the one-dimensional onecomponent log–gas. Journal of Statistical Physics, 102(3–4): 737–780, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004846818738.

Frachebourg:2001:IBA

[FJM01b] L. Frachebourg, V. Jacquemet, and Ph. A. Martin. Inhomogeneous ballistic aggregation. Journal of Statistical Physics, 105(5– 6):745–769, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013540925139.

Foias:2002:SEN

[FJMR02] C. Foias, M. S. Jolly, O. P. Manley, and R. Rosa. Statistical estimates for the Navier–Stokes equations and the Kraichnan theory of 2-D fully developed turbulence. Journal of Statistical Physics, 108(3–4):591–645, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1015782025005.

Foias:2003:LLD

[FJMR03] C. Foias, M. S. Jolly, O. P. Manley, and R. Rosa. On the Landau– Lifschitz degrees of freedom in 2-D turbulence. Journal of Statistical Physics, 111(3–4):1017–1019, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022814702548.

Fantoni:2003:POC

[FJT03] R. Fantoni, B. Jancovici, and G. Téllez. Pressures for a onecomponent plasma on a pseudosphere. Journal of Statistical Physics, 112(1–2):27–57, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023671419021.

Fannjiang:2000:FBM

[FK00] Albert Fannjiang and Tomasz Komorowski. Fractional Brownian motions and enhanced diffusion in a unidirectional wavelike turbulence. Journal of Statistical Physics, 100(5–6):1071– 1095, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018738009970.

Fel:2002:ECL

[FK02] Leonid G. Fel and Konstantin M. Khanin. On effective conductivity on Z^{d} lattice. Journal of Statistical Physics, 108(5–6): 1015–1031, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019787330168.

Fiala:2004:TFF

[FK04] Jan Fiala and Peter Kleban. Thermodynamics of the Farey fraction spin chain. Journal of Statistical Physics, 116(5–6): 1471–1490, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000041727.66775.d3.

Fiala:2005:GNT

[FK05] Jan Fiala and Peter Kleban. Generalized number theoretic spin chain-connections to dynamical systems and expectation values. Journal of Statistical Physics, 121(3–4):553–577, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-7579-8.

Finkelshtein:2009:RMS

[FK09a] Dmitri Finkelshtein and Yuri Kondratiev. Regulation mechanisms in spatial stochastic development models. Journal of Statistical Physics, 136(1):103–115, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9764-7.

Fleurke:2009:SRP

[FK09b] S. R. Fleurke and C. Külske. A second-row parking paradox. Journal of Statistical Physics, 136(2):285–295, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9775-4.pdf.

Friedland:2008:VAM

[FKLM08] S. Friedland, E. Krop, P. H. Lundow, and K. Markström. On the validations of the asymptotic matching conjectures. Journal of Statistical Physics, 133(3):513–533, November 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9550-y.

Frisch:2005:MFA

[FKM05] U. Frisch, K. Khanin, and T. Matsumoto. Multifractality of the Feigenbaum attractor and fractional derivatives. Journal of Statistical Physics, 121(5–6):671–695, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7011-4.

Fiala:2003:PTS

[FKÖ03] Jan Fiala, Peter Kleban, and Ali Özlük. The phase transition in statistical models defined on Farey fractions. Journal of Statistical Physics, 110(1–2):73–86, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021014627403.

Fishman:2008:PDL

[FKS08] Shmuel Fishman, Yevgeny Krivolapov, and Avy Soffer. On the problem of dynamical localization in the nonlinear Schrödinger equation with a random potential. Journal of Statistical Physics, 131(5):843–865, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9472-0.

Felderhof:2003:RBM

[FL03a] B. U. Felderhof and C. Löhn. Rotational Brownian motion of a pair of dipoles coupled via a classical Heisenberg interaction. Journal of Statistical Physics, 110(3–6):1099–1115, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022196810741.

Franz:2003:RBO

[FL03b] Silvio Franz and Michele Leone. Replica bounds for optimization problems and diluted spin systems. Journal of Statistical Physics, 111(3–4):535–564, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022885828956.

Feng:2006:NSP

[FL06] Z. C. Feng and Y. Charles Li. Nanoscale spontaneous patterning. Journal of Statistical Physics, 123(4):741–751, May 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9083-1.

Fischer:2000:SLG

[FLM00] Werner Fischer, Hajo Leschke, and Peter Müller. Spectral localization by Gaussian random potentials in multi-dimensional continuous space. Journal of Statistical Physics, 101(5–6): 935–985, December 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026425621261.

Fernandez:2003:VPA

[FLR03] Roberto Fernández, Arnaud Le Ny, and Frank Redig. Variational principle and almost quasilocality for renormalized measures. Journal of Statistical Physics, 111(1–2):465–478, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022281730113.

Ferrari:2007:CFN

[FLS07] Pablo A. Ferrari, Claudio Landim, and Valentin V. Sisko. Condensation for a fixed number of independent random variables. Journal of Statistical Physics, 128(5):1153–1158, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9356-3.

Fedorenko:2008:FTC

[FLW08] Andrei A. Fedorenko, Pierre Le Doussal, and Kay Jörg Wiese. Field theory conjecture for loop-erased random walks. Journal of Statistical Physics, 133(5):805–812, December 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9642-8.

Fabricius:2001:BEI

[FM01a] Klaus Fabricius and Barry M. McCoy. Bethe’s equation is incomplete for the XXZ model at roots of unity. Journal of Statistical Physics, 103(5–6):647–678, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010380116927.

Fabricius:2001:CBE

[FM01b] Klaus Fabricius and Barry M. McCoy. Completing Bethe’s equations at roots of unity. Journal of Statistical Physics, 104(3– 4):573–587, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010372504158.

Fournier:2001:MPA

[FM01c] Nicolas Fournier and Sylvie Méléard. A Markov process associated with a Boltzmann equation without cutoff and for non– Maxwell molecules. Journal of Statistical Physics, 104(1–2):359– 385, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010322130480.

Fort:2002:ASW

[FM02a] Joaquim Fort and Vicenç Méndez. On approximate solutions to the wavefront speed problem. Journal of Statistical Physics, 107 (3–4):805–820, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014590230955.

Frisch:2002:MFD

[FM02b] U. Frisch and T. Matsumoto. On multifractality and fractional derivatives. Journal of Statistical Physics, 108(5–6):1181– 1202, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019843616965.

Fabricius:2003:NDE

[FM03] Klaus Fabricius and Barry M. McCoy. New developments in the eight vertex model. Journal of Statistical Physics, 111(1– 2):323–337, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022213209641.

Fabricius:2005:NDE

[FM05a] Klaus Fabricius and Barry M. McCoy. New developments in the eight vertex model II. Chains of odd length. Journal of Statistical Physics, 120(1–2):37–70, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-4410-5.

Fernandez:2005:CCC

[FM05b] Roberto Fernández and Grégory Maillard. Chains with complete connections: General theory, uniqueness, loss of memory and mixing properties. Journal of Statistical Physics, 118(3–4): 555–588, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-8821-5.

Fontbona:2007:PCE

[FM07] Joaquin Fontbona and Miguel Martinez. Paths clustering and an existence result for stochastic vortex systems. Journal of Statistical Physics, 128(3):699–719, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9339-4.

Falco:2008:RGA

[FM08] Pierluigi Falco and Vieri Mastropietro. Renormalization group and asymptotic spin-charge separation for chiral Luttinger liquids. Journal of Statistical Physics, 131(1):79–116, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9474-y.

Fabricius:2009:NMT

[FM09a] Klaus Fabricius and Barry M. McCoy. New Q matrices and their functional equations for the eight vertex model at elliptic roots of unity. Journal of Statistical Physics, 134(4):643–668, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9692-6.

Forrester:2009:MCC

[FM09b] Peter J. Forrester and Anthony Mays. A method to calculate correlation functions for β = 1 random matrices of odd size. Journal of Statistical Physics, 134(3):443–462, February 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9684-6.

Frisch:2003:SEF

[FMB03] U. Frisch, T. Matsumoto, and J. Bec. Singularities of Euler flow? Not out of the blue! Journal of Statistical Physics, 113(5–6): 761–781, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1027308602344.

Frohlich:2004:IAT

[FMS04] J. Fröhlich, M. Merkli, and I. M. Sigal. Ionization of atoms in a thermal field. Journal of Statistical Physics, 116(1–4):311–359, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037226.16493.5e.

Franosch:2000:PDN

[FN00] T. Franosch and David R. Nelson. Population dynamics near an oasis with time-dependent convection. Journal of Statistical Physics, 99(3–4):1021–1030, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018607932123.

Fritz:2006:EFS

[FNO06] József Fritz, Katalin Nagy, and Stefano Olla. Equilibrium fluctuations for a system of harmonic oscillators with conservative noise. Journal of Statistical Physics, 122(3):399–415, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8070-2.

Ferrari:2007:HPN

[FNP07] Pablo A. Ferrari, Beat M. Niederhauser, and Eugene A. Pechersky. Harness processes and non-homogeneous crystals. Journal of Statistical Physics, 128(5):1159–1176, September 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9343-8.

Fournier:2000:SPS

[Fou00] Nicolas Fournier. Strict positivity of a solution to a onedimensional Kac equation without cutoff. Journal of Statistical Physics, 99(3–4):725–749, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018683226672.

Fournier:2006:UCS

[Fou06] Nicolas Fournier. Uniqueness for a class of spatially homogeneous Boltzmann equations without angular cutoff. Journal of Statistical Physics, 125(4):923–942, November 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9208-6.

Friedli:2004:NAV

[FP04] S. Friedli and C.-E. Pfister. Non-analyticity and the van der Waals limit. Journal of Statistical Physics, 114(3–4):665–734, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012506.98828.dd.

Feigenbaum:2001:DFF

[FPD01] Mitchell J. Feigenbaum, Itamar Procaccia, and Benny Davidovich. Dynamics of finger formation in Laplacian growth without surface tension. Journal of Statistical Physics, 103(5– 6):973–1007, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010313106449.

Fitzgerald:2000:MCT

[FPS00] M. Fitzgerald, R. R. Picard, and R. N. Silver. Monte Carlo transition dynamics and variance reduction. Journal of Statistical Physics, 98(1–2):321–345, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018635108073.

Fellner:2004:ECE

[FPS04] Klemens Fellner, Frederic Poupaud, and Christian Schmeiser. Existence and convergence to equilibrium of a kinetic model for cometary flows. Journal of Statistical Physics, 114(5–6): 1481–1499, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013956.08390.b1.

Fernandez:2007:ARH

[FPS07] Roberto Fernández, Aldo Procacci, and Benedetto Scoppola. The analyticity region of the hard sphere gas. Improved bounds. Journal of Statistical Physics, 128(5):1139–1143, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9352-7.

Frohlich:2001:UQH

[FPSW01] Jürg Fröhlich, Bill Pedrini, Christoph Schweigert, and Johannes Walcher. Universality in quantum Hall systems: Coset construction of incompressible states. Journal of Statistical Physics, 103 (3–4):527–567, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010389232079.

Feng:2000:SAE

[FR00] Y. Feng and H. J. Ruskin. Shell analysis and effective disorder in a 2D froth. Journal of Statistical Physics, 99(1–2):263–272, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018600824582.

Forrester:2007:SML

[FR07] Peter J. Forrester and Eric M. Rains. Symmetrized models of last passage percolation and non-intersecting lattice paths. Journal of Statistical Physics, 129(5–6):833–855, December 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9413-y.

Franceschetti:2007:WRW

[Fra07a] Massimo Franceschetti. When a random walk of fixed length can lead uniformly anywhere inside a hypersphere. Journal of Statistical Physics, 127(4):813–823, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9305-1.

Franz:2007:DKS

[Fra07b] Silvio Franz. On the dynamics of Kac p-spin glasses. Journal of Statistical Physics, 126(4–5):765–780, March 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9196-6.

Freidlin:2001:SOE

[Fre01] Mark I. Freidlin. On stable oscillations and equilibriums induced by small noise. Journal of Statistical Physics, 103(1–2):283–300, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004827921214.

Freidlin:2003:DPP

[Fre03] Mark Freidlin. Deterministic 3D-Perturbations of planar incompressible flow lead to stochasticity. Journal of Statistical Physics, 111(5–6):1209–1218, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1023052216159.

Freidlin:2004:SRS

[Fre04] Mark Freidlin. Some remarks on the Smoluchowski–Kramers approximation. Journal of Statistical Physics, 117(3–4):617– 634, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2273-9.

Freidlin:2006:RCI

[Fre06] Mark Freidlin. Reaction-convection in incompressible 3D-Fluid: A homogenization problem. Journal of Statistical Physics, 122 (2):333–349, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8073-z.

Freilikher:2008:JPF

[Fre08a] Valentin Freilikher. J.-P. Fouque, J. Garnier, G. Papanicolaou, K. Sølna: Wave Propagation and Time Reversal in Randomly Layered Media. Journal of Statistical Physics, 131(1):169–171, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9489-z.

Freund:2008:JWG

[Fre08b] Isaac Freund. Joseph w. Goodman: Speckle phenomena in optics: Theory and applications. Journal of Statistical Physics, 130 (2):413–414, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9440-8.

Farnell:2009:HOC

[FRZB09] D. J. J. Farnell, J. Richter, R. Zinke, and R. F. Bishop. Highorder coupled cluster method (CCM) calculations for quantum magnets with valence-bond ground states. Journal of Statistical Physics, 135(1):175–198, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9703-7.

Ferrari:2003:SFF

[FS03] Patrik L. Ferrari and Herbert Spohn. Step fluctuations for a faceted crystal. Journal of Statistical Physics, 113(1–2):1–46, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1025703819894.

Figotin:2005:STT

[FS05] A. Figotin and J. H. Schenker. Spectral theory of time dispersive and dissipative systems. Journal of Statistical Physics, 118(1– 2):199–263, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-8783-7.

Figotin:2006:OSV

[FS06] Alexander Figotin and Stephen P. Shipman. Open systems viewed through their conservative extensions. Journal of Statistical Physics, 125(2):359–409, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9195-7.

Fellner:2007:CES

[FS07a] Klemens Fellner and Christian Schmeiser. Classification of equilibrium solutions of the cometary flow equation and explicit solutions of the Euler equations for monatomic ideal gases. Journal of Statistical Physics, 129(3):493–507, November 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9396-8.

Figotin:2007:HSD

[FS07b] Alexander Figotin and Jeffrey H. Schenker. Hamiltonian structure for dispersive and dissipative dynamical systems. Journal of Statistical Physics, 128(4):969–1056, August 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9321-1.

Fontes:2008:BPH

[FS08] L. R. G. Fontes and R. H. Schonmann. Bootstrap percolation on homogeneous trees has 2 phase transitions. Journal of Statistical Physics, 132(5):839–861, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9583-2.

Feldman:2008:KPO

[FSG08] E. P. Feldman, L. I. Stefanovich, and K. V. Gumennyk. Kinetics of polydomain ordering at second-order phase transitions (by the example of the AuCu_{3} alloy). Journal of Statistical Physics, 132 (3):501–517, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9559-2.

Fan:2006:CAR

[FST06] Ai Hua Fan, Károly Simon, and Hajnal R. Tóth. Contracting on average random IFS with repelling fixed point. Journal of Statistical Physics, 122(1):169–193, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8013-4.

Fridkin:2001:FSX

[FSZ01] V. Fridkin, Yu. Stroganov, and D. Zagier. Finite size XXZ spin chain with anisotropy parameter Δ = \frac{1}{2}. Journal of Statistical Physics, 102(3–4):781–794, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004898802809.

Franz:2006:NGT

[FT06] Silvio Franz and Francesca Tria. A note on the Guerra and Talagrand theorems for mean field spin glasses: The simple case of spherical models. Journal of Statistical Physics, 122(2):313–332, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8007-9.

Ferrero:2007:TDT

[FT07] Alejandro Ferrero and Gabriel Téllez. Two-dimensional twocomponent plasma with adsorbing impurities. Journal of Statistical Physics, 129(4):759–786, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9401-2.

Fantoni:2008:TDO

[FT08] Riccardo Fantoni and Gabriel Téllez. Two-dimensional onecomponent plasma on Flamm’s paraboloid. Journal of Statistical Physics, 133(3):449–489, November 2008. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9616-x.

Frohlich:2005:HRM

[FU05] Jürg Fröhlich and Daniel Ueltschi. Hund’s rule and metallic ferromagnetism. Journal of Statistical Physics, 118(5–6):973–978, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2174-y.

Fukuyama:2000:DSB

[Fuk00] M. Fukuyama. Discrete symmetry breaking for certain shortrange interactions. Journal of Statistical Physics, 98(5–6): 1049–1061, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018607526690.

Fukushima:2003:NMH

[Fuk03] Noboru Fukushima. A new method of the high temperature series expansion. Journal of Statistical Physics, 111(5– 6):1049–1090, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023091930271.

Fukushima:2008:AWS

[Fuk08] Ryoki Fukushima. Asymptotics for the Wiener sausage among Poissonian obstacles. Journal of Statistical Physics, 133(4): 639–657, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9629-5.

Franosch:2002:CMS

[FV02] T. Franosch and Th. Voigtmann. Completely monotone solutions of the mode-coupling theory for mixtures. Journal of Statistical Physics, 109(1–2):237–259, October 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019991729106.

Fatkullin:2003:SDC

[FVE03] Ibrahim Fatkullin and Eric Vanden-Eijnden. Statistical description of contact-interacting Brownian walkers on the line. Journal of Statistical Physics, 112(1–2):155–163, July 2003.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023627604000.

Fannjiang:2003:NID

[FW03] Albert Fannjiang and Lech Wolowski. Noise induced dissipation in Lebesgue-measure preserving maps on d-dimensional torus. Journal of Statistical Physics, 113(1–2):335–378, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025787124437.

Freidlin:2006:LTB

[FW06] M. I. Freidlin and A. D. Wentzell. Long-time behavior of weakly coupled oscillators. Journal of Statistical Physics, 123 (6):1311–1337, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9133-8.

Fyodorov:2007:RSB

[FW07] Yan V. Fyodorov and Ian Williams. Replica symmetry breaking condition exposed by random matrix calculation of landscape complexity. Journal of Statistical Physics, 129(5–6):1081– 1116, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9386-x.

Farmer:2006:CRT

[FY06] David W. Farmer and Mark Yerrington. Crystallization of random trigonometric polynomials. Journal of Statistical Physics, 123(6):1219–1230, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9126-7.

Fonseca:2003:IFT

[FZ03] P. Fonseca and A. Zamolodchikov. Ising field theory in a magnetic field: Analytic properties of the free energy. Journal of Statistical Physics, 110(3–6):527–590, March 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022147532606.

Foffi:2000:KAO

[FZS^{+}00] G. Foffi, E. Zaccarelli, F. Sciortino, P. Tartaglia, and K. A. Dawson. Kinetic arrest originating in competition between attractive interaction and packing force. Journal of Statistical Physics, 100 (1–2):363–376, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018660200206.

Garzo:2005:TCI

[GA05] Vicente Garzó and Antonio Astillero. Transport coefficients for inelastic Maxwell mixtures. Journal of Statistical Physics, 118 (5–6):935–971, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2006-0.

Gacs:2001:RCA

[Gác01] Peter Gács. Reliable cellular automata with self-organization. Journal of Statistical Physics, 103(1–2):45–267, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004823720305.

Goldenfeld:2006:RGA

[GAD06] Nigel Goldenfeld, Badrinarayan P. Athreya, and Jonathan A. Dantzig. Renormalization group approach to multiscale modelling in materials science. Journal of Statistical Physics, 125 (5–6):1015–1023, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9013-7.

Gallavotti:2001:LMR

[Gal01] Giovanni Gallavotti. The Luttinger model: Its role in the RG–Theory of one dimensional many body Fermi systems. Journal of Statistical Physics, 103(3–4):459–483, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010381030262.

Galatolo:2006:HTD

[Gal06] Stefano Galatolo. Hitting time and dimension in Axiom A systems, generic interval exchanges and an application to Birkoff sums. Journal of Statistical Physics, 123(1):111–124, April 2006.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9041-y.

Galvan:2008:QMI

[Gal08] Bruno Galvan. Quantum mechanics and imprecise probability. Journal of Statistical Physics, 131(6):1155–1167, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9530-2.

Gallavotti:2009:TCO

[Gal09] Giovanni Gallavotti. Thermostats, chaos and Onsager reciprocity. Journal of Statistical Physics, 134(5–6):1121–1131, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9655-3.

Gamba:2003:FTL

[Gam03] Andrea Gamba. Finite-time Lyapunov exponents for products of random transformations. Journal of Statistical Physics, 112 (1–2):193–218, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023631704909.

Garnier:2001:LTD

[Gar01] J. Garnier. Long-time dynamics of Korteweg–de Vries solitons driven by random perturbations. Journal of Statistical Physics, 105(5–6):789–833, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013549126956.

Garzo:2003:NTI

[Gar03] Vicente Garzó. Nonlinear transport in inelastic Maxwell mixtures under simple shear flow. Journal of Statistical Physics, 112 (3–4):657–683, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023828109434.

Garbaczewski:2006:DED

[Gar06] Piotr Garbaczewski. Differential entropy and dynamics of uncertainty. Journal of Statistical Physics, 123(2):315–355, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9058-2.

Garriga:2008:LMN

[Gar08] Adan Garriga. L. Leuzzi and T.M. Nieuwenhuizen: Thermodynamics of the Glassy State. Journal of Statistical Physics, 133(6):1185–1186, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9650-8.

Gaspard:2002:TFN

[Gas02] Pierre Gaspard. Trace formula for noisy flows. Journal of Statistical Physics, 106(1–2):57–96, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013167928166.

Gaspard:2004:TRD

[Gas04] Pierre Gaspard. Time-reversed dynamical entropy and irreversibility in Markovian random processes. Journal of Statistical Physics, 117(3–4):599–615, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

004-3455-1. See erratum [Gas07].

Gaspard:2007:ETR

[Gas07] Pierre Gaspard. Erratum: Time-reversed dynamical entropy and irreversibility in Markovian random processes. Journal of Statistical Physics, 126(4–5):1109, March 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

006-9198-4. See [Gas04].

Gates:2000:ESS

[Gat00] D. J. Gates. Exact stationary states of a two-dimensional transport model. Journal of Statistical Physics, 99(3–4):1031–1035, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018659916194.

Giaquinta:2000:SGE

[GAV00] A. Giaquinta, M. Argentina, and M. G. Velarde. A simple generalized excitability model mimicking salient features of neuron dynamics. Journal of Statistical Physics, 101(1–2):665–678, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026401819366.

Garcia-Colin:2005:BRB

[GC05] L. S. García-Colín. Book review: How Fluids Unmix. Journal of Statistical Physics, 118(3–4):791–792, February 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8830-4.

Garcia-Colin:2008:CCE

[GC08a] L. S. García-Colín. Carlo Cercignani and Ester Gabetta (eds): Transport Phenomena and Kinetic Theory. Applications to Gases, Semiconductors, Photons and Biological Systems. Journal of Statistical Physics, 132(6):1145–1146, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9599-7.

Garcia-Colin:2008:JOS

[GC08b] Leopoldo S. García-Colín. J. Ortiz de Sarate, J.V. Sengers: Hydrodynamic Fluctuations. Journal of Statistical Physics, 130(6): 1217–1218, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9485-3.

Garcia-Colin:2002:BRB

[GCU02] L. S. Garcia-Colin and F. J. Uribe. Book review: Equations of State for Fluid and Fluid Mixtures, Parts I and II. J. V. Sengers, R. F. Kayser, and H. J. White, Jr. Elsevier Science B.V., The Netherlands, 2000. Journal of Statistical Physics, 106(1– 2):403–404, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013148818597.

Garcia-Colin:2003:E

[GCU03] L. S. Garcia-Colin and F. J. Uribe. Erratum. Journal of Statistical Physics, 112(3–4):885, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023892428998.

Gambassi:2006:CDT

[GD06] A. Gambassi and S. Dietrich. Critical dynamics in thin films. Journal of Statistical Physics, 123(5):929–1005, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9089-8.

Granovsky:2009:TDC

[GE09] Boris L. Granovsky and Michael M. Erlihson. On time dynamics of coagulation–fragmentation processes. Journal of Statistical Physics, 134(3):567–588, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9671-3.

Gentile:2004:PPS

[Gen04] Guido Gentile. Pure point spectrum for two-level systems in a strong quasi-periodic field. Journal of Statistical Physics, 115 (5–6):1605–1620, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000028070.11031.0c.

Gentile:2006:RPS

[Gen06] Guido Gentile. Resummation of perturbation series and reducibility for Bryuno skew-product flows. Journal of Statistical Physics, 125(2):317–357, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9127-6.

Giroux:2008:GSG

[GF08] Gaston Giroux and René Ferland. Global spectral gap for Dirichlet–Kac random motions. Journal of Statistical Physics, 132(3):561–567, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9571-6.

Gintautas:2008:RFC

[GFH08] Vadas Gintautas, Glenn Foster, and Alfred W. Hübler. Resonant forcing of chaotic dynamics. Journal of Statistical Physics, 130 (3):617–629, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9444-4.

Gonze:2000:EVC

[GG00] Didier Gonze and Albert Goldbeter. Entrainment versus chaos in a model for a circadian oscillator driven by light–dark cycles. Journal of Statistical Physics, 101(1–2):649–663, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1026410121183.

Gray:2001:ETT

[GG01] Lawrence Gray and David Griffeath. The ergodic theory of traffic jams. Journal of Statistical Physics, 105(3–4):413– 452, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012202706850.

Gielis:2002:RIP

[GG02] Guy Gielis and Geoffrey Grimmett. Rigidity of the interface in percolation and random-cluster models. Journal of Statistical Physics, 109(1–2):1–37, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019950525471.

Graham:2006:RCR

[GG06] B. T. Graham and G. R. Grimmett. Random-cluster representation of the Blume–Capel model. Journal of Statistical Physics, 125(2):283–316, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9135-6.

Godreche:2009:FTF

[GGL09] C. Godrèche, H. Grandclaude, and J. M. Luck. Finite-time fluctuations in the degree statistics of growing networks. Journal of Statistical Physics, 137(5–6):1117–1146, December 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9847-5.

Glimm:2002:NSA

[GGZD02] James Glimm, John Grove, Yongmin Zhang, and Srabasti Dutta. Numerical study of axisymmetric Richtmyer–Meshkov instability and azimuthal effect on spherical mixing. Journal of Statistical Physics, 107(1–2):241–260, April 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014518906518.

Gawedzki:2004:SBF

[GH04] Krzysztof Gawedzki and Péter Horvai. Sticky behavior of fluid particles in the compressible Kraichnan model. Journal of Statistical Physics, 116(5–6):1247–1300, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041740.90705.d5.

Gherardi:2009:WPS

[Ghe09] Marco Gherardi. Whole-plane self-avoiding walks and radial Schramm–Loewner evolution: A numerical study. Journal of Statistical Physics, 136(5):864–874, September 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9797-y.

Ganesh:2000:PBI

[GHO^{+}00] A. Ganesh, B. M. Hambly, Neil O’Connell, Dudley Stark, and P. J. Upton. Poissonian behavior of Ising spin systems in an external field. Journal of Statistical Physics, 99(1–2):613–626, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018669312287.

Ghomeshi:2005:EUS

[Gho05] S. Ghomeshi. Existence and uniqueness of solutions for the Couette problem. Journal of Statistical Physics, 118(1–2):265–300, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8784-6.

Gibilisco:2008:DUP

[GI08] Paolo Gibilisco and Tommaso Isola. A dynamical uncertainty principle in von Neumann algebras by operator monotone functions. Journal of Statistical Physics, 132(5):937–944, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9582-3.

Gibilisco:2008:VIQ

[GII08] Paolo Gibilisco, Daniele Imparato, and Tommaso Isola. A volume inequality for quantum Fisher information and the uncertainty principle. Journal of Statistical Physics, 130(3):545–559, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9454-2.

Giraud:2001:CSG

[Gir01] Christophe Giraud. Clustering in a self-gravitating onedimensional gas at zero temperature. Journal of Statistical Physics, 105(3–4):585–604, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1012227809576.

Giraud:2003:SFB

[Gir03] Christophe Giraud. On a shock front in Burgers turbulence. Journal of Statistical Physics, 111(1–2):387–402, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022221511458.

Gitterman:2007:RKB

[Git07] Moshe Gitterman. R. Kerner: Models of Agglomeration and Glass Transition. Journal of Statistical Physics, 128(5):1261– 1262, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9342-9.

Gitterman:2008:IHB

[Git08] Moshe Gitterman. Igor Herbut: A Modern Approach to Critical Phenomena. Journal of Statistical Physics, 131(1):173, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9490-6.

Gitterman:2009:TUS

[Git09] Moshe Gitterman. Y. Tsori, U. Steiner: Polymers, Liquids and Colloids in Electric Fields. Journal of Statistical Physics, 135 (4):775–776, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9762-9.

Giuliani:2009:LRO

[Giu09] Alessandro Giuliani. Long range order for lattice dipoles. Journal of Statistical Physics, 134(5–6):1059–1070, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9596-x.

Ge:2008:GJE

[GJ08a] Hao Ge and Da-Quan Jiang. Generalized Jarzynski’s equality of inhomogeneous multidimensional diffusion processes. Journal of Statistical Physics, 131(4):675–689, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9520-4.

Goncalves:2008:SLG

[GJ08b] Patrícia Gonçalves and Milton Jara. Scaling limits for gradient systems in random environment. Journal of Statistical Physics, 131(4):691–716, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9509-z.

Goncalves:2008:SLT

[GJ08c] Patrícia Gonçalves and Milton Jara. Scaling limits of a tagged particle in the exclusion process with variable diffusion coefficient. Journal of Statistical Physics, 132(6):1135–1143, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9595-y.

Ge:2006:SDM

[GJQ06] Hao Ge, Da-Quan Jiang, and Min Qian. A simple discrete model of Brownian motors: Time-periodic Markov chains. Journal of Statistical Physics, 123(4):831–859, May 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9099-6.

Galtbayar:2004:LTD

[GJY04] A. Galtbayar, A. Jensen, and K. Yajima. Local time-decay of solutions to Schrödinger equations with time-periodic potentials. Journal of Statistical Physics, 116(1–4):231–282, August 2004.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037203.79298.ec.

Groeneveld:2002:NNR

[GK02] J. Groeneveld and R. Klages. Negative and nonlinear response in an exactly solved dynamical model of particle transport. Journal of Statistical Physics, 109(3–4):821–861, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020487219455.

Gruzberg:2004:LEM

[GK04] Ilya A. Gruzberg and Leo P. Kadanoff. The Loewner equation: Maps and shapes. Journal of Statistical Physics, 114(5–6): 1183–1198, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013973.40984.3b.

Germinet:2006:NCR

[GK06] François Germinet and Abel Klein. New characterizations of the region of complete localization for random Schrödinger operators. Journal of Statistical Physics, 122(1):73–94, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8068-9.

Gamarnik:2009:SCM

[GK09a] David Gamarnik and Dmitriy Katz. Sequential cavity method for computing free energy and surface pressure. Journal of Statistical Physics, 137(2):205–232, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9849-3.

Giardina:2009:TTT

[GK09b] Cristian Giardinà and Jorge Kurchan. Thinking transport as a twist. Journal of Statistical Physics, 135(5–6):895–914, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9736-y.

Garriga:2002:SSE

[GKR02] A. Garriga, J. Kurchan, and F. Ritort. Strong Soret effect in one dimension. Journal of Statistical Physics, 106(1–2):109–123, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013182909209.

Giardina:2009:DHS

[GKRV09] Cristian Giardinà, Jorge Kurchan, Frank Redig, and Kiamars Vafayi. Duality and hidden symmetries in interacting particle systems. Journal of Statistical Physics, 135(1):25–55, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9716-2.pdf.

Gandolfo:2001:RPB

[GL01a] D. Gandolfo and L. Laanait. Retrieval properties of bidirectional associative memories. Journal of Statistical Physics, 104(5–6): 1389–1396, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010426400738.

Godreche:2001:SOT

[GL01b] C. Godrèche and J. M. Luck. Statistics of the occupation time of renewal processes. Journal of Statistical Physics, 104(3–4): 489–524, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010364003250.

Graffi:2004:LIB

[GL04] Sandro Graffi and Marco Lenci. Localization in infinite billiards: A comparison between quantum and classical ergodicity. Journal of Statistical Physics, 116(1–4):821–830, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000037218.05161.f3.

Gallavotti:2002:LDR

[GLM02] G. Gallavotti, J. L. Lebowitz, and V. Mastropietro. Large deviations in rarefied quantum gases. Journal of Statistical Physics, 108(5–6):831–861, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1019766826534.

Gandolfo:2007:LML

[GLMSR07] Daniel Gandolfo, Lahoussine Laanait, Salvador Miracle-Sole, and Jean Ruiz. A lattice model for the line tension of a sessile drop. Journal of Statistical Physics, 126(1):133–156, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9152-5.

Goldstein:2006:DWF

[GLTZ06] Sheldon Goldstein, Joel L. Lebowitz, Roderich Tumulka, and Nino Zanghì. On the distribution of the wave function for systems in thermal equilibrium. Journal of Statistical Physics, 125(5–6):1193–1221, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9210-z.

Guo:2000:MCS

[GM00] Hongyu Guo and Bruce N. Miller. Monte Carlo study of localization on a one-dimensional lattice. Journal of Statistical Physics, 98(1–2):347–374, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018687124911.

Gruber:2003:BEA

[GM03a] Christian Gruber and Gary P. Morriss. A Boltzmann equation approach to the dynamics of the simple piston. Journal of Statistical Physics, 113(1–2):297–333, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025735107599.

Guerin:2003:CBL

[GM03b] Hélène Guérin and Sylvie Méléard. Convergence from Boltzmann to Landau processes with soft potential and particle approximations. Journal of Statistical Physics, 111(3–4):931–966, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022858517569.

Golinelli:2005:SDT

[GM05] O. Golinelli and K. Mallick. Spectral degeneracies in the totally asymmetric exclusion process. Journal of Statistical Physics, 120(5–6):779–798, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-6972-7.

Garzo:2007:NST

[GM07a] Vicente Garzó and José María Montanero. Navier–Stokes transport coefficients of d-dimensional granular binary mixtures at low density. Journal of Statistical Physics, 129(1):27–58, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9357-2.

Gobron:2007:FOP

[GM07b] T. Gobron and I. Merola. First-order phase transition in Potts models with finite-range interactions. Journal of Statistical Physics, 126(3):507–583, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-006-9230-8.

Gunaratne:2005:VSR

[GMNT05] Gemunu H. Gunaratne, Joseph L. McCauley, Matthew Nicol, and Andrei Török. Variable step random walks and selfsimilar distributions. Journal of Statistical Physics, 121(5–6): 887–899, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5474-y.

Grun:2006:TFF

[GMR06] Günther Grün, Klaus Mecke, and Markus Rauscher. Thinfilm flow influenced by thermal noise. Journal of Statistical Physics, 122(6):1261–1291, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-006-9028-8.

Ghirardi:2002:EPC

[GMW02] GianCarlo Ghirardi, Luca Marinatto, and Tullio Weber. Entanglement and properties of composite quantum systems: A conceptual and mathematical analysis. Journal of Statistical Physics, 108(1–2):49–122, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015439502289.

Gotoh:2003:RPT

[GN03] Toshiyuki Gotoh and Tohru Nakano. Role of pressure in turbulence. Journal of Statistical Physics, 113(5–6):855–874, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027316804161.

Gangardt:2008:WTO

[GN08] D. M. Gangardt and S. K. Nechaev. Wetting transition on a one-dimensional disorder. Journal of Statistical Physics, 130(3): 483–502, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9433-7.

Galka:2006:WTE

[GOBY06] Andreas Galka, Tohru Ozaki, Jorge Bosch Bayard, and Okito Yamashita. Whitening as a tool for estimating mutual information in spatiotemporal data sets. Journal of Statistical Physics, 124(5):1275–1315, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9131-x.

Grimmett:2008:EQI

[GOS08] Geoffrey R. Grimmett, Tobias J. Osborne, and Petra F. Scudo. Entanglement in the quantum Ising model. Journal of Statistical Physics, 131(2):305–339, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9502-6.

Garcia-Ojalvo:2000:ESS

[GOSG00] J. García-Ojalvo and L. Schimansky-Geier. Excitable structures in stochastic bistable media. Journal of Statistical Physics, 101(1–2):473–481, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026493517549.

Gottlieb:2005:EBF

[Got05] Alex D. Gottlieb. Examples of bosonic de Finetti states over finite dimensional Hilbert spaces. Journal of Statistical Physics, 121(3–4):497–509, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-7005-2.

Groppi:2004:TKM

[GP04] Maria Groppi and Jacek Polewczak. On two kinetic models for chemical reactions: Comparisons and existence results. Journal of Statistical Physics, 117(1–2):211–241, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044059.59066.a9.

Garcia-Pelayo:2008:TSP

[GP08] Ricardo García-Pelayo. Twice scattered particles in a plane are uniformly distributed. Journal of Statistical Physics, 133(2): 401–404, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9612-1.

Gruber:2002:DMC

[GPL02] Christian Gruber, Séverine Pache, and Annick Lesne. Deterministic motion of the controversial piston in the thermodynamic limit. Journal of Statistical Physics, 108(3–4):669–701, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015786125914.

Gruber:2003:TTS

[GPL03] Christian Gruber, Séverine Pache, and Annick Lesne. Twotime-scale relaxation towards thermal equilibrium of the enigmatic piston. Journal of Statistical Physics, 112(5–6):1177– 1206, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1024671710343.

Gruber:2004:SLT

[GPL04] Christian Gruber, Séverine Pache, and Annick Lesne. On the second law of thermodynamics and the piston problem. Journal of Statistical Physics, 117(3–4):739–772, November 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2271-y.

Glass:2005:CDE

[GPM^{+}05] Leon Glass, Theodore J. Perkins, Jonathan Mason, Hava T. Siegelmann, and Roderick Edwards. Chaotic dynamics in an electronic model of a genetic network. Journal of Statistical Physics, 121(5–6):969–994, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-7009-y.

Gayrard:2002:DOI

[GPT02] V. Gayrard, E. Presutti, and L. Triolo. Density oscillations at the interface between vapor and liquid. Journal of Statistical Physics, 108(5–6):863–884, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019723010604.

Gyorgyi:2000:BSC

[GR00] G. Györgyi and P. Reimann. Beyond storage capacity in a single model neuron: Continuous replica symmetry breaking. Journal of Statistical Physics, 101(1–2):679–702, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026441500710.

Gabetta:2006:SNR

[GR06] Ester Gabetta and Eugenio Regazzini. Some new results for McKean’s graphs with applications to Kac’s equation. Journal of Statistical Physics, 125(4):943–970, November 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9187-7.

Gupta:2009:SMC

[GR09] Shalabh Gupta and Asok Ray. Statistical mechanics of complex systems for pattern identification. Journal of Statistical Physics, 134(2):337–364, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9679-3.

Graham:2000:LEC

[Gra00] Robert Graham. Langevin equation of collective modes of Bose– Einstein condensates in traps. Journal of Statistical Physics, 101(1–2):243–257, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026453803436.

Gray:2001:RGG

[Gra01] Lawrence F. Gray. A reader’s guide to Gacs’s ‘Positive Rates’ paper. Journal of Statistical Physics, 103(1–2):1–44, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004824203467.

Grassberger:2009:SLE

[Gra09] Peter Grassberger. Scaling of loop-erased walks in 2 to 4 dimensions. Journal of Statistical Physics, 136(2):399–404, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9787-0.

Griffiths:2000:CQR

[Gri00] Robert B. Griffiths. Consistent quantum realism: A reply to Bassi and Ghirardi. Journal of Statistical Physics, 99(5–6): 1409–1425, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018601225280.

Grmela:2008:ECH

[Grm08] Miroslav Grmela. Extensions of classical hydrodynamics. Journal of Statistical Physics, 132(3):581–602, August 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9558-3.

Gandolfo:2007:MRC

[GRU07] Daniel Gandolfo, Jean Ruiz, and Daniel Ueltschi. On a model of random cycles. Journal of Statistical Physics, 129(4):663– 676, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9410-1.

Grynberg:2001:GDO

[Gry01] M. D. Grynberg. Growth diversity in one dimensional fluctuating interfaces. Journal of Statistical Physics, 103(1–2):395–408, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004836107101.

Gonchenko:2000:TDA

[GS00] S. V. Gonchenko and L. P. Shilnikov. On two-dimensional areapreserving diffeomorphisms with infinitely many elliptic islands. Journal of Statistical Physics, 101(1–2):321–356, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026418323000.

Grecchi:2001:CMD

[GS01] Vincenzo Grecchi and Andrea Sacchetti. Critical metastability and destruction of the splitting in non-autonomous systems. Journal of Statistical Physics, 103(1–2):339–368, April 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004832022123.

Gaveau:2003:CDR

[GS03a] B. Gaveau and L. S. Schulman. Creation, dissipation and recycling of resources in non-equilibrium systems. Journal of Statistical Physics, 110(3–6):1317–1367, March 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022165431214.

Ginoza:2003:AES

[GS03b] Mitsuaki Ginoza and Moises Silbert. Analytic equation of state of a quasi one-dimensional model lipid monolayer. Journal of Statistical Physics, 110(1–2):419–429, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021087116017.

Gesztesy:2004:CIM

[GS04] Fritz Gesztesy and Barry Simon. Connectedness of the isospectral manifold for one-dimensional half-line Schrödinger operators. Journal of Statistical Physics, 116(1–4):361–365, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000037217.89500.b3.

Giardina:2007:VBG

[GS07a] Cristian Giardinà and Shannon Starr. Variational bounds for the generalized random energy model. Journal of Statistical Physics, 127(1):1–20, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-007-9296-y.pdf.

Goldstein:2007:UQE

[GS07b] S. Goldstein and W. Struyve. On the uniqueness of quantum equilibrium in Bohmian mechanics. Journal of Statistical Physics, 128(5):1197–1209, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9354-5.

Grima:2007:MSA

[GS07c] Ramon Grima and Santiago Schnell. A mesoscopic simulation approach for modeling intracellular reactions. Journal of Statistical Physics, 128(1–2):139–164, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9202-z.

Garnier:2008:RBP

[GS08a] Josselin Garnier and Knut Sølna. Random backscattering in the parabolic scaling. Journal of Statistical Physics, 131(3):445–486, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9488-0.

Gouet:2008:BDP

[GS08b] Raul Gouet and Aidan Sudbury. Blocking and dimer processes on the Cayley tree. Journal of Statistical Physics, 130 (5):935–955, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9451-5.

Grosskinsky:2008:DCTa

[GS08c] Stefan Großkinsky and Gunter M. Schütz. Discontinuous condensation transition and nonequivalence of ensembles in a zerorange process. Journal of Statistical Physics, 132(1):77–108, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9541-z.

Grosskinsky:2008:DCTb

[GS08d] Stefan Großkinsky and Gunter M. Schütz. Discontinuous condensation transition and nonequivalence of ensembles in a zero-range process. Journal of Statistical Physics, 133(4):801– 803, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9628-6.

Giuliani:2009:GSE

[GS09] Alessandro Giuliani and Robert Seiringer. The ground state energy of the weakly interacting Bose gas at high density. Journal of Statistical Physics, 135(5–6):915–934, June 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9718-0.

Grosskinsky:2003:CZR

[GSS03] Stefan Großkinsky, Gunter M. Schütz, and Herbert Spohn. Condensation in the zero range process: Stationary and dynamical properties. Journal of Statistical Physics, 113(3–4): 389–410, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026008532442.

Grosskinsky:2007:RRS

[GSW07] Stefan Großkinsky, Gunter M. Schütz, and Richard D. Willmann. Rigorous results on spontaneous symmetry breaking in a one-dimensional driven particle system. Journal of Statistical Physics, 128(3):587–606, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9341-x.

Gurevich:2002:HDS

[GT02] B. M. Gurevich and A. A. Tempelman. Hausdorff dimension of sets of generic points for Gibbs measures. Journal of Statistical Physics, 108(5–6):1281–1301, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019760018782.

Grill:2004:ODF

[GT04a] Karl Grill and Christian Tutschka. One-dimensional falling bodies. Journal of Statistical Physics, 117(5–6):1015–1022, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5713-7.

Guerra:2004:HTR

[GT04b] Francesco Guerra and Fabio Lucio Toninelli. The high temperature region of the Viana–Bray diluted spin glass model. Journal of Statistical Physics, 115(1–2):531–555, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019815.11115.54.

Grill:2006:LGF

[GT06] Karl Grill and Christian Tutschka. Lattice gas with finite-range interaction under gravity. Journal of Statistical Physics, 125(3): 717–726, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9201-0.

Gonzalez:2008:WSD

[GT08] Diego Luis González and Gabriel Téllez. Wigner surmise for domain systems. Journal of Statistical Physics, 132(1):187–205, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9548-5.

Ganikhodjaev:2009:MPP

[GTA09] N. Ganikhodjaev, S. Temir, and H. Akin. Modulated phase of a Potts model with competing binary interactions on a Cayley tree. Journal of Statistical Physics, 137(4):701–715, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9869-z.

Gravner:2001:LTH

[GTW01] Janko Gravner, Craig A. Tracy, and Harold Widom. Limit theorems for height fluctuations in a class of discrete space and time growth models. Journal of Statistical Physics, 102(5–6): 1085–1132, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004879725949.

Gruber:2002:BKT

[GTZ02] Christian Gruber, Hiroshi Tamura, and Valentin A. Zagrebnov. Berezinskii–Kosterlitz–Thouless order in two-dimensional O(2)-ferrofluid. Journal of Statistical Physics, 106(5–6):875–893, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014021400408.

Gaspard:2002:WDT

[GvB02] Pierre Gaspard and Henk van Beijeren. When do tracer particles dominate the Lyapunov spectrum? Journal of Statistical Physics, 109(3–4):671–704, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1020466732292.

Gough:2007:SPQ

[GvH07] John Gough and Ramon van Handel. Singular perturbation of quantum stochastic differential equations with coupling through an oscillator mode. Journal of Statistical Physics, 127(3):575– 607, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9284-2.

Garcia:2000:LKE

[GW00] Alejandro L. Garcia and Wolfgang Wagner. The limiting kinetic equation of the consistent Boltzmann algorithm for dense gases. Journal of Statistical Physics, 101(5–6):1065–1086, December 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026429722169.

Gaiotto:2009:SBC

[GW09] Davide Gaiotto and Edward Witten. Supersymmetric boundary conditions in N = 4 super Yang–Mills theory. Journal of Statistical Physics, 135(5–6):789–855, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9687-3.

Georgii:2005:CIG

[GY05] Hans-Otto Georgii and Hyun Jae Yoo. Conditional intensity and Gibbsianness of determinantal point processes. Journal of Statistical Physics, 118(1–2):55–84, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-8777-5.

Georgii:2001:EDP

[GZ01] Hans-Otto Georgii and Valentin Zagrebnov. Entropy-driven phase transitions in multitype lattice gas models. Journal of Statistical Physics, 102(1–2):35–67, January 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026556508220.

Giuliani:2005:FRB

[GZG05] A. Giuliani, F. Zamponi, and G. Gallavotti. Fluctuation relation beyond linear response theory. Journal of Statistical Physics, 119 (3–4):909–944, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3021-5.

Huepe:2002:DPT

[HAG02] Cristián Huepe and Maximino Aldana-González. Dynamical phase transition in a neural network model with noise: An exact solution. Journal of Statistical Physics, 108(3–4):527–540, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015777824097.

Hagele:2007:PCL

[Häg07] Daniel Hägele. Proof of the cases p ≤ 7 of the Lieb–Seiringer formulation of the Bessis–Moussa–Villani conjecture. Journal of Statistical Physics, 127(6):1167–1171, June 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9327-8.

Hansen:2002:BRB

[Han02] Per Lyngs Hansen. Book review: Theory of Critical Phenomena in Finite-Size Systems, Scaling and Quantum Effects. J. G. Brankov, D. M. Danchev, and N. S. Tonchev, Series in Condensed Matter Physics, Vol. 9, World Scientific, 2000. Journal of Statistical Physics, 107(5–6):1299–1302, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015174128018.

Hansen:2006:ELC

[Han06] Frank Hansen. Extensions of Lieb’s concavity theorem. Journal of Statistical Physics, 124(1):87–101, July 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9155-2.

Hansen:2007:WYE

[Han07] Frank Hansen. The Wigner–Yanase entropy is not subadditive. Journal of Statistical Physics, 126(3):643–648, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9265-x.

Hao:2005:CSO

[Hao05] Bailin Hao. Critical slowing down in one-dimensional maps and beyond. Journal of Statistical Physics, 121(5–6):749–757, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8669-3.

Hartmann:2008:ESS

[Har08] Carsten Hartmann. An ergodic sampling scheme for constrained Hamiltonian systems with applications to molecular dynamics. Journal of Statistical Physics, 130(4):687–711, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9470-2.

Hastings:2000:GSS

[Has00] M. B. Hastings. Ground state and spin-glass phase of the large-N infinite-range spin glass via supersymmetry. Journal of Statistical Physics, 99(1–2):171–217, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018696623673.

Hastings:2001:EDS

[Has01] M. B. Hastings. Eigenvalue distribution in the self-dual non– Hermitian ensemble. Journal of Statistical Physics, 103(5– 6):903–913, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010356821471.

Hastings:2002:SIB

[Has02] M. B. Hastings. Scale-invariant branch distribution from a soluble stochastic model. Journal of Statistical Physics, 107(5– 6):1031–1042, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015105607546.

Hattori:2007:FPG

[Hat07] Tetsuya Hattori. The fixed point of a generalization of the renormalization group maps for self-avoiding paths on gaskets. Journal of Statistical Physics, 127(3):609–627, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9283-3.

Hoye:2000:CPS

[HB00] J. S. Høye and I. Brevik. The Casimir problem of spherical dielectrics: A solution in terms of quantum statistical mechanics. Journal of Statistical Physics, 100(1–2):223–232, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018695813410.

Horhammer:2008:IEQ

[HB08] Christian Hörhammer and Helmut Büttner. Information and entropy in quantum Brownian motion. Journal of Statistical Physics, 133(6):1161–1174, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9640-x.

Hilhorst:2008:RLT

[HC08] H. J. Hilhorst and P. Calka. Random line tessellations of the plane: Statistical properties of many-sided cells. Journal of Statistical Physics, 132(4):627–647, August 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9577-0.

He:2002:TFK

[HD02] Xiaoyi He and Gary D. Doolen. Thermodynamic foundations of kinetic theory and lattice Boltzmann models for multiphase flows. Journal of Statistical Physics, 107(1–2):309–328, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014527108336.

Harbola:2003:DTB

[HD03] Upendra Harbola and Shankar P. Das. Dynamic transition in a binary liquid and its dependence on the mass–ratio: Results from a self consistent mode coupling model. Journal of Statistical Physics, 112(5–6):1109–1125, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024627902591.

Hierro:2009:SBF

[HD09] Juan Hierro and César Dopazo. Singular boundaries in the forward Chapman–Kolmogorov differential equation. Journal of Statistical Physics, 137(2):305–329, October 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9842-x.

Heinrichs:2004:STD

[HDMF04] S. Heinrichs, W. Dieterich, P. Maass, and H. L. Frisch. Static and time dependent density functional theory with internal degrees of freedom: Merits and limitations demonstrated for the Potts model. Journal of Statistical Physics, 114(3–4):1115–1125, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012518.27400.2a.

Hanney:2003:ERN

[HE03] T. Hanney and M. R. Evans. Einstein relation for nonequilibrium steady states. Journal of Statistical Physics, 111(5– 6):1377–1390, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023068619793.

Heilmann:2004:ESA

[Hei04] Ole J. Heilmann. Exact solution of 1D asymmetric exclusion model with variable cluster size. Journal of Statistical Physics, 116(1–4):855–879, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000037248.12745.
fa.

Hemmer:2000:DIB

[Hem00] P. C. Hemmer. Demixing in isotropic binary mixtures of rodlike macromolecules. Journal of Statistical Physics, 100(1–2):3–11, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018671107958.

Hepp:2009:EMP

[Hep09] Klaus Hepp. The eye of a mathematical physicist. Journal of Statistical Physics, 134(5–6):1033–1057, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9719-z.

Henderson:2000:OAF

[HGST00] Douglas Henderson, Ian Garcia, Stefan Sokolowski, and Andrij Trokhymchuk. Overlapping associating fluids with directional bonds in a bulk and near a hard wall: Monte Carlo study. Journal of Statistical Physics, 100(1–2):153–166, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018687611593.

Hut:2002:ODR

[HH02] P. Hut and D. C. Heggie. Orbital divergence and relaxation in the gravitational N-body problem. Journal of Statistical Physics, 109(5–6):1017–1025, December 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020472526203.

Hoover:2009:SSM

[HHB09] W. G. Hoover, Carol G. Hoover, and Marcus N. Bannerman. Single-speed molecular dynamics of hard parallel squares and cubes. Journal of Statistical Physics, 136(4):715–732, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9795-0.pdf.

Homma:2004:FDW

[HI04] Makoto Homma and Chigak Itoi. Ferromagnetic domain wall and spiral ground states in one-dimensional deformed flat-band Hubbard model. Journal of Statistical Physics, 117(3–4):477– 519, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2266-8.

Homma:2005:GEA

[HI05] Makoto Homma and Chigak Itoi. Gapless excitation above a domain wall ground state in a flat-band Hubbard model. Journal of Statistical Physics, 119(1–2):391–425, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2723-4.

Hirayama:2005:SVF

[Hir05] Michihiro Hirayama. Second variational formulae for dimension spectra. Journal of Statistical Physics, 118(1–2):103–118, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8779-3.

Hillar:2005:PCC

[HJ05] Christopher J. Hillar and Charles R. Johnson. On the positivity of the coefficients of a certain polynomial defined by two positive definite matrices. Journal of Statistical Physics, 118(3–4): 781–789, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-8829-x.

Horowitz:2009:EFC

[HJ09] Jordan M. Horowitz and Christopher Jarzynski. Exact formula for currents in strongly pumped diffusive systems. Journal of Statistical Physics, 136(5):917–925, September 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9818-x.

Hinrichsen:2000:FSP

[HJDRD00] Haye Hinrichsen, Andrea Jiménez-Dalmaroni, Yadin Rozov, and Eytan Domany. Flowing sand — a possible physical realization of directed percolation. Journal of Statistical Physics, 98(5–6): 1149–1168, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018667712578.

Haller:2001:PGS

[HK01] Karl Haller and Tom Kennedy. Periodic ground states in the neutral Falicov–Kimball model in two dimensions. Journal of Statistical Physics, 102(1–2):15–34, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026504524149.

Hald:2002:ANA

[HK02a] Ole H. Hald and Raz Kupferman. Asymptotic and numerical analyses for mechanical models of heat baths. Journal of Statistical Physics, 106(5–6):1121–1184, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014093921790.

Hryniv:2002:STO

[HK02b] Ostap Hryniv and Roman Kotecký. Surface tension and the Ornstein–Zernike behaviour for the 2D Blume–Capel model. Journal of Statistical Physics, 106(3–4):431–476, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013797920029.

Hatayama:2001:FCM

[HKT01] Goro Hatayama, Atsuo Kuniba, and Taichiro Takagi. Factorization of combinatorial R matrices and associated cellular automata. Journal of Statistical Physics, 102(3–4):843–863, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004803003717.

Horvai:2007:FTA

[HKW07] Peter Horvai, Tomasz Komorowski, and Jan Wehr. Finite time approach to equilibrium in a fractional Brownian velocity field. Journal of Statistical Physics, 127(3):553–565, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9270-0.

Haggkvist:2002:IPF

[HL02] Roland Häggkvist and Per Håkan Lundow. The Ising partition function for 2D grids with periodic boundary: Computation and analysis. Journal of Statistical Physics, 108(3–4):429–457, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015721706350.

Hagendorf:2008:SDC

[HL08] Christian Hagendorf and Pierre Le Doussal. SLE on doublyconnected domains and the winding of loop-erased random walks. Journal of Statistical Physics, 133(2):231–254, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9614-z.

Hildebrand:2000:NWP

[HM00] M. Hildebrand and A. S. Mikhailov. Nanoscale wave patterns in reactive adsorbates with attractive lateral interactions. Journal of Statistical Physics, 101(1–2):599–620, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026497618457.

Hurtado:2008:DNN

[HMG08] Pablo I. Hurtado, J. Marro, and P. L. Garrido. Demagnetization via nucleation of the nonequilibrium metastable phase in a model of disorder. Journal of Statistical Physics, 133(1):29–58, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9602-3.

Hartmann:2005:SDP

[HMH05] Michael Hartmann, Günter Mahler, and Ortwin Hess. Spectral densities and partition functions of modular quantum systems as derived from a Central Limit Theorem. Journal of Statistical Physics, 119(5–6):1139–1151, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-4298-5.

Hu:2007:TAA

[HMO07] Jifeng Hu, Anastasios Matzavinos, and Hans G. Othmer. A theoretical approach to actin filament dynamics. Journal of Statistical Physics, 128(1–2):111–138, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9204-x.

Hochberg:2000:EPR

[HMPPMV00] David Hochberg, Carmen Molina-París, Juan Pérez-Mercader, and Matt Visser. Effective potential for the reaction–diffusion– decay system. Journal of Statistical Physics, 99(3–4):903–941, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018699630306.

Horbacz:2005:SSG

[HMS05] Katarzyna Horbacz, Józef Myjak, and Tomasz Szarek. On stability of some general random dynamical system. Journal of Statistical Physics, 119(1–2):35–60, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2045-6.

Howard:2003:PTZ

[HN03] C. Douglas Howard and Charles M. Newman. The percolation transition for the zero-temperature stochastic Ising model on the hexagonal lattice. Journal of Statistical Physics, 111(1–2):57–62, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022296706006.

Honig:2005:FOCa

[HNO05a] Andreas Hönig, Barbara Niethammer, and Felix Otto. On firstorder corrections to the LSW theory i: Infinite systems. Journal of Statistical Physics, 119(1–2):61–122, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2057-2.

Honig:2005:FOCb

[HNO05b] Andreas Hönig, Barbara Niethammer, and Felix Otto. On firstorder corrections to the LSW theory II: Finite systems. Journal of Statistical Physics, 119(1–2):123–164, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2058-1.

Hivert:2007:DSE

[HNOV07] F. Hivert, S. Nechaev, G. Oshanin, and O. Vasilyev. On the distribution of surface extrema in several one- and two-dimensional random landscapes. Journal of Statistical Physics, 126(2): 243–279, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9231-7.

Hauert:2004:DFD

[HNS04] Ch. Hauert, J. Nagler, and H. G. Schuster. Of dogs and fleas: The dynamics of N uncoupled two-state systems. Journal of Statistical Physics, 116(5–6):1453–1469, September 2004.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041725.70622.c4.

Horvai:2008:CPD

[HNS08] P. Horvai, S. V. Nazarenko, and T. H. M. Stein. Coalescence of particles by differential sedimentation. Journal of Statistical Physics, 130(6):1177–1195, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9466-y.

Horita:2009:NLG

[HO09] Vanderlei Horita and Krerley Oliveira. Non-lacunary Gibbs measures for certain fractal repellers. Journal of Statistical Physics, 136(5):842–863, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9811-4.

Hoddeson:2001:JBT

[Hod01] Lillian Hoddeson. John Bardeen and the theory of superconductivity: A late revision of a homework assignment for J. M. Luttinger. Journal of Statistical Physics, 103(3–4):625–640, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010301602037.

Holovatch:2002:BRB

[Hol02a] Yurij Holovatch. Book review: Critical Properties of ϕ^{4}- Theories. Hagen Kleinert, Verena Schulte–Frohlinde, World Scientific, Singapore, 2001. Journal of Statistical Physics, 107(5– 6):1303–1304, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015126212089.

Holroyd:2002:IEP

[Hol02b] Alexander E. Holroyd. Inequalities in entanglement percolation. Journal of Statistical Physics, 109(1–2):317–323, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019952014993.

Holroyd:2002:KPP

[Hol02c] Alexander E. Holroyd. Knotted paths in percolation. Journal of Statistical Physics, 109(1–2):325–330, October 2002.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019904131832.

Holcman:2007:MDV

[Hol07] D. Holcman. Modeling DNA and virus trafficking in the cell cytoplasm. Journal of Statistical Physics, 127(3):471–494, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9282-4.

Horbacz:2006:PRD

[Hor06] Katarzyna Horbacz. Pointwise and Renyi dimensions of an invariant measure of random dynamical systems with jumps. Journal of Statistical Physics, 122(5):1041–1059, March 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-9017-3.

Hosokawa:2000:SWI

[Hos00] Iwao Hosokawa. Statistics of ‘worms’ in isotropic turbulence treated on the multifractal basis. Journal of Statistical Physics, 99(3–4):783–798, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018639411651.

Hosaka:2006:THM

[Hos06] Kenshi Hosaka. Triviality of hierarchical models with small negative ϕ^{4} coupling in four dimensions. Journal of Statistical Physics, 122(2):237–253, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8010-1.

Hairer:2004:PHH

[HP04] M. Hairer and G. A. Pavliotis. Periodic homogenization for hypoelliptic diffusions. Journal of Statistical Physics, 117(1–2): 261–279, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000044055.59822.20.

Hairer:2008:BDB

[HP08] M. Hairer and G. A. Pavliotis. From ballistic to diffusive behavior in periodic potentials. Journal of Statistical Physics, 131 (1):175–202, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9493-3.

Hoover:2000:CSI

[HPCH00] Wm. G. Hoover, H. A. Posch, V. M. Castillo, and C. G. Hoover. Computer simulation of irreversible expansions via molecular dynamics, smooth particle applied mechanics, Eulerian, and Lagrangian continuum mechanics. Journal of Statistical Physics, 100(1–2):313–326, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018656132065.

Hoover:2002:LMT

[HPF^{+}02] Wm. G. Hoover, Harald A. Posch, Christina Forster, Christoph Dellago, and Mary Zhou. Lyapunov modes of two-dimensional many-body systems; soft disks, hard disks, and rotors. Journal of Statistical Physics, 109(3–4):765–776, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020474901341.

Hat:2007:HNA

[HPK^{+}07] Beata Hat, Pawel Paszek, Marek Kimmel, Kazimierz Piechor, and Tomasz Lipniacki. How the number of alleles influences gene expression. Journal of Statistical Physics, 128(1–2):511– 533, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9218-4.

Holm:2003:BEE

[HPWW03] Darryl D. Holm, Vakhtang Putkaradze, Patrick D. Weidman, and Beth A. Wingate. Boundary effects on exact solutions of the Lagrangian-averaged Navier–Stokes-α equations. Journal of Statistical Physics, 113(5–6):841–854, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1027364720090.

Hammond:2006:KLS

[HR06] Alan Hammond and Fraydoun Rezakhanlou. Kinetic limit for a system of coagulating planar Brownian particles. Journal of Statistical Physics, 124(2–4):997–1040, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-0105-1.

Haggkvist:2004:MCS

[HRA^{+}04] Roland Häggkvist, Anders Rosengren, Daniel Andrén, Petras Kundrotas, Per Håkan Lundow, and Klas Markström. A Monte Carlo sampling scheme for the Ising model. Journal of Statistical Physics, 114(1–2):455–480, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003116.17579.5d.

Hyvaluoma:2005:LBS

[HRK^{+}05] J. Hyväluoma, P. Raiskinmäki, A. Koponen, M. Kataja, and J. Timonen. Lattice-Boltzmann simulation of particle suspensions in shear flow. Journal of Statistical Physics, 121(1–2): 149–161, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-4314-4.

Hoye:2003:DHT

[HRS03] Johan Høye, Fernando O. Raineri, and George Stell. Debye– Hückel theory for charged aligned needles and for polyelectrolyte solutions. Journal of Statistical Physics, 110(3–6):835–860, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022168004381.

Hiemer:1998:PBS

[HS98] Philipp Hiemer and Vadim Snurnikov. Polygonal billiards with small obstacles. Journal of Statistical Physics, 90(1–2):453–466, January 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1023/A%3A1023228423836. See counter-example [Mon04].

Hara:2000:SLI

[HS00] Takashi Hara and Gordon Slade. The scaling limit of the incipient infinite cluster in high-dimensional percolation. I. Critical exponents. Journal of Statistical Physics, 99(5–6):1075–1168, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018628503898.

Holcman:2004:ETS

[HS04] D. Holcman and Z. Schuss. Escape through a small opening: Receptor trafficking in a synaptic membrane. Journal of Statistical Physics, 117(5–6):975–1014, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-5712-8.

Hanyga:2008:MFC

[HS08] A. Hanyga and M. Seredyńska. On a mathematical framework for the constitutive equations of anisotropic dielectric relaxation. Journal of Statistical Physics, 131(2):269–303, April 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9501-7.

Haskovec:2009:DLK

[HS09a] Jan Haskovec and Christian Schmeiser. Diffusive limit of a kinetic model for cometary flows. Journal of Statistical Physics, 136(1):179–194, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9784-3.

Haskovec:2009:SPA

[HS09b] Jan Haskovec and Christian Schmeiser. Stochastic particle approximation for measure valued solutions of the 2D Keller–Segel system. Journal of Statistical Physics, 135(1):133–151, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9717-1.

Hellen:2000:MCS

[HSA00] E. K. O. Hellén, P. Szelestey, and M. J. Alava. Multilayer cooperative sequential adsorption. Journal of Statistical Physics, 98(1– 2):265–280, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018678923094.

Haggstrom:2001:SIR

[HT01] Olle Häggström and Tatyana Turova. A strict inequality for the random triangle model. Journal of Statistical Physics, 104 (1–2):471–482, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010378215459.

Hattori:2002:RGA

[HT02] Tetsuya Hattori and Toshiro Tsuda. Renormalization group analysis of the self-avoiding paths on the d-dimensional Sierpiński gaskets. Journal of Statistical Physics, 109(1–2): 39–66, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019927309542.

Huang:2009:GVP

[Hua09] M. Huang. Gamow vectors in a periodically perturbed quantum system. Journal of Statistical Physics, 137(3):569–592, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9853-7.

Horita:2001:HDN

[HV01] Vanderlei Horita and Marcelo Viana. Hausdorff dimension of non-hyperbolic repellers. I: Maps with holes. Journal of Statistical Physics, 105(5–6):835–862, December 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013501211027.

Hilhorst:2009:MSO

[HvdHMO09] D. Hilhorst, R. van der Hout, M. Mimura, and I. Ohnishi. A mathematical study of the one-dimensional Keller and Rubinow model for Liesegang bands. Journal of Statistical Physics, 135 (1):107–132, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9701-9.pdf.

Heydenreich:2008:MFB

[HvdHS08] Markus Heydenreich, Remco van der Hofstad, and Akira Sakai. Mean-field behavior for long- and finite range Ising model, percolation and self-avoiding walk. Journal of Statistical Physics, 132(6):1001–1049, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9580-5.

Hermisson:2001:FSQ

[HWB01] Joachim Hermisson, Holger Wagner, and Michael Baake. Fourstate quantum chain as a model of sequence evolution. Journal of Statistical Physics, 102(1–2):315–343, January 2001.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026577012763.

Hauksson:2004:AEM

[HY04] Bergthór Hauksson and Jakob Yngvason. Asymptotic exactness of magnetic Thomas–Fermi theory at nonzero temperature. Journal of Statistical Physics, 116(1–4):523–546, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037223.74597.4e.

Iarotski:2001:FEM

[Iar01] D. A. Iarotski. ‘Free’ evolution of multi-particle excitations in the Glauber dynamics at high temperature. Journal of Statistical Physics, 104(5–6):1091–1111, September 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010453728055.

Inoue:2002:DSM

[ICO02] Yasuhiro Inoue, Yu Chen, and Hirotada Ohashi. Development of a simulation model for solid objects suspended in a fluctuating fluid. Journal of Statistical Physics, 107(1–2):85–100, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014550318814.

Iommi:2009:LSA

[IK09a] Godofredo Iommi and Jan Kiwi. The Lyapunov spectrum is not always concave. Journal of Statistical Physics, 135(3):535–546, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9750-0.

Its:2009:FHF

[IK09b] A. R. Its and V. E. Korepin. The Fisher–Hartwig formula and entanglement entropy. Journal of Statistical Physics, 137(5–6): 1014–1039, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9835-9.

Ioffe:2002:NQW

[Iof02] D. Ioffe. A note on the quantum Widom–Rowlison model. Journal of Statistical Physics, 106(1–2):375–384, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013192600850.

Imamura:2007:DTP

[IS07] T. Imamura and T. Sasamoto. Dynamics of a tagged particle in the asymmetric exclusion process with the step initial condition. Journal of Statistical Physics, 128(4):799–846, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9326-9.

Iovanella:2007:SSG

[ISS07] Antonio Iovanella, Benedetto Scoppola, and Elisabetta Scoppola. Some spin glass ideas applied to the clique problem. Journal of Statistical Physics, 126(4–5):895–915, March 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9255-z.

Sasa:2006:SST

[iST06] Shin ichi Sasa and Hal Tasaki. Steady state thermodynamics. Journal of Statistical Physics, 125(1):125–224, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-9021-7.

Ito:2002:RGR

[Ito02] K. R. Ito. Renormalization group recursion formulas and flows of 2D O(N) spin models. Journal of Statistical Physics, 107 (3–4):821–856, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014542315025.

Ioffe:2006:EDP

[IVZ06] D. Ioffe, Y. Velenik, and M. Zahradník. Entropy-driven phase transition in a polydisperse hard-rods lattice system. Journal of Statistical Physics, 122(4):761–786, February 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8085-8.

Jabin:2001:VPS

[Jab01] Pierre-Emmanuel Jabin. The Vlasov–Poisson system with infinite mass and energy. Journal of Statistical Physics, 103(5– 6):1107–1123, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010321308267.

Jacquet:2009:TTP

[Jac09] Philippe A. Jacquet. ThermoElectric transport properties of a chain of quantum dots with self-consistent reservoirs. Journal of Statistical Physics, 134(4):709–748, February 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9697-1.

Jafarpour:2003:AYL

[Jaf03] Farhad H. Jafarpour. The application of the Yang–Lee theory to study a phase transition in a non-equilibrium system. Journal of Statistical Physics, 113(1–2):269–281, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1025731006690.

Jancovici:2000:PMT

[Jan00a] B. Jancovici. Pressure and Maxwell tensor in a Coulomb fluid. Journal of Statistical Physics, 99(5–6):1281–1295, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018640806624.

Jancovici:2000:SRT

[Jan00b] B. Jancovici. A sum rule for the two-dimensional two-component plasma. Journal of Statistical Physics, 100(1–2):201–207, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018691712501.

Jancovici:2001:PST

[Jan01a] B. Jancovici. Pressure and stress tensor in a Yukawa fluid. Journal of Statistical Physics, 102(5–6):1315–1330, March 2001.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004896513654.

Janssen:2001:DPC

[Jan01b] Hans-Karl Janssen. Directed percolation with colors and flavors. Journal of Statistical Physics, 103(5–6):801–839, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010300703724.

Jancovici:2003:CFF

[Jan03] B. Jancovici. Charge fluctuations in finite Coulomb systems. Journal of Statistical Physics, 110(3–6):879–902, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022172105290.

Janzing:2006:CPM

[Jan06a] Dominik Janzing. On the computational power of molecular heat engines. Journal of Statistical Physics, 122(3):531–556, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8015-9.

Janzing:2006:QTM

[Jan06b] Dominik Janzing. Quantum thermodynamics with missing reference frames: Decompositions of free energy into nonincreasing components. Journal of Statistical Physics, 125(3): 761–776, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9220-x.

Jarzynski:2000:HDD

[Jar00] C. Jarzynski. Hamiltonian derivation of a detailed fluctuation theorem. Journal of Statistical Physics, 98(1–2):77–102, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018670721277.

Jeon:2008:PVM

[JAWC08] Junhwan Jeon, Nelson R. Alexander, Alissa M. Weaver, and Peter T. Cummings. Protrusion of a virtual model Lamellipodium by actin polymerization: A coarse-grained Langevin dynamics model. Journal of Statistical Physics, 133(1):79–100, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9600-5.

Jafarizadeh:2001:HCM

[JBKN01] M. A. Jafarizadeh, S. Behnia, S. Khorram, and H. Nagshara. Hierarchy of chaotic maps with an invariant measure. Journal of Statistical Physics, 104(5–6):1013–1028, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010449627146.

Jensen:2001:ELA

[Jen01] Iwan Jensen. Enumerations of lattice animals and trees. Journal of Statistical Physics, 102(3–4):865–881, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004855020556.

Just:2003:EFC

[JGB^{+}03] Wolfram Just, Katrin Gelfert, Nilüfer Baba, Anja Riegert, and Holger Kantz. Elimination of fast chaotic degrees of freedom: On the accuracy of the Born approximation. Journal of Statistical Physics, 112(1–2):277–292, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023635805818.

Jayaprakash:2003:EDR

[JH03] C. Jayaprakash and F. Hayot. An equation for the dissipation rate correlation and its implications for the intermittency exponent μ in turbulence. Journal of Statistical Physics, 111 (1–2):371–386, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022269427387.

Joo:2007:ENE

[JHA07] Jaewook Joo, Eric Harvill, and Réka Albert. Effects of noise on ecological invasion processes: Bacteriophage-mediated competition in bacteria. Journal of Statistical Physics, 128(1–2):229– 256, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9182-z.

Jiang:2009:CPE

[JHW09] Jianguo Jiang, Yineng Huang, and Jichun Wu. Can the prunedenriched method be used for the simulation of fluids? Journal of Statistical Physics, 136(5):984–988, September 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9812-3.

Jiang:2003:SRB

[Jia03] Miaohua Jiang. Sinai-Ruelle-Bowen measures for lattice dynamical systems. Journal of Statistical Physics, 111(3–4):863–902, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022854416660.

Jiang:2007:GSR

[Jia07] Zhenglu Jiang. Global solution to the relativistic Enskog equation with near-vacuum data. Journal of Statistical Physics, 127 (4):805–812, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9269-6.

Jiang:2008:GEP

[Jia08] Zhenglu Jiang. Global existence proof for relativistic Boltzmann equation with hard interactions. Journal of Statistical Physics, 130(3):535–544, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-007-9453-3.

Janke:2001:SFS

[JK01] W. Janke and R. Kenna. The strength of first and second order phase transitions from partition function zeroes. Journal of Statistical Physics, 102(5–6):1211–1227, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004836227767.

Jin:2004:QSC

[JK04] B.-Q. Jin and V. E. Korepin. Quantum spin chain, Toeplitz determinants and the Fisher–Hartwig conjecture. Journal of Statistical Physics, 116(1–4):79–95, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037230.37166.42.

Jimbo:2001:FFC

[JKO^{+}01] Michio Jimbo, Hitoshi Konno, Satoru Odake, Yaroslav Pugai, and Jun’ichi Shiraishi. Free field construction for the ABF models in regime II. Journal of Statistical Physics, 102(3–4):883–921, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004807104626.

Jansons:2000:ENS

[JL00a] Kalvis M. Jansons and G. D. Lythe. Efficient numerical solution of stochastic differential equations using exponential timestepping. Journal of Statistical Physics, 100(5–6):1097– 1109, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018711024740.

Jitomirskaya:2000:ZDS

[JL00b] S. Jitomirskaya and M. Landrigan. Zero-dimensional spectral measures for quasi-periodic operators with analytic potential. Journal of Statistical Physics, 100(3–4):791–796, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018635811535.

Jancovici:2001:BFT

[JL01] B. Jancovici and J. L. Lebowitz. Bounded fluctuations and translation symmetry breaking: A solvable model. Journal of Statistical Physics, 103(3–4):619–624, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010349517967.

Jedrzejewski:2000:GSL

[JM00] Janusz Jȩdrzejewski and Jacek Miȩkisz. Ground states of lattice gases with ‘almost’ convex repulsive interactions. Journal of Statistical Physics, 98(3–4):589–620, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018663106236.

Jenkinson:2005:ZTL

[JMU05] O. Jenkinson, R. D. Mauldin, and M. Urbański. Zero temperature limits of Gibbs-equilibrium states for countable alphabet subshifts of finite type. Journal of Statistical Physics, 119 (3–4):765–776, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3035-z.

Jiang:2009:PSB

[JNH^{+}09] Rui Jiang, Katsuhiro Nishinari, Mao-Bin Hu, Yong-Hong Wu, and Qing-Song Wu. Phase separation in a bidirectional two-lane asymmetric exclusion process. Journal of Statistical Physics, 136(1):73–88, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9770-9.

Jaksic:2006:LRT

[JOP06] V. Jaksić, Y. Ogata, and C.-A. Pillet. Linear response theory for thermally driven quantum open systems. Journal of Statistical Physics, 123(3):547–569, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9075-1.

Jaksic:2001:NEL

[JP01] V. Jaksić and C.-A. Pillet. A note on eigenvalues of Liouvilleans. Journal of Statistical Physics, 105(5–6):937–941, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013561529682.

Jaksic:2002:MTN

[JP02] V. Jaksić and C.-A. Pillet. Mathematical theory of nonequilibrium quantum statistical mechanics. Journal of Statistical Physics, 108(5–6):787–829, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019818909696.

Jaeck:2009:NBE

[JPZ09] Thomas Jaeck, Joseph V. Pulé, and Valentin A. Zagrebnov. On the nature of Bose–Einstein condensation in disordered systems. Journal of Statistical Physics, 137(1):19–55, October 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9825-y.

Jana:2006:GRE

[JR06] N. K. Jana and B. V. Rao. Generalized random energy model. Journal of Statistical Physics, 123(5):1033–1058, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9043-9.

Jana:2007:GRE

[JR07] N. K. Jana and B. V. Rao. Generalized random energy model II. Journal of Statistical Physics, 127(4):841–850, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9288-y.

Jaroszewska:2008:HDI

[JR08] Joanna Jaroszewska and Michal Rams. On the Hausdorff dimension of invariant measures of weakly contracting on average measurable IFS. Journal of Statistical Physics, 132(5): 907–919, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9566-3.

Jacobsen:2001:TMP

[JS01a] Jesper Lykke Jacobsen and Jesús Salas. Transfer matrices and partition-function zeros for antiferromagnetic Potts models. II. Extended results for square-lattice chromatic polynomial. Journal of Statistical Physics, 104(3–4):701–723, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010328721905.

Jancovici:2001:CCC

[JS01b] B. Jancovici and L. Samaj. Charge correlations in a Coulomb system along a plane wall: A relation between asymptotic behavior and dipole moment. Journal of Statistical Physics, 105(1– 2):193–209, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012286127824.

Jancovici:2001:CSI

[JS01c] B. Jancovici and L. Samaj. Coulomb systems with ideal dielectric boundaries: Free fermion point and universality. Journal of Statistical Physics, 104(3–4):753–775, August 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010332822814.

Jin:2001:RBE

[JS01d] Shi Jin and Marshall Slemrod. Regularization of the Burnett equations via relaxation. Journal of Statistical Physics, 103(5– 6):1009–1033, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010365123288.

Johnson:2001:ERV

[JS01e] Oliver Johnson and Yurii Suhov. Entropy and random vectors. Journal of Statistical Physics, 104(1–2):145–165, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010353526846.

Jancovici:2004:MCD

[JS04] B. Jancovici and L. Samaj. Microscopic calculation of the dielectric susceptibility tensor for Coulomb fluids II. Journal of Statistical Physics, 114(5–6):1211–1234, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013972.61656.65.

Jacobsen:2006:TMP

[JS06] Jesper Lykke Jacobsen and Jesús Salas. Transfer matrices and partition-function zeros for antiferromagnetic Potts models. Journal of Statistical Physics, 122(4):705–760, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-8077-8.

Jacobsen:2008:BCP

[JS08a] Jesper Lykke Jacobsen and Hubert Saleur. Boundary chromatic polynomial. Journal of Statistical Physics, 132(4):707–719, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9585-0.

Jancovici:2008:GCP

[JS08b] Bernard Jancovici and Ladislav Samaj. Guest charge and potential fluctuations in two-dimensional classical Coulomb systems. Journal of Statistical Physics, 131(4):613–629, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9515-1.

Jeng:2008:SJP

[JS08c] M. Jeng and J. M. Schwarz. On the study of Jamming percolation. Journal of Statistical Physics, 131(4):575–595, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9514-2.

Jacobsen:2003:TMP

[JSS03] Jesper Lykke Jacobsen, Jesús Salas, and Alan D. Sokal. Transfer matrices and partition-function zeros for antiferromagnetic Potts models. III. Triangular-lattice chromatic polynomial. Journal of Statistical Physics, 112(5–6):921–1017, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024611424456.

Jacobsen:2005:SFS

[JSS05] Jesper Lykke Jacobsen, Jesús Salas, and Alan D. Sokal. Spanning forests and the q-state Potts model in the limit q → 0. Journal of Statistical Physics, 119(5–6):1153–1281, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-4409-y.

Jancovici:2004:CFC

[JT04] B. Jancovici and G. Téllez. Charge fluctuations for a Coulomb fluid in a disk on a pseudosphere. Journal of Statistical Physics, 116(1–4):205–230, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037207.05672.5b.

Jung:2003:ERM

[Jun03] Paul Jung. Extremal reversible measures for the exclusion process. Journal of Statistical Physics, 112(1–2):165–191, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023679620839.

Just:2001:EPT

[Jus01] Wolfram Just. Equilibrium phase transitions in coupled map lattices: A pedestrian approach. Journal of Statistical Physics, 105(1–2):133–142, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012230010077.

Josic:2000:DRD

[JW00] Kresimir Josić and C. Eugene Wayne. Dynamics of a ring of diffusively coupled Lorenz oscillators. Journal of Statistical Physics, 98(1–2):1–30, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018600203530.

Jackson:2004:GAB

[JW04] R. K. Jackson and M. I. Weinstein. Geometric analysis of bifurcation and symmetry breaking in a Gross–Pitaevskii equation. Journal of Statistical Physics, 116(1–4):881–905, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000037238.94034.75.

Junk:2005:AAL

[JY05] Michael Junk and Zhaoxia Yang. Asymptotic analysis of lattice Boltzmann boundary conditions. Journal of Statistical Physics, 121(1–2):3–35, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-8321-2.

Kadanoff:2003:IE

[Kad03a] Leo P. Kadanoff. Imports and exports. Journal of Statistical Physics, 111(5–6):1391–1396, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023020703863.

Kadanoff:2003:IDC

[Kad03b] Leo P. Kadanoff. Intelligent design and complexity research. An essay and book review of the book: No Free Lunch: Why Specified Complexity Cannot be Purchased without Intelligence, William A. Dembski, 429 pp., Roman and Littlefield, 2002. Journal of Statistical Physics, 110(1–2):451–455, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021043400996.

Kadanoff:2006:PFR

[Kad06] Leo P. Kadanoff. Pulled fronts and the reproductive individual. Journal of Statistical Physics, 122(6):1293–1296, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9073-3.

Kadanoff:2009:MSP

[Kad09] Leo P. Kadanoff. More is the same; phase transitions and mean field theories. Journal of Statistical Physics, 137(5–6): 777–797, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9814-1.

Kanter:2001:BRB

[Kan01] Ido Kanter. Book review: Statistical Mechanics of Learning. By A. Engel and C. Van den Broeck, Cambridge University Press. Journal of Statistical Physics, 105(3–4):719–720, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012240212301.

Kaplan:2006:CP

[Kap06] T. A. Kaplan. The chemical potential. Journal of Statistical Physics, 122(6):1237–1260, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8067-x.

Karwowski:2007:GME

[Kar07] Andrzej Karwowski. Grad’s 13 moment equations in a modified form. Journal of Statistical Physics, 128(3):667–698, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9216-6.

Kashaev:2001:NCQ

[Kas01] R. M. Kashaev. The non-compact quantum dilogarithm and the Baxter equations. Journal of Statistical Physics, 102(3–4): 923–936, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004859121464.

Kastner:2002:EOP

[Kas02] Michael Kastner. Existence and order of the phase transition of the Ising model with fixed magnetization. Journal of Statistical Physics, 109(1–2):133–142, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019983427289.

Kim:2003:CCB

[KAS03] Young C. Kim, Mikhail A. Anisimov, and Jan V. Sengers. Crossover critical behavior in the three-dimensional Ising model. Journal of Statistical Physics, 110(3–6):591–609, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022199516676.

Kawasaki:2003:DVW

[Kaw03] Kyozi Kawasaki. Dynamical van der Waals model of glassy behavior. Journal of Statistical Physics, 110(3–6):1249–1304, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022161330306.

Kawasaki:2006:PME

[Kaw06] Kyozi Kawasaki. Principle of maximum entropy and reduced dynamics. Journal of Statistical Physics, 123(4):711–740, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9121-z.

Kirkpatrick:2002:LTT

[KBS02] T. R. Kirkpatrick, D. Belitz, and J. V. Sengers. Long-time tails, weak localization, and classical and quantum critical behavior. Journal of Statistical Physics, 109(3–4):373–405, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020485809093.

Kincaid:2002:NPS

[KC02] John M. Kincaid and E. G. D. Cohen. Nano- and pico-scale transport phenomena in fluids. Journal of Statistical Physics, 109(3–4):361–371, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020433725023.

Kurzynski:2003:MFP

[KC03] M. Kurzyński and P. Chelminiak. Mean first-passage time in the stochastic theory of biochemical processes. Application to actomyosin molecular motor. Journal of Statistical Physics, 110(1– 2):137–181, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021018728312.

Keskin:2007:DDQ

[KCK07] Mustafa Keskin, Osman Canko, and Muharrem Kirak. Dynamic dipole and quadrupole phase transitions in the kinetic spin-3/2 model. Journal of Statistical Physics, 127(2):359–380, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9236-2.

Kalmykov:2008:MEP

[KCT08] Yuri P. Kalmykov, William T. Coffey, and Serguey V. Titov. Master equation in phase space for a uniaxial spin system. Journal of Statistical Physics, 131(5):969–987, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9522-2.

Klages:2000:DDD

[KD00] R. Klages and Christoph Dellago. Density-dependent diffusion in the periodic Lorentz gas. Journal of Statistical Physics, 101(1– 2):145–159, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026445601619.

Kehrwald:2005:LBS

[Keh05] Dirk Kehrwald. Lattice Boltzmann simulation of shear-thinning fluids. Journal of Statistical Physics, 121(1–2):223–237, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-5963-z.

Kennedy:2002:FIP

[Ken02] Tom Kennedy. A faster implementation of the pivot algorithm for self-avoiding walks. Journal of Statistical Physics, 106(3–4): 407–429, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013750203191.

Kennedy:2004:CIS

[Ken04] Tom Kennedy. Conformal invariance and stochastic Loewner evolution predictions for the 2D self-avoiding walk — Monte Carlo tests. Journal of Statistical Physics, 114(1–2):51–78, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003104.35024.f9.

Kennedy:2007:FAS

[Ken07] Tom Kennedy. A fast algorithm for simulating the chordal Schramm–Loewner evolution. Journal of Statistical Physics, 128(5):1125–1137, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9358-1.

Kennedy:2008:CLD

[Ken08] Tom Kennedy. Computing the Loewner driving process of random curves in the half plane. Journal of Statistical Physics, 131(5):803–819, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9535-x.

Kennedy:2009:NCS

[Ken09] Tom Kennedy. Numerical computations for the Schramm– Loewner evolution. Journal of Statistical Physics, 137(5–6): 839–856, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9866-2.

Krug:2000:APS

[KG00] J. Krug and J. García. Asymmetric particle systems on R. Journal of Statistical Physics, 99(1–2):31–55, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018688421856.

Krattenthaler:2003:VWF

[KG03] Christian Krattenthaler and Anthony J. Guttmann. Vicious walkers, friendly walkers, and Young tableaux. III. Between two walls. Journal of Statistical Physics, 110(3–6):1069–1086, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022192709833.

Kalyuzhnyi:2000:HAE

[KHV00] Yu. V. Kalyuzhnyi, M. F. Holovko, and V. Vlachy. Highly asymmetric electrolytes in the associative mean-spherical approximation. Journal of Statistical Physics, 100(1–2):243–265, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018699914319.

Kralj-Iglic:2006:QOP

[KIBG^{+}06] Veronika Kralj-Iglic, Blaz Babnik, Dorit R. Gauger, Sylvio May, and Ales Iglic. Quadrupolar ordering of phospholipid molecules in narrow necks of phospholipid vesicles. Journal of Statistical Physics, 125(3):727–752, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9051-9.

Kiessling:2004:EFTa

[Kie04a] Michael K.-H. Kiessling. Electromagnetic field theory without divergence problems 1. The Born legacy. Journal of Statistical Physics, 116(1–4):1057–1122, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037250.72634.2a.

Kiessling:2004:EFTb

[Kie04b] Michael K.-H. Kiessling. Electromagnetic field theory without divergence problems 2. A least invasively quantized theory.

Journal of Statistical Physics, 116(1–4):1123–1159, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037251.24558.5c.

Kiessling:2009:MQG

[Kie09a] Michael K.-H. Kiessling. Monotonicity of quantum ground state energies: Bosonic atoms and stars. Journal of Statistical Physics, 137(5–6):1063–1078, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9843-9.

Kiessling:2009:NCG

[Kie09b] Michael K.-H. Kiessling. A note on classical ground state energies. Journal of Statistical Physics, 136(2):275–284, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-009-9769-2.

Kiessling:2009:RCT

[Kie09c] Michael K.-H. Kiessling. On Ruelle’s construction of the thermodynamic limit for the classical microcanonical entropy. Journal of Statistical Physics, 134(1):19–25, January 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9638-4.

Katori:2008:TBB

[KIK08] Makoto Katori, Minami Izumi, and Naoki Kobayashi. Two Bessel bridges conditioned never to collide, double Dirichlet series, and Jacobi theta function. Journal of Statistical Physics, 131(6):1067–1083, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9524-0.

Khorrami:2000:EDP

[KK00a] M. Khorrami and V. Karimipour. Exact determination of the phase structure of a multi-species asymmetric exclusion process. Journal of Statistical Physics, 100(5–6):999–1030, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018758907902.

Konno:2000:EAE

[KK00b] Norio Konno and Makoto Katori. Extension of the Arrowsmith– Essam formula to the Domany–Kinzel model. Journal of Statistical Physics, 101(3–4):747–774, November 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026421520352.

Kawasaki:2002:EDT

[KK02a] Kyozi Kawasaki and Bongsoo Kim. Out of equilibrium dynamics of the toy model with mode coupling and trivial Hamiltonian. Journal of Statistical Physics, 109(3–4):591–606, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020458430474.

Komorowski:2002:EEI

[KK02b] T. Komorowski and G. Krupa. Erratum: ‘On the Existence of Invariant Measure for Lagrangian Velocity in Compressible Environments,’ Journal of Statistical Physics, Vol. 106, Nos. 3/ 4, February 2002. Journal of Statistical Physics, 109(1–2):341, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1019908200881. See [KK02c].

Komorowski:2002:EIM

[KK02c] T. Komorowski and G. Krupa. On the existence of invariant measure for Lagrangian velocity in compressible environments. Journal of Statistical Physics, 106(3–4):635–651, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/

A%3A1013762406825. See erratum [KK02b].

Kim:2008:CNR

[KK08] Jin Min Kim and Dae Ho Kim. Conserved noise restricted curvature model. Journal of Statistical Physics, 131(6):1179–1184, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9525-z.

Kuznetsov:2005:MCS

[KKS05] S. P. Kuznetsov, A. P. Kuznetsov, and I. R. Sataev. Multiparameter critical situations, universality and scaling in twodimensional period–doubling maps. Journal of Statistical Physics, 121(5–6):697–748, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-6973-6.

Katori:2000:SPD

[KKT00] Makoto Katori, Norio Konno, and Hideki Tanemura. Survival probabilities for discrete-time models in one dimension. Journal of Statistical Physics, 99(1–2):603–612, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018617328216.

Klumpp:2003:TMM

[KL03] Stefan Klumpp and Reinhard Lipowsky. Traffic of molecular motors through tube-like compartments. Journal of Statistical Physics, 113(1–2):233–268, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1025778922620.

Kiessling:2006:LFP

[KL06] Michael Kiessling and Carlo Lancellotti. The linear Fokker– Planck equation for the Ornstein–Uhlenbeck process as an (almost) nonlinear kinetic equation for an isolated N-particle system. Journal of Statistical Physics, 123(3):525–546, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9057-3.

Kiessling:2008:LFP

[KL08] Michael Kiessling and Carlo Lancellotti. The linear Fokker– Planck equation for the Ornstein–Uhlenbeck process as an (almost) nonlinear kinetic equation for an isolated N-particle system. Journal of Statistical Physics, 130(4):827, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9447-1.

Keller:2009:REE

[KL09] Gerhard Keller and Carlangelo Liverani. Rare events, escape rates and quasistationarity: Some exact formulae. Journal of Statistical Physics, 135(3):519–534, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9747-8.

Klauder:2004:ONP

[Kla04] John R. Klauder. Overcoming Nonrenormalizability.Part 2. Journal of Statistical Physics, 116(1–4):1181–1187, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037237.78375.e8.

Krafczyk:2005:PMM

[KLL05] Manfred Krafczyk, Anthony J. C. Ladd, and Li-Shi Luo. Preface: Mesoscopic methods in engineering and science. Journal of Statistical Physics, 121(1–2):1–2, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-9107-2.

Kuna:2007:RPP

[KLS07] T. Kuna, J. L. Lebowitz, and E. R. Speer. Realizability of point processes. Journal of Statistical Physics, 129(3):417– 439, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9393-y.

Kurka:2000:LSC

[KM00] Petr Kůrka and Alejandro Maass. Limit sets of cellular automata associated to probability measures. Journal of Statistical Physics, 100(5–6):1031–1047, September 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018706923831.

Kotecky:2001:FSS

[KM01] R. Kotecký and I. Medved’. Finite-size scaling for the 2D Ising model with minus boundary conditions. Journal of Statistical Physics, 104(5–6):905–943, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010495725329.

Klein:2006:SEA

[KM06] Abel Klein and Stanislav Molchanov. Simplicity of eigenvalues in the Anderson model. Journal of Statistical Physics, 122 (1):95–99, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8009-7.

Kuster:2007:IPT

[KM07] Jens Küster and Gernot Münster. The interfacial profile in two-loop order. Journal of Statistical Physics, 129(3):441– 451, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9404-z.

Kopf:2008:IRF

[KM08a] Michael H. Köpf and Gernot Münster. Interfacial roughening in field theory. Journal of Statistical Physics, 132(3):417–430, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9572-5.

Kung:2008:MLD

[KM08b] William Kung and M. Cristina Marchetti. Mode-locking in driven disordered systems as a boundary-value problem. Journal of Statistical Physics, 132(5):945–955, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9573-4.

Klenke:2009:MIE

[KM09a] Achim Klenke and Peter Mörters. Multiple intersection exponents for planar Brownian motion. Journal of Statistical Physics, 136(2):373–397, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9780-7.

Korada:2009:ESG

[KM09b] Satish Babu Korada and Nicolas Macris. Exact solution of the gauge symmetric p-spin glass model on a complete graph. Journal of Statistical Physics, 136(2):205–230, July 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9781-6.

Kupiainen:2007:SRS

[KMG07] A. Kupiainen and P. Muratore-Ginanneschi. Scaling, renormalization and statistical conservation laws in the Kraichnan model of turbulent advection. Journal of Statistical Physics, 126(3):

669–724, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9205-9.

Krstic:2007:MPO

[KMP^{+}07] Vladimir Krstić, Zeljka Maglica, Hana Cipcić Paljetak, Boris Podobnik, and Nenad Pavin. Min-protein oscillations in E. coli: Three-dimensional off-lattice stochastic reaction–diffusion model. Journal of Statistical Physics, 128(1–2):5–20, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9189-5.

Kuznetsov:2008:BNC

[KMS08] S. P. Kuznetsov, A. A. Mailybaev, and I. R. Sataev. Birth of a new class of period-doubling scaling behavior as a result of bifurcation in the renormalization equation. Journal of Statistical Physics, 130(3):599–616, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9442-6.

Klumper:2001:SXC

[KMSS01] A. Klümper, J. R. Reyes Martínez, C. Scheeren, and M. Shiroishi. The spin-1/2 XXZ chain at finite magnetic field: Crossover phenomena driven by temperature. Journal of Statistical Physics, 102(3–4):937–951, February 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004811305534.

Kalinay:2000:SMS

[KMST00] P. Kalinay, P. Markos, L. Samaj, and I. Travenec. The sixthmoment sum rule for the pair correlations of the two-dimensional one-component plasma: Exact result. Journal of Statistical Physics, 98(3–4):639–666, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018667207145.

Kramer:2003:CAP

[KMVE03] Peter R. Kramer, Andrew J. Majda, and Eric Vanden-Eijnden. Closure approximations for passive scalar turbulence: A comparative study on an exactly solvable model with complex features. Journal of Statistical Physics, 111(3–4):565–679, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022837913026.

Kaminaga:2003:LRQ

[KN03] Masahiro Kaminaga and Fumihiko Nakano. The Landauer resistivity on quantum wires. Journal of Statistical Physics, 111 (1–2):339–353, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022265226479.

Kager:2004:GSL

[KN04] Wouter Kager and Bernard Nienhuis. A guide to stochastic Löwner evolution and its applications. Journal of Statistical Physics, 115(5–6):1149–1229, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028058.87266.be.

Kager:2004:ESL

[KNK04] Wouter Kager, Bernard Nienhuis, and Leo P. Kadanoff. Exact solutions for Loewner evolutions. Journal of Statistical Physics, 115(3–4):805–822, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000022380.93241.24.

Keldermann:2007:PRD

[KNP07] R. H. Keldermann, M. P. Nash, and A. V. Panfilov. Pacemakers in a reaction–diffusion mechanics system. Journal of Statistical Physics, 128(1–2):375–392, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-006-9219-3.pdf.

Komatsu:2009:RNS

[KNST09] Teruhisa S. Komatsu, Naoko Nakagawa, Shin-Ichi Sasa, and Hal Tasaki. Representation of nonequilibrium steady states in large mechanical systems. Journal of Statistical Physics, 134 (2):401–423, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9678-4.pdf.

Komorowski:2002:SBP

[KO02] Tomasz Komorowski and Stefano Olla. On the superdiffusive behavior of passive tracer with a Gaussian drift. Journal of Statistical Physics, 108(3–4):647–668, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1015734109076.

Komorowski:2005:MER

[KO05] Tomasz Komorowski and Stefano Olla. On mobility and Einstein relation for tracers in time-mixing random environments. Journal of Statistical Physics, 118(3–4):407–435, February 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8815-3.

Kohn:2001:PT

[Koh01] Walter Kohn. Periodic thermodynamics. Journal of Statistical Physics, 103(3–4):417–423, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010327828445.

Koibuchi:2007:PTM

[Koi07a] Hiroshi Koibuchi. Phase transition of meshwork models for spherical membranes. Journal of Statistical Physics, 129(4): 605–621, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9385-y.

Koibuchi:2007:PTT

[Koi07b] Hiroshi Koibuchi. Phase transition of triangulated spherical surfaces with elastic skeletons. Journal of Statistical Physics, 127 (3):457–470, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9287-z.

Kolokoltsov:2004:HLC

[Kol04] Vassili N. Kolokoltsov. Hydrodynamic limit of coagulation– fragmentation type models of k-nary interacting particles. Journal of Statistical Physics, 115(5–6):1621–1653, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028071.96950.12.

Kolokoltsov:2007:NMS

[Kol07] Vassili N. Kolokoltsov. Nonlinear Markov semigroups and interacting Lévy type processes. Journal of Statistical Physics, 126 (3):585–642, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9211-y.

Kolesnik:2008:RMF

[Kol08] Alexander D. Kolesnik. Random motions at finite speed in higher dimensions. Journal of Statistical Physics, 131(6): 1039–1065, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9532-0.

Koma:2000:IQH

[Kom00a] Tohru Koma. Insensitivity of quantized Hall conductance to disorder and interactions. Journal of Statistical Physics, 99(1– 2):383–459, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018657009561.

Koma:2000:SGQ

[Kom00b] Tohru Koma. Spectral gaps of quantum Hall systems with interactions. Journal of Statistical Physics, 99(1–2):313–381, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018604925491.

Komarova:2007:LGF

[Kom07] Natalia L. Komarova. Loss- and gain-of-function mutations in cancer: Mass-action, spatial and hierarchical models. Journal of Statistical Physics, 128(1–2):413–446, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9238-0.

Koma:2008:WHC

[Kom08] Tohru Koma. Widths of the Hall conductance plateaus. Journal of Statistical Physics, 130(5):843–934, March 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9432-8.

Konno:2002:DCF

[Kon02a] Norio Konno. Dualities for a class of finite range probabilistic cellular automata in one dimension. Journal of Statistical Physics, 106(5–6):915–922, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1014025501317.

Konno:2002:SDM

[Kon02b] Norio Konno. Self-duality for multi-state probabilistic cellular automata with finite range interactions. Journal of Statistical Physics, 106(5–6):923–930, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1014077518155.

Koralov:2000:TVF

[Kor00] Leonid Koralov. Transport by vector fields with Kolmogorov spectrum. Journal of Statistical Physics, 98(1–2):405–418, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018691325820.

Kozak:2000:RWF

[Koz00] John J. Kozak. Random walks on a fractal solid. Journal of Statistical Physics, 101(1–2):405–414, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026474407979.

Kastner:2001:CCB

[KP01a] M. Kastner and M. Promberger. Comment on ‘A Comparison Between Broad Histogram and Multicanonical Methods’. Journal of Statistical Physics, 104(3–4):893–894, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/

A%3A1010393008701. See [LdOP00, LPdO01].

Kastner:2001:CMC

[KP01b] M. Kastner and M. Promberger. Consistency of microcanonical and canonical finite-size scaling. Journal of Statistical Physics, 103(5–6):893–902, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010304804632.

Komorowski:2004:TPT

[KP04] Tomasz Komorowski and Szymon Peszat. Transport of a passive tracer by an irregular velocity field. Journal of Statistical Physics, 115(5–6):1361–1388, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028063.58764.68.

Kuksin:2005:FBR

[KP05] Sergei Kuksin and Oliver Penrose. A family of balance relations for the two-dimensional Navier–Stokes equations with random forcing. Journal of Statistical Physics, 118(3–4):437–449, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8816-2.

Kalinay:2006:EDR

[KP06] Pavol Kalinay and Jerome K. Percus. Exact dimensional reduction of linear dynamics: Application to confined diffusion. Journal of Statistical Physics, 123(5):1059–1069, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9081-3.

Kozitsky:2007:EGM

[KP07] Yuri Kozitsky and Tatiana Pasurek. Euclidean Gibbs measures of interacting quantum anharmonic oscillators. Journal of Statistical Physics, 127(5):985–1047, June 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

006-9274-9. See addendum and corrigendum [KP08b].

Klopp:2008:LQG

[KP08a] Frédéric Klopp and Konstantin Pankrashkin. Localization on quantum graphs with random vertex couplings. Journal of Statistical Physics, 131(4):651–673, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9517-z.

Kozitsky:2008:ACE

[KP08b] Yuri Kozitsky and Tatiana Pasurek. Addendum and corrigendum to ‘Euclidean Gibbs Measures of Interacting Quantum Anharmonic Oscillators’. Journal of Statistical Physics, 132(4):

755–757, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9531-1. See [KP07].

Kastner:2000:MFS

[KPH00] Michael Kastner, Michael Promberger, and Alfred Hüller. Microcanonical finite-size scaling. Journal of Statistical Physics, 99 (5–6):1251–1264, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018636705716.

Katsoulakis:2005:MMC

[KPT05] Markos A. Katsoulakis, Petr Plechác, and Dimitrios K. Tsagkarogiannis. Mesoscopic modeling for continuous spin lattice systems: Model problems and micromagnetics applications. Journal of Statistical Physics, 119(1–2):347–389, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-2126-6.

Kruis:2006:SDE

[KPvB06] H. V. Kruis, Debabrata Panja, and Henk van Beijeren. Systematic density expansion of the Lyapunov exponents for a two-dimensional random Lorentz gas. Journal of Statistical Physics, 124(2–4):823–842, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-005-9001-y.

Kruglikov:2006:DEZ

[KR06] B. Kruglikov and M. Rypdal. Dynamics and entropy in the Zhang model of self-organized criticality. Journal of Statistical Physics, 122(5):975–1039, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9011-9.

Kirsch:2007:WFB

[KR07] R. Kirsch and S. Rjasanow. A weak formulation of the Boltzmann equation based on the Fourier transform. Journal of Statistical Physics, 129(3):483–492, November 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9374-1.

Kramer:2003:TDR

[Kra03] Peter R. Kramer. Two different rapid decorrelation in time limits for turbulent diffusion. Journal of Statistical Physics, 110(1– 2):87–136, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021066611474.

Krause:2009:DKF

[Kra09] Paul Krause. The diffusion kernel filter. Journal of Statistical Physics, 134(2):365–380, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9673-1.

Kierlik:2000:SCO

[KRT00] E. Kierlik, M. L. Rosinberg, and G. Tarjus. A self-consistent Ornstein–Zernike approximation for the Edwards–Anderson spin-glass model. Journal of Statistical Physics, 100(1–2):423– 443, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018612317044.

Kaufmann:2000:TCC

[KS00] Z. Kaufmann and P. Szépfalusy. Transient chaos and critical states in generalized Baker maps. Journal of Statistical Physics, 101(1–2):107–124, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026459432086.

Kalinay:2002:TPT

[KS02] P. Kalinay and L. Samaj. Thermodynamic properties of the two-dimensional Coulomb gas in the low-density limit. Journal of Statistical Physics, 106(5–6):857–874, March 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014088716338.

Kastner:2006:MFS

[KS06] Michael Kastner and Oliver Schnetz. On the mean-field spherical model. Journal of Statistical Physics, 122(6):1195–1214, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8031-9.

Kessler:2007:ERF

[KS07] David A. Kessler and Nadav M. Shnerb. Extinction rates for fluctuation-induced metastabilities: A real-space WKB approach. Journal of Statistical Physics, 127(5):861–886, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9312-2.

Klep:2008:SHS

[KS08] Igor Klep and Markus Schweighofer. Sums of Hermitian squares and the BMV conjecture. Journal of Statistical Physics, 133(4): 739–760, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9632-x.

Kang:2009:DWP

[KS09] Yang Kang and Jeffrey Schenker. Diffusion of wave packets in a Markov random potential. Journal of Statistical Physics, 134(5– 6):1005–1022, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9714-4.

Khain:2007:RCC

[KSSM07] Evgeniy Khain, Leonard M. Sander, and Casey M. Schneider-Mizell. The role of Cell–Cell adhesion in wound healing. Journal of Statistical Physics, 128(1–2):209–218, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9194-8.

Katsoulakis:2006:ILC

[KT06] Markos A. Katsoulakis and José Trashorras. Information loss in coarse-graining of stochastic particle dynamics. Journal of Statistical Physics, 122(1):115–135, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8063-1.

Katori:2007:NBM

[KT07] Makoto Katori and Hideki Tanemura. Noncolliding Brownian motion and determinantal processes. Journal of Statistical Physics, 129(5–6):1233–1277, December 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9421-y.

Katori:2009:ZAF

[KT09] Makoto Katori and Hideki Tanemura. Zeros of Airy function and relaxation process. Journal of Statistical Physics, 136(6): 1177–1204, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9829-7.

Korner:2005:LBM

[KTH^{+}05] C. Körner, M. Thies, T. Hofmann, N. Thürey, and U. Rüde. Lattice Boltzmann model for free surface flow for modeling foaming. Journal of Statistical Physics, 121(1–2):179–196, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8879-8.

Kuchanov:2006:CES

[KTZ06] Semion I. Kuchanov, Konstantin V. Tarasevich, and Timur V. Zharnikov. Configurational effects in statistical theory of branched non-random polycondensation. Journal of Statistical Physics, 122(5):875–908, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9016-4.

Kuksin:2004:ELS

[Kuk04] Sergei B. Kuksin. The Eulerian limit for 2D statistical hydrodynamics. Journal of Statistical Physics, 115(1–2):469–492, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000019830.64243.a2.

Kuksin:2006:RBR

[Kuk06] Sergei B. Kuksin. Remarks on the balance relations for the twodimensional Navier–Stokes equation with random forcing. Journal of Statistical Physics, 122(1):101–114, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8084-9.

Kulske:2001:GNR

[Kül01] Christof Külske. On the Gibbsian nature of the random field Kac model under block-averaging. Journal of Statistical Physics, 104(5–6):991–1012, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010497510308.

Kulske:2003:ANG

[Kül03] Christof Külske. Analogues of non–Gibbsianness in joint measures of disordered mean field models. Journal of Statistical Physics, 112(5–6):1079–1108, September 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024615625364.

Kupferman:2004:FKK

[Kup04] Raz Kupferman. Fractional kinetics in Kac–Zwanzig heat bath models. Journal of Statistical Physics, 114(1–2):291–326, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003113.22621.f0.

Kurkova:2003:TDG

[Kur03] I. Kurkova. Temperature dependence of the Gibbs state in the random energy model. Journal of Statistical Physics, 111 (1–2):35–56, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022244721936.

Kruse:2000:DAI

[KVM^{+}00] Norbert Kruse, Christian Voss, Valentin Medvedev, Christopher Bodenstein, David Hanon, and Jean Pierre Boon. Dynamics of adsorbate islands with nanoscale resolution. Journal of Statistical Physics, 101(1–2):621–629, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026407405872.

Katsov:2000:DLS

[KW00a] Kirill Katsov and John D. Weeks. Determining liquid structure from the tail of the direct correlation function. Journal of Statistical Physics, 100(1–2):107–134, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018683410684.

Koprucki:2000:NEG

[KW00b] Thomas Koprucki and Heinz-Jürgen Wagner. New exact ground states for one-dimensional quantum many-body systems. Journal of Statistical Physics, 100(3–4):779–790, August 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018683727464.

King:2002:NCD

[KW02] C. King and F. Y. Wu. New correlation duality relations for the planar Potts model. Journal of Statistical Physics, 107(3–4):919– 940, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014550516842.

Kytola:2006:CFT

[Kyt06] Kalle Kytölä. On conformal field theory of SLE(κ, ρ). Journal of Statistical Physics, 123(6):1169–1181, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9138-3.

Kleban:2003:CPM

[KZ03a] Peter Kleban and Don Zagier. Crossing probabilities and modular forms. Journal of Statistical Physics, 113(3–4):431– 454, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026012600583.

Kuchanov:2003:SBP

[KZ03b] Semion I. Kuchanov and Timur V. Zharnikov. Stochastic branching process for description of non-random irreversible polycondensation. Journal of Statistical Physics, 111(5–6): 1273–1298, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023008401138.

Kondratiev:2007:SAS

[KZ07] Yu. Kondratiev and E. Zhizhina. Spectral analysis of a stochastic Ising model in continuum. Journal of Statistical Physics, 129 (1):121–149, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9363-4.

Langer:2006:SBP

[Lan06] Steve Langer. The Santa Barbara physicist (to the tune of The Modern Major General, with apologies to Gilbert & Sullivan).

Journal of Statistical Physics, 125(5–6):1267–1268, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-0150-9.

Lancellotti:2009:FAV

[Lan09] Carlo Lancellotti. On the fluctuations about the Vlasov limit for n-particle systems with mean-field interactions. Journal of Statistical Physics, 136(4):643–665, August 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9800-7.

Lecomte:2007:TFS

[LARvW07] V. Lecomte, C. Appert-Rolland, and F. van Wijland. Thermodynamic formalism for systems with Markov dynamics. Journal of Statistical Physics, 127(1):51–106, April 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9254-0.

Latz:2002:NEP

[Lat02] A. Latz. Non-equilibrium projection-operator for a quenched thermostatted system. Journal of Statistical Physics, 109(3–4): 607–622, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020410514545.

Lawler:2009:PFL

[Law09] Gregory F. Lawler. Partition functions, loop measure, and versions of SLE. Journal of Statistical Physics, 134(5–6):813–837, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9704-6.

Le:2001:EDN

[LB01] K. C. Le and V. L. Berdichevsky. Energy distribution in a neutral gas of point vortices. Journal of Statistical Physics, 104(3– 4):881–892, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010341024631.

Li:2004:MTM

[LB04] Xiangting Li and David J. Bergman. A mixed-transfer-matrix method for simulating normal Conductor/ perfect Insulator/ perfect conductor random networks. Journal of Statistical Physics, 117(3–4):427–452, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2272-x.

Lipsmeier:2009:RES

[LB09] Florian Lipsmeier and Ellen Baake. Rare event simulation for T-cell activation. Journal of Statistical Physics, 134(3):537–566, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9672-2.

Lynden-Bell:2004:RPP

[LBLB04] D. Lynden-Bell and R. M. Lynden-Bell. Relaxation to a perpetually pulsating equilibrium. Journal of Statistical Physics, 117(1– 2):199–209, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000044068.53435.eb.

Lombardo:2008:NLS

[LCS08] Maria Carmela Lombardo, Russel E. Caflisch, and Marco Sammartino. Non-local scattering kernel and the hydrodynamic limit. Journal of Statistical Physics, 130(1):69–82, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9360-7.

Lubeck:2001:CVE

[LD01] S. Lübeck and D. Dhar. Continuously varying exponents in a sandpile model with dissipation near surface. Journal of Statistical Physics, 102(1–2):1–14, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026538607311.

Lee-Dadswell:2008:DET

[LDNG08] G. R. Lee-Dadswell, B. G. Nickel, and C. G. Gray. Detailed examination of transport coefficients in cubic-plus-quartic oscillator chains. Journal of Statistical Physics, 132(1):1–33, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9551-x.

Lima:2003:UCS

[LdO03] Marcus V. Lima and César R. de Oliveira. Uniform Cantor singular continuous spectrum for nonprimitive Schrödinger operators. Journal of Statistical Physics, 112(1–2):357–374, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023644007635.

Lima:2000:CBB

[LdOP00] A. R. Lima, P. M. C. de Oliveira, and T. J. P. Penna. A comparison between broad histogram and multicanonical methods. Journal of Statistical Physics, 99(3–4):691–705, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/

A%3A1018679025763. See comments [KP01a, LPdO01].

Lebowitz:2002:Pa

[Leb02a] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 107(1–2):1, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014538916088.

Lebowitz:2002:Pb

[Leb02b] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 108(5– 6):721–722, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019753606970.

Lebowitz:2002:Pc

[Leb02c] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 109(3– 4):353–354, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020466324114.

Lebowitz:2003:P

[Leb03] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 110 (3–6):465–466, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1022154929880.

Lebowitz:2004:PSM

[Leb04a] Joel Lebowitz. Program of the 91st Statistical Mechanics Meeting, Rutgers University, May 16–18, 2004. Journal of Statistical Physics, 117(3–4):775–781, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2270-z.

Lebowitz:2004:LEL

[Leb04b] Joel L. Lebowitz. A laudatio for Elliott Lieb on his receiving the Poincaré Medal at the International Congress on Mathematical Physics in Lisbon, July 30, 2003. Journal of Statistical Physics, 116(1–4):9–11, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000037241.49804.8a.

Lebowitz:2004:Pa

[Leb04c] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 115(1–2):1, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000020049.14806.92.

Lebowitz:2004:Pb

[Leb04d] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 116 (1–4):1, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000037252.48115.02.

Lebowitz:2005:P

[Leb05a] Joel Lebowitz. Preface. Journal of Statistical Physics, 121(5– 6):609, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8672-8.

Lebowitz:2005:PSMa

[Leb05b] Joel L. Lebowitz. Program of the 92nd Statistical Mechanics Meeting Rutgers University, December 19–21, 2004. Journal of Statistical Physics, 118(5–6):1265–1271, March 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3024-2.

Lebowitz:2005:PSMb

[Leb05c] Joel L. Lebowitz. Program of the 93rd Statistical Mechanics Meeting — Rutgers University, May 15–17, 2005. Journal of Statistical Physics, 120(5–6):1165–1170, September 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8715-1.

Lebowitz:2006:P

[Leb06a] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 125 (5–6):1013, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9188-6.

Lebowitz:2006:PSMb

[Leb06b] Joel L. Lebowitz. Program of the 94th Statistical Mechanics Meeting Rutgers University, December 18–20, 2005. Journal of Statistical Physics, 125(1):277–282, October 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9084-0.

Lebowitz:2006:PSMa

[Leb06c] Joel L. Lebowitz. Program of the 95th Statistical Mechanics Meeting Rutgers University, May 7–9, 2006. Journal of Statistical Physics, 124(1):261–267, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9150-7.

Lebowitz:2007:P

[Leb07a] Joel Lebowitz. Preface. Journal of Statistical Physics, 129(5– 6):807, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9437-3.

Lebowitz:2007:PSMa

[Leb07b] Joel L. Lebowitz. Program of the 96th Statistical Mechanics Conference Rutgers University, Hill Center, Room 114 Sunday, Monday and Tuesday, December 17–19, 2006. Journal of Statistical Physics, 127(4):853–859, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9306-0.

Lebowitz:2007:PSMb

[Leb07c] Joel L. Lebowitz. Program of the 97th Statistical Mechanics Conference Rutgers University, Hill Center, Room 114 Sunday, Monday and Tuesday, May 6–8, 2007. Journal of Statistical Physics, 129(2):411–415, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9395-9.

Lebowitz:2008:PSMa

[Leb08a] Joel L. Lebowitz. Program of the 98th Statistical Mechanics Conference. Journal of Statistical Physics, 131(3):559–565, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9492-4.

Lebowitz:2008:PSMb

[Leb08b] Joel L. Lebowitz. Program of the 99th Statistical Mechanics Conference Honoring Edouard Brezin & Giorgio Parisi. Journal of Statistical Physics, 132(6):1147–1151, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9592-1.

Lebowitz:2009:P

[Leb09a] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 137(5– 6):775–776, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9890-2.

Lebowitz:2009:PSMa

[Leb09b] Joel L. Lebowitz. Program of the 100th Statistical Mechanics Conference/ DIMACS Workshop. Journal of Statistical Physics, 134(4):797–805, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-009-9690-8.

Lebowitz:2009:PSMb

[Leb09c] Joel L. Lebowitz. Program of the 101st Statistical Mechanics Conference. Journal of Statistical Physics, 136(1):199–203, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9767-4.

Ledoux:2000:DOS

[Led00] M. Ledoux. On the distribution of overlaps in the Sherrington– Kirkpatrick spin glass model. Journal of Statistical Physics, 100(5–6):871–892, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018771210627.

Lee:2005:NAS

[Lee05] T. D. Lee. A new approach to solve the low-lying states of the Schrödinger equation. Journal of Statistical Physics, 121(5–6): 1015–1071, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5476-9.

Lee:2009:GSE

[Lee09] Ji Oon Lee. Ground state energy of dilute Bose gas in small negative potential case. Journal of Statistical Physics, 134(1):1–18, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9661-5.

Leggett:2003:BEC

[Leg03] A. J. Leggett. Bose–Einstein condensation in a harmonic trap: Effect of interactions on T_{c}. Journal of Statistical Physics, 110 (3–6):903–910, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022124222128.

Lemarchand:2000:SAC

[Lem00] A. Lemarchand. Selection of an attractor in a continuum of stable solutions: Descriptions of a wave front at different scales. Journal of Statistical Physics, 101(1–2):579–598, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026430625726.

Lendi:2000:HOQ

[Len00] K. Lendi. Higher order quantum Onsager coefficients from dynamical invariants. Journal of Statistical Physics, 99(3–4): 1037–1043, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018612000264.

Lerman:2000:DPN

[Ler00] L. M. Lerman. Dynamical phenomena near a saddle-focus homoclinic connection in a Hamiltonian system. Journal of Statistical Physics, 101(1–2):357–372, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026411506781.

Levin:2003:TST

[Lev03] Yan Levin. Thermodynamics of surface tension: Application to electrolyte solutions. Journal of Statistical Physics, 110(3– 6):825–834, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022116020311.

Lemanski:2004:CSD

[LFB04] R. Lemański, J. K. Freericks, and G. Banach. Charge stripes due to electron correlations in the two-dimensional spinless Falicov– Kimball model. Journal of Statistical Physics, 116(1–4):699–718, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037213.25834.33.

Liboff:2001:HQB

[LG01] Richard L. Liboff and Joseph Greenberg. The hexagon quantum billiard. Journal of Statistical Physics, 105(1–2):389–402, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012298530550.

Li:2006:QES

[Li06] Gu-Qiang Li. Quantum entropy of spin fields in the Schwarzschild–anti-de Sitter black hole with a global monopole. Journal of Statistical Physics, 125(3):753–760, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9147-2.

Li:2007:RCE

[Li07] Dong Li. On the rate of convergence to equilibrium of the Andersen thermostat in molecular dynamics. Journal of Statistical Physics, 129(2):265–287, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9391-0.

Li:2009:NRR

[Li09] Dong Li. On a nonlinear recurrent relation. Journal of Statistical Physics, 134(4):681–700, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9691-7.

Liang:2006:TFH

[Lia06] Faming Liang. A theory on flat histogram Monte Carlo algorithms. Journal of Statistical Physics, 122(3):511–529, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-8016-8.

Lindner:2004:MFP

[Lin04] Benjamin Lindner. Moments of the first passage time under external driving. Journal of Statistical Physics, 117(3–4): 703–737, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2269-5.

Lindner:2008:DCB

[Lin08] Benjamin Lindner. Diffusion coefficient of a Brownian particle with a friction function given by a power law. Journal of Statistical Physics, 130(3):523–533, February 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-007-9438-2.pdf.

Liu:2009:GTP

[Liu09] Da-Jiang Liu. Generic two-phase coexistence and nonequilibrium criticality in a lattice version of Schlögl’s second model for autocatalysis. Journal of Statistical Physics, 135(1):77–85, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9708-2.

Lipowsky:2003:MMC

[LJ03] Reinhard Lipowsky and Nicole Jaster. Molecular motor cycles: From ratchets to networks. Journal of Statistical Physics, 110(3– 6):1141–1167, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022101011650.

Lee:2006:PSV

[LK06a] Ha Youn Lee and Mehran Kardar. Patterns and symmetries in the visual cortex and in natural images. Journal of Statistical Physics, 125(5–6):1243–1266, December 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9107-x.

Lobaskin:2006:VFD

[LK06b] Dmitry Lobaskin and Stefan Kehrein. Violation of the fluctuation–dissipation theorem and heating effects in the timedependent Kondo model. Journal of Statistical Physics, 123 (2):301–313, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9055-5.

Lohmar:2009:DLR

[LK09] Ingo Lohmar and Joachim Krug. Diffusion-limited reactions and mortal random walkers in confined geometries. Journal of Statistical Physics, 134(2):307–336, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9680-x.

Lam:2009:OET

[LKL09] Khanh-Dang Nguyen Thu Lam, Jorge Kurchan, and Dov Levine. Order in extremal trajectories. Journal of Statistical Physics, 137(5–6):1079–1093, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9828-8.

Lerma:2007:SAC

[LKMGG07] Claudia Lerma, Trine Krogh-Madsen, Michael Guevara, and Leon Glass. Stochastic aspects of cardiac arrhythmias. Journal of Statistical Physics, 128(1–2):347–374, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9191-y.

Lippi:2000:HCT

[LL00] Andrea Lippi and Roberto Livi. Heat conduction in twodimensional nonlinear lattices. Journal of Statistical Physics, 100(5–6):1147–1172, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018721525900.

Lieb:2002:BBE

[LL02] Elliott H. Lieb and Michael Loss. A bound on binding energies and mass renormalization in models of quantum electrodynamics. Journal of Statistical Physics, 108(5–6):1057–1069, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1019870931077.

Lan:2008:DNS

[LL08a] Yueheng Lan and Y. Charles Li. On the dynamics of Navier– Stokes and Euler equations. Journal of Statistical Physics, 132(1):35–76, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9555-6.

Lipowsky:2008:CCM

[LL08b] Reinhard Lipowsky and Steffen Liepelt. Chemomechanical coupling of molecular motors: Thermodynamics, network representations, and balance conditions. Journal of Statistical Physics, 130(1):39–67, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-007-9425-7.pdf.

Langmann:2009:FGR

[LL09a] Edwin Langmann and Göran Lindblad. Fermi’s golden rule and exponential decay as a RG fixed point. Journal of Statistical Physics, 134(4):749–768, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9700-x.

Lipowsky:2009:CCM

[LL09b] Reinhard Lipowsky and Steffen Liepelt. Chemomechanical coupling of molecular motors: Thermodynamics, network representations, and balance conditions. Journal of Statistical Physics, 135(4):777–778, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9746-9.pdf.

Larralde:2003:TPM

[LLMM03] H. Larralde, F. Leyvraz, and C. Mejía-Monasterio. Transport properties of a modified Lorentz gas. Journal of Statistical Physics, 113(1–2):197–231, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1025726905782.

Lockard:2002:EPA

[LLMS02] David P. Lockard, Li-Shi Luo, Seth D. Milder, and Bart A. Singer. Evaluation of PowerFLOW for aerodynamic applications. Journal of Statistical Physics, 107(1–2):423–478, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014539411062.

Langlands:2000:UCI

[LLSA00] Robert P. Langlands, Marc-André Lewis, and Yvan Saint-Aubin. Universality and conformal invariance for the Ising model in domains with boundary. Journal of Statistical Physics, 98(1–2): 131–244, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018674822185.

Lipowsky:2009:ECM

[LLV09] Reinhard Lipowsky, Steffen Liepelt, and Angelo Valleriani. Energy conversion by molecular motors coupled to nucleotide hydrolysis. Journal of Statistical Physics, 135(5–6):951–975, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9754-9.pdf.

Lorinczi:2001:GMB

[LM01] József Lörinczi and Robert A. Minlos. Gibbs measures for Brownian paths under the effect of an external and a small pair potential. Journal of Statistical Physics, 105(3–4):605–647, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1012279826414.

Laurencot:2002:BDL

[LM02] Philippe Laurençot and Stéphane Mischler. From the Becker– Döring to the Lifshitz–Slyozov–Wagner equations. Journal of Statistical Physics, 106(5–6):957–991, March 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014081619064.

Lemaitre:2006:SRQ

[LM06] Anaël Lemaître and Craig Maloney. Sum rules for the quasistatic and visco-elastic response of disordered solids at zero temperature. Journal of Statistical Physics, 123(2):415–453, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-9015-5.

Lebensztayn:2005:IUB

[LMP05] Élcio Lebensztayn, Fábio P. Machado, and Serguei Popov. An improved upper bound for the critical probability of the frog model on homogeneous trees. Journal of Statistical Physics, 119 (1–2):331–345, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2051-8.

Levine:2005:ZRP

[LMS05] E. Levine, D. Mukamel, and G. M. Schütz. Zero-range process with open boundaries. Journal of Statistical Physics, 120(5–6): 759–778, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-7000-7.

Liverani:2007:RCF

[LMV07] C. Liverani, Ph. Marie, and S. Vaienti. Random classical fidelity. Journal of Statistical Physics, 128(4):1079–1091, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9338-5.

Lorenz:2000:SPT

[LMZ00] Christian D. Lorenz, Raechelle May, and Robert M. Ziff. Similarity of percolation thresholds on the HCP and FCC lattices. Journal of Statistical Physics, 98(3–4):961–970, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018648130343.

Landa:2006:CEP

[LNM06] P. S. Landa, Yu. I. Neimark, and P. V. E. McClintock. Changes in the effective parameters of averaged motion in nonlinear systems subject to noise. Journal of Statistical Physics, 125(3): 593–620, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9209-5.

Leimkuhler:2009:GST

[LNT09] Ben Leimkuhler, Emad Noorizadeh, and Florian Theil. A gentle stochastic thermostat for molecular dynamics. Journal of Statistical Physics, 135(2):261–277, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9734-0.

Lo:2008:WLM

[Lo08] Assane Lo. Witten Laplacian method for the decay of correlations. Journal of Statistical Physics, 132(2):355–396, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9547-6.

Los:2005:HGM

[Los05] Victor F. Los. Homogeneous generalized master equations. Journal of Statistical Physics, 119(1–2):241–271, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2722-5.

Lou:2003:SLL

[Lou03] Bendong Lou. Singular limit of a p-Laplacian reaction–diffusion equation with a spatially inhomogeneous reaction term. Journal of Statistical Physics, 110(1–2):377–383, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/

A%3A1021083015108. See erratum [Lou08a].

Loulakis:2005:SDC

[Lou05] Michail Loulakis. On the symmetry of the diffusion coefficient in asymmetric simple exclusion. Journal of Statistical Physics, 119 (3–4):853–860, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3018-0.

Lou:2007:SLS

[Lou07] Bendong Lou. Singular limits of spatially inhomogeneous convection-reaction-diffusion equations. Journal of Statistical Physics, 129(3):509–516, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9400-3.

Lou:2008:ESL

[Lou08a] Bendong Lou. Erratum on ‘Singular Limit of a p-Laplacian Reaction–Diffusion Equation with a Spatially Inhomogeneous Reaction Term’. Journal of Statistical Physics, 133(1):203, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9611-2. See [Lou03].

Loulidi:2008:AAO

[Lou08b] M. Loulidi. Analytical approach to the one-dimensional disordered exclusion process with open boundaries and random sequential dynamics. Journal of Statistical Physics, 132(1):109– 127, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9538-7.

Landim:2004:VFD

[LOV04] C. Landim, S. Olla, and S. R. S. Varadhan. On viscosity and fluctuation–dissipation in exclusion processes. Journal of Statistical Physics, 115(1–2):323–363, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019814.73545.28.

Leuzzi:2001:KSP

[LP01] Luca Leuzzi and Giorgio Parisi. The K-SAT problem in a simple limit. Journal of Statistical Physics, 103(5–6):679–695, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010388300998.

Lytova:2008:ABM

[LP08] A. Lytova and L. Pastur. On asymptotic behavior of multilinear eigenvalue statistics of random matrices. Journal of Statistical Physics, 133(5):871–882, December 2008. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9644-6.

Lytova:2009:FME

[LP09] A. Lytova and L. Pastur. Fluctuations of matrix elements of regular functions of Gaussian random matrices. Journal of Statistical Physics, 134(1):147–159, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9665-1.

Lafitte:2008:NTP

[LPD08] P. Lafitte, P. E. Parris, and S. De Bièvre. Normal transport properties in a metastable stationary state for a classical particle coupled to a non–Ohmic bath. Journal of Statistical Physics, 132(5): 863–879, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9590-3.

Lima:2001:RCC

[LPdO01] A. R. Lima, T. J. P. Penna, and P. M. C. de Oliveira. Reply to a comment on ‘A Comparison Between Broad Histogram and Multicanonical Methods’. Journal of Statistical Physics, 104(3– 4):895–899, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010345125540. See [LdOP00,

KP01a].

Liu:2004:CBM

[LPE04] Da-Jiang Liu, N. Pavlenko, and J. W. Evans. Crossover between mean-field and Ising critical behavior in a lattice-gas reaction–diffusion model. Journal of Statistical Physics, 114(1– 2):101–114, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000003105.50683.c6.

Loskutov:2000:PSC

[LR00] A. Loskutov and A. B. Ryabov. Properties of some chaotic billiards with time-dependent boundaries. Journal of Statistical Physics, 101(1–2):705, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026438927543.

LeNy:2002:STC

[LR02a] Arnaud Le Ny and Frank Redig. Short time conservation of Gibbsianness under local stochastic evolutions. Journal of Statistical Physics, 109(5–6):1073–1090, December 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020428711181.

Loskutov:2002:PDT

[LR02b] Alexander Loskutov and Alexei Ryabov. Particle dynamics in time-dependent stadium-like billiards. Journal of Statistical Physics, 108(5–6):995–1014, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019735313330.

Lochmann:2006:ATZ

[LR06] Andreas Lochmann and Manfred Requardt. An analysis of the transition zone between the various scaling regimes in the smallworld model. Journal of Statistical Physics, 122(2):255–278, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8083-x.

Lepri:2000:GCF

[LRB00] S. Lepri, L. Rondoni, and G. Benettin. The Gallavotti–Cohen fluctuation theorem for a nonchaotic model. Journal of Statistical Physics, 99(3–4):857–872, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018695529398.

Lenci:2005:LDQ

[LRB05] Marco Lenci and Luc Rey-Bellet. Large deviations in quantum lattice systems: One-phase region. Journal of Statistical Physics, 119(3–4):715–746, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3015-3.

Leonenko:2006:SLM

[LRM06] N. N. Leonenko and M. D. Ruiz-Medina. Scaling laws for the multidimensional Burgers equation with quadratic external potential. Journal of Statistical Physics, 124(1):191–205, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9136-5.

Landim:2005:STD

[LRY05] Claudio Landim, José A. Ramírez, and Horng-Tzer Yau. Superdiffusivity of two dimensional lattice gas models. Journal of Statistical Physics, 119(5–6):963–995, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-4297-1.

LaCour:2000:MDN

[LS00a] Brian R. La Cour and William C. Schieve. Macroscopic determinism in noninteracting systems using large deviation theory. Journal of Statistical Physics, 99(5–6):1225–1249, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018684621645.

Loutsenko:2000:SSI

[LS00b] Igor Loutsenko and Vyacheslav Spiridonov. Soliton solutions of integrable hierarchies and Coulomb plasmas. Journal of Statistical Physics, 99(3–4):751–767, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018635310742.

LaCour:2002:MDI

[LS02] Brian R. La Cour and William C. Schieve. Macroscopic determinism in interacting systems using large deviation theory. Journal of Statistical Physics, 107(3–4):729–756, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014582013208.

Lefevere:2004:PAA

[LS04a] R. Lefevere and A. Schenkel. Perturbative analysis of anharmonic chains of oscillators out of equilibrium. Journal of Statistical Physics, 115(5–6):1389–1421, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028064.17281.bb.

Lieb:2004:EFB

[LS04b] Elliott H. Lieb and Robert Seiringer. Equivalent forms of the Bessis–Moussa–Villani conjecture. Journal of Statistical Physics, 115(1–2):185–190, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000019811.15510.
27.

Lewin:2009:SCP

[LS09a] Mathieu Lewin and Robert Seiringer. Strongly correlated phases in rapidly rotating Bose gases. Journal of Statistical Physics, 137(5–6):1040–1062, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9833-y.

Lukkarinen:2009:NON

[LS09b] Jani Lukkarinen and Herbert Spohn. Not to normal order — notes on the kinetic limit for weakly interacting quantum fluids. Journal of Statistical Physics, 134(5–6):1133–1172, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9682-8.

Ledrappier:2003:RVD

[LSSW03] François Ledrappier, Michael Shub, Carles Simó, and Amie Wilkinson. Random versus deterministic exponents in a Rich family of diffeomorphisms. Journal of Statistical Physics, 113 (1–2):85–149, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025770720803.

Lods:2004:DLB

[LT04] Bertrand Lods and Giuseppe Toscani. Dissipative linear Boltzmann equation for hard spheres. Journal of Statistical Physics, 117(3–4):635–664, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-2267-7.

Liu:2002:MZS

[LTWW02] Qing-Hui Liu, Bo Tan, Zhi-Xiong Wen, and Jun Wu. Measure zero spectrum of a class of Schrödinger operators. Journal of Statistical Physics, 106(3–4):681–691, February 2002.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013718624572.

Lu:2000:MBE

[Lu00] Xuguang Lu. A modified Boltzmann equation for Bose– Einstein particles: Isotropic solutions and long-time behavior. Journal of Statistical Physics, 98(5–6):1335–1394, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018628031233.

Lu:2001:SHS

[Lu01] Xuguang Lu. On spatially homogeneous solutions of a modified Boltzmann equation for Fermi–Dirac particles. Journal of Statistical Physics, 105(1–2):353–388, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1012282516668.

Lu:2004:IDS

[Lu04] Xuguang Lu. On isotropic distributional solutions to the Boltzmann equation for Bose–Einstein particles. Journal of Statistical Physics, 116(5–6):1597–1649, September 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041750.11320.9c.

Lu:2005:BEB

[Lu05] Xuguang Lu. The Boltzmann equation for Bose–Einstein particles: Velocity concentration and convergence to equilibrium. Journal of Statistical Physics, 119(5–6):1027–1067, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-3767-9.

Lu:2006:BEF

[Lu06] Xuguang Lu. On the Boltzmann equation for Fermi–Dirac particles with very soft potentials: Averaging compactness of weak solutions. Journal of Statistical Physics, 124(2–4):517–547, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9039-5.

Lubeck:2006:TDP

[Lüb06] S. Lübeck. Tricritical directed percolation. Journal of Statistical Physics, 123(1):193–221, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9000-z.

Lucarini:2008:SBP

[Luc08a] Valerio Lucarini. From symmetry breaking to Poisson point process in 2D Voronoi tessellations: the generic nature of hexagons. Journal of Statistical Physics, 130(6):1047–1062, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9475-x.

Lucarini:2008:RTE

[Luc08b] Valerio Lucarini. Response theory for equilibrium and nonequilibrium statistical mechanics: Causality and generalized Kramers–Kronig relations. Journal of Statistical Physics, 131 (3):543–558, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9498-y.

Lucarini:2009:EDR

[Luc09a] Valerio Lucarini. Evidence of dispersion relations for the nonlinear response of the Lorenz 63 system. Journal of Statistical Physics, 134(2):381–400, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9675-z.

Lucarini:2009:TDR

[Luc09b] Valerio Lucarini. Three-dimensional random Voronoi tessellations: From cubic crystal lattices to Poisson point processes. Journal of Statistical Physics, 134(1):185–206, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9668-y.

Luo:2001:FIM

[Luo01] Shunlong Luo. Fisher information matrix of Husimi distribution. Journal of Statistical Physics, 102(5–6):1417–1428, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1004856832310.

Luo:2007:NSW

[Luo07] Shunlong Luo. Notes on superadditivity of Wigner–Yanase– Dyson information. Journal of Statistical Physics, 128(5):1177– 1188, September 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9349-2.

Ladd:2001:LBS

[LV01] A. J. C. Ladd and R. Verberg. Lattice-Boltzmann simulations of particle–fluid suspensions. Journal of Statistical Physics, 104 (5–6):1191–1251, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010414013942.

Lauwers:2002:FBG

[LV02] J. Lauwers and A. Verbeure. Fluctuations in the Bose gas with attractive boundary conditions. Journal of Statistical Physics, 108(1–2):123–168, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015491519127.

Lauwers:2003:BEC

[LVZ03] J. Lauwers, A. Verbeure, and V. A. Zagrebnov. Bose–Einstein condensation for homogeneous interacting systems with a oneparticle spectral gap. Journal of Statistical Physics, 112(1– 2):397–420, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023648208543.

Laurencot:2001:DCE

[LW01a] Philippe Laurençot and Dariusz Wrzosek. The discrete coagulation equations with collisional breakage. Journal of Statistical Physics, 104(1–2):193–220, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010309727754.

Lu:2001:DFZ

[LW01b] Wentao T. Lu and F. Y. Wu. Density of the Fisher zeroes for the Ising model. Journal of Statistical Physics, 102(3–4):953–970, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004863322373.

Lee:2004:EDB

[LW04a] Jihoon Lee and Ming-Yih Wu. Ergodicity for the dissipative Boussinesq equations with random forcing. Journal of Statistical Physics, 117(5–6):929–973, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-5711-9.

Lubeck:2004:USB

[LW04b] S. Lübeck and R. D. Willmann. Universal scaling behavior of directed percolation around the upper critical dimension. Journal of Statistical Physics, 115(5–6):1231–1250, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028059.24904.3b.

Langmann:2007:MFM

[LW07] Edwin Langmann and Mats Wallin. Mean field magnetic phase diagrams for the two dimensional t–t′–U Hubbard model. Journal of Statistical Physics, 127(4):825–840, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9308-y.

Levesque:2000:CFT

[LWL00] D. Levesque, J.-J. Weis, and J. L. Lebowitz. Charge fluctuations in the two-dimensional one-component plasma. Journal of Statistical Physics, 100(1–2):209–222, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018643829340.

Lieb:2001:GSE

[LY01] Elliott H. Lieb and Jakob Yngvason. The ground state energy of a dilute two-dimensional Bose gas. Journal of Statistical Physics, 103(3–4):509–526, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010337215241.

Lin:2007:CNS

[LY07] Kevin K. Lin and Lai-Sang Young. Correlations in nonequilibrium steady states of random halves models. Journal of Statistical Physics, 128(3):607–639, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9318-9.

Lyberg:2005:FVC

[Lyb05] I. Lyberg. The fourth virial coefficient of a fluid of hard spheres in odd dimensions. Journal of Statistical Physics, 119(3–4):747– 764, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-3020-6.

Luo:2004:ICS

[LZ04] Shunlong Luo and Zhengmin Zhang. An informational characterization of Schrödinger’s uncertainty relations. Journal of Statistical Physics, 114(5–6):1557–1576, March 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013971.75667.c8.

Luo:2008:SWY

[LZ08] Shunlong Luo and Qiang Zhang. Superadditivity of Wigner– Yanase–Dyson information revisited. Journal of Statistical Physics, 131(6):1169–1177, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9534-y.

Luo:2009:OCT

[LZ09] Shunlong Luo and Qiang Zhang. Observable correlations in twoqubit states. Journal of Statistical Physics, 136(1):165–177, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9779-0.

Levine:2004:PTT

[LZGM04] E. Levine, G. Ziv, L. Gray, and D. Mukamel. Phase transitions in traffic models. Journal of Statistical Physics, 117(5–6): 819–830, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-5706-6.

Machta:2002:SCT

[Mac02] J. Machta. Sampling chaotic trajectories quickly in parallel. Journal of Statistical Physics, 109(3–4):863–873, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020431119088.

Macdonald:2009:DCZ

[Mac09] Brian Macdonald. Density of complex zeros of a system of real random polynomials. Journal of Statistical Physics, 136(5): 807–833, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9810-5.

Magdziarz:2009:BSF

[Mag09a] Marcin Magdziarz. Black–Scholes formula in subdiffusive regime. Journal of Statistical Physics, 136(3):553–564, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9791-4.

Magdziarz:2009:LPS

[Mag09b] Marcin Magdziarz. Langevin picture of subdiffusion with infinitely divisible waiting times. Journal of Statistical Physics, 135(4):763–772, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9751-z.

Maier:2003:CEF

[Mai03] Robert S. Maier. On crossing event formulas in critical twodimensional percolation. Journal of Statistical Physics, 111(5– 6):1027–1048, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023006413433.

Malyshev:2005:MMC

[Mal05] V. A. Malyshev. Microscopic models for chemical thermodynamics. Journal of Statistical Physics, 119(5–6):997–1026, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-4408-z.

Malecki:2007:TDK

[Mal07] Justin Malecki. The two dimensional Kondo model with Rashba spin-orbit coupling. Journal of Statistical Physics, 129(4):741– 757, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9414-x.

Mandelbrot:2003:MPL

[Man03] Benoît B. Mandelbrot. Multifractal power law distributions: Negative and critical dimensions and other ‘anomalies,’ explained by a simple example. Journal of Statistical Physics, 110 (3–6):739–774, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022159802564;   http://www.
springerlink.com/content/j20k474mu4063439/. Special is-

sue in honor of Michael E. Fisher’s 70th birthday (Piscataway, NJ, 2001).

Melik-Alaverdian:2001:QPM

[MAOB01] V. Melik-Alaverdian, G. Ortiz, and N. E. Bonesteel. Quantum projector method on curved manifolds. Journal of Statistical Physics, 104(1–2):449–470, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010326231389.

Masoliver:2000:BRB

[Mas00] Jaume Masoliver. Book review: An Introduction to Econophysics, Correlations, and Complexity in Finance, N. Rosario, H. Mantegna, and H. E. Stanley, Cambridge University Press, Cambridge, 2000. Journal of Statistical Physics, 100(3–4): 801–802, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018639912443.

Masoliver:2002:BRB

[Mas02a] Jaume Masoliver. Book reviews: Theory of Financial Risks: From Statistical Mechanics to Risk Management. Jean-Philippe Bouchaud and Marc Potters, Cambridge University Press, Cambridge, 2001. Stochastic Processes from Physics to Finance. Wolfang Paul and Jörg Baschnagel, Springer, Berlin, 1999. Journal of Statistical Physics, 109(1–2):339–340, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020060216811.

Mason:2002:MSC

[Mas02b] R. J. Mason. A multi-speed compressible lattice-Boltzmann model. Journal of Statistical Physics, 107(1–2):385–400, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014535310153.

Masoliver:2003:BRB

[Mas03a] Jaume Masoliver. Book review: Patterns of Speculation. A Study in Observational Econophysics. Bertrand M. Roehner, Cambridge University Press, Cambridge, 2002. Journal of Statistical Physics, 112(3–4):881–882, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023888328089.

Mastropietro:2003:NUI

[Mas03b] V. Mastropietro. Non-universality in Ising models with four spin interaction. Journal of Statistical Physics, 111(1–2):201–259, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022257024662.

Masoliver:2005:BRB

[Mas05a] Jaume Masoliver. Book review: Quantum Finance, Path Integrals and Hamiltonians for Options and Interest Rates. Journal of Statistical Physics, 120(1–2):417–418, July 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5473-z.

Mastropietro:2005:RPL

[Mas05b] Vieri Mastropietro. Rigorous proof of Luttinger liquid behavior in the 1D Hubbard model. Journal of Statistical Physics, 121(3– 4):373–432, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-7007-0.

Masoliver:2006:BRB

[Mas06] Jaume Masoliver. Book review: The Noisy Oscillator. Journal of Statistical Physics, 123(3):703, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9094-y.

Masoliver:2007:CD

[Mas07] Jaume Masoliver. Chaotic dynamics. Journal of Statistical Physics, 127(3):655–656, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9302-4.

Masoliver:2008:MLW

[Mas08] Jaume Masoliver. Melvin Lax, Wei cai, min xu: Random processes in physics and finance. Journal of Statistical Physics, 130(4):821, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9468-9.

Masoliver:2009:FBH

[Mas09a] Jaume Masoliver. Floyd b. Hanson: Applied stochastic processes and control for jump-diffusions. Journal of Statistical Physics, 134(1):207, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-008-9669-x.pdf.

Masoliver:2009:GCM

[Mas09b] Jaume Masoliver. S. Gianfausto, C. De Michele, N.T. Kottegoda, R. Renzo: Extremes in Nature: An Approach Using Copulas. Journal of Statistical Physics, 134(2):425, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-008-9666-0.pdf.

Mattis:2001:LMM

[Mat01] Daniel C. Mattis. Luttinger’s model and the matter of dispersion. Journal of Statistical Physics, 103(3–4):503–507, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010385131170.

Mattingly:2002:DSS

[Mat02] Jonathan C. Mattingly. The dissipative scale of the stochastics Navier–Stokes equation: Regularization and analyticity. Journal of Statistical Physics, 108(5–6):1157–1179, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019799700126.

Mattis:2004:DN

[Mat04] Daniel C. Mattis. Design of a nanomagnet. Journal of Statistical Physics, 116(1–4):773–782, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037212.88196.7c.

Mathieu:2008:QIP

[Mat08] P. Mathieu. Quenched invariance principles for random walks with random conductances. Journal of Statistical Physics, 130 (5):1025–1046, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9465-z.

Maudlin:2008:DRM

[Mau08] Tim Maudlin. David Ruelle: The Mathematician’s brain. Journal of Statistical Physics, 130(4):823–825, February 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9469-8.

Meyer:2002:PGL

[MB02] David A. Meyer and Heather Blumer. Parrondo games as lattice gas automata. Journal of Statistical Physics, 107(1–2):225–239, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014566822448.

Manoj:2003:PSD

[MB03] G. Manoj and Mustansir Barma. Phase separation driven by a fluctuating two-dimensional self-affine potential field. Journal of Statistical Physics, 110(3–6):1305–1316, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022113414376.

Margolin:2006:NTS

[MB06] Gennady Margolin and Eli Barkai. Nonergodicity of a time series obeying Lévy statistics. Journal of Statistical Physics, 122(1): 137–167, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8076-9.

Majumdar:2008:EME

[MBL08] Satya N. Majumdar, Oriol Bohigas, and Arul Lakshminarayan. Exact minimum eigenvalue distribution of an entangled random pure state. Journal of Statistical Physics, 131(1):33–49, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9491-5.

Majumdar:2005:ADF

[MC05] Satya N. Majumdar and Alain Comtet. Airy distribution function: From the area under a Brownian excursion to the maximal height of fluctuating interfaces. Journal of Statistical Physics, 119(3–4):777–826, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3022-4.

McCoy:2001:BR

[McC01] Barry M. McCoy. The Baxter revolution. Journal of Statistical Physics, 102(3–4):375–384, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004894530125.

McDonald:2001:GAD

[McD01] Elaine T. McDonald. General asymptotics of the density of a restricted coalescing random walk system. Journal of Statistical Physics, 102(1–2):97–113, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026560609128.

Marchetti:2008:HSM

[MCG08] Domingos H. U. Marchetti, William R. P. Conti, and Leonardo F. Guidi. Hierarchical spherical model from a geometric point of view. Journal of Statistical Physics, 132(5): 811–838, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9568-1.

Majumdar:2006:USE

[MCZ06] Satya N. Majumdar, Alain Comtet, and Robert M. Ziff. Unified solution of the expected maximum of a discrete time random walk and the discrete flux to a spherical trap. Journal of Statistical Physics, 122(5):833–856, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-9002-x.

McGuire:2001:EBA

[MD01] J. B. McGuire and Charlotte Dirk. Extending the Bethe ansatz: The quantum three-particle ring. Journal of Statistical Physics, 102(3–4):971–980, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004815406443.

Mansour:2000:MSL

[MDB00] M. Malek Mansour, J. Dethier, and F. Baras. Microscopic simulation of limit cycle behavior in spatially extended systems. Journal of Statistical Physics, 101(1–2):425–441, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026451230269.

Mordant:2003:LVF

[MDL^{+}03] N. Mordant, J. Delour, E. Léveque, O. Michel, A. Arnéodo, and J.-F. Pinton. Lagrangian velocity fluctuations in fully developed turbulence: Scaling, intermittency, and dynamics. Journal of Statistical Physics, 113(5–6):701–717, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1027352417365.

Mei:2009:ETD

[MDW09] Dong Cheng Mei, Lu Chun Du, and Can Jun Wang. The effects of time delay on stochastic resonance in a bistable system with correlated noises. Journal of Statistical Physics, 137(4): 625–638, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9864-4.

Mecke:2001:EMC

[Mec01] Klaus R. Mecke. Exact moments of curvature measures in the Boolean model. Journal of Statistical Physics, 102(5–6): 1343–1381, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004800714563.

Mermin:2003:MLF

[Mer03] N. David Mermin. My life with Fisher. Journal of Statistical Physics, 110(3–6):467–473, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022187113951.

Messager:2002:QPTa

[Mes02a] Alain Messager. On quantum phase transition. I. Spinless electrons strongly correlated with ions. Journal of Statistical Physics, 106(3–4):723–783, February 2002. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013722725481.

Messager:2002:QPTb

[Mes02b] Alain Messager. On quantum phase transition. II. The Falicov– Kimball model. Journal of Statistical Physics, 106(3–4):785–810, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013774709551.

Mezzasalma:2001:PCS

[Mez01] Stefano A. Mezzasalma. Polymer chain size from geodesic path and geometrical Bolyai–Lobachevskij partition function. Application to swelling of macromolecules in solution and micellar growth. Journal of Statistical Physics, 102(5–6):1331–1341, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004848630492.

Morriss:2002:SWD

[MG02] G. P. Morriss and Ch. Gruber. Strong and weak damping in the adiabatic motion of the simple piston. Journal of Statistical Physics, 109(3–4):549–568, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1020454329566.

Mobilia:2007:PTS

[MGT07] Mauro Mobilia, Ivan T. Georgiev, and Uwe C. Täuber. Phase transitions and spatio-temporal fluctuations in stochastic lattice Lotka–Volterra models. Journal of Statistical Physics, 128 (1–2):447–483, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9146-3.

Medved:2007:SMS

[MH07] Igor Medved’ and Dale A. Huckaby. A statistical mechanical study of current spikes due to phase transitions at electrode– electrolyte interfaces. Journal of Statistical Physics, 129(2): 335–376, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9359-0.

Miekisz:2004:SSS

[Mie04] Jacek Miekisz. Stochastic stability in spatial games. Journal of Statistical Physics, 117(1–2):99–110, October 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044065.65866.bc.

Misumi:2006:CVL

[Mis06] Jun Misumi. Critical values in a long-range percolation on spaces like fractals. Journal of Statistical Physics, 125(4): 873–883, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9243-3.

Mirin:2003:EDA

[MK03] Nikolay Mirin and Anatoly B. Kolomeisky. Effect of detachments in asymmetric simple exclusion processes. Journal of Statistical Physics, 110(3–6):811–823, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022163903473.

Molchan:2004:SVM

[MK04] G. Molchan and A. Khokhlov. Small values of the maximum for the integral of fractional Brownian motion. Journal of Statistical Physics, 114(3–4):923–946, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012512.18060.a5.

Mohanty:2007:DDS

[MK07] P. K. Mohanty and K. Kruse. Driven diffusive systems of active filament bundles. Journal of Statistical Physics, 128(1–2):95– 110, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9125-8.

Majumdar:2000:NPT

[MKB00] Satya N. Majumdar, Supriya Krishnamurthy, and Mustansir Barma. Nonequilibrium phase transition in a model of diffusion, aggregation, and fragmentation. Journal of Statistical Physics, 99(1–2):1–29, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018632005018.

Muller:2008:MSM

[MKL08] Melanie J. I. Müller, Stefan Klumpp, and Reinhard Lipowsky. Motility states of molecular motors engaged in a stochastic tugof-war. Journal of Statistical Physics, 133(6):1059–1081, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-008-9651-7.pdf.

Moore:2000:IDL

[MM00] Cristopher Moore and Jonathan Machta. Internal diffusionlimited aggregation: Parallel algorithms and complexity. Journal of Statistical Physics, 99(3–4):661–690, May 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018627008925.

Mera:2002:DFT

[MM02] M. Eugenia Mera and Manuel Morán. Degrees of freedom of a time series. Journal of Statistical Physics, 106(1–2):125–145, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013172129075.

Mallick:2005:AOD

[MM05a] Kirone Mallick and Philippe Marcq. Anharmonic oscillator driven by additive Ornstein–Uhlenbeck noise. Journal of Statistical Physics, 119(1–2):1–33, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2135-5.

Mondaini:2005:SGC

[MM05b] Leonardo Mondaini and E. C. Marino. Sine–Gordon/ Coulomb gas soliton correlation functions and an exact evaluation of the Kosterlitz–Thouless critical exponent. Journal of Statistical Physics, 118(3–4):767–779, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-8828-y.

Muller:2005:PWR

[MM05c] Melanie Müller and Gernot Münster. Profile and width of rough interfaces. Journal of Statistical Physics, 118(3–4):669–686, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8824-2.

Messager:2006:MSF

[MM06a] Alain Messager and Bruno Nachtergaele Molchanov. A model with simultaneous first and second order phase transitions. Journal of Statistical Physics, 122(1):1–14, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8071-1.

Mischler:2006:CPIb

[MM06b] S. Mischler and C. Mouhot. Cooling process for inelastic Boltzmann equations for hard spheres,Part II: Self-similar solutions and tail behavior. Journal of Statistical Physics, 124(2–4): 703–746, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9097-8.

Miao:2008:GWC

[MM08] Dong Miao and Fuming Ma. On guided waves created by line defects. Journal of Statistical Physics, 130(6):1197–1215, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9480-8.

Mischler:2006:CPIa

[MMR06] S. Mischler, C. Mouhot, and M. Rodriguez Ricard. Cooling process for inelastic Boltzmann equations for hard spheres, part i: The Cauchy problem. Journal of Statistical Physics, 124(2– 4):655–702, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9096-9.

Mejia-Monasterio:2008:FRN

[MMR08a] Carlos Mejía-Monasterio and Lamberto Rondoni. On the fluctuation relation for Nosé-Hoover boundary thermostated systems. Journal of Statistical Physics, 133(4):617–637, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9618-8.

Mitra:2008:ABI

[MMR08b] Mithun K. Mitra, Gautam I. Menon, and R. Rajesh. Asymptotic behavior of inflated lattice polygons. Journal of Statistical Physics, 131(3):393–404, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9512-4.

Moore:2000:HRC

[MN00] Cristopher Moore and M. E. J. Newman. Height representation, critical exponents, and ergodicity in the four-state triangular Potts antiferromagnet. Journal of Statistical Physics, 99 (3–4):629–660, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018638624854.

Meurice:2002:NSF

[MN02] Y. Meurice and S. Niermann. From nonlinear scaling fields to critical amplitudes. Journal of Statistical Physics, 108(1–2):213– 246, July 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015495620036.

Maes:2003:TRE

[MN03] Christian Maes and Karel Netocný. Time-reversal and entropy. Journal of Statistical Physics, 110(1–2):269–310, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021026930129.

Maltsev:2004:DST

[MN04] A. Ya. Maltsev and S. P. Novikov. Dynamical systems, topology, and conductivity in normal metals. Journal of Statistical Physics, 115(1–2):31–46, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019835.01125.92.

Mitter:2005:IEF

[MN05] Sanjoy K. Mitter and Nigel J. Newton. Information and entropy flow in the Kalman–Bucy filter. Journal of Statistical Physics, 118(1–2):145–176, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-8781-9.

Michalowicz:2009:ITM

[MNBO09] J. V. Michalowicz, J. M. Nichols, F. Bucholtz, and C. C. Olson. An Isserlis’ theorem for mixed Gaussian variables: Application to the auto-bispectral density. Journal of Statistical Physics, 136(1):89–102, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9768-3.

McKane:2007:ABO

[MNNS07] A. J. McKane, J. D. Nagy, T. J. Newman, and M. O. Stefanini. Amplified biochemical oscillations in cellular systems. Journal of Statistical Physics, 128(1–2):165–191, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9221-9.

Manzo:2004:EFM

[MNOS04] F. Manzo, F. R. Nardi, E. Olivieri, and E. Scoppola. On the essential features of metastability: Tunnelling time and critical configurations. Journal of Statistical Physics, 115(1–2):591–642, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000019822.45867.ec.

Mamasakhlisov:2008:PAD

[MNP08] Yevgeni S. Mamasakhlisov, Ali Naji, and Rudolf Podgornik. Partially annealed disorder and collapse of like–charged macroions. Journal of Statistical Physics, 133(4):659–681, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9635-7.

Machta:2008:PSS

[MNS08] J. Machta, C. M. Newman, and D. L. Stein. The percolation signature of the spin glass transition. Journal of Statistical Physics, 130(1):113–128, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9446-2.

Maes:2003:HCN

[MNV03] Christian Maes, Karel Netocný, and Michel Verschuere. Heat conduction networks. Journal of Statistical Physics, 111(5–6): 1219–1244, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023004300229.

Manzo:2001:DBC

[MO01] F. Manzo and E. Olivieri. Dynamical Blume–Capel model: Competing metastable states at infinite volume. Journal of Statistical Physics, 104(5–6):1029–1090, September 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010401711216.

Mashkevich:2009:ADT

[MO09a] Stefan Mashkevich and Stéphane Ouvry. Area distribution of two-dimensional random walks on a square lattice. Journal of Statistical Physics, 137(1):71–78, October 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9827-9.

Mourragui:2009:LGM

[MO09b] Mustapha Mourragui and Enza Orlandi. Lattice gas model in random medium and open boundaries: Hydrodynamic and relaxation to the steady state. Journal of Statistical Physics, 136 (4):685–714, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9796-z.

Molchan:2002:MCM

[Mol02] G. Molchan. Mandelbrot cascade measures independent of branching parameter. Journal of Statistical Physics, 107(5– 6):977–988, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015149422567.

Molchan:2004:UBP

[Mol04] G. Molchan. On the uniqueness of the branching parameter for a random cascade measure. Journal of Statistical Physics, 115 (3–4):855–868, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000022382.88228.fd.

Monteil:2004:CET

[Mon04] Thierry Monteil. A counter-example to the theorem of Hiemer and Snurnikov. Journal of Statistical Physics, 114(5–6):1619– 1623, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013974.81162.20. See

[HS98].

Morris:2007:EZT

[Mor07] I. D. Morris. Entropy for zero-temperature limits of Gibbsequilibrium states for countable-alphabet subshifts of finite type. Journal of Statistical Physics, 126(2):315–324, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9215-7.

Macris:2001:LRO

[MP01] N. Macris and C.-A. Piguet. Long-range orders in models of itinerant electrons interacting with heavy quantum fields. Journal of Statistical Physics, 105(5–6):909–935, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013509512844.

Mezard:2003:CMZ

[MP03] Marc Mézard and Giorgio Parisi. The cavity method at zero temperature. Journal of Statistical Physics, 111(1–2):1–34, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022221005097.

Marder:2006:GES

[MP06] M. Marder and N. Papanicolaou. Geometry and elasticity of strips and flowers. Journal of Statistical Physics, 125(5–6): 1065–1092, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9087-x.

Morais:2009:APT

[MP09] Thiago Morais and Aldo Procacci. Absence of phase transitions in a class of integer spin systems. Journal of Statistical Physics, 136(4):677–684, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9799-9.

Munoz:2009:SCK

[MPD09] Enrique Muñoz, Jeong-Man Park, and Michael W. Deem. Solution of the Crow–Kimura and eigen models for alphabets of arbitrary size by Schwinger spin coherent states. Journal of Statistical Physics, 135(3):429–465, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-009-9732-2.pdf.

Marinari:2000:RSB

[MPRT^{+}00] Enzo Marinari, Giorgio Parisi, Federico Ricci-Tersenghi, Juan J. Ruiz-Lorenzo, and Francesco Zuliani. Replica symmetry breaking in short-range spin glasses: Theoretical foundations and numerical evidences. Journal of Statistical Physics, 98(5–6): 973–1074, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018607809852.

Mazel:2002:GPA

[MPS02] Alexander Mazel, Aldo Procacci, and Benedetto Scoppola. Gas phase of asymmetric nearest neighbor Ising model. Journal of Statistical Physics, 106(5–6):1241–1248, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014002323607.

Maier-Paape:2000:SDM

[MPSW00] Stanislaus Maier-Paape, Barbara Stoth, and Thomas Wanner. Spinodal decomposition for multicomponent Cahn–Hilliard systems. Journal of Statistical Physics, 98(3–4):871–896, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018687811688.

Milanovic:2006:GCE

[MPT06] Lj. Milanović, H. A. Posch, and W. Thirring. Gravitational collapse and ergodicity in confined gravitational systems. Journal of Statistical Physics, 124(2–4):843–858, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9095-x.

Maes:2000:PEP

[MR00] Christian Maes and Frank Redig. Positivity of entropy production. Journal of Statistical Physics, 101(1–2):3–15, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1026434726635.

Morozov:2001:NMQ

[MR01] V. G. Morozov and Gerd Röpke. Non-Markovian quantum kinetics and conservation laws. Journal of Statistical Physics, 102(1– 2):285–313, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026525028692.

Mukhamedov:2004:GMM

[MR04] Farruh Mukhamedov and Utkir Rozikov. On Gibbs measures of models with competing ternary and binary interactions and corresponding von Neumann algebras. Journal of Statistical Physics, 114(3–4):825–848, February 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012509.10642.83.

Mukhamedov:2005:GMM

[MR05] Farruh Mukhamedov and Utkir Rozikov. On Gibbs measures of models with competing ternary and binary interactions and corresponding von Neumann algebras II. Journal of Statistical Physics, 119(1–2):427–446, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-2056-3.

Maes:2007:FTN

[MRS07] C. Maes, F. Redig, and E. Saada. Freezing transitions in non– Fellerian particle systems. Journal of Statistical Physics, 127 (1):171–189, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9279-z.

Mezard:2003:TSD

[MRTZ03] M. Mézard, F. Ricci-Tersenghi, and R. Zecchina. Two solutions to diluted p-spin models and XORSAT problems. Journal of Statistical Physics, 111(3–4):505–533, May 2003. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022886412117.

Maes:2002:NCH

[MRV02] Christian Maes, Frank Redig, and Michel Verschuere. No current without heat. Journal of Statistical Physics, 106(3–4):569–587, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013706321846.

Markowich:2002:ENS

[MRW02] Peter A. Markowich, Gerhard Rein, and Gershon Wolansky. Existence and nonlinear stability of stationary states of the Schrödinger–Poisson system. Journal of Statistical Physics, 106(5–6):1221–1239, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014050206769.

Myjak:2001:LEH

[MS01] Józef Myjak and Tomasz Szarek. A lower estimation of the Hausdorff dimension for attractors with overlaps. Journal of Statistical Physics, 105(3–4):649–657, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1012231910484.

Miller:2002:LBM

[MS02] W. Miller and S. Succi. A lattice Boltzmann model for anisotropic crystal growth from melt. Journal of Statistical Physics, 107(1–2):173–186, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014510704701.

Mazilu:2003:HTE

[MS03] I. Mazilu and B. Schmittmann. High temperature expansion for a driven bilayer system. Journal of Statistical Physics, 113(3–4): 505–525, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026016701492.

Maioli:2005:TLS

[MS05a] Marco Maioli and Andrea Sacchetti. Two level systems driven by a stochastic perturbation. Journal of Statistical Physics, 119 (5–6):1383–1396, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5106-6.

Mattingly:2005:SSS

[MS05b] Jonathan C. Mattingly and Toufic M. Suidan. The small scales of the stochastic Navier–Stokes equations under rough forcing. Journal of Statistical Physics, 118(1–2):343–364, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-8787-3.

Montanari:2006:DGT

[MS06a] Andrea Montanari and Guilhem Semerjian. On the dynamics of the glass transition on Bethe lattices. Journal of Statistical Physics, 124(1):103–189, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9103-1.

Montanari:2006:RIB

[MS06b] Andrea Montanari and Guilhem Semerjian. Rigorous inequalities between length and time scales in glassy systems. Journal of Statistical Physics, 125(1):23–54, October 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9175-y.

Mitter:2008:GRG

[MS08a] P. K. Mitter and B. Scoppola. The global renormalization group trajectory in a critical supersymmetric field theory on the lattice Z^{3}. Journal of Statistical Physics, 133(5):921–1011, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9626-8.

Morters:2008:CWS

[MS08b] Peter Mörters and Nadia Sidorova. A class of weakly selfavoiding walks. Journal of Statistical Physics, 133(2):255–269, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9619-7.

Merkli:2009:RTO

[MS09] Marco Merkli and Shannon Starr. A resonance theory for open quantum systems with time-dependent dynamics. Journal of Statistical Physics, 134(5–6):871–898, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9645-5.

Mattingly:2007:ADS

[MSVE07] Jonathan C. Mattingly, Toufic M. Suidan, and Eric Vanden-Eijnden. Anomalous dissipation in a stochastically forced infinite-dimensional system of coupled oscillators. Journal of Statistical Physics, 128(5):1145–1152, September 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9351-8.

Merchan:2004:CCS

[MT04] Lina Merchán and Gabriel Téllez. Confined Coulomb systems with adsorbing boundaries: The two-dimensional twocomponent plasma. Journal of Statistical Physics, 114(3–4): 735–761, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000012507.14488.a3.

Matthes:2008:SDK

[MT08] Daniel Matthes and Giuseppe Toscani. On steady distributions of kinetic models of conservative economies. Journal of Statistical Physics, 130(6):1087–1117, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9462-2.

Moessner:2004:PPQ

[MTS04] R. Moessner, Oleg Tchernyshyov, and S. L. Sondhi. Planar pyrochlore, quantum ice and sliding ice. Journal of Statistical Physics, 116(1–4):755–772, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037247.54022.62.

Mezard:2008:GTR

[MTT08] M. Mézard, M. Tarzia, and C. Toninelli. Group testing with random pools: Phase transitions and optimal strategy. Journal of Statistical Physics, 131(5):783–801, June 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9528-9.

Mungan:2007:SML

[Mun07] Muhittin Mungan. String matching and 1D lattice gases. Journal of Statistical Physics, 126(1):207–242, January 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9247-z.

Murmann:2001:HLD

[Mür01] Michael G. Mürmann. The hydrodynamic limit of a deterministic particle system with conservation of mass and momentum. Journal of Statistical Physics, 105(3–4):483–510, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012219607758.

Murugan:2008:MSP

[Mur08] Rajamanickam Murugan. Multiple stochastic point processes in gene expression. Journal of Statistical Physics, 131(1):153–165, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9484-4.

Michoel:2001:ICB

[MV01] T. Michoel and A. Verbeure. Interferencing in coupled Bose– Einstein condensates. Journal of Statistical Physics, 102(5–6): 1383–1405, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004852731401.

Mogilner:2003:SDP

[MV03] A. Mogilner and D. W. Verzi. A simple 1-D physical model for the crawling nematode sperm cell. Journal of Statistical Physics, 110(3–6):1169–1189, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022153028488.

Meson:2007:SMA

[MV07] Alejandro Mesón and Fernando Vericat. A statistical mechanics approach for a rigidity problem. Journal of Statistical Physics, 126(2):391–417, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9259-8.

Muratov:2004:BUG

[MVE04] Cyrill B. Muratov and Eric Vanden-Eijnden. Breakup of universality in the generalized spinodal nucleation theory. Journal of Statistical Physics, 114(3–4):605–623, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012504.46631.89.

Maes:2003:MMK

[MvW03] Christian Maes and Maarten H. van Wieren. A Markov model for kinesin. Journal of Statistical Physics, 112(1–2):329–355, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023691923564.

Menshikov:2008:ABR

[MVW08] M. V. Menshikov, M. Vachkovskaia, and A. R. Wade. Asymptotic behaviour of randomly reflecting billiards in unbounded tubular domains. Journal of Statistical Physics, 132(6):1097– 1133, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9578-z.

Mariz:2001:SOD

[MvWH^{+}01] A. M. Mariz, F. van Wijland, H. J. Hilhorst, S. R. Gomes Júnior, and C. Tsallis. Statistics of the one-dimensional Riemann walk. Journal of Statistical Physics, 102(1–2):259–283, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026572911854.

Meester:2002:NTD

[MZ02a] Ronald Meester and Dmitri Znamenski. Non-triviality of a discrete Bak–Sneppen evolution model. Journal of Statistical Physics, 109(5–6):987–1004, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1020468325294.

Minlos:2002:LBT

[MZ02b] R. Minlos and E. Zhizhina. Leading branches of the transfermatrix spectrum for lattice spin systems (quasi-particles of different species). Journal of Statistical Physics, 108(5–6):885– 904, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019775027443.

Mora:2008:RST

[MZ08] Thierry Mora and Lenka Zdeborová. Random subcubes as a toy model for constraint satisfaction problems. Journal of Statistical Physics, 131(6):1121–1138, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9543-x.

Nadiga:2000:SPI

[Nad00] B. T. Nadiga. Scaling properties of an inviscid mean-motion fluid model. Journal of Statistical Physics, 98(3–4):935–948, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018644029435.

Nagahata:2003:FDE

[Nag03] Yukio Nagahata. Fluctuation dissipation equation for lattice gas with energy. Journal of Statistical Physics, 110(1–2):219–246, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021022829221.

Nagle:2004:RED

[Nag04] John F. Nagle. Regarding the entropy of distinguishable particles. Journal of Statistical Physics, 117(5–6):1047–1062, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5715-5.

Nagao:2007:PER

[Nag07] Taro Nagao. Pfaffian expressions for random matrix correlation functions. Journal of Statistical Physics, 129(5–6):1137– 1158, December 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9415-9.

Najar:2004:ABI

[Naj04] Hatem Najar. Asymptotic behavior of the integrated density of states of acoustic operators with random long range perturbations. Journal of Statistical Physics, 115(3–4):977–996, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000022377.63297.c6.

Najar:2007:LTA

[Naj07] Hatem Najar. Lifshitz tails for acoustic waves in random quantum waveguide. Journal of Statistical Physics, 128(4): 1093–1112, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9333-x.

Najar:2008:NLT

[Naj08] Hatem Najar. Non-Lifshitz tails at the spectral bottom of some random operators. Journal of Statistical Physics, 130(4): 713–725, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9467-x.

Nakano:2006:RBL

[Nak06] Fumihiko Nakano. The repulsion between localization centers in the Anderson model. Journal of Statistical Physics, 123(4):803– 810, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9086-y.

Nambu:2004:SBL

[Nam04] Yoichiro Nambu. Spontaneous breaking of Lie and current algebras. Journal of Statistical Physics, 115(1–2):7–17, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019827.74407.2d.

Nagy:2002:RTA

[NAS02] Zoltán Nagy, Cécile Appert, and Ludger Santen. Relaxation times in the ASEP model using a DMRG method. Journal of Statistical Physics, 109(3–4):623–639, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020462531383.

Nelson:2003:EDI

[ND03] Kevin Nelson and Dean J. Driebe. Ensemble dynamics of intermittency and power-law decay. Journal of Statistical Physics, 111(5–6):1183–1207, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023000232088.

Nishino:2008:ADE

[ND08] Akinori Nishino and Tetsuo Deguchi. An algebraic derivation of the eigenspaces associated with an Ising-like spectrum of the superintegrable chiral Potts model. Journal of Statistical Physics, 133(4):587–615, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9624-x.

Nie:2002:LBS

[NDC02] Xiaobo Nie, Gary D. Doolen, and Shiyi Chen. Lattice-Boltzmann simulations of fluid flows in MEMS. Journal of Statistical Physics, 107(1–2):279–289, April 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014523007427.

Nepomechie:2003:FRB

[Nep03] Rafael I. Nepomechie. Functional relations and Bethe ansatz for the XXZ chain. Journal of Statistical Physics, 111(5–6): 1363–1376, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023016602955.

Newman:2000:MSW

[New00] M. E. J. Newman. Models of the small world. Journal of Statistical Physics, 101(3–4):819–841, November 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026485807148.

Newton:2008:ISM

[New08] Nigel J. Newton. Interactive statistical mechanics and nonlinear filtering. Journal of Statistical Physics, 133(4):711–737, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9622-z.

Neuhaus:2003:CSD

[NH03] Thomas Neuhaus and Johannes S. Hager. 2D crystal shapes, droplet condensation, and exponential slowing down in simulations of first-order phase transitions. Journal of Statistical Physics, 113(1–2):47–83, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025718703965.

Nicholls:2001:SMP

[Nic01] Geoff K. Nicholls. Spontaneous magnetization in the plane. Journal of Statistical Physics, 102(5–6):1229–1251, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004888211837.

Nienhuis:2001:TC

[Nie01] Bernard Nienhuis. Tiles and colors. Journal of Statistical Physics, 102(3–4):981–996, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1004867423281.

Nishimori:2007:DFD

[Nis07] Hidetoshi Nishimori. Duality in finite-dimensional spin glasses. Journal of Statistical Physics, 126(4–5):977–986, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9156-1.

Nielsen:2000:NAD

[NKC00] Steve Nielsen, Raymond Kapral, and Giovanni Ciccotti. Nonadiabatic dynamics in mixed quantum–classical systems. Journal of Statistical Physics, 101(1–2):225–242, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026458004345.

Najnudel:2009:SLR

[NNR09] Joseph Najnudel, Ashkan Nikeghbali, and Felix Rubin. Scaled limit and rate of convergence for the largest eigenvalue from the generalized Cauchy random matrix ensemble. Journal of Statistical Physics, 137(2):373–406, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9854-6.

Nechaev:2000:APT

[NOB00] S. Nechaev, G. Oshanin, and A. Blumen. Anchoring of polymers by traps randomly placed on a line. Journal of Statistical Physics, 98(1–2):281–303, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018631007164.

Nolan:2004:BRB

[Nol04] John P. Nolan. Book review: Theory and Applications of Long-Range Dependence. Paul Doukhan, George Oppenheim, and Murad Taqqu, Birkhäuser Press, Boston, 2003. Journal of Statistical Physics, 114(5–6):1625–1626, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000014280.37342.2c.

Normand:2009:MCM

[Nor09] Raoul Normand. A model for coagulation with mating. Journal of Statistical Physics, 137(2):343–371, October 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9855-5.

Nardi:2005:AEN

[NOS05] F. R. Nardi, E. Olivieri, and E. Scoppola. Anisotropy effects in nucleation for conservative dynamics. Journal of Statistical Physics, 119(3–4):539–595, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-3247-7.

Nachtergaele:2006:PCQ

[NOS06] Bruno Nachtergaele, Yoshiko Ogata, and Robert Sims. Propagation of correlations in quantum lattice systems. Journal of Statistical Physics, 124(1):1–13, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9143-6.

Nouri:2002:EGR

[Nou02] A. Nouri. The evolution of a gas in a radiation field from a kinetic point of view. Journal of Statistical Physics, 106(3–4): 589–622, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013758305917.

Nouri:2004:EBC

[NOV04] A. Nouri, A. Omrane, and J. P. Vila. Erratum on ‘Boundary Conditions for Scalar Conservation Laws, from a Kinetic Point of View’. Journal of Statistical Physics, 115(5–6):1755–1756, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028246.81983.ba.

Nozieres:2004:CMI

[Noz04] Ph. Nozières. Can a metal–insulator transition induce s-wave superconductivity? Journal of Statistical Physics, 115(1–2):19– 30, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000019826.53628.83.

Niethammer:2001:LMD

[NP01] Barbara Niethammer and Robert L. Pego. The LSW model for domain coarsening: Asymptotic behavior for conserved total mass. Journal of Statistical Physics, 104(5–6):1113– 1144, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010405812125.

Netocny:2004:LDQ

[NR04] K. Netocný and F. Redig. Large deviations for quantum spin systems. Journal of Statistical Physics, 117(3–4):521–547, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-3452-4.

Newman:2002:SRS

[NS02] C. M. Newman and D. L. Stein. The state(s) of replica symmetry breaking: Mean field theories vs. short-ranged spin glasses. Journal of Statistical Physics, 106(1–2):213–244, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013128314054.

Neishtadt:2004:APD

[NS04] A. I. Neishtadt and Y. G. Sinai. Adiabatic piston as a dynamical system. Journal of Statistical Physics, 116(1–4):815–820, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000037222.64432.62.

Newell:2005:PF

[NS05] Alan C. Newell and Patrick D. Shipman. Plants and Fibonacci. Journal of Statistical Physics, 121(5–6):937–968, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8665-7.

Niu:2004:ISH

[NSCW04] X. D. Niu, C. Shu, Y. T. Chew, and T. G. Wang. Investigation of stability and hydrodynamics of different lattice Boltzmann models. Journal of Statistical Physics, 117(3–4):665–680, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-2264-x.

Nachtergaele:2004:FOE

[NSS04] Bruno Nachtergaele, Wolfgang Spitzer, and Shannon Starr. Ferromagnetic ordering of energy levels. Journal of Statistical Physics, 116(1–4):719–738, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/B%3AJOSS.0000037227.
24460.e5.

Narnhofer:2004:RBS

[NT04] Heide Narnhofer and Walter Thirring. On the relation between strong subadditivity and entanglement. Journal of Statistical Physics, 116(1–4):1189–1197, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037210.51757.65.

Niethammer:2008:SIF

[NV08] B. Niethammer and J. J. L. Velázquez. On screening induced fluctuations in Ostwald ripening. Journal of Statistical Physics, 130(3):415–453, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9449-z.

Oshanin:2003:ESM

[OBB03] G. Oshanin, O. Bénichou, and A. Blumen. Exactly solvable model of reactions on a random catalytic chain. Journal of Statistical Physics, 112(3–4):541–586, August 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023871824455.

OCarroll:2001:ANE

[O’C01] Michael O’Carroll. Absence of negative energy spectrum for Nparticle Hamiltonians. Journal of Statistical Physics, 105(3–4): 711–717, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012288128231.

Ouared:2005:LBS

[OC05] Rafik Ouared and Bastien Chopard. Lattice Boltzmann simulations of blood flow: Non–Newtonian rheology and clotting processes. Journal of Statistical Physics, 121(1–2):209–221, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8415-x.

OConnell:2006:DTL

[O’C06] R. F. O’Connell. Does the Third Law of Thermodynamics hold in the quantum regime? Journal of Statistical Physics, 124(1):15–23, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9151-6.

Owczarek:2001:SAA

[OEB01] A. L. Owczarek, J. W. Essam, and R. Brak. Scaling analysis for the adsorption transition in a watermelon network of n directed non-intersecting walks. Journal of Statistical Physics, 102(3–4): 997–1017, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004819507352.

Orrick:2001:SSL

[ONGP01] W. P. Orrick, B. Nickel, A. J. Guttmann, and J. H. H. Perk. The susceptibility of the square lattice Ising model: New developments. Journal of Statistical Physics, 102(3–4):795–841, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004850919647.

Oono:2003:RT

[Oon03] Y. Oono. Renormalization and taxonomy. Journal of Statistical Physics, 110(3–6):1369–1374, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022117515284.

Onofrio:2004:UAF

[OP04] Roberto Onofrio and Carlo Presilla. Ultracold atomic Fermi– Bose mixtures in bichromatic optical dipole traps: A novel route to study fermion superfluidity. Journal of Statistical Physics, 115(1–2):57–89, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000019829.71660.40.

Oppenheim:2001:BRB

[Opp01] Irwin Oppenheim. Book review: Statistical Physics Statics, Dynamics, and Renormalization. Leo P. Kadanoff, World Scientific, Singapore, 2000. Journal of Statistical Physics, 105(5–6): 943–944, December 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013513613752.

Oppenheim:2002:BRBa

[Opp02a] Irwin Oppenheim. Book review: Statistical Mechanics — A Survival Guide. A. M. Glazer and J. S. Wark, Oxford University Press, Oxford, 2001. Journal of Statistical Physics, 107(3–4): 941, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014502617751.

Oppenheim:2002:BRBb

[Opp02b] Irwin Oppenheim. Book review: Thermal Physics — Entropy and Free Energies. J. C. Lee, World Scientific, Singapore, 2002. Journal of Statistical Physics, 109(5–6):1121–1122, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1020484828928.

Oppenheim:2003:BRBa

[Opp03a] Irwin Oppenheim. Book review: Methods of Statistical Physics. Tomoyasu Tanaka, Cambridge University Press, Cambridge, 2002. Journal of Statistical Physics, 110(1–2):457–458, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1021095417834.

Oppenheim:2003:BRBb

[Opp03b] Irwin Oppenheim. Book review: Statistical Mechanics, F. Schwabl, Trans. W. Brewer, Springer, New York, 2002. Journal of Statistical Physics, 112(5–6):1219–1220, September 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024675811252.

Oppenheim:2004:BRBb

[Opp04a] Irwin Oppenheim. Book review: Entropy. A. Greven, G. Keller, and G. Warnecke (eds.), Princeton University Press, Princeton, 2003. Journal of Statistical Physics, 115(5–6):1767–1768, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000028245.25527.45.

Oppenheim:2004:BRBd

[Opp04b] Irwin Oppenheim. Book review: Equilibrium and Nonequilibrium Statistical Thermodynamics. Journal of Statistical Physics, 117(5–6):1071–1072, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-5717-3.

Oppenheim:2004:BRBa

[Opp04c] Irwin Oppenheim. Book review: Statistical Mechanics Made Simple. Daniel C. Mattis, World Scientific Publishing Co., Singapore, 2003. Journal of Statistical Physics, 114(3–4):1179–1180, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012706.82083.6f.

Oppenheim:2004:BRBc

[Opp04d] Irwin Oppenheim. Book review: The Kinetic Theory of Gases: An Anthology of Classic Papers with Historical Commentary.

S. G. Brush, World Scientific Publishing Co., River Edge, NJ, 2003. Journal of Statistical Physics, 117(1–2):377–379, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000044204.75830.ee.

Oppenheim:2005:BRP

[Opp05] Irwin Oppenheim. Book review: Phase transitions: A brief account with modern applications. Journal of Statistical Physics, 119(5–6):1419–1420, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-4406-1.

Oppenheim:2006:BET

[Opp06a] Irwin Oppenheim. Beyond equilibrium thermodynamics. Journal of Statistical Physics, 122(2):371–372, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8003-0.

Oppenheim:2006:BRB

[Opp06b] Irwin Oppenheim. Book review: Statistical Physics: Including Applications to Condensed Matter. Journal of Statistical Physics, 123(3):701–702, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9093-z.

Oppenheim:2006:STF

[Opp06c] Irwin Oppenheim. Statistical thermodynamics: Fundamentals and applications. Journal of Statistical Physics, 125(1):245–276, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9088-9.

Oppenheim:2007:BRB

[Opp07] Irwin Oppenheim. Book review: Nonequilibrium Statistical Mechanics. Journal of Statistical Physics, 127(4):851–852, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9314-0.

Oppenheim:2009:PCM

[Opp09] Irwin Oppenheim. P. Castiglione, M. Falcioni, A. Lesne, A. Vulpiani: Chaos and Coarse Graining in Statistical Mechanics.

Journal of Statistical Physics, 135(1):199–200, April 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9712-6.

Ostrovsky:2004:LLM

[Ost04] Dmitry Ostrovsky. Limit lognormal multifractal as an exponential functional. Journal of Statistical Physics, 116(5–6):1491– 1520, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000041726.07161.46.

Ostrovsky:2007:FFK

[Ost07] Dmitry Ostrovsky. Functional Feynman–Kac equations for limit lognormal multifractals. Journal of Statistical Physics, 127 (5):935–965, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9315-z.

Orlandini:2004:LFK

[OSv04] E. Orlandini, A. L. Stella, and C. vanderzande. Loose, flat knots in collapsed polymers. Journal of Statistical Physics, 115 (1–2):681–700, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000019820.70798.ed.

Olson:2003:DMC

[OT03] Eric Olson and Edriss S. Titi. Determining modes for continuous data assimilation in 2D turbulence. Journal of Statistical Physics, 113(5–6):799–840, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1027312703252.

Oerding:2000:FIF

[OvWLH00] Klaus Oerding, Frédéric van Wijland, Jean-Pierre Leroy, and Hendrik Jan Hilhorst. Fluctuation-induced first-order transition in a nonequilibrium steady state. Journal of Statistical Physics, 99(5–6):1365–1395, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018697024371.

Provata:2000:SDC

[PA00] A. Provata and Y. Almirantis. Statistical dynamics of clustering in the genome structure. Journal of Statistical Physics, 101 (1–2):709, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026490911614.

Provata:2002:SDC

[PA02a] A. Provata and Y. Almirantis. Statistical dynamics of clustering in the genome structure. Journal of Statistical Physics, 106(1– 2):23–56, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013115911328.

Proville:2002:MPS

[PA02b] Laurent Proville and Serge Aubry. Many-polaron states in the Holstein–Hubbard model. Journal of Statistical Physics, 106(5– 6):1185–1195, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014046105860.

Pajevic:2007:PAC

[Paj07] Sinisa Pajevic. Performance analysis of communications networks and systems. Journal of Statistical Physics, 126(3): 725–727, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9268-7.

Panja:2002:EPL

[Pan02] Debabrata Panja. An elementary proof of Lyapunov exponent pairing for hard-sphere systems at constant kinetic energy. Journal of Statistical Physics, 109(3–4):705–727, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020418816362.

Panchenko:2008:ECO

[Pan08] Dmitry Panchenko. Exponential control of overlap in the replica method for p-spin Sherrington–Kirkpatrick model. Journal of Statistical Physics, 130(5):831–842, March 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9463-1.

Parisi:2008:MCR

[Par08] Giorgio Parisi. On the most compact regular lattices in large dimensions: A statistical mechanical approach. Journal of Statistical Physics, 132(2):207–234, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9539-6.

Patrick:2007:EWC

[Pat07a] A. E. Patrick. Euler walk on a Cayley tree. Journal of Statistical Physics, 127(3):629–653, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9281-5.

Patrick:2007:SMR

[Pat07b] A. E. Patrick. Spherical model in a random field. Journal of Statistical Physics, 128(5):1211–1235, September 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9353-6.

Perk:2009:NRC

[PAY09] Jacques H. H. Perk and Helen Au-Yang. New results for the correlation functions of the Ising model and the transverse Ising chain. Journal of Statistical Physics, 135(4):599–619, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-009-9758-5.

Piaud:2005:ECL

[PBFC05] Benjamin Piaud, Stéphane Blanco, Richard Fournier, and Michael J. Clifton. Energy-conserving lattice Boltzmann thermal model in two dimensions. Journal of Statistical Physics, 121(1–2):119–131, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-5256-6.

Phillips:2005:TDF

[PD05] James M. Phillips and J. G. Dash. Thermal disorder, fluctuations, growth and fragmentation of finite one-dimensional atomic chains. Journal of Statistical Physics, 120(3–4):721–735, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-5252-x.

Park:2006:SBF

[PD06] Jeong-Man Park and Michael W. Deem. Schwinger Boson formulation and solution of the Crow–Kimura and eigen models of quasispecies theory. Journal of Statistical Physics, 125(4): 971–1011, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9190-z.

Pessoa:2009:ZPV

[PdKV09] Renato Pessoa, Maurice de Koning, and Silvio Antonio Vitiello. Zero-point vacancy concentration in a model quantum solid: A reversible-work approach. Journal of Statistical Physics, 134(4): 769–780, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9702-8.

Palla:2007:CPC

[PDV07] Gergely Palla, Imre Derényi, and Tamás Vicsek. The critical point of k-clique percolation in the Erdős–Rényi graph. Journal of Statistical Physics, 128(1–2):219–227, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9184-x.

Panja:2000:LTT

[PDvB00] D. Panja, J. R. Dorfman, and Henk van Beijeren. Long-time-tail effects on Lyapunov exponents of a random, two-dimensional field-driven Lorentz gas. Journal of Statistical Physics, 100 (1–2):279–311, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018604115227.

Penrose:2001:LTM

[Pen01] Mathew D. Penrose. Limit theorems for monolayer ballistic deposition in the continuum. Journal of Statistical Physics, 105(3– 4):561–583, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012275725505.

Penrose:2008:GRI

[Pen08] Mathew D. Penrose. Growth and roughness of the interface for ballistic deposition. Journal of Statistical Physics, 131 (2):247–268, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9507-1.

Percus:2002:HII

[Per02] J. K. Percus. Homopolymers with intrachain interactions. Journal of Statistical Physics, 106(1–2):357–368, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013188532709.

Peyre:2009:SIA

[Pey09] Rémi Peyre. Some ideas about quantitative convergence of collision models to their mean field limit. Journal of Statistical Physics, 136(6):1105–1130, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9820-3.

Pauls:2007:BTM

[PF07] W. Pauls and U. Frisch. A Borel transform method for locating singularities of Taylor and Fourier series. Journal of Statistical Physics, 127(6):1095–1119, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9307-z.

Prellberg:2006:CAF

[PFK06] Thomas Prellberg, Jan Fiala, and Peter Kleban. Cluster approximation for the Farey fraction spin chain. Journal of Statistical Physics, 123(2):455–471, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9034-x.

Prestipino:1999:SEL

[PG99] Santi Prestipino and Paolo V. Giaquinta. Statistical entropy of a lattice-gas model: Multiparticle correlation expansion. Journal of Statistical Physics, 96(1–2):135–167, July 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004520432275. See errata [PG00].

Prestipino:2000:ESE

[PG00] Santi Prestipino and Paolo V. Giaquinta. Errata: Statistical Entropy of a Lattice-Gas Model: Multiparticle Correlation Expansion. Journal of Statistical Physics, 98(1–2):507–509, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018603728546. See [PG99].

Prestipino:2003:CEE

[PG03] Santi Prestipino and Paolo V. Giaquinta. The concavity of entropy and extremum principles in thermodynamics. Journal of Statistical Physics, 111(1–2):479–493, April 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022233814184.

Pleimling:2001:CCL

[PH01] Michel Pleimling and Alfred Hüller. Crossing the coexistence line at constant magnetization. Journal of Statistical Physics, 104(5–6):971–989, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010445526237.

PhD:2007:LBM

[PhD07] Irina Ginzburg PhD. Lattice Boltzmann modeling with discontinuous collision components: Hydrodynamic and advection– diffusion equations. Journal of Statistical Physics, 126(1): 157–206, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9234-4.

Piasecki:2000:SAW

[Pia00] Jaroslaw Piasecki. Stochastic adiabatic wall. Journal of Statistical Physics, 101(1–2):703, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026459422711.

Piasecki:2001:DVI

[Pia01] J. Piasecki. Drift velocity induced by collisions. Journal of Statistical Physics, 104(5–6):1145–1154, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010457828963.

Posazhennikova:2003:CBF

[PIRB03] A. I. Posazhennikova, J. O. Indekeu, D. Ross, and D. Bonn. Cross-over between first-order and critical wetting at the liquid– vapour interface of n-alkane/methanol mixtures: Tricritical wetting and critical prewetting. Journal of Statistical Physics, 110 (3–6):611–658, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022151600747.

Pivato:2003:MCA

[Piv03] Marcus Pivato. Multiplicative cellular automata on nilpotent groups: Structure, entropy, and asymptotics. Journal of Statistical Physics, 110(1–2):247–267, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021074813291.

Pruessner:2004:WCP

[PM04] Gunnar Pruessner and Nicholas R. Moloney. Winding clusters in percolation on the torus and the Möbius strip. Journal of Statistical Physics, 115(3–4):839–853, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022376.25660.7b.

Pulkkinen:2005:PTM

[PM05] Otto Pulkkinen and Juha Merikoski. Phase transitions on Markovian bipartite graphs — an application of the zero-range process. Journal of Statistical Physics, 119(3–4):881–907, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-3011-7.

Povolotsky:2006:BAS

[PM06] A. M. Povolotsky and J. F. F. Mendes. Bethe ansatz solution of discrete time stochastic processes with fully parallel update. Journal of Statistical Physics, 123(1):125–166, April 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9023-0.

Prodan:2003:KSE

[PN03] E. Prodan and P. Nordlander. On the Kohn–Sham equations with periodic background potentials. Journal of Statistical Physics, 111(3–4):967–992, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022810601639.

Pavan:2007:NME

[PO07] V. Pavan and L. Oxarango. A new momentum equation for gas flow in porous media: The Klinkenberg effect seen through the kinetic theory. Journal of Statistical Physics, 126(2):355–389, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9110-2.

Podgornik:2001:BRBb

[Pod01a] Rudi Podgornik. Book review: An Introduction to Chaos in Nonequilibrium Statistical Mechanics Cambridge. J. R. Dorfman, Cambridge Lecture Notes in Physics 14, Cambridge University 1999. Journal of Statistical Physics, 104(1–2):483–485, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010330332297.

Podgornik:2001:BRBa

[Pod01b] Rudi Podgornik. Book review: Quantum Phase Transitions. S. Sachdev, Cambridge University Press, 1999. Journal of Statistical Physics, 103(5–6):1139–1141, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010377426014.

Podgornik:2002:BRBa

[Pod02a] Rudi Podgornik. Book review: Fluctuations and Scaling in Biology. Tamas Vicsek. Oxford University Press, 2001, 246 p. Journal of Statistical Physics, 106(3–4):851–852, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013730927298.

Podgornik:2002:BRBb

[Pod02b] Rudi Podgornik. Book review: Statistical Physics of Spin Glasses and Information Processing. Hidetoshi Nishimori, Oxford Science Publications, Oxford University Press, 2001. Journal of Statistical Physics, 109(1–2):335–337, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020091132740.

Podgornik:2003:BRB

[Pod03] Rudi Podgornik. Book review: Phase Transition Dynamics. Akira Onuki, 714 pp., Cambridge University Press, 2002. Journal of Statistical Physics, 112(1–2):429–432, July 2003.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023652309452.

Podgornik:2004:BRBb

[Pod04a] Rudi Podgornik. Book review: Polymer Physics. M. Rubinshtein and R. H. Colby, Oxford University Press, 2003. Journal of Statistical Physics, 115(5–6):1757–1761, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028242.25527.8b.

Podgornik:2004:BRBa

[Pod04b] Rudi Podgornik. Book review: The Physics of Phase Transitions. P. Papon, J. Leblond and P. H. E., Meijer, Springer– Verlag, New York, 2002. Journal of Statistical Physics, 115(3– 4):1135–1137, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000022575.57136.89.

Podgornik:2005:BRB

[Pod05] Rudi Podgornik. Book review: Structured Fluids, Polymers, Colloids, Surfactants: Structured Fluids, Polymers, Colloids, Surfactants. Journal of Statistical Physics, 119(1–2):453–455, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2102-1.

Podgornik:2006:SPP

[Pod06a] R. Podgornik. Statistical physics of polymers. Journal of Statistical Physics, 122(2):377–380, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8021-y.

Podgornik:2006:BRB

[Pod06b] Rudi Podgornik. Book review: Statistics of Linear Polymers in Disordered Media. Journal of Statistical Physics, 123(2): 499–501, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9068-0.

Podgornik:2007:BLA

[Pod07a] Rudi Podgornik. Energy Landscapes: Applications to Clusters, Biomolecules and Glasses (Cambridge Molecular Science). Journal of Statistical Physics, 126(2):423–427, January 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9264-y.

Podgornik:2007:CDG

[Pod07b] Rudi Podgornik. C. de Dominicis and I. Giardina: Random fields and spin glasses: a field theory approach. Journal of Statistical Physics, 129(1):191–192, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9388-8.

Podgornik:2007:AVJ

[Pod07c] Rudi Podgornik. S. Albeverio, V. Jentsch, H. Kantz (eds.): Extreme events in nature and society. (The Frontiers Collection). Journal of Statistical Physics, 129(1):189–190, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9389-7.

Podgornik:2008:IHM

[Pod08] Rudi Podgornik. Igor herbut: A modern approach to critical phenomena. Journal of Statistical Physics, 131(2):381–383, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9494-2.

Polewczak:2000:KTS

[Pol00] Jacek Polewczak. The kinetic theory of simple reacting spheres: I. Global existence result in a dilute-gas case. Journal of Statistical Physics, 100(1–2):327–362, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018608216136.

Pomeau:2005:FSP

[Pom05] Yves Pomeau. Friction on a spinning piece of matter. Journal of Statistical Physics, 121(5–6):1083–1095, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-5966-9.

Popov:2001:FRE

[Pop01] S. Yu. Popov. Frogs in random environment. Journal of Statistical Physics, 102(1–2):191–201, January 2001. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026516826875.

Porra:2004:BRB

[Por04] Josep M. Porrà. Book review: Why Stock Markets Crash: Critical Events in Complex Financial Systems. Didier Sornette, Princeton University Press, Princeton, 2003. Journal of Statistical Physics, 117(3–4):773–774, November 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-3456-0.

Porra:2006:MBM

[Por06] Josep M. Porrà. The (mis)behavior of markets. Journal of Statistical Physics, 122(2):373–375, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8004-Z.

Patsahan:2001:DOP

[PP01] O. V. Patsahan and T. M. Patsahan. Determination of the order parameter in binary fluid mixtures. Journal of Statistical Physics, 105(1–2):285–307, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1012242312803.

Povolotsky:2003:AAP

[PPH03] A. M. Povolotsky, V. B. Priezzhev, and Chin-Kun Hu. The asymmetric avalanche process. Journal of Statistical Physics, 111(5–6):1149–1182, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023048115250.

Pini:2000:FIA

[PPR00] D. Pini, A. Parola, and L. Reatto. Fluctuations-inclusive approach to phase transitions in binary mixtures. Journal of Statistical Physics, 100(1–2):13–38, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018657307050.

Piacquadio:2007:MSE

[PR07] M. Piacquadio and M. Rosen. Multifractal spectrum of an experimental (video feedback) Farey tree. Journal of Statistical Physics, 127(4):783–804, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9217-5.

Pretti:2003:NCT

[Pre03] Marco Pretti. A note on cactus trees: Variational vs. recursive approach. Journal of Statistical Physics, 111(3–4):993–1015, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022862618478.

Pretti:2005:CKN

[Pre05] Marco Pretti. On the convergence of Kikuchi’s natural iteration method. Journal of Statistical Physics, 119(3–4):659–675, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-4426-x.

Pretti:2007:AVA

[Pre07] M. Pretti. Alternative variational approach to cactus lattices. Journal of Statistical Physics, 127(6):1237–1253, June 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9324-y.

Priezzhev:2000:UCD

[Pri00] V. B. Priezzhev. The upper critical dimension of the Abelian sandpile model. Journal of Statistical Physics, 98(3–4):667–684, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018619323983.

Privman:2003:IDO

[Pri03] Vladimir Privman. Initial decoherence of open quantum systems. Journal of Statistical Physics, 110(3–6):957–970, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022180407107.

Procacci:2007:APM

[Pro07] Aldo Procacci. Abstract polymer models with general pair interactions. Journal of Statistical Physics, 129(1):171–188, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/s10955-007-9378-x. See erratum and addendum

[Pro09a].

Procacci:2009:EAA

[Pro09a] Aldo Procacci. Erratum and addendum: ‘Abstract Polymer Models with General Pair Interactions’. Journal of Statistical Physics, 135(4):779–786, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10955-
009-9753-x.pdf. See [Pro07].

Proville:2009:DDE

[Pro09b] Laurent Proville. Depinning of a discrete elastic string from a random array of weak pinning points with finite dimensions. Journal of Statistical Physics, 137(4):717–727, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-009-9860-8.

Pruisken:2001:DUT

[Pru01] A. M. M. Pruisken. At the dawn of a unifying theory. Journal of Statistical Physics, 103(3–4):569–573, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010341316149.

Penrose:2000:CCP

[PS00] Oliver Penrose and George Stell. Close to close packing. Journal of Statistical Physics, 100(1–2):89–95, July 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018679309775.

Procacci:2001:DCU

[PS01] Aldo Procacci and Benedetto Scoppola. On decay of correlations for unbounded spin systems with arbitrary boundary conditions. Journal of Statistical Physics, 105(3–4):453–482, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012267523688.

Patrascioiu:2002:PES

[PS02a] Adrian Patrascioiu and Erhard Seiler. Percolation and the existence of a soft phase in the classical Heisenberg model.

Journal of Statistical Physics, 106(3–4):811–826, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013726826390.

Polewczak:2002:TCS

[PS02b] Jacek Polewczak and George Stell. Transport coefficients in some stochastic models of the revised Enskog equation. Journal of Statistical Physics, 109(3–4):569–590, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020406413636.

Prahofer:2002:SIP

[PS02c] Michael Prähofer and Herbert Spohn. Scale invariance of the PNG droplet and the Airy process. Journal of Statistical Physics, 108(5–6):1071–1106, September 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019791415147.

Popkov:2003:SED

[PS03a] Vladislav Popkov and Gunter M. Schütz. Shocks and excitation dynamics in a driven diffusive two-channel system. Journal of Statistical Physics, 112(3–4):523–540, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023819807616.

Pumir:2003:LPA

[PS03b] Alain Pumir and Boris I. Shraiman. Lagrangian particle approach to large eddy simulations of hydrodynamic turbulence. Journal of Statistical Physics, 113(5–6):693–700, December 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1027300400526.

Prahofer:2004:ESF

[PS04a] Michael Prähofer and Herbert Spohn. Exact scaling functions for one-dimensional stationary KPZ growth. Journal of Statistical Physics, 115(1–2):255–279, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019810.21828.fc.

Procacci:2004:IGN

[PS04b] Aldo Procacci and Benedetto Scoppola. Infinite graphs with a nontrivial bond percolation threshold: Some sufficient conditions. Journal of Statistical Physics, 115(3–4):1113–1127, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000022369.76414.33.

Procacci:2006:AMP

[PS06] Aldo Procacci and Benedetto Scoppola. Analyticity and mixing properties for random cluster model with q > 0 on Z^{d}. Journal of Statistical Physics, 123(6):1285–1310, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9117-8.

Pavliotis:2007:PEM

[PS07a] G. A. Pavliotis and A. M. Stuart. Parameter estimation for multiscale diffusions. Journal of Statistical Physics, 127(4):741– 781, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9300-6.

Peche:2007:WRM

[PS07b] Sandrine Péché and Alexander Soshnikov. Wigner random matrices with non-symmetrically distributed entries. Journal of Statistical Physics, 129(5–6):857–884, December 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9340-y.

Pastur:2008:BUR

[PS08a] L. Pastur and M. Shcherbina. Bulk universality and related properties of Hermitian matrix models. Journal of Statistical Physics, 130(2):205–250, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9434-6.

Pavlyukevich:2008:ODS

[PS08b] Ilya Pavlyukevich and Igor M. Sokolov. One-dimensional spacediscrete transport subject to Lévy perturbations. Journal of Statistical Physics, 133(1):205–215, October 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9607-y.

Prager:2006:DDP

[PSG06] T. Prager and L. Schimansky-Geier. Drift and diffusion in periodically driven renewal processes. Journal of Statistical Physics, 123(2):391–413, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9029-7.

Pugh:2004:PDI

[PSW04] Charles Pugh, Michael Shub, and Amie Wilkinson. Partial differentiability of invariant splittings. Journal of Statistical Physics, 114(3–4):891–921, February 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012511.80422.33.

Peppin:2009:ORP

[PSW09] S. S. L. Peppin, M. J. Spannuth, and J. S. Wettlaufer. Onsager reciprocity in premelting solids. Journal of Statistical Physics, 134(4):701–708, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-009-9699-z.

Pulvirenti:2004:API

[PT04] Ada Pulvirenti and Giuseppe Toscani. Asymptotic properties of the inelastic Kac model. Journal of Statistical Physics, 114(5– 6):1453–1480, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013964.98706.00.

Pareschi:2006:SSP

[PT06] Lorenzo Pareschi and Giuseppe Toscani. Self-similarity and power-like tails in nonconservative kinetic models. Journal of Statistical Physics, 124(2–4):747–779, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9025-y.

Pakonski:2003:FLG

[PTZ03] Prot Pakoński, Gregor Tanner, and Karol Zyczkowski. Families of line-graphs and their quantization. Journal of Statistical Physics, 111(5–6):1331–1352, June 2003. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1023012502046.

Planes:2002:EFS

[PV02] Antoni Planes and Eduard Vives. Entropic formulation of statistical mechanics. Journal of Statistical Physics, 106(3–4): 827–850, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013778810460.

Pavlo:2002:PRU

[PVV02] Pavol Pavlo, George Vahala, and Linda Vahala. Preliminary results in the use of energy-dependent octagonal lattices for thermal lattice Boltzmann simulations. Journal of Statistical Physics, 107(1–2):499–519, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014543611970.

Pule:2005:DTM

[PVZ05] Joseph V. Pulé, André F. Verbeure, and Valentin A. Zagrebnov. A Dicke type model for equilibrium BEC superradiance. Journal of Statistical Physics, 119(1–2):309–329, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2007-z.

Pagonabarraga:2002:BFD

[PWC02] Ignacio Pagonabarraga, Alexander J. Wagner, and M. E. Cates. Binary fluid demixing: The crossover region. Journal of Statistical Physics, 107(1–2):39–52, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1014594101067.

Pechersky:2009:PPN

[PY09a] E. Pechersky and A. Yambartsev. Percolation properties of the non-ideal gas. Journal of Statistical Physics, 137(3):501– 520, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9856-4.

Putz:2009:HSM

[PY09b] V. B. Putz and J. M. Yeomans. Hydrodynamic synchronisation of model microswimmers. Journal of Statistical Physics, 137(5–6):1001–1013, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9826-x.

Parisi:2006:HDB

[PZ06a] Giorgio Parisi and Francesco Zamponi. On the high density behavior of Hamming codes with fixed minimum distance. Journal of Statistical Physics, 123(6):1145–1167, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9142-7.

Pesin:2006:PTU

[PZ06b] Yakov Pesin and Ke Zhang. Phase transitions for uniformly expanding maps. Journal of Statistical Physics, 122(6):1095–1110, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-9005-7.

Qu:2002:NTS

[QAK02] X. Qu, M. Aldana, and Leo P. Kadanoff. Numerical and theoretical studies of noise effects in the Kauffman model. Journal of Statistical Physics, 109(5–6):967–986, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020416308456.

Qi:2002:LMO

[QLAS02] Dewei Qi, Lishi Luo, Raja Aravamuthan, and William Strieder. Lateral migration and orientation of elliptical particles in Poiseuille flows. Journal of Statistical Physics, 107(1–2):101– 120, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1014502402884.

Qian:2002:TGD

[QQT02] Hong Qian, Min Qian, and Xiang Tang. Thermodynamics of the general diffusion process: Time-reversibility and entropy production. Journal of Statistical Physics, 107(5–6):1129–1141, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015109708454.

Racz:2000:PEF

[Rác00] Z. Rácz. Presence of energy flux in quantum spin chains: An experimental signature. Journal of Statistical Physics, 101(1– 2):273–281, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026478508888.

Radin:2008:RCP

[Rad08a] Charles Radin. Random close packing of granular matter. Journal of Statistical Physics, 131(4):567–573, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9523-1.

Radulescu:2008:CTE

[Rad08b] Anca Radulescu. Computing topological entropy in a space of quartic polynomials. Journal of Statistical Physics, 130(2): 373–385, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9428-4.

Rajabpour:2009:BPC

[Raj09] M. A. Rajabpour. Bessel process and conformal quantum mechanics. Journal of Statistical Physics, 136(4):785–805, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9806-1.

Ramm:2007:DPW

[Ram07] A. G. Ramm. Distribution of particles which produces a ‘Smart’ material. Journal of Statistical Physics, 127(5):915–934, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9303-3.

Rapaport:2001:BRB

[Rap01] D. C. Rapaport. Book review: A Guide to Monte Carlo Simulations in Statistical Physics. D. P. Landau and K. Binder, Cambridge University Press, 2000, pp. 384 + xiii. Journal of Statistical Physics, 103(5–6):1143–1145, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010329510084.

Rath:2009:MFF

[Rát09] Balázs Ráth. Mean field frozen percolation. Journal of Statistical Physics, 137(3):459–499, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9863-5.

Rawlings:2003:MGR

[Raw03] Philip K. Rawlings. Modes of a Gaussian random walk. Journal of Statistical Physics, 111(3–4):769–788, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022846114843.

Rebenshtok:2008:WNE

[RB08] A. Rebenshtok and E. Barkai. Weakly non-ergodic statistical physics. Journal of Statistical Physics, 133(3):565–586, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9610-3.

Raimbault:2001:SGT

[RC01] Jean-Luc Raimbault and Jean-Michel Caillol. Sine-Gordon theory for the equation of state of classical hard-core Coulomb systems. II. High-temperature expansion. Journal of Statistical Physics, 103(5–6):777–799, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010348619653.

Roy:2008:HTO

[RD08] Dibyendu Roy and Abhishek Dhar. Heat transport in ordered harmonic lattices. Journal of Statistical Physics, 131(3):535–541, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9487-1.

Reuse:2004:CDS

[RdCM04] F. A. Reuse, V. de Coulon, and K. Maschke. Coherent and dissipative spin dynamics in n-electron systems. Journal of Statistical Physics, 114(1–2):361–453, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003115.75917.08.

Reiss:2000:MRB

[Rei00] Howard Reiss. ‘Magic relation’ between the structures of coexisting phases at a first-order phase transition in a hard sphere system. Journal of Statistical Physics, 100(1–2):73–87, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018627325705.

Reiner:2005:ICSa

[Rei05] Albert Reiner. Infinite compressibility states in the hierarchical reference theory of fluids. I. Analytical considerations. Journal of Statistical Physics, 118(5–6):1107–1127, March 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2017-x.

Reimann:2008:TPS

[Rei08] Peter Reimann. Typicality of pure states randomly sampled according to the Gaussian adjusted projected measure. Journal of Statistical Physics, 132(5):921–935, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9576-1.

Rouet:2002:GCP

[RF02] J. L. Rouet and M. R. Feix. A generalization of the Collatz problem. Building cycles and a stochastic approach. Journal of Statistical Physics, 107(5–6):1283–1298, June 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015122111180.

Richard:2004:PSM

[RG04] C. Richard and A. J. Guttmann. Poland–Scheraga models and the DNA denaturation transition. Journal of Statistical Physics, 115(3–4):925–947, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000022370.48118.8b.

Risau-Gusman:2006:TPT

[RGRTS06] Sebastián Risau-Gusman, Ana C. Ribeiro-Teixeira, and Daniel A. Stariolo. Topology and phase transitions: The case of the short range spherical model. Journal of Statistical Physics, 124(5):

1231–1253, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9178-8.

Rheinlander:2005:CGC

[Rhe05] Martin Rheinländer. A consistent grid coupling method for lattice-Boltzmann schemes. Journal of Statistical Physics, 121 (1–2):49–74, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8412-0.

Rice:2000:ACM

[Ric00] Stuart A. Rice. Active control of molecular dynamics: Coherence versus chaos. Journal of Statistical Physics, 101(1–2):187–212, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026422423909.

Richard:2002:SBT

[Ric02] Christoph Richard. Scaling behaviour of two-dimensional polygon models. Journal of Statistical Physics, 108(3–4):459–493, August 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1015773723188.

Richards:2004:TPP

[Ric04] Donald St. P. Richards. Total positivity properties of generalized hypergeometric functions of matrix argument. Journal of Statistical Physics, 116(1–4):907–922, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037249.50382.92.

Rider:2003:FTL

[Rid03] B. Rider. Fluctuations in the thermodynamic limit of focussing cubic Schrödinger. Journal of Statistical Physics, 113(3–4): 575–594, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026072819239.

Rider:2004:OSG

[Rid04] B. Rider. Order statistics and Ginibre’s ensembles. Journal of Statistical Physics, 114(3–4):1139–1148, February 2004.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012520.37908.07.

Ritort:2001:BRB

[Rit01] Felix Ritort. Book review: Introduction to the Replica Theory of Disordered Statistical Systems. Victor Dotsenko, Cambridge University Press, 2000. Journal of Statistical Physics, 105(1– 2):403–404, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012250614620.

Ritort:2003:BRB

[Rit03] Felix Ritort. Book review: Defects and Geometry in Condensed Matter Physics. D. R. Nelson, Cambridge University Press, Cambridge, 2002. Journal of Statistical Physics, 112(3– 4):883–884, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023840412159.

Ritchie:2006:CTJ

[Rit06] Thomas Logan Ritchie. Construction of the thermodynamic jamming limit for the parking process and other exclusion schemes on Z^{d}. Journal of Statistical Physics, 122(3):381–398, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8025-7.

Rivasseau:2002:TDH

[Riv02] V. Rivasseau. The two dimensional Hubbard model at halffilling. I. Convergent contributions. Journal of Statistical Physics, 106(3–4):693–722, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013770608643.

Rivoire:2004:PAG

[Riv04] Olivier Rivoire. Properties of atypical graphs from negative complexities. Journal of Statistical Physics, 117(3–4):453– 476, November 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-2265-9.

Rikvold:2000:AAV

[RK00] Per Arne Rikvold and M. Kolesik. Analytic approximations for the velocity of field-driven Ising interfaces. Journal of Statistical Physics, 100(1–2):377–403, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018664301115.

Reiner:2005:ICSb

[RK05] Albert Reiner and Gerhard Kahl. Infinite compressibility states in the hierarchical reference theory of fluids. II. Numerical evidence. Journal of Statistical Physics, 118(5–6):1129–1149, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2022-0.

Rateitschak:2000:NHT

[RKH00] K. Rateitschak, R. Klages, and Wm. G. Hoover. The Nosé-Hoover thermostated Lorentz gas. Journal of Statistical Physics, 101(1–2):61–77, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026447620778.

Rateitschak:2000:TDS

[RKN00] K. Rateitschak, R. Klages, and G. Nicolis. Thermostating by deterministic scattering: The periodic Lorentz gas. Journal of Statistical Physics, 99(5–6):1339–1364, June 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018645007533.

Rajesh:2000:CMM

[RM00] R. Rajesh and Satya N. Majumdar. Conserved mass models and particle systems in one dimension. Journal of Statistical Physics, 99(3–4):943–965, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018651714376.

Robinson:2008:LMD

[RM08] D. J. Robinson and G. P. Morriss. Lyapunov mode dynamics in hard-disk systems. Journal of Statistical Physics, 131(1):1–31, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9473-z.

Roma:2004:GSE

[RNVRP04] F. Romá, F. Nieto, E. E. Vogel, and A. J. Ramirez-Pastor. Ground-state entropy of ±J Ising lattices by Monte Carlo simulations. Journal of Statistical Physics, 114(5–6):1325–1341, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000013967.52237.6e.

Robledo:2000:RGO

[Rob00] A. Robledo. The renormalization group and optimization of entropy. Journal of Statistical Physics, 100(1–2):475–487, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018620618862.

Rodriguez:2002:BRB

[Rod02] Rosalio Rodriguez. Book review: Selected Papers of N. G. van Kampen. Paul H. E. Meijer, ed. World Scientific, 2000. Journal of Statistical Physics, 106(5–6):1249–1253, March 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014054307677.

Roepstorff:2003:BRB

[Roe03] G. Roepstorff. Book review: Path Integrals in Physics, Vol. I: Stochastic Processes and Quantum Mechanics (336 p.), Vol. II: Quantum Field Theory, Statistical Physics and Other Modern Applications (345 p.). M. Chaichian and A. Demichev, Institute of Physics Publishing Ltd 2001, Bristol and Philadelphia. Journal of Statistical Physics, 111(1–2):495–496, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022285831022.

Romito:2004:EFD

[Rom04] Marco Romito. Ergodicity of the finite dimensional approximation of the 3D Navier–Stokes equations forced by a degenerate noise. Journal of Statistical Physics, 114(1–2):155–177, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000003108.92097.5c.

Romito:2008:AES

[Rom08] Marco Romito. Analysis of equilibrium states of Markov solutions to the 3D Navier–Stokes equations driven by additive noise. Journal of Statistical Physics, 131(3):415–444, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9477-8.

Rottenstreich:2006:SPA

[Rot06] Sivan Rottenstreich. A stationary phase approach to the weak coupling Schrödinger equation. Journal of Statistical Physics, 124(1):47–86, July 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9157-0.

Rozikov:2006:CDG

[Roz06] U. A. Rozikov. A constructive description of ground states and Gibbs measures for Ising model with two-step interactions on Cayley tree. Journal of Statistical Physics, 122(2):217–235, January 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8029-3.

Rozikov:2008:CMC

[Roz08] U. A. Rozikov. A contour method on Cayley trees. Journal of Statistical Physics, 130(4):801–813, February 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9455-1.

Ramakrishnan:2001:QMP

[RR01] T. V. Ramakrishnan and A. K. Rajagopal. Quasiparticles in the mixed phase of superconducting cuprates: A semiclassical Green’s function approach. Journal of Statistical Physics, 103 (3–4):441–457, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010329013424.

Reiterer:2000:NDM

[RRS^{+}00] P. Reiterer, C. Reitshammer, F. Schürrer, F. Hanser, and T. Eitzenberger. New discrete model Boltzmann equations for arbitrary partitions of the velocity space. Journal of Statistical Physics, 98(1–2):419–440, January 2000. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018643409890.

Rakos:2004:ESM

[RS04a] A. Rákos and G. M. Schütz. Exact shock measures and steadystate selection in a driven diffusive system with two conserved densities. Journal of Statistical Physics, 117(1–2):55–76, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000044064.62295.29.

Reich:2004:RDF

[RS04b] Hendrik Reich and Matthias Schmidt. Replica density functional study of one-dimensional hard core fluids in porous media. Journal of Statistical Physics, 116(5–6):1683–1702, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041752.55138.0a.

Rakos:2005:CDR

[RS05] A. Rákos and G. M. Schütz. Current distribution and random matrix ensembles for an integrable asymmetric fragmentation process. Journal of Statistical Physics, 118(3–4):511–530, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8819-z.

Rousset:2006:ESN

[RS06] Mathias Rousset and Gabriel Stoltz. Equilibrium sampling from nonequilibrium dynamics. Journal of Statistical Physics, 123 (6):1251–1272, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9090-2.

Rajabpour:2008:ERG

[RS08] M. A. Rajabpour and R. Sepehrinia. Explicit renormalization group for D = 2 random bond Ising model with long-range correlated disorder. Journal of Statistical Physics, 130(4):815–820, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9443-5.

Raz:2009:ELR

[RS09] Hillel Raz and Robert Sims. Estimating the Lieb–Robinson velocity for classical anharmonic lattice systems. Journal of Statistical Physics, 137(1):79–108, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9839-5.

Raiskinmaki:2002:LBS

[RSMJ^{+}02] P. Raiskinmäki, A. Shakib-Manesh, A. Jäsberg, A. Koponen, J. Merikoski, and J. Timonen. Lattice-Boltzmann simulation of capillary rise dynamics. Journal of Statistical Physics, 107 (1–2):143–158, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014506503793.

Rybko:2009:SRC

[RSV09] Alexander Rybko, Senya Shlosman, and Alexander Vladimirov. Spontaneous resonances and the coherent states of the queuing networks. Journal of Statistical Physics, 134(1):67–104, January 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9658-0.

Ramos:2008:CMC

[RT08a] A. D. Ramos and A. Toom. Chaos and Monte Carlo approximations of the flip-annihilation process. Journal of Statistical Physics, 133(4):761–771, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9625-9.

Ramos:2008:ECL

[RT08b] A. D. Ramos and A. Toom. An error correction. Letter to the Editor. Journal of Statistical Physics, 131(1):167–168, April 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9483-5.

Rath:2008:TPM

[RT08c] Balázs Ráth and Bálint Tóth. Triangle percolation in mean field random graphs-with PDE. Journal of Statistical Physics, 131(3):385–391, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9499-x.

Rubi:2004:BRB

[Rub04] J. M. Rubi. Book review: Bose–Einstein Condensation. Lev Pitaevskii and Sandro Stringari, Oxford University Press, Oxford, 2003. Journal of Statistical Physics, 115(5–6):1763–1764, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028243.07395.b3.

Rubi:2006:BRB

[Rub06a] J. M. Rubi. Book review: Statistical Physics for Cosmic Structures. Journal of Statistical Physics, 123(3):699–700, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9092-0.

Rubi:2006:TFS

[Rub06b] J. M. Rubi. Theory of fluctuations in superconductors. Journal of Statistical Physics, 123(2):497–498, April 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-9007-5.

Rubi:2009:JZJ

[Rub09a] J. M. Rubi. J. Zinn–Justin: Phase Transitions and Renormalization Group. Journal of Statistical Physics, 134(4):795–796, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9711-7.

Rubi:2009:VMB

[Rub09b] J. M. Rubi. Vieri Mastroprieto: Non-Perturbative Renormalization. Journal of Statistical Physics, 134(4):793–794, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9707-3.

Ruelle:2000:NNS

[Rue00a] David Ruelle. Natural nonequilibrium states in quantum statistical mechanics. Journal of Statistical Physics, 98(1–2):57–75, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018618704438.

Ruelle:2000:REI

[Rue00b] David Ruelle. A remark on the equivalence of isokinetic and isoenergetic thermostats in the thermodynamic limit. Journal of Statistical Physics, 100(3–4):757–763, August 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018679626556.

Ruelle:2002:SIF

[Rue02] David Ruelle. Some ill-formulated problems on regular and messy behavior in statistical mechanics and smooth dynamics for which I would like the advice of Yasha Sinai. Journal of Statistical Physics, 108(5–6):723–728, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019758523808.

Rugh:2008:JRC

[Rug08] Hans Henrik Rugh. J.-R. Chazottes and B. Fernandez (eds): Dynamics of Coupled Map Lattices and of Related Spatially Extended Systems. Journal of Statistical Physics, 130(3):631, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9439-1.

Rutkevich:2001:AVD

[Rut01] S. B. Rutkevich. Analytic verification of the droplet picture in the two-dimensional Ising model. Journal of Statistical Physics, 104(3–4):589–608, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010324620997.

Rutkevich:2008:ESB

[Rut08] S. B. Rutkevich. Energy spectrum of bound-spinons in the quantum Ising spin-chain ferromagnet. Journal of Statistical Physics, 131(5):917–939, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9495-1.

Ruzmaikina:2000:SIH

[Ruz00] A. A. Ruzmaikina. Stieltjes integrals of Hölder continuous functions with applications to fractional Brownian motion. Journal of Statistical Physics, 100(5–6):1049–1069, September 2000.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018754806993.

Radicchi:2007:PTB

[RVMO07] Filippo Radicchi, Daniele Vilone, and Hildegard Meyer-Ortmanns. Phase transition between synchronous and asynchronous updating algorithms. Journal of Statistical Physics, 129(3):593–603, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9416-8.

Ravishankar:2007:LTE

[RY07] K. Ravishankar and Lai-Sang Young. Local thermodynamic equilibrium for some stochastic models of Hamiltonian origin. Journal of Statistical Physics, 128(3):641–665, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9335-8.

Ryckman:2008:LSP

[Ryc08] E. Ryckman. Linear statistics of point processes via orthogonal polynomials. Journal of Statistical Physics, 132(3):473–486, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9564-5.

Santen:2002:AEP

[SA02] Ludger Santen and Cécile Appert. The asymmetric exclusion process revisited: Fluctuations and dynamics in the domain wall picture. Journal of Statistical Physics, 106(1–2):187–199, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013176229983.

Saakian:2007:NMS

[Saa07] David B. Saakian. A new method for the solution of models of biological evolution: Derivation of exact steady–state distributions. Journal of Statistical Physics, 128(3):781–798, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9334-9.

Sabelfeld:2008:SFU

[Sab08] K. K. Sabelfeld. Stokes flows under random boundary velocity excitations. Journal of Statistical Physics, 133(6):1107– 1136, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9654-4.

Sachdev:2004:QIM

[Sac04] Subir Sachdev. Quantum impurity in a magnetic environment. Journal of Statistical Physics, 115(1–2):47–56, April 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019819.67227.f8.

Sacchetti:2005:NDW

[Sac05] Andrea Sacchetti. Nonlinear double-well Schrödinger equations in the semiclassical limit. Journal of Statistical Physics, 119(5– 6):1347–1382, June 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-3766-x.

Sahimi:2002:BRB

[Sah02] Muhammad Sahimi. Book review: Random Heterogeneous Materials. Salvatore Torquato, Springer–Verlag, New York, 2002, 701 pages. Journal of Statistical Physics, 109(1–2):331–333, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1019956115902.

Saied:2000:BNP

[Sai00] Effat A. Saied. On the Brazil nuts problem: Is it of relevance to hydrodynamics? Journal of Statistical Physics, 98(5–6): 1395–1407, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018680015304.

Sakagawa:2000:EBC

[Sak00] Hironobu Sakagawa. Equivalence between canonical Gibbs measures and stationary measures for stochastic lattice-gas model. Journal of Statistical Physics, 98(3–4):949–959, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018696013505.

Sakai:2001:MFC

[Sak01] Akira Sakai. Mean-field critical behavior for the contact process. Journal of Statistical Physics, 104(1–2):111–143, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010320523031.

Sakai:2002:HIC

[Sak02] Akira Sakai. Hyperscaling inequalities for the contact process and oriented percolation. Journal of Statistical Physics, 106(1– 2):201–211, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013197011935.

Sakagawa:2004:ERT

[Sak04a] Hironobu Sakagawa. Entropic repulsion for two dimensional multi-layered harmonic crystals. Journal of Statistical Physics, 114(1–2):37–49, January 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOSS.0000003103.86191.
24.

Sakai:2004:MFB

[Sak04b] Akira Sakai. Mean-field behavior for the survival probability and the percolation point-to-surface connectivity. Journal of Statistical Physics, 117(1–2):111–130, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044061.83860.62. See erratum [Sak05].

Sakai:2005:EMF

[Sak05] Akira Sakai. Erratum on ‘Mean-field Behavior for the Survival Probability and the Percolation Point-to-Surface Connectivity’. Journal of Statistical Physics, 119(1–2):447–448, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-004-2049-2. See [Sak04b].

Sakagawa:2006:ERH

[Sak06] Hironobu Sakagawa. Entropic repulsion for the high dimensional Gaussian lattice field between two walls. Journal of Statistical Physics, 124(5):1255–1274, September 2006. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9049-3.

Sakagawa:2007:BMH

[Sak07] Hironobu Sakagawa. Bounds on the mass for the high dimensional Gaussian lattice field between two hard walls. Journal of Statistical Physics, 129(3):537–553, November 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9394-x.

Sakagawa:2009:ERM

[Sak09] Hironobu Sakagawa. Entropic repulsion of the massless field with a class of self-potentials. Journal of Statistical Physics, 135(3):467–481, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9745-x.

Salmhofer:2009:CFT

[Sal09] Manfred Salmhofer. Clustering of fermionic truncated expectation values via functional integration. Journal of Statistical Physics, 134(5–6):941–952, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9698-0.

Samaj:2000:MCD

[Sam00] L. Samaj. Microscopic calculation of the dielectric susceptibility tensor for Coulomb fluids. Journal of Statistical Physics, 100(5– 6):949–967, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018779412445.

Samaj:2001:STI

[Sam01a] L. Samaj. Surface tension of an ideal dielectric–electrolyte boundary: Exactly solvable model. Journal of Statistical Physics, 103(5–6):737–752, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010344518745.

Samaj:2001:TPO

[Sam01b] L. Samaj. Thermodynamic properties of the one-dimensional two-component log–gas. Journal of Statistical Physics, 105(1– 2):173–191, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012234110986.

Samaj:2003:ESC

[Sam03] L. Samaj. Exact solution of a charge-asymmetric twodimensional Coulomb gas. Journal of Statistical Physics, 111 (1–2):261–290, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022209108732.

Samaj:2004:TDO

[Sam04] L. Samaj. Is the two-dimensional one-component plasma exactly solvable? Journal of Statistical Physics, 117(1–2):131–158, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000044056.19438.2c.

Samaj:2005:AEG

[Sam05a] L. Samaj. Anomalous effects of ‘Guest’ charges immersed in electrolyte: Exact 2D results. Journal of Statistical Physics, 120 (1–2):125–146, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-5477-8.

Samaj:2005:SEP

[Sam05b] L. Samaj. Saturation of electrostatic potential: Exactly solvable 2D Coulomb models. Journal of Statistical Physics, 119 (3–4):459–478, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-4416-z.

Samaj:2006:RHC

[Sam06] L. Samaj. Renormalization of a hard-core guest charge immersed in a two-dimensional electrolyte. Journal of Statistical Physics, 124(5):1179–1206, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9122-y.

Samaj:2007:THT

[Sam07] Ladislav Samaj. A trickiness of the high-temperature limit for number density correlation functions in classical Coulomb fluids. Journal of Statistical Physics, 128(3):569–586, August 2007.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9337-6.

Samaj:2009:CCS

[Sam09] Ladislav Samaj. Charge and current sum rules in quantum media coupled to radiation. Journal of Statistical Physics, 137(1):1–17, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9845-7.

Santoro:2005:HPM

[San05] Manuel Santoro. On the Helmholtz potential metric: The isotherm length–work theorem. Journal of Statistical Physics, 120(3–4):737–755, August 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-7006-1.

Stratford:2005:LBB

[SAPD05] K. Stratford, R. Adhikari, I. Pagonabarraga, and J.-C. Desplat. Lattice Boltzmann for binary fluids with suspended colloids. Journal of Statistical Physics, 121(1–2):163–178, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-8411-1.

Seaton:2001:IRD

[SB01a] Katherine A. Seaton and Murray T. Batchelor. The inversion relation and the dilute a3, 4, 6 eigenspectrum. Journal of Statistical Physics, 102(3–4):1019–1027, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004871524190.

Spehner:2001:KMQ

[SB01b] D. Spehner and J. Bellissard. A kinetic model of quantum jumps. Journal of Statistical Physics, 104(3–4):525–572, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010320520088.

Sreenivasan:2006:CPT

[SB06] K. R. Sreenivasan and A. Bershadskii. Clustering properties in turbulent signals. Journal of Statistical Physics, 125(5–6):

1141–1153, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9112-0.

Stanley:2003:ASP

[SBG03] H. E. Stanley, S. V. Buldyrev, and N. Giovambattista. Application of statistical physics to understand static and dynamic anomalies in liquid water. Journal of Statistical Physics, 110(3– 6):1039–1054, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022188608924.

Starruss:2007:NMC

[SBSAD07] J. Starruß, Th. Bley, L. Søgaard-Andersen, and A. Deutsch. A new mechanism for collective migration in Myxococcus xanthus. Journal of Statistical Physics, 128(1–2):269–286, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9298-9.

Shalizi:2001:CMP

[SC01] Cosma Rohilla Shalizi and James P. Crutchfield. Computational mechanics: Pattern and prediction, structure and simplicity. Journal of Statistical Physics, 104(3–4):817–879, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010388907793.

Shargel:2009:FTE

[SC09] Benjamin Hertz Shargel and Tom Chou. Fluctuation theorems for entropy production and heat dissipation in periodically driven Markov chains. Journal of Statistical Physics, 137(1): 165–188, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9836-8.pdf.

Schutz:2000:ETD

[Sch00] Gunter M. Schütz. Exact tracer diffusion coefficient in the asymmetric random average process. Journal of Statistical Physics, 99 (3–4):1045–1049, May 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018664117102.

Schmeling:2001:EPU

[Sch01] Jörg Schmeling. Entropy preservation under Markov coding. Journal of Statistical Physics, 104(3–4):799–815, August 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010336923722.

Schreiber:2006:DKS

[Sch06] Tomasz Schreiber. Dobrushin–Kotecký-Shlosman theorem for polygonal Markov fields in the plane. Journal of Statistical Physics, 123(3):631–684, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9053-7.

Schinazi:2007:SSM

[Sch07] Rinaldo B. Schinazi. A spatial stochastic model for virus dynamics. Journal of Statistical Physics, 128(3):771–779, August 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9323-z.

Schreiber:2008:NHP

[Sch08] Tomasz Schreiber. Non-homogeneous polygonal Markov fields in the plane: Graphical representations and geometry of higher order correlations. Journal of Statistical Physics, 132(4):669–705, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9584-1.

Semerjian:2004:SDD

[SCM04] G. Semerjian, L. F. Cugliandolo, and A. Montanari. On the stochastic dynamics of disordered spin models. Journal of Statistical Physics, 115(1–2):493–530, April 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000019821.08230.72.

Sumedha:2005:EIE

[SD05] Sumedha and Deepak Dhar. Efficiency of the incomplete enumeration algorithm for Monte–Carlo simulation of linear and branched polymers. Journal of Statistical Physics, 120(1–2):71– 100, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-5462-2. See correction [SD23].

Sumedha:2006:QAS

[SD06] Sumedha and Deepak Dhar. Quenched averages for self-avoiding walks and polygons on deterministic fractals. Journal of Statistical Physics, 125(1):55–76, October 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9098-7.

Simon:2008:QSR

[SD08] Damien Simon and Bernard Derrida. Quasi-stationary regime of a branching random walk in presence of an absorbing wall. Journal of Statistical Physics, 131(2):203–233, April 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9504-4.

Sadhu:2009:SSS

[SD09] Tridib Sadhu and Deepak Dhar. Steady state of stochastic sandpile models. Journal of Statistical Physics, 134(3):427–441, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9683-7.

Sumedha:2023:CEI

[SD23] Sumedha and Deepak Dhar. Correction: Efficiency of the incomplete enumeration algorithm for Monte–Carlo simulation of linear and branched polymers. Journal of Statistical Physics, 190(11):??, November 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03197-1.

Stattenberger:2007:NST

[SDBS07] Günther Stattenberger, Markus Dankesreiter, Florian Baumgartner, and Johannes J. Schneider. On the neighborhood structure of the traveling salesman problem generated by local search moves. Journal of Statistical Physics, 129(4):623–648, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-007-9382-1.

Stapleton:2004:SIC

[SDC04] Matthew Stapleton, Martin Dingler, and Kim Christensen. Sensitivity to initial conditions in self-organized critical systems. Journal of Statistical Physics, 117(5–6):891–900, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-5709-3.

Seaton:2002:UAR

[Sea02] Katherine A. Seaton. A universal amplitude ratio for the q ≤ 4 Potts model from a solvable lattice model. Journal of Statistical Physics, 107(5–6):1255–1265, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1015118010271.

Semerjian:2008:FVR

[Sem08] Guilhem Semerjian. On the freezing of variables in random constraint satisfaction problems. Journal of Statistical Physics, 130 (2):251–293, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9417-7.

Seppalainen:2001:HPT

[Sep01] Timo Seppäläinen. Hydrodynamic profiles for the totally asymmetric exclusion process with a slow bond. Journal of Statistical Physics, 102(1–2):69–96, January 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026508625058.

Sergeev:2006:TLS

[Ser06] S. M. Sergeev. Thermodynamic limit for a spin lattice. Journal of Statistical Physics, 123(6):1231–1250, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9128-5.

Sethuraman:2006:SOT

[Set06] Sunder Sethuraman. Superdiffusivity of occupation-time variance in 2-dimensional asymmetric exclusion processes with density ρ = 1/2. Journal of Statistical Physics, 123(4):787–802, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9061-7.

Serrano:2005:VFP

[SEZ05] Mar Serrano, Pep Español, and Ignacio Zúñiga. Voronoi fluid particle model for Euler equations. Journal of Statistical Physics, 121(1–2):133–147, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8414-y.

Salem:2007:CTP

[SF07a] Walid K. Abou Salem and Jürg Fröhlich. Cyclic thermodynamic processes and entropy production. Journal of Statistical Physics, 126(3):431–466, February 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9111-1.

Salem:2007:SFL

[SF07b] Walid K. Abou Salem and Jürg Fröhlich. Status of the fundamental laws of thermodynamics. Journal of Statistical Physics, 126(4–5):1045–1068, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9222-8.

Succi:2002:TRL

[SFCO02] S. Succi, O. Filippova, H. Chen, and S. Orszag. Towards a renormalized lattice Boltzmann equation for fluid turbulence. Journal of Statistical Physics, 107(1–2):261–278, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014570923357.

Seeger:2005:DHC

[SH05] S. Seeger and K. H. Hoffmann. On the domain of hyperbolicity of the cumulant equations. Journal of Statistical Physics, 121 (1–2):75–90, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-6969-2.

Schutz:2007:HZR

[SH07] G. M. Schütz and R. J. Harris. Hydrodynamics of the zerorange process in the condensation regime. Journal of Statistical Physics, 127(2):419–430, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9280-6.

Shankar:2001:LRR

[Sha01] R. Shankar. Luttinger revisited — the renormalization group approach. Journal of Statistical Physics, 103(3–4):485–502, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010333114332.

Shankar:2006:DD

[Sha06] R. Shankar. Dots for dummies. Journal of Statistical Physics, 125(5–6):1173–1182, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-005-9018-2.

Sharipov:2010:CSL

[Sha10] Felix Sharipov. Comments on ‘Symmetry of the Linearized Boltzmann Equation’ by S. Takata. Journal of Statistical Physics, 139(3):536–537, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9955-2. See

[Tak09a, Tak09b].

Shcherbina:2009:EUO

[Shc09] M. Shcherbina. Edge universality for orthogonal ensembles of random matrices. Journal of Statistical Physics, 136(1):35–50, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9766-5.

Schlesener:2003:CCF

[SHD03] F. Schlesener, A. Hanke, and S. Dietrich. Critical Casimir forces in colloidal suspensions. Journal of Statistical Physics, 110(3– 6):981–1013, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022184508016.

Shirai:2006:LDF

[Shi06a] Tomoyuki Shirai. Large deviations for the fermion point process associated with the exponential kernel. Journal of Statistical Physics, 123(3):615–629, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9026-x.

Shiwa:2006:ERG

[Shi06b] Y. Shiwa. Exact renormalization group for the Brazovskii model of striped patterns. Journal of Statistical Physics, 124(5):1207– 1229, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9080-4.

Shiozawa:2009:CLT

[Shi09] Yuichi Shiozawa. Central limit theorem for branching Brownian motions in random environment. Journal of Statistical Physics, 136(1):145–163, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9774-5.

Shlesinger:2004:BRB

[Shl04] Michael F. Shlesinger. Book review: Dynamics and Thermodynamics of Systems with Long-Range Interactions. T. Dauxois, S. Ruffo, E. Arimondo, and M. Wilkens (eds.). Springer–Verlag, Berlin, 2002. Journal of Statistical Physics, 115(5–6):1765–1766, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000028244.59318.e3.

Shlesinger:2005:BRB

[Shl05] Michael F. Shlesinger. Book review: Modeling Complex Systems. Journal of Statistical Physics, 119(3–4):949–950, May 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-2132-8.

Shlesinger:2007:CAS

[Shl07] Michael F. Shlesinger. Complex adaptive systems: An introduction to computational models of social life. Journal of Statistical Physics, 129(2):409–410, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-007-9384-z.

Shneidman:2003:LEN

[Shn03] Vitaly A. Shneidman. On the lowest energy nucleation path in a supersaturated lattice gas. Journal of Statistical Physics, 112 (1–2):293–318, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023687822656.

Sasamoto:2004:FOD

[SI04] T. Sasamoto and T. Imamura. Fluctuations of the onedimensional polynuclear growth model in half-space. Journal of Statistical Physics, 115(3–4):749–803, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022374.73462.85.

Siders:2005:EFO

[Sid05] Paul D. Siders. Effects of field orientation on the driven lattice gas. Journal of Statistical Physics, 119(3–4):861–880, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-005-4427-9.

Sigal:2009:GSR

[Sig09] Israel Michael Sigal. Ground state and resonances in the standard model of the non-relativistic QED. Journal of Statistical Physics, 134(5–6):899–939, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9721-5.

Simon:2008:HDR

[Sim08] Thomas Simon. On the Hausdorff dimension of regular points of inviscid Burgers equation with stable initial data. Journal of Statistical Physics, 131(4):733–747, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9508-0.

Simon:2009:CJT

[Sim09] Barry Simon. A celebration of Jürg and Tom. Journal of Statistical Physics, 134(5–6):809–812, March 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9627-7.

Sinai:2002:WMO

[Sin02] Ya. G. Sinai. What, in my opinion, David Ruelle should do in the coming years? Journal of Statistical Physics, 108(5–6): 729–732, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019710607878.

Sinai:2005:PSS

[Sin05] Yakov Sinai. Power series for solutions of the 3D–Navier– Stokes system on R^{3}. Journal of Statistical Physics, 121(5–6): 779–803, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8670-x.

Singh:2008:OMT

[Sin08] Navinder Singh. Onsager–Machlup theory and work fluctuation theorem for a harmonically driven Brownian particle. Journal of Statistical Physics, 131(3):405–414, May 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9503-5.

Sinclair:2009:CFE

[Sin09] Christopher D. Sinclair. Correlation functions for β = 1 ensembles of matrices of odd size. Journal of Statistical Physics, 136(1):17–33, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9771-8.

Samaj:2001:STM

[SJ01a] L. Samaj and B. Jancovici. Surface tension of a metal– electrolyte boundary: Exactly solvable model. Journal of Statistical Physics, 103(5–6):717–735, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010392401906.

Schmuser:2001:NEB

[SJ01b] Frank Schmüser and Wolfram Just. Non-equilibrium behaviour in unidirectionally coupled map lattices. Journal of Statistical Physics, 105(3–4):525–559, November 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1012223708667.

Stenull:2001:RFT

[SJ01c] Olaf Stenull and Hans-Karl Janssen. Renormalized field theory of resistor diode percolation. Journal of Statistical Physics, 104(5–6):1273–1297, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010418214851.

Samaj:2002:DCT

[SJ02a] L. Samaj and B. Jancovici. Density correlations in the twodimensional Coulomb gas. Journal of Statistical Physics, 106(1– 2):323–355, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013136515871.

Samaj:2002:LDB

[SJ02b] L. Samaj and B. Jancovici. Large-distance behavior of particle correlations in the two-dimensional two-component plasma. Journal of Statistical Physics, 106(1–2):301–321, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013184431800.

Sapin:2007:UQL

[SJW07] O. Sapin, H. R. Jauslin, and Stefan Weigert. Upper quantum Lyapunov exponent and Anosov relations for quantum systems driven by a classical flow. Journal of Statistical Physics, 127 (4):699–719, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9310-4.

Sharipov:2006:OCR

[SK06] Felix Sharipov and Denize Kalempa. Onsager–Casimir reciprocal relations based on the Boltzmann equation and gas–surface interaction. Gaseous mixtures. Journal of Statistical Physics, 125(3):661–675, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9200-1.

Seferoglu:2007:CBB

[SK07] N. Seferoglu and B. Kutlu. Critical behavior of the Blume– Emery–Griffiths model for a simple cubic lattice on the cellular automaton. Journal of Statistical Physics, 129(3):453– 468, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9392-z.

Straley:2004:BMO

[SKM04] Joseph P. Straley, Eugene B. Kolomeisky, and Stephen C. Milne. The Bose molecule in one dimension. Journal of Statistical Physics, 116(5–6):1579–1596, September 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041749.74054.b9.

Skrypnik:2000:LLS

[Skr00] W. I. Skrypnik. LRO in lattice systems of linear classical and quantum oscillators. Strong nearest-neighbor pair quadratic interaction. Journal of Statistical Physics, 100(5–6):853–870, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1023/A%3A1018767109719.

Skrypnik:2003:LRO

[Skr03] W. I. Skrypnik. Long-range order in nonequilibrium systems of interacting Brownian linear oscillators. Journal of Statistical Physics, 111(1–2):291–321, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022261125570.

Subramanian:2009:SAS

[SKT09] Natarajan Subramanian, Louise H. Kellogg, and Donald L. Turcotte. Statistics of advective stretching in three-dimensional incompressible flows. Journal of Statistical Physics, 136(5): 926–944, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9807-0.pdf.

Sridhar:2002:SBR

[SL02] S. Sridhar and W. T. Lu. Sinai billiards, Ruelle zeta-functions and Ruelle resonances: Microwave experiments. Journal of Statistical Physics, 108(5–6):755–765, September 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1019714808787.

Stilck:2002:TPS

[SLA02] Jürgen F. Stilck, Yan Levin, and Jeferson J. Arenzon. Thermodynamic properties of a simple model of like-charged attracting rods. Journal of Statistical Physics, 106(1–2):287–299, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1013132414962.

Slanina:2009:EIB

[Sla09a] Frantisek Slanina. Efficiency of interacting Brownian motors: Improved mean-field treatment. Journal of Statistical Physics, 135(5–6):935–950, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9743-z.

Slawny:2009:SDT

[Sla09b] Joseph Slawny. Self-dual transitive spin 1/2 models. Journal of Statistical Physics, 135(4):639–650, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9659-z.

Seshadri:2000:CWP

[SLB00] S. Seshadri, S. Lakshmibala, and V. Balakrishnan. Control of wave packet revivals using geometric phases. Journal of Statistical Physics, 101(1–2):213–223, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026488815731.

Szulga:2001:WMP

[SM01] Jerzy Szulga and Fred Molz. The Weierstrass–Mandelbrot process revisited. Journal of Statistical Physics, 104(5–6):1317– 1348, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010422315759.

Sengupta:2006:CMC

[SM06] A. Mayukh Sengupta and Partha Pratim Mitra. Capacity of multivariate channels with multiplicative noise: Random matrix techniques and large-n expansions (2). Journal of Statistical Physics, 125(5–6):1223–1242, December 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9076-0.

Schehr:2008:RRR

[SM08] Grégory Schehr and Satya N. Majumdar. Real roots of random polynomials and zero crossing properties of diffusion equation. Journal of Statistical Physics, 132(2):235–273, July 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9574-3.

Schehr:2009:CRR

[SM09] Grégory Schehr and Satya N. Majumdar. Condensation of the roots of real random polynomials on the real axis. Journal of Statistical Physics, 135(4):587–598, May 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9755-8.

Smereka:2008:LTB

[Sme08] Peter Smereka. Long time behavior of a modified Becker-Döring system. Journal of Statistical Physics, 132(3):519–533, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9552-9.

Shakib-Manesh:2002:SSC

[SMRK^{+}02] A. Shakib-Manesh, P. Raiskinmäki, A. Koponen, M. Kataja, and J. Timonen. Shear stress in a Couette flow of liquid–particle suspensions. Journal of Statistical Physics, 107(1–2):67–84, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014598201975.

Schneider-Mizell:2009:GVM

[SMS09] Casey M. Schneider-Mizell and Leonard M. Sander. A generalized voter model on complex networks. Journal of Statistical Physics, 136(1):59–71, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9757-6.

Shahbazi:2003:STS

[SMT03] F. Shahbazi, A. A. Masoudi, and M. Reza Rahimi Tabar. Singularity time scale of the Kardar–Parisi–Zhang equation in 2+1 dimensions. Journal of Statistical Physics, 112(3–4):437–456, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023821922637.

Shu:2005:LBK

[SNC05] C. Shu, X. D. Niu, and Y. T. Chew. A lattice Boltzmann kinetic model for microflow and heat transfer. Journal of Statistical Physics, 121(1–2):239–255, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8413-z.

Schor:2000:DBS

[SO00a] Ricardo S. Schor and Michael O’Carroll. Decay of the Bethe– Salpeter kernel and bound states for lattice classical ferromagnetic spin systems at high temperature. Journal of Statistical Physics, 99(5–6):1265–1279, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018688722554.

Schor:2000:TMS

[SO00b] Ricardo S. Schor and Michael O’Carroll. Transfer matrix spectrum and bound states for lattice classical ferromagnetic spin systems at high temperature. Journal of Statistical Physics, 99 (5–6):1207–1223, June 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018632604807.

Schor:2002:TMS

[SO02] Ricardo S. Schor and Michael O’Carroll. Transfer matrix spectrum for lattice classical O(N) ferromagnetic spin systems at high temperature. Journal of Statistical Physics, 109(1–2): 279–288, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019947914085.

Shreif:2008:CAA

[SO08] Z. Shreif and P. Ortoleva. Curvilinear all-atom multiscale (CAM) theory of macromolecular dynamics. Journal of Statistical Physics, 130(4):669–685, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9452-4.

Sodin:2009:TWL

[Sod09] Sasha Sodin. The Tracy–Widom law for some sparse random matrices. Journal of Statistical Physics, 136(5):834–841, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9813-2. See erratum [Sod17].

Sodin:2017:ETW

[Sod17] Sasha Sodin. Erratum to: The Tracy–Widom Law for Some Sparse Random Matrices. Journal of Statistical Physics, 166(5): 1343, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-017-1715-0.pdf. See [Sod09].

Sonigo:2009:SIT

[Son09] Nicky Sonigo. Semi-infinite TASEP with a complex boundary mechanism. Journal of Statistical Physics, 136(6):1069– 1094, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9823-0.

Soshnikov:2000:GFN

[Sos00] Alexander B. Soshnikov. Gaussian fluctuation for the number of particles in Airy, Bessel, sine, and other determinantal random point fields. Journal of Statistical Physics, 100(3–4):491–522, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018672622921.

Soshnikov:2002:NUD

[Sos02] Alexander Soshnikov. A note on universality of the distribution of the largest eigenvalues in certain sample covariance matrices. Journal of Statistical Physics, 108(5–6):1033–1056, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1019739414239.

Soshnikov:2003:JDI

[Sos03] Alexander Soshnikov. Janossy densities. II. Pfaffian ensembles. Journal of Statistical Physics, 113(3–4):611–622, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026077020147.

Sotirov:2007:BSZ

[Sot07] Alexander Sotirov. The boundary structure of zero-temperature driven hard spheres. Journal of Statistical Physics, 126(1): 95–116, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9253-1.

Spohn:2003:BRB

[Spo03] Herbert Spohn. Book review: Quantum Mechanics and Its Emergent Macrophysics. Geoffrey L. Sewell, Princeton University Press, 2002. Journal of Statistical Physics, 111(3–4): 1021–1023, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022866719386.

Spohn:2006:CIP

[Spo06a] Herbert Spohn. Collisional invariants for the phonon Boltzmann equation. Journal of Statistical Physics, 124(5):1131– 1135, September 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9180-1.

Spohn:2006:EPB

[Spo06b] Herbert Spohn. Erratum on ‘The Phonon Boltzmann Equation, Properties and Link to Weakly Anharmonic Lattice Dynamics’. Journal of Statistical Physics, 123(3):707, May 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

006-9144-5. See [Spo06c].

Spohn:2006:PBE

[Spo06c] Herbert Spohn. The phonon Boltzmann equation, properties and link to weakly anharmonic lattice dynamics. Journal of Statistical Physics, 124(2–4):1041–1104, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

005-8088-5. See erratum [Spo06b].

Sondergaard:2000:AHO

[SPVV00] Niels Søndergaard, Gergely Palla, Gábor Vattay, and André Voros. Asymptotics of high order noise corrections. Journal of Statistical Physics, 101(1–2):385–395, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026403314340.

Salas:2000:UAR

[SS00a] Jesús Salas and Alan D. Sokal. Universal amplitude ratios in the critical two-dimensional Ising model on a torus.

Journal of Statistical Physics, 98(3–4):551–588, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018611122166.

Stoyan:2000:RSA

[SS00b] Dietrich Stoyan and Martin Schlather. Random sequential adsorption: Relationship to dead leaves and characterization of variability. Journal of Statistical Physics, 100(5–6):969– 979, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018769422266.

Salas:2001:TMP

[SS01] Jesús Salas and Alan D. Sokal. Transfer matrices and partitionfunction zeros for antiferromagnetic Potts models. I. General theory and square-lattice chromatic polynomial. Journal of Statistical Physics, 104(3–4):609–699, August 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010376605067.

Sanchez:2004:LTD

[SS04] Óscar Sánchez and Juan Soler. Long-time dynamics of the Schrödinger–Poisson–Slater system. Journal of Statistical Physics, 114(1–2):179–204, January 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000003109.97208.53.

Scott:2005:RLG

[SS05] Alexander D. Scott and Alan D. Sokal. The repulsive lattice gas, the independent-set polynomial, and the Lovász local lemma. Journal of Statistical Physics, 118(5–6):1151–1261, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-2055-4.

Sabelfeld:2008:EHP

[SS08] K. K. Sabelfeld and I. A. Shalimova. Elastic half-plane under random displacement excitations on the boundary. Journal of Statistical Physics, 132(6):1071–1095, September 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9597-9.

Sabelfeld:2009:EHP

[SS09a] Karl Sabelfeld and Irina Shalimova. Elastostatics of a halfplane under random boundary excitations. Journal of Statistical Physics, 137(3):521–537, November 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9857-3.

Salas:2009:TMP

[SS09b] Jesús Salas and Alan D. Sokal. Transfer matrices and partitionfunction zeros for antiferromagnetic Potts models. Journal of Statistical Physics, 135(2):279–373, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9725-1.

Sasamoto:2009:SLK

[SS09c] Tomohiro Sasamoto and Herbert Spohn. Superdiffusivity of the 1D lattice Kardar–Parisi–Zhang equation. Journal of Statistical Physics, 137(5–6):917–935, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9831-0.

Sengers:2009:ECE

[SS09d] Jan V. Sengers and Joseph G. Shanks. Experimental criticalexponent values for fluids. Journal of Statistical Physics, 137(5– 6):857–877, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9840-z.

Shalimova:2009:SAE

[SS09e] I. A. Shalimova and K. K. Sabelfeld. Stochastic analysis of an elastic 3D half-space respond to random boundary displacements: Exact results and Karhunen-Loéve expansion. Journal of Statistical Physics, 135(3):547–569, May 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9737-x.

Stoop:2004:CDV

[SSB04] Ruedi Stoop, Norbert Stoop, and Leonid Bunimovich. Complexity of dynamics as variability of predictability. Journal of Statistical Physics, 114(3–4):1127–1137, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012519.93677.15.

Sabhapandit:2000:DAS

[SSD00] Sanjib Sabhapandit, Prabodh Shukla, and Deepak Dhar. Distribution of avalanche sizes in the hysteretic response of the random-field Ising model on a Bethe lattice at zero temperature. Journal of Statistical Physics, 98(1–2):103–129, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1018622805347.

Singer:2005:AEP

[SSE05] A. Singer, Z. Schuss, and R. S. Eisenberg. Attenuation of the electric potential and field in disordered systems. Journal of Statistical Physics, 119(5–6):1397–1418, June 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-3025-1.

Singer:2006:NEPb

[SSH06a] A. Singer, Z. Schuss, and D. Holcman. Narrow escape,Part II: The circular disk. Journal of Statistical Physics, 122(3):465–489, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8027-5.

Singer:2006:NEPc

[SSH06b] A. Singer, Z. Schuss, and D. Holcman. Narrow escape,Part III: Non-smooth domains and Riemann surfaces. Journal of Statistical Physics, 122(3):491–509, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8028-4.

Singer:2006:NEPa

[SSHE06] A. Singer, Z. Schuss, D. Holcman, and R. S. Eisenberg. Narrow escape, Part I. Journal of Statistical Physics, 122(3):437–463, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-8026-6.

Succi:2002:LBS

[SSK02] S. Succi, G. Smith, and E. Kaxiras. Lattice Boltzmann simulation of reactive microflows over catalytic surfaces. Journal of Statistical Physics, 107(1–2):343–366, April 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014531209244.

Scala:2000:ASH

[SSLG^{+}00] A. Scala, M. R. Sadr-Lahijany, N. Giovambattista, S. V. Buldyrev, and H. E. Stanley. Applications of the Stell–Hemmer potential to understanding second critical points in real systems. Journal of Statistical Physics, 100(1–2):97–106, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018631426614.

Schmiedl:2007:EPM

[SSS07] Tim Schmiedl, Thomas Speck, and Udo Seifert. Entropy production for mechanically or chemically driven biomolecules. Journal of Statistical Physics, 128(1–2):77–93, July 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9148-1.

Stillinger:2000:TDE

[SST^{+}00] Dorothea K. Stillinger, Frank H. Stillinger, Salvatore Torquato, Thomas M. Truskett, and Pablo G. Debenedetti. Triangle distribution and equation of state for classical rigid disks. Journal of Statistical Physics, 100(1–2):49–72, July 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018675208867.

Samaj:2000:TPT

[ST00] L. Samaj and I. Travenec. Thermodynamic properties of the two-dimensional two-component plasma. Journal of Statistical Physics, 101(3–4):713–730, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026489924895.

Seta:2002:NSA

[ST02] Takeshi Seta and Ryoichi Takahashi. Numerical stability analysis of FDLBM. Journal of Statistical Physics, 107(1–2):557–572, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1014599729717.

Stein:2004:CBK

[Ste04] D. L. Stein. Critical behavior of the Kramers escape rate in asymmetric classical field theories. Journal of Statistical Physics, 114(5–6):1537–1556, March 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013968.89846.1c.

Struchtrup:2006:SEM

[Str06] Henning Struchtrup. Scaling and expansion of moment equations in kinetic theory. Journal of Statistical Physics, 125(3): 569–591, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9199-3.

Strahov:2008:MKM

[Str08] Eugene Strahov. Matrix kernels for measures on partitions. Journal of Statistical Physics, 133(5):899–919, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9641-9.

Saussol:2002:RDL

[STV02] B. Saussol, S. Troubetzkoy, and S. Vaienti. Recurrence, dimensions, and Lyapunov exponents. Journal of Statistical Physics, 106(3–4):623–634, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013710422755.

Sudbury:2009:RSA

[Sud09] Aidan Sudbury. Random sequential adsorption on random trees. Journal of Statistical Physics, 136(1):51–58, July 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9776-3.

Suidan:2000:ADL

[Sui00] Toufic Suidan. Adhesion dynamics on the line: The mass aggregation process. Journal of Statistical Physics, 101(3–4): 893–903, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026494008966.

Sumedha:2005:RST

[Sum05] Sumedha. Rooted spiral trees on hyper-cubic lattices. Journal of Statistical Physics, 120(1–2):101–123, July 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7773-8.

Suto:2002:EEC

[Süt02] András Sütö. Exact eigenstates for contact interactions. Journal of Statistical Physics, 109(5–6):1051–1072, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020476627111.

Suto:2003:TLP

[Süt03] András Sütö. Thermodynamic limit and proof of condensation for trapped bosons. Journal of Statistical Physics, 112(1–2):375– 396, July 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023696124473.

Suto:2004:NGB

[Süt04]  András Sütö.  Normal and generalized Bose condensa-

tion in traps: One dimensional examples. Journal of Statistical Physics, 117(1–2):301–341, October 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044057.55220.51.

Suwanna:2009:FTA

[Suw09] S. Suwanna. Finite Trotter approximation to the averaged mean square distance in the Anderson model. Journal of Statistical Physics, 136(6):1131–1175, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9817-y.

Suzuki:2001:FRS

[Suz01] J. Suzuki. Functional relations in Stokes multipliers — fun with x^{6} + αx^{2} potential. Journal of Statistical Physics, 102(3–4): 1029–1047, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004823608260.

Suzuki:2003:MAC

[Suz03] Masuo Suzuki. Methodology of analytic and computational studies on quantum systems. Journal of Statistical Physics, 110(3– 6):945–956, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022128423037.

Szasz:2007:LLR

[SV07] Domokos Szász and Tamás Varjú. Limit laws and recurrence for the planar Lorentz process with infinite horizon. Journal of Statistical Physics, 129(1):59–80, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9367-0.

Sinai:2009:SSQ

[SV09] Y. G. Sinai and I. Vinogradov. Separating solution of a quadratic recurrent equation. Journal of Statistical Physics, 136 (4):603–613, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9801-6.

Snoeijer:2002:FRM

[SvL02] Jacco H. Snoeijer and J. M. J. van Leeuwen. Force relaxation in the q-model for granular media. Journal of Statistical Physics, 109(3–4):449–469, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020442026840.

Salmhofer:2000:PCF

[SW00] Manfred Salmhofer and Christian Wieczerkowski. Positivity and convergence in fermionic quantum field theory. Journal of Statistical Physics, 99(1–2):557–586, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018661110470.

Sunden:2008:BAM

[SW08] Mattias Sundén and Bernt Wennberg. Brownian approximation and Monte Carlo simulation of the non-cutoff Kac equation. Journal of Statistical Physics, 130(2):295–312, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9424-8.

Swendsen:2002:SMC

[Swe02] Robert H. Swendsen. Statistical mechanics of classical systems with distinguishable particles. Journal of Statistical Physics, 107 (5–6):1143–1166, June 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015161825292.

Swendsen:2004:RNC

[Swe04] Robert H. Swendsen. Response to Nagle’s criticism of my proposed definition of the entropy. Journal of Statistical Physics, 117(5–6):1063–1070, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-5716-4.

Straube:2007:RRS

[SWF07] Ronny Straube, Michael J. Ward, and Martin Falcke. Reaction rate of small diffusing molecules on a cylindrical membrane. Journal of Statistical Physics, 129(2):377–405, October 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9371-4.

Samaj:2004:TSB

[SWK04] L. Samaj, J. Wagner, and P. Kalinay. Translation symmetry breaking in the one-component plasma on the cylinder. Journal of Statistical Physics, 117(1–2):159–178, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044066.98352.12.

Schmidt:2009:NEW

[SWK09] John R. Schmidt, Jost O. L. Wendt, and Alan R. Kerstein. Non-equilibrium wall deposition of inertial particles in turbulent flow. Journal of Statistical Physics, 137(2):233–257, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9844-8.pdf.

Seiler:2003:CBQ

[SY03] Erhard Seiler and Karim Yildirim. Critical behavior in a quasi d dimensional spin model. Journal of Statistical Physics, 112(3– 4):457–495, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023815706708.

Seiringer:2008:GSE

[SY08] Robert Seiringer and Jun Yin. Ground state energy of the low density Hubbard model. Journal of Statistical Physics, 131 (6):1139–1154, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9527-x.

Spohn:2003:LTB

[SZ03] H. Spohn and E. Zhizhina. Long-time behavior for the 1-D stochastic Ising model with unbounded random couplings. Journal of Statistical Physics, 111(1–2):419–431, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022225612366.

Shlosman:2004:MT

[SZ04] Senya Shlosman and Valentin Zagrebnov. Magnetostriction transition. Journal of Statistical Physics, 114(3–4):563–574, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000012502.75889.09.

Szabo:2006:BRB

[Sza06] Gábor Szabó. Book review: Evolution and Structure of the Internet: A Statistical Physics Approach. Journal of Statistical Physics, 122(6):1297–1298, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-006-9060-8.

Tahmasbi:2009:PWS

[Tah09] Nader Tahmasbi. Partial-wave scattering and statistical mechanics via the l-wave non-local separable potential of rank-two. Journal of Statistical Physics, 136(5):989–1003, September 2009.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-009-9815-0.

Takata:2009:SLBa

[Tak09a] Shigeru Takata. Symmetry of the linearized Boltzmann equation and its application. Journal of Statistical Physics, 136 (4):751–784, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9793-2. See comments

[Sha10].

Takata:2009:SLBb

[Tak09b] Shigeru Takata. Symmetry of the linearized Boltzmann equation II. Journal of Statistical Physics, 136(5):945–983, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9805-2.

Takata:2010:RCS

[Tak10] Shigeru Takata. Reply to the comments of Sharipov on ‘Symmetry of the Linearized Boltzmann Equation’. Journal of Statistical Physics, 139(3):538–539, May 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

010-9953-4. See [Tak09a, Tak09b, Sha10].

Taguchi:2006:VFA

[TAL06] Satoshi Taguchi, Kazuo Aoki, and Vladimir Latocha. Vapor flows along a plane condensed phase with weak condensation in the presence of a noncondensable gas. Journal of Statistical Physics, 124(2–4):321–369, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-0001-8.

Talagrand:2007:LDG

[Tal07] Michel Talagrand. Large deviations, Guerra’s and A.S.S. schemes, and the Parisi Hypothesis. Journal of Statistical Physics, 126(4–5):837–894, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-006-9108-9.

Tao:2001:PIA

[Tao01] R. Tao. Path-integral approach to the statistical physics of onedimensional random systems. Journal of Statistical Physics, 103 (3–4):575–588, May 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1010393300220.

Taqqu:2004:ERB

[Taq04] Murad S. Taqqu. An essay and review of the book: Self Similar Processes. Paul Embrechts and Makoto Maejima, Princeton University Press, 2003. Journal of Statistical Physics, 114(3–4): 1171–1177, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000012705.64638.ae.

Tasaki:2006:CRD

[Tas06] Hal Tasaki. The coefficient of restitution does not exceed unity. Journal of Statistical Physics, 123(6):1361–1374, June 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9129-4.

Toninelli:2004:DAT

[TB04] Cristina Toninelli and Giulio Biroli. Dynamical arrest, tracer diffusion and kinetically constrained lattice gases. Journal of Statistical Physics, 117(1–2):27–54, October 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044063.86539.19.

Touchette:2006:NEM

[TB06] Hugo Touchette and Christian Beck. Nonconcave entropies in multifractals and the thermodynamic formalism. Journal of Statistical Physics, 125(2):455–471, October 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9174-z.

Toninelli:2007:JPG

[TB07] Cristina Toninelli and Giulio Biroli. Jamming percolation and glassy dynamics. Journal of Statistical Physics, 126(4– 5):731–763, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9177-9.

Toninelli:2008:NCC

[TB08] Cristina Toninelli and Giulio Biroli. A new class of cellular automata with a discontinuous glass transition. Journal of Statistical Physics, 130(1):83–112, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9420-z.

Toninelli:2005:CBK

[TBF05] Cristina Toninelli, Giulio Biroli, and Daniel S. Fisher. Cooperative behavior of kinetically constrained lattice gas models of glassy dynamics. Journal of Statistical Physics, 120(1–2):167– 238, July 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-005-5250-z.

Tutschka:2003:OFE

[TC03] Christian Tutschka and José A. Cuesta. Overcomplete free energy functional for D = 1 particle systems with next neighbor interactions. Journal of Statistical Physics, 111(5–6):1125–1148, June 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023096031180.

Taniguchi:2007:OMT

[TC07] Tooru Taniguchi and E. G. D. Cohen. Onsager–Machlup theory for nonequilibrium steady states and fluctuation theorems. Journal of Statistical Physics, 126(1):1–41, January 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9252-2.

Taniguchi:2008:IEN

[TC08a] Tooru Taniguchi and E. G. D. Cohen. Inertial effects in nonequilibrium work fluctuations by a path integral approach. Journal of Statistical Physics, 130(1):1–26, January 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9398-6.

Taniguchi:2008:NSS

[TC08b] Tooru Taniguchi and E. G. D. Cohen. Nonequilibrium steady state thermodynamics and fluctuations for stochastic systems. Journal of Statistical Physics, 130(4):633–667, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9471-1.

Taniguchi:2002:LSP

[TDM02] Tooru Taniguchi, Carl P. Dettmann, and Gary P. Morriss. Lyapunov spectra of periodic orbits for a many-particle system. Journal of Statistical Physics, 109(3–4):747–764, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020422917270.

Tuzel:2000:DDS

[TE00] Erkan Tüzel and Ayse Erzan. Dissipative dynamics and the statistics of energy states of a Hookean model for protein folding. Journal of Statistical Physics, 100(1–2):405–422, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018616417953.

Tellez:2001:TDC

[Tél01] Gabriel Téllez. Two-dimensional Coulomb systems in a disk with ideal dielectric boundaries. Journal of Statistical Physics, 104(5–6):945–970, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010493409399.

Telcs:2006:ERR

[Tel06a] András Telcs. The Einstein relation for random walks on graphs. Journal of Statistical Physics, 122(4):617–645, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-8002-1.

Tellez:2006:GCE

[Tél06b] Gabriel Téllez. Guest charges in an electrolyte: Renormalized charge, long- and short-distance behavior of the electric potential and density profiles. Journal of Statistical Physics, 122(4): 787–798, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8069-8.

Tellez:2007:ESF

[Tél07] Gabriel Téllez. Equation of state in the fugacity format for the two-dimensional Coulomb gas. Journal of Statistical Physics, 126(2):281–298, January 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9251-3.

Tasaki:2000:EPT

[TG00] S. Tasaki and P. Gaspard. Entropy production and transports in a conservative multibaker map with energy. Journal of Statistical Physics, 101(1–2):125–144, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1026443028452.

Tasaki:2002:SPP

[TG02] S. Tasaki and P. Gaspard. Spectral properties of a piecewise linear intermittent map. Journal of Statistical Physics, 109(3– 4):803–820, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020479002249.

Tian:2004:LSR

[Tia04] Guang-Shan Tian. Lieb’s spin-reflection-positivity method and its applications to strongly correlated electron systems. Journal of Statistical Physics, 116(1–4):629–680, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037214.70064.78.

Tidriri:2001:NMS

[Tid01a] M. Tidriri. New models for the solution of intermediate regimes in transport theory and radiative transfer: Existence theory, positivity, asymptotic analysis, and approximations. Journal of Statistical Physics, 104(1–2):291–325, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010365812733.

Tidriri:2001:RDA

[Tid01b] M. Tidriri. Rigorous derivation and analysis of coupling of kinetic equations and their hydrodynamic limits for a simplified Boltzmann model. Journal of Statistical Physics, 104(1–2):255– 290, July 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010313828663.

Tidriri:2004:HLB

[Tid04] M. Tidriri. Hydrodynamic limit of a B.G.K. like model on domains with boundaries and analysis of kinetic boundary conditions for scalar multidimensional conservation laws. Journal of Statistical Physics, 115(5–6):1715–1754, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028079.51072.06.

Tindel:2003:QLD

[Tin03] Samy Tindel. Quenched large deviation principle for the overlap of a p-spins system. Journal of Statistical Physics, 110(1– 2):51–72, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021062510565.

Taguchi:2008:TSP

[TJ08] Satoshi Taguchi and Ansgar Jüngel. A two-surface problem of the electron flow in a semiconductor on the basis of kinetic theory. Journal of Statistical Physics, 130(2):313–342, January 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9426-6.

Tsutahara:2002:DET

[TK02] Michihisa Tsutahara and Ho Keun Kang. A discrete effect of the thermal lattice BGK model. Journal of Statistical Physics, 107 (1–2):479–498, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014591527900.

Tolke:2002:MSD

[TKR02] Jonas Tölke, Manfred Krafczyk, and Ernst Rank. A multigridsolver for the discrete Boltzmann equation. Journal of Statistical Physics, 107(1–2):573–591, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014551813787.

Tellez:2002:SME

[TM02] Gabriel Téllez and Lina Merchán. Solvable model for electrolytic soap films: The two-dimensional two-component plasma. Journal of Statistical Physics, 108(3–4):495–525, August 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015725807258.

Taillefumier:2008:HLC

[TM08] Thibaud Taillefumier and Marcelo O. Magnasco. A Haar-like construction for the Ornstein Uhlenbeck process. Journal of Statistical Physics, 132(2):397–415, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9545-8.

Tanase-Nicola:2004:MST

[TNK04] Sorin Tanase-Nicola and Jorge Kurchan. Metastable states, transitions, basins and borders at finite temperatures. Journal of Statistical Physics, 116(5–6):1201–1245, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000041739.53068.6a.

Toninelli:2007:CPF

[Ton07] Fabio Lucio Toninelli. Critical properties and finite-size estimates for the depinning transition of directed random polymers. Journal of Statistical Physics, 126(4–5):1025–1044, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9123-x.

Toninelli:2008:ABB

[Ton08a] Fabio Toninelli. Anton Bovier: Statistical mechanics of disordered systems. A mathematical perspective. Journal of Statistical Physics, 132(5):957–958, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9581-4.

Toninelli:2008:GGB

[Ton08b] Fabio Toninelli. Giambattista Giacomin: Random Polymer Models. Journal of Statistical Physics, 130(6):1219–1220, March 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9478-7.

Toom:2000:AUE

[Too00] André Toom. Algorithmical unsolvability of the ergodicity problem for locally interacting processes with continuous time. Journal of Statistical Physics, 98(1–2):495–501, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018699527637.

Toom:2002:LIR

[Too02] André Toom. On large isolated regions in supercritical percolation. Journal of Statistical Physics, 109(5–6):1091–1108, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020480728020.

Toom:2004:NED

[Too04] André Toom. Non-ergodicity in a 1-D particle process with variable length. Journal of Statistical Physics, 115(3–4):895–924, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOSS.0000022371.44066.f6.

Toom:2007:ECO

[Too07] André Toom. Every continuous operator has an invariant measure. Journal of Statistical Physics, 129(3):555–566, November 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9407-9.

Toom:2008:LLN

[Too08] André Toom. Law of large numbers for non-local functions on probabilistic cellular automata. Journal of Statistical Physics, 133(5):883–897, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9643-7.

Toral:2004:NLR

[Tor04] Raúl Toral. On the nonextensivity of the long range X–Y model. Journal of Statistical Physics, 114(5–6):1393–1398, March 2004.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000013963.16180.a3.

Toldin:2009:SDP

[TPV09] Francesco Parisen Toldin, Andrea Pelissetto, and Ettore Vicari. Strong-disorder paramagnetic–ferromagnetic fixed point in the square-lattice ±J Ising model. Journal of Statistical Physics, 135 (5–6):1039–1061, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9705-5.pdf.

Turaev:2003:SBC

[TRK03] D. Turaev and V. Rom-Kedar. Soft billiards with corners. Journal of Statistical Physics, 112(3–4):765–813, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023884227180.

Tij:2004:PFH

[TS04] Mohamed Tij and Andrés Santos. Poiseuille flow in a heated granular gas. Journal of Statistical Physics, 117(5–6):901– 928, December 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-004-5710-x.

Torres:2005:GCF

[TT05] Aldemar Torres and Gabriel Téllez. General considerations on the finite-size corrections for Coulomb systems in the Debye-Hückel regime. Journal of Statistical Physics, 118(3–4):735–765, February 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8827-z.

Takahashi:2001:SHF

[TTK01] Sarato Takahashi, Alex Yu. Tretyakov, and Norio Konno. On some Harris–FKG type correlation inequalities for a non– attractive model. Journal of Statistical Physics, 102(5–6): 1429–1438, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1004808916380.

Tij:2001:NCF

[TTM^{+}01] M. Tij, E. E. Tahiri, J. M. Montanero, V. Garzó, A. Santos, and J. W. Dufty. Nonlinear Couette flow in a low density granular gas. Journal of Statistical Physics, 103(5–6):1035–1068, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010317207358.

Tailleur:2006:KES

[TTNK06] Julien Tailleur, Sorin Tanase-Nicola, and Jorge Kurchan. Kramers equation and supersymmetry. Journal of Statistical Physics, 122(4):557–595, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-005-8059-x.

Tumulka:2006:RVG

[Tum06] Roderich Tumulka. A relativistic version of the Ghirardi– Rimini–Weber model. Journal of Statistical Physics, 125(4): 821–840, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9227-3.

Turova:2003:LPC

[Tur03] Tatyana S. Turova. Long paths and cycles in dynamical graphs. Journal of Statistical Physics, 110(1–2):385–417, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021035131946.

Turova:2006:PTD

[Tur06] Tatyana S. Turova. Phase transitions in dynamical random graphs. Journal of Statistical Physics, 123(5):1007–1032, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9101-3.

Toscani:2000:TES

[TV00] G. Toscani and C. Villani. On the trend to equilibrium for some dissipative systems with slowly increasing a priori bounds. Journal of Statistical Physics, 98(5–6):1279–1309, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018623930325.

Toth:2002:BEF

[TV02] Bálint Tóth and Benedek Valkó. Between equilibrium fluctuations and Eulerian scaling: Perturbation of equilibrium for a class of deposition models. Journal of Statistical Physics, 109(1– 2):177–205, October 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019987628197.

Toth:2003:ORE

[TV03] Bálint Tóth and Benedek Valkó. Onsager relations and Eulerian hydrodynamic limit for systems with several conservation laws. Journal of Statistical Physics, 112(3–4):497–521, August 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023867723546.

Tzeng:2003:DSQ

[TW03] W.-J. Tzeng and F. Y. Wu. Dimers on a simple-quartic net with a vacancy. Journal of Statistical Physics, 110(3–6):671–689, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022155701655.

Tracy:2008:FDR

[TW08] Craig A. Tracy and Harold Widom. A Fredholm determinant representation in ASEP. Journal of Statistical Physics, 132 (2):291–300, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9562-7.

Tracy:2009:ASB

[TW09] Craig A. Tracy and Harold Widom. On ASEP with step Bernoulli initial condition. Journal of Statistical Physics, 137(5– 6):825–838, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-009-9867-1.pdf.

Ugawa:2007:EHE

[UC07] Hideaki Ugawa and Patricio Cordero. Extended hydrodynamics from Enskog’s equation for a two-dimensional system general formalism. Journal of Statistical Physics, 127(2):339–358, April 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-006-9257-x.

Uchiyama:2004:EFZ

[Uch04] Kôhei Uchiyama. Equilibrium fluctuations for zero-range-exclusion processes. Journal of Statistical Physics, 115(5–6): 1423–1460, June 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000028065.88090.af.

Ueltschi:2004:SAH

[Uel04] Daniel Ueltschi. Segregation in the asymmetric Hubbard model. Journal of Statistical Physics, 116(1–4):681–697, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037231.88815.04.

Umantsev:2009:TST

[Uma09] Alex Umantsev. Thermodynamic stability of transition states in nanosystems. Journal of Statistical Physics, 136(1):117–130, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-009-9765-6.

Urrutia:2008:THS

[Urr08] Ignacio Urrutia. Two hard spheres in a spherical pore: Exact analytic results in two and three dimensions. Journal of Statistical Physics, 131(4):597–611, May 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9513-3.

Valdinoci:2000:ENR

[Val00] Enrico Valdinoci. Estimates for non-resonant normal forms in Hamiltonian perturbation theory. Journal of Statistical Physics, 101(3–4):905–919, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026428203514.

Valageas:2009:SSP

[Val09a] Patrick Valageas. Some statistical properties of the Burgers equation with white-noise initial velocity. Journal of Statistical Physics, 137(4):729–764, November 2009. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9809-y.

Valageas:2009:SPB

[Val09b] Patrick Valageas. Statistical properties of the Burgers equation with Brownian initial velocity. Journal of Statistical Physics, 134 (3):589–640, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9685-5.

Varandas:2008:CDR

[Var08a] Paulo Varandas. Correlation decay and recurrence asymptotics for some robust nonuniformly hyperbolic maps. Journal of Statistical Physics, 133(5):813–839, December 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9639-3.

Vargas:2008:RIQ

[Var08b] Rodrigo Vargas. Repeated interaction quantum systems: Van Hove limits and asymptotic states. Journal of Statistical Physics, 133(3):491–511, November 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9605-0.

vanBeijeren:2003:UTR

[vB03] Henk van Beijeren. The uphill turtle race; on short time nucleation probabilities. Journal of Statistical Physics, 110(3–6): 1397–1410, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022121616193.

Volchenkov:2007:NDT

[VB07] D. Volchenkov and Ph. Blanchard. Nonlinear diffusion through large complex networks containing regular subgraphs. Journal of Statistical Physics, 127(4):677–697, May 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9313-1.

Volchenkov:2008:MCM

[VB08] D. Volchenkov and P. Blanchard. Markov chain methods for analyzing urban networks. Journal of Statistical Physics, 132(6):

1051–1069, September 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9591-2.

vanBeijeren:2002:NRP

[vBD02] H. van Beijeren and J. R. Dorfman. A note on the Ruelle pressure for a dilute disordered Sinai billiard. Journal of Statistical Physics, 108(5–6):767–785, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1019860925625.

Volkov:2006:NES

[VBM06] Igor Volkov, Jayanth R. Banavar, and Amos Maritan. A novel ensemble in statistical physics. Journal of Statistical Physics, 123(1):167–180, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9032-z.

Vannimenus:2001:SMI

[VD01] Jean Vannimenus and Bernard Derrida. A solvable model of interface depinning in random media. Journal of Statistical Physics, 105(1–2):1–23, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012278408260.

Vidigal:2005:ABO

[VD05] Ronaldo Vidigal and Ronald Dickman. Asymptotic behavior of the order parameter in a stochastic sandpile. Journal of Statistical Physics, 118(1–2):1–25, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-8775-7.

vandenEsker:2008:UDF

[vdEvdHH08] Henri van den Esker, Remco van der Hofstad, and Gerard Hooghiemstra. Universality for the distance in finite variance random graphs. Journal of Statistical Physics, 133(1):169–202, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-008-9594-z.pdf.

vanderHofstad:2004:IIC

[vdHJ04] Remco van der Hofstad and Antal A. Járai. The incipient infinite cluster for high-dimensional unoriented percolation. Journal of Statistical Physics, 114(3–4):625–663, February 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000012505.39213.6a.

vanderHofstad:2001:SOD

[vdHK01] Remco van der Hofstad and Wolfgang König. A survey of one-dimensional random polymers. Journal of Statistical Physics, 103(5–6):915–944, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1010309005541.

vanderHofstad:2008:PTR

[vdHK08] Remco van der Hofstad and Wouter Kager. Pattern theorems, ratio limit theorems and Gumbel maximal clusters for random fields. Journal of Statistical Physics, 130(3):503–522, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-007-9435-5.

vanderHofstad:2002:CAR

[vdHKK02] Remco van der Hofstad, Achim Klenke, and Wolfgang König. The critical attractive random polymer in dimension one. Journal of Statistical Physics, 106(3–4):477–520, February 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013750004100.

Hofstad:2006:MCN

[vDHR06] Remco van Der Hofstad and Frank Redig. Maximal clusters in non-critical percolation and related models. Journal of Statistical Physics, 122(4):671–703, February 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-8012-z.

vanEnter:2000:RNR

[vE00] Aernout C. D. van Enter. A remark on the notion of robust phase transitions. Journal of Statistical Physics, 98(5–6):1409–1416, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018632132142.

vanEnter:2007:TCB

[vEK07] Aernout C. D. van Enter and Christof Külske. Two connections between random systems and non–Gibbsian measures. Journal of Statistical Physics, 126(4–5):1007–1024, March 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-006-9185-9.pdf.

Velazquez:2000:ESF

[Vel00] J. J. L. Velázquez. On the effect of stochastic fluctuations in the dynamics of the Lifshitz–Slyozov–Wagner model. Journal of Statistical Physics, 99(1–2):57–113, April 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018640505926.

Velazquez:2008:SSB

[Vel08] J. J. L. Velázquez. Self-similar behaviour for noncompactly supported solutions of the LSW model. Journal of Statistical Physics, 130(4):757–799, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9456-0.

vanEnter:2005:IMR

[vENS05] A. C. D. van. Enter, K. Netocný, and H. G. Schaap. On the Ising model with random boundary condition. Journal of Statistical Physics, 118(5–6):997–1056, March 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-004-2138-2.

vanEnter:2007:CTD

[vER07] A. C. D. van Enter and W. M. Ruszel. Chaotic temperature dependence at zero temperature. Journal of Statistical Physics, 127(3):567–573, May 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-006-9260-2.pdf.

Vanden-Eijnden:2008:RES

[VEW08] Eric Vanden-Eijnden and Maria G. Westdickenberg. Rare events in stochastic partial differential equations on large spatial domains. Journal of Statistical Physics, 131(6):1023–1038, June 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-008-9537-8.

Vincent:2009:RSG

[VHO09] E. Vincent, J. Hammann, and M. Ocio. Real spin glasses relax slowly in the shade of hierarchical trees. Journal of Statistical Physics, 135(5–6):1105–1120, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9730-4.

Villani:2006:ICS

[Vil06a] Cédric Villani. Introduction for Carlo’s special issue. Journal of Statistical Physics, 124(2–4):271–273, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9118-7.

Villani:2006:MGM

[Vil06b] Cédric Villani. Mathematics of granular materials. Journal of Statistical Physics, 124(2–4):781–822, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9038-6.

Vincenzi:2002:KKD

[Vin02] D. Vincenzi. The Kraichnan–Kazantsev dynamo. Journal of Statistical Physics, 106(5–6):1073–1091, March 2002. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014089820881.

vanKampen:2002:RMO

[vK02] N. G. van Kampen. The road from molecules to Onsager. Journal of Statistical Physics, 109(3–4):471–481, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020494010910.

vanKampen:2004:NAN

[vK04] N. G. van Kampen. A new approach to noise in quantum mechanics. Journal of Statistical Physics, 115(3–4):1057–1072, May 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000022383.06086.6c.

Valli:2002:SWF

[VKVT02] A. Valli, A. Koponen, T. Vesala, and J. Timonen. Simulations of water flow through bordered pits of conifer xylem. Journal of Statistical Physics, 107(1–2):121–142, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014554419722.

Veerman:2005:FF

[VLCW05] J. J. P. Veerman, G. Lafferriere, J. S. Caughman, and A. Williams. Flocks and formations. Journal of Statistical Physics, 121(5–6):901–936, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-6999-9.

Vollmer:2002:ERF

[VMT02] Jürgen Vollmer, László Mátyás, and Tamás Tél. Escape-rate formalism, decay to steady states, and divergences in the entropyproduction rate. Journal of Statistical Physics, 109(3–4):875– 893, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020483103158.

Visco:2006:FPI

[VPB^{+}06] Paolo Visco, Andrea Puglisi, Alain Barrat, Emmanuel Trizac, and Frédéric van Wijland. Fluctuations of power injection in randomly driven granular gases. Journal of Statistical Physics, 125 (3):533–568, November 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9161-4.

vanRensburg:2001:TI

[vR01] E. J. Janse van Rensburg. Trees at an interface. Journal of Statistical Physics, 102(5–6):1177–1209, March 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004884110928.

vanRensburg:2001:ACD

[vRR01] E. J. Janse van Rensburg and A. Rechnitzer. Adsorbing and collapsing directed animals. Journal of Statistical Physics, 105 (1–2):49–91, October 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012225909169.

vonSolms:2000:MP

[vSC00] N. von Solms and Y. C. Chiew. A model for polyelectrolytes. Journal of Statistical Physics, 100(1–2):267–277, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018652031157.

Veerman:2009:ATF

[VST09] J. J. P. Veerman, B. D. Stosić, and F. M. Tangerman. Automated traffic and the finite size resonance. Journal of Statistical Physics, 137(1):189–203, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9816-z.

Vollmer:2000:MTE

[VTM00] Jürgen Vollmer, Tamás Tél, and László Mátyás. Modeling thermostating, entropy currents, and cross effects by dynamical systems. Journal of Statistical Physics, 101(1–2):79–105, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1026499113431.

vanWonderen:2000:VLM

[vWL00] A. J. van Wonderen and K. Lendi. Virtues and limitations of Markovian master equations with a time-dependent generator. Journal of Statistical Physics, 100(3–4):633–658, August 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018671424739.

vanWijland:2005:LDW

[vWR05] Frédéric van Wijland and Zoltán Rácz. Large deviations in weakly interacting boundary driven lattice gases. Journal of Statistical Physics, 118(1–2):27–54, January 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-8776-6.

vanZon:2006:TDS

[vZC06] R. van Zon and E. G. D. Cohen. Theorem on the distribution of short-time particle displacements with physical applications.

Journal of Statistical Physics, 123(1):1–37, April 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9027-9.

vanZon:2002:FPT

[vZvB02] R. van Zon and Henk van Beijeren. Front propagation techniques to calculate the largest Lyapunov exponent of dilute hard disk gases. Journal of Statistical Physics, 109(3–4):641–669, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020414615453.

Wreszinski:2009:PFT

[WA09] Walter F. Wreszinski and Elcio Abdalla. A precise formulation of the Third Law of Thermodynamics. Journal of Statistical Physics, 134(4):781–792, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-009-9693-5.

Wagner:2000:LIH

[Wag00] Christoph Wagner. Lyapunov instability for a hard-disk fluid in equilibrium and nonequilibrium thermostated by deterministic scattering. Journal of Statistical Physics, 98(3–4):723–742, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018675408962.

Warnaar:2001:RTC

[War01] S. Ole Warnaar. Refined q-trinomial coefficients and character identities. Journal of Statistical Physics, 102(3–4):1065–1081, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004827709169.

Watanabe:2004:THS

[Wat04] Hiroshi Watanabe. Triviality of hierarchical O ( N ) spin model in four dimensions with large N. Journal of Statistical Physics, 115(5–6):1669–1713, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028073.49354.9d.

Wreszinski:2004:SAO

[WB04] Walter F. Wreszinski and Oscar Bolina. A self-averaging ‘order parameter’ for the Sherrington–Kirkpatrick spin glass model. Journal of Statistical Physics, 116(5–6):1389–1404, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000041743.24497.63.

Wakou:2002:TLG

[WBE02] J. Wakou, R. Brito, and M. H. Ernst. Towards a Landau– Ginzburg-type theory for granular fluids. Journal of Statistical Physics, 107(1–2):3–22, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014590000158.

Wedagedera:2001:PDS

[WD01] J. R. Wedagedera and T. C. Dorlas. The phase diagram of a spin glass model. Journal of Statistical Physics, 103(5–6):697–716, June 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1010340417836.

Widom:2005:RTH

[WDMB05] M. Widom, N. Destainville, R. Mosseri, and F. Bailly. Random tilings of high symmetry: II. Boundary conditions and numerical studies. Journal of Statistical Physics, 120(5–6): 837–873, September 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-6998-x.

Wang:2007:MCM

[WE07] Hongyun Wang and Timothy C. Elston. Mathematical and computational methods for studying energy Transduction in protein motors. Journal of Statistical Physics, 128(1–2):35–76, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9169-9.

Weeks:2003:EFD

[Wee03] John D. Weeks. External fields, density functionals, and the Gibbs inequality. Journal of Statistical Physics, 110(3–6): 1209–1218, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022157229397.

Wehr:1997:SLL

[Weh97] Jan Wehr. A strong law of large numbers for iterated functions of independent random variables. Journal of Statistical Physics, 86(5–6):1373–1384, March 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02183629. See er-

ratum [Ano01d].

Weiss:2001:BRB

[Wei01] George H. Weiss. Book review: Diffusion and Reaction in Fractals and Disordered Systems. Daniel ben-Avraham and Shlomo Havlin. Cambridge University Press, Cambridge England, 2000. Journal of Statistical Physics, 104(5–6):1407–1408, September 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1010430517577.

Weiss:2002:BRB

[Wei02] George Weiss. Book review: Brownian Motion: Fluctuations, Dynamics, and Applications. Robert M. Mazo, Clarendon Press, 2002. Journal of Statistical Physics, 109(5–6):1123–1124, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020436912999.

Weiss:2005:BRB

[Wei05] George H. Weiss. Book review: Elements of the Random Walk: An Introduction for Advanced Students and Researchers. Journal of Statistical Physics, 119(1–2):449–451, April 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
004-1994-0.

Weiss:2007:FBP

[Wei07] George H. Weiss. Fractal-based point processes. Journal of Statistical Physics, 127(6):1287–1288, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-007-9328-7.

Weiss:2009:RKG

[Wei09] George H. Weiss. R. Klages, G. Radons, I.M. Sokolov (eds.): Anomalous Transport: Foundations and Applications. Journal of Statistical Physics, 135(2):389–391, April 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9713-5.

West:2004:BRB

[Wes04] Bruce J. West. Book review: Brownian Agents and Active Particles. F. Schweitzer, Springer, Berlin, 2003. Journal of Statistical Physics, 117(1–2):381–383, October 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000044205.26242.f8.

West:2009:BPT

[Wes09] Bruce J. West. Benoît perthame: Transport equations in biology. Journal of Statistical Physics, 134(3):641–642, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9686-4.

Willemsen:2002:MSP

[WHI02] S. M. Willemsen, H. C. J. Hoefsloot, and P. D. Iedema. Mesoscopic simulation of polymers in fluid dynamics problems. Journal of Statistical Physics, 107(1–2):53–65, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014546217905.

Widom:2004:AAP

[Wid04] Harold Widom. On asymptotics for the Airy process. Journal of Statistical Physics, 115(3–4):1129–1134, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022384.58696.61.

Wiese:2000:PPP

[Wie00] Kay Jörg Wiese. The passive polymer problem. Journal of Statistical Physics, 101(3–4):843–891, November 2000. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026473504422.

Winkel:2002:LCI

[Win02] Matthias Winkel. Limit clusters in the inviscid Burgers turbulence with certain random initial velocities. Journal of Statistical Physics, 107(3–4):893–917, May 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014598400004.

Witte:2003:BRB

[Wit03] F. M. C. Witte. Book review: Path Integrals in Quantum Mechanics, Statistics, Polymer Physics and Financial Markets. Prof. Dr. Hagen Kleinert, 3rd extended edition, World Scientific Publishing, Singapore. Journal of Statistical Physics, 111 (1–2):497–499, April 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022237915092.

Wojtkiewicz:2000:DET

[WK00] Jacek Wojtkiewicz and Andreas Klümper. Distribution of energies for the two-dimensional Ising model. Journal of Statistical Physics, 98(5–6):1063–1073, March 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018659510760.

Wu:2004:OEV

[WK04] F. Y. Wu and H. Kunz. The odd eight-vertex model. Journal of Statistical Physics, 116(1–4):67–78, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000037206.47155.58.

Wen:2007:SSF

[WLTH07] Qi Wen, Guanglai Li, Jay X. Tang, and Greg Huber. Switching statistics of a flagellar motor: First-passage time and dynamic binding. Journal of Statistical Physics, 128(1–2):257–267, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-006-9249-x.

Wilkinson:2007:SDR

[WM07] Michael Wilkinson and Bernhard Mehlig. Strength distribution of repeatedly broken chains. Journal of Statistical Physics, 127 (6):1279–1286, June 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-007-9320-2.

Widom:2002:AOT

[WMDB02] M. Widom, R. Mosseri, N. Destainville, and F. Bailly. Arctic octahedron in three-dimensional rhombus tilings and related integer solid partitions. Journal of Statistical Physics, 109(5–6): 945–965, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020464224385.

Wadati:2001:SNS

[WNUK01] Miki Wadati, Akinori Nishino, Hideaki Ujino, and Yasushi Komori. Symmetric and non-symmetric bases of quantum integrable particle systems with long-range interactions. Journal of Statistical Physics, 102(3–4):1049–1064, February 2001. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1004875625099.

Wojtkiewicz:2003:SGS

[Woj03] Jacek Wojtkiewicz. Stability of ground states of 2d strongly asymmetric correlated–hopping Hubbard model. Journal of Statistical Physics, 112(5–6):1127–1151, September 2003. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1024667609435.

Wojtkiewicz:2006:PDT

[Woj06] Jacek Wojtkiewicz. Phase diagram of the two-dimensional t– t′ Falicov–Kimball model. Journal of Statistical Physics, 123 (3):585–600, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9054-6.

Wojtkiewicz:2009:EFE

[Woj09] Jacek Wojtkiewicz. Estimations of the free energy for the Hubbard model. Journal of Statistical Physics, 135(2):375–387, April 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9720-6.

Wolf:2006:GPS

[Wol06] Christian Wolf. Generalized physical and SRB measures for hyperbolic diffeomorphisms. Journal of Statistical Physics, 122(6): 1111–1138, March 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8024-8.

Wagner:2002:LEB

[WP02] Alexander J. Wagner and Ignacio Pagonabarraga. Lees–Edwards boundary conditions for lattice Boltzmann. Journal of Statistical Physics, 107(1–2):521–537, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014595628808.

Wu:2003:BPS

[WP03] Xian-Yuan Wu and S. Yu. Popov. On AB bond percolation on the square lattice and AB site percolation on its line graph. Journal of Statistical Physics, 110(1–2):443–449, January 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021091316925.

Wagner:2002:RAO

[WS02a] C. Wagner and R. Stoop. Renormalization approach to optimal limiter control in 1-D chaotic systems. Journal of Statistical Physics, 106(1–2):97–107, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013120112236.

Wang:2002:TMM

[WS02b] Jian-Sheng Wang and Robert H. Swendsen. Transition matrix Monte Carlo method. Journal of Statistical Physics, 106(1–2): 245–285, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013180330892.

Wu:2000:IMH

[Wu00] C. Chris Wu. Ising models on hyperbolic graphs II. Journal of Statistical Physics, 100(5–6):893–904, September 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018763008810.

Wu:2002:ZTD

[Wu02] C. Chris Wu. Zero-temperature dynamics of Ising models on the triangular lattice. Journal of Statistical Physics, 106(1–2): 369–373, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013140616779.

Wu:2009:SGS

[Wu09] Zhigang Wu. Stability of global solution for the relativistic Enskog equation near vacuum. Journal of Statistical Physics, 137 (1):149–164, October 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9848-4.

Wehr:2001:CLT

[WW01] Jan Wehr and Jung M. Woo. Central limit theorems for nonlinear hierarchical sequences of random variables. Journal of Statistical Physics, 104(3–4):777–797, August 2001. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1010384806884.

Widom:2005:BRB

[WW05] A. Widom and F. Y. Wu. Book review: Lectures on the Kinetic Theory of Gases, Non-equilibrium Thermodynamics and Statistical Theories. Journal of Statistical Physics, 119(3–4):945–948, May 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-2112-z.

Wennberg:2006:KET

[WW06] B. Wennberg and Y. Wondmagegne. The Kac equation with a thermostatted force field. Journal of Statistical Physics, 124(2– 4):859–880, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-9020-8.

Wehr:2006:SLW

[WX06] Jan Wehr and Jack Xin. Scaling limits of waves in convex scalar conservation laws under random initial perturbations. Journal of Statistical Physics, 122(2):361–370, January 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-8006-x.

Wang:2009:LTA

[WZ09] W.-M. Wang and Zhifei Zhang. Long time Anderson localization for the nonlinear random Schrödinger equation. Journal of Statistical Physics, 134(5–6):953–968, March 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
008-9649-1.

Xing:2009:TGS

[Xin09] Xiangjun Xing. Topology and geometry of smectic order on compact curved substrates. Journal of Statistical Physics, 134 (3):487–536, February 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9681-9.

Xu:2008:EDM

[XKHK08] Ying Xu, Hosho Katsura, Takaaki Hirano, and Vladimir E. Korepin. Entanglement and density matrix of a block of spins in AKLT model. Journal of Statistical Physics, 133(2):347–377, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-008-9617-9.

Yan:2003:HIM

[Yan03] Jun Yan. Harmonic interaction model and its applications in Bose–Einstein condensation. Journal of Statistical Physics, 113(3–4):623–634, November 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026029104217.

Yarotsky:2005:UGS

[Yar05] D. A. Yarotsky. Uniqueness of the ground state in weak perturbations of non-interacting gapped quantum lattice systems. Journal of Statistical Physics, 118(1–2):119–144, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-004-8780-x.

Yarotsky:2008:RWA

[Yar08] D. A. Yarotsky. Random walk analysis of the commensurate– incommensurate transition in the isotropic spin-1 chain. Journal of Statistical Physics, 130(5):957–981, March 2008. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
007-9458-y.

Yong:2006:URW

[YBMS06] Zhang Yong, David A. Benson, Mark M. Meerschaert, and Hans-Peter Scheffler. On using random walks to solve the spacefractional advection–dispersion equations. Journal of Statistical Physics, 123(1):89–110, April 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9042-x.

Yi:2007:MYC

[YCCN07] Tau-Mu Yi, Shanqin Chen, Ching-Shan Chou, and Qing Nie. Modeling yeast cell polarization induced by pheromone gradients. Journal of Statistical Physics, 128(1–2):193–207, July 2007. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9285-1.

Yepez:2002:QLG

[Yep02] Jeffrey Yepez. Quantum lattice-gas model for the Burgers equation. Journal of Statistical Physics, 107(1–2):203–224, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1014514805610.

Yukhimets:2000:MFH

[YFP00] A. Yukhimets, H. L. Frisch, and J. K. Percus. Molecular fluids at high dimensionality. Journal of Statistical Physics, 100 (1–2):135–151, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018635527522.

Yamamoto:2002:SCF

[YHD02] Kazuhiro Yamamoto, Xiaoyi He, and Gary D. Doolen. Simulation of combustion field with lattice Boltzmann method. Journal of Statistical Physics, 107(1–2):367–383, April 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014583226083.

Yip:2006:SMS

[Yip06] Sidney Yip. Soft-mode scenarios of shear localization: Atomiclevel landscapes. Journal of Statistical Physics, 125(5–6):1109– 1120, December 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9030-1.

Yngvason:2004:ELA

[Yng04] Jakob Yngvason. Elliott Lieb and the art of mathematical physics. Journal of Statistical Physics, 116(1–4):13–16, August 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000037242.87441.1c.

Yoo:2007:VPD

[Yoo07] Hyun Jae Yoo. A variational principle in the dual pair of reproducing kernel Hilbert spaces and an application. Journal of Statistical Physics, 126(2):325–354, January 2007. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9258-9.

Yoshida:2008:PTG

[Yos08] Nobuo Yoshida. Phase transitions for the growth rate of linear stochastic evolutions. Journal of Statistical Physics, 133(6): 1033–1058, December 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-008-9646-4.

Young:2002:WSM

[You02] Lai-Sang Young. What are SRB measures, and which dynamical systems have them? Journal of Statistical Physics, 108(5–6): 733–754, September 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1019762724717.

Yethiraj:2000:IET

[YS00] Arun Yethiraj and George Stell. An integral equation theory for the Widom–Rowlinson mixture. Journal of Statistical Physics, 100(1–2):39–47, July 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018623224797.

Yakhot:2005:ASS

[YS05] Victor Yakhot and Katepalli R. Sreenivasan. Anomalous scaling of structure functions and dynamic constraints on turbulence simulations. Journal of Statistical Physics, 121(5–6): 823–841, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8666-6.

Yu:2006:DGF

[Yu06] Shih-Hsien Yu. The development of the Green’s function for the Boltzmann equation. Journal of Statistical Physics, 124(2– 4):301–320, August 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-006-9064-4.

Yau:2009:SOU

[YY09] Horng-Tzer Yau and Jun Yin. The second order upper bound for the ground energy of a Bose gas. Journal of Statistical Physics, 136(3):453–503, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9792-3.

Zaburdaev:2006:RWM

[Zab06] Vasily Yu. Zaburdaev. Random walk model with waiting times depending on the preceding jump length. Journal of Statistical Physics, 123(4):871–881, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-006-9104-0.

Zaburdaev:2008:MAR

[Zab08] Vasily Yu. Zaburdaev. Microscopic approach to random walks. Journal of Statistical Physics, 133(1):159–167, October 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-008-9598-8.

Zahran:2002:OFF

[Zah02] M. A. Zahran. 1/2-Order fractional Fokker–Planck equation on comblike model. Journal of Statistical Physics, 109(5–6): 1005–1016, December 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020420409364.

Zhu:2009:SPS

[ZF09] Ping Zhu and Yang Fu. Stationary properties and stochastic resonance for a saturation laser model with cross-correlation between quantum noise terms. Journal of Statistical Physics, 136(1):131–143, July 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-009-9773-6.

Zheng:2002:CKT

[ZGA02] Yihao Zheng, Alejandro L. Garcia, and Berni J. Alder. Comparison of kinetic theory and hydrodynamics for Poiseuille flow. Journal of Statistical Physics, 109(3–4):495–505, November 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020498111819.

Zhang:2000:SAB

[Zha00a] Jian Zhang. Stability of attractive Bose–Einstein condensates. Journal of Statistical Physics, 101(3–4):731–746, November 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1026437923987.

Zhang:2000:CBM

[Zha00b] Yu Zhang. Critical behavior for maximal flows on the cubic lattice. Journal of Statistical Physics, 98(3–4):799–811, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1018631726709.

Zhang:2008:GEC

[Zha08] Xianwen Zhang. Global existence of classical solutions to the Fokker–Planck-BGK equation. Journal of Statistical Physics, 132(3):535–550, August 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9546-7.

Zhang:2009:EMM

[Zha09] Yunxin Zhang. The efficiency of molecular motors. Journal of Statistical Physics, 134(4):669–679, February 2009. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
009-9695-3.

Zhizhina:2000:LTR

[Zhi00] Elena Zhizhina. The Lifshitz tail and relaxation to equilibrium in the one-dimensional disordered Ising model. Journal of Statistical Physics, 98(3–4):701–721, February 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018623424891.

Zinn-Justin:2000:DPM

[ZJ00] P. Zinn-Justin. The dilute Potts model on random surfaces. Journal of Statistical Physics, 98(1–2):245–264, January 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1018626906256.

Zhdanov:2000:SRK

[ZK00] V. P. Zhdanov and B. Kasemo. Surface restructuring, kinetic oscillations, and chaos in heterogeneous catalytic reactions. Journal of Statistical Physics, 101(1–2):631–647, October 2000. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1026470307071.

Zaliapin:2003:BDEa

[ZKBG03a] Ilya Zaliapin, Vladimir Keilis-Borok, and Michael Ghil. A Boolean delay equation model of colliding cascades. Part I: Multiple seismic regimes. Journal of Statistical Physics, 111 (3–4):815–837, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022850215752.

Zaliapin:2003:BDEb

[ZKBG03b] Ilya Zaliapin, Vladimir Keilis-Borok, and Michael Ghil. A Boolean delay equation model of colliding Cascades.Part II: Prediction of critical transitions. Journal of Statistical Physics, 111 (3–4):839–861, May 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022802432590.

Zuk:2005:EBH

[ZKD05] Or Zuk, Ido Kanter, and Eytan Domany. The entropy of a binary hidden Markov process. Journal of Statistical Physics, 121(3–4): 343–360, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-7576-y.

Zhang:2005:GFM

[ZQS^{+}05] Y. H. Zhang, R. S. Qin, Y. H. Sun, R. W. Barber, and D. R. Emerson. Gas flow in microchannels — a lattice Boltzmann method approach. Journal of Statistical Physics, 121(1–2): 257–267, October 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-005-8416-9.

Zamponi:2004:FEP

[ZRA04] F. Zamponi, G. Ruocco, and L. Angelani. Fluctuations of entropy production in the isokinetic ensemble. Journal of Statistical Physics, 115(5–6):1655–1668, June 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000028072.34588.32.

Zielen:2002:EMF

[ZS02] Frank Zielen and Andreas Schadschneider. Exact mean-field solutions of the asymmetric random average process. Journal of Statistical Physics, 106(1–2):173–185, January 2002. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013145026048.

Zuckerman:2004:SFS

[ZW04] Daniel M. Zuckerman and Thomas B. Woolf. Systematic finite-sampling inaccuracy in free energy differences and other nonlinear quantities. Journal of Statistical Physics, 114(5–6): 1303–1323, March 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOSS.0000013961.84860.5b.

Zweifel:2005:MPM

[Zwe05] P. F. Zweifel. The mathematical physics of music. Journal of Statistical Physics, 121(5–6):1097–1104, December 2005. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
005-7581-1.

Zhang:2006:SMS

[ZZ06] Xicheng Zhang and Xianwen Zhang. Supports of measure solutions for spatially homogeneous Boltzmann equations. Journal of Statistical Physics, 124(2–4):485–495, August 2006. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
006-9067-1.